A Bibliography of Papers in Lecture Notes in Computer Science (2012): Volumes 6121–7125

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.01

Title word cross-reference

(Δ + 1) [1577]. (ρ, G) [266]. (r|p) [781]. 1 [1022]. 1^{st} [1342]. 2 [27, 1294, 1138, 432, 1028, 281, 758, 272, 1440, 546, 861, 867, 1352, 578, 561]. 3 [579, 1293, 1381, 176, 1355, 1623, 1294, 1012, 1358, 341, 1370, 1028, 157, 160, 978, 1440, 861, 1385, 279, 995, 1340, 1400, 1433, 1352, 173, 1295, 1343, 1560, 1409, 662]. 4 [1349]. [0, 1]^{d} [660]. ^{+} [204]. ^{2} [608, 1012]. ^{3} [1012, 622]. _{p} [647]. A∗ [1264]. B [623]. β [217]. C^{1} [673]. C^{2} [656]. ℓ_{0} [268]. ϵ [324, 1470]. G^{2} [649]. GM(1, 1) [536]. H_{∞} [392]. K [1026, 909, 1433, 1516, 930, 1033]. L_{1} [673]. μ [1709]. p [526, 240, 1089]. P_{0} [103]. q [683]. R [297, 1012]. ρ [1643, 1626]. τ [522].

-Algorithm [1626]. -AntWars [758]. -Based [662]. -Bernstein-Type [683]. -Centroid [781]. -Coloring [1577]. -Cycles [526]. -D [867, 561]. -elastica [240]. -Equivariant [266]. -Induction [930]. -Matrix [103]. -Optimal [324].

-Planarity [1022]. -Quasi-Planar [1033]. -Root [1026]. -Space [1433]. -Trees [1012]. -Values [1089].

/or [97].

066313 [1619].

1 [1804]. 1-Form [662]. 10th [1823]. 12th [1866]. 13th [1849, 1835, 1836]. 15th [1854]. 19th [1843].

2.0 [1224]. 2008-2010 [1840]. 2009 [1681]. 2010 [1319, 1685]. 2011 [1008]. 27th [1815]. 2DPCA [455].

31- [1818].

4 [1408]. 454 [585]. 4th [1870, 1830, 1818, 1837, 1847].

5th [1856, 1857].

65nm [1140]. 6th [1870, 1864].

7th [1833, 1821, 1832, 1860, 1863, 1826, 1827, 1828, 1848, 1869].

8th [1821, 1822, 1867].

9th [1838].

A4MMC [1816]. AAM [1377]. AAMAS [1851, 1865]. ABC [574]. ABC-Based [574]. Abdominal [987, 1001, 978, 995, 994, 1004, 1000, 1842]. Abelian [845]. Ablation [989]. Abnormal [1359]. above [1586, 1595]. ABS [925, 922, 923]. Abstract [1650, 1065, 922, 57]. Abstraction [747]. ACC [464]. Accelerate [761, 1732]. Accelerated [1074, 1019, 171, 1376]. Accelerating [1634, 22, 177]. Accelerator [30, 26]. Access [4, 1058, 1423, 621, 1213, 622, 1451, 1794, 623, 620, 1084, 55, 1632]. Access-Control [621, 623]. Account [1829]. Accuracy [767]. Accurate [166, 105, 811, 111, 788, 840]. Accurately [516]. Achieving [40]. Acid [559, 563]. ACO [1668, 782]. ACO-Algorithm [1668]. Acoustic [453]. Acquisition [1104, 1456]. Across [1484, 1418]. ACSI [1630]. Action [81, 266, 200, 122, 831, 123, 131, 126, 207, 148, 133, 1492, 339]. Action-Rule [200]. Actions [128, 1722, 717, 1841]. Activation [1303, 1304]. Active [4, 1300, 429, 247, 352, 121, 257]. Active-Training [1300]. Activities [1229, 1271, 127]. Activity [701, 462, 278]. Actor [428]. Actor-Critic [428]. Acyclic [1044]. Ad [1863, 1580, 1072, 873]. Adaptable [1457]. Adaptation [205, 1274, 488, 1232, 1453, 1237, 1726, 1750]. Adapting [1253, 1175].

Adaption [698, 1310]. Adaptive [242, 1303, 1304, 784, 1265, 245, 1437, 715, 463, 1161, 351, 659, 489, 664, 1176, 64, 1159, 764, 243, 1215, 1726, 1397, 697, 1468, 1155, 752, 434, 353, 317, 1411, 1865]. Adaptively [346]. AdaRank [369]. Adding [239]. Additive [1648]. Addressing [1065]. Adipose [999]. Adjacent [1040, 810, 1414]. Adjustable [1569]. Adjustment [176, 1370, 1382, 165]. ADMI [1860]. Admits [1039]. Advanced [1637, 1851, 1826, 1827]. Advances [1017, 1845, 856, 804, 691, 1856, 1857, 1847]. Aerial [163]. Aerodynamics [1673]. AES [1140]. AETROS [868]. Affine [987, 997, 134]. affinity [146]. Affordance [1171]. Affordance-Based [1171]. against [614, 1798, 1172, 515]. Age [352, 461, 1550]. Ageing [1450, 1452]. Agent [1160, 1174, 208, 1481, 191, 1493, 1183, 1181, 1822, 1483, 1177, 1186, 1616, 1851, 1849, 1250, 1195, 1173, 1804, 1252, 1164, 203, 1271, 1502, 1272, 22, 188, 468, 1809, 1482, 1811, 1484, 1180, 857, 198, 1251, 1249, 207, 187, 1269, 1158, 1168, 1275, 1278, 1479, 1498, 1273, 1167, 1192, 1871, 1165, 637, 1492, 1187, 1151, 1810, 1806, 1485, 1196, 1179, 1480, 1178, 1267, 1185, 184, 1612, 1248, 1258, 1807, 433]. Agent-Based [1160, 1181, 1483, 1177, 1250, 1195, 1271, 1272, 22, 188, 1809, 1180, 857, 1251, 1249, 187, 1269, 1278, 1479, 1498, 1273, 1167, 1165, 1187, 1151, 1810, 1806, 1485]. Agent-Enhanced [1482]. Agent-Environment [208]. Agent-Oriented [1186, 203]. Agents [1190, 1860, 1615, 1480, 968, 194, 1488, 1491, 749, 831, 1176, 1184, 1191, 1194, 1253, 1812, 193, 515, 1188, 1865, 1731, 201, 1279, 1821, 1860]. AgentStore [1252]. Agglomeration [1643]. Aggregates [1466]. Aggregating [1784]. Aggregation [1643, 790]. Aggregations [1708]. Aging [471]. Agreements [1782]. AI [1115]. Aided [1835, 1836, 975, 198, 838, 1292, 473]. AIME [1834]. Air [1680, 1685]. Aircraft [658]. al [1786, 717]. ALA [1865]. Alert [1453]. Alerts [1027]. Algebra [107]. Algebraic [644, 650, 685, 1818]. Algebras [710]. Algorithm [350, 475, 1666, 1362, 1668, 1489, 242, 333, 412, 260, 761, 60, 318, 554, 428, 411, 371, 526, 1437, 471, 806, 1138, 429, 388, 351, 413, 409, 349, 335, 763, 556, 7, 1589, 420, 90, 830, 538, 439, 1719, 461, 91, 470, 575, 387, 574, 469, 410, 1344, 523, 279, 1376, 529, 577, 815, 1095, 327, 1094, 376, 358, 1674, 802, 1675, 771, 1626, 386, 326, 492, 496, 1689, 552, 415, 476, 564, 406, 572, 385, 558, 402, 1264, 408, 348, 570, 380, 462, 457, 384, 438, 571, 1409, 1651, 522, 367, 390, 332, 527, 518, 555, 346, 405, 314, 1427, 893, 1552]. Algorithmic [701, 924]. Algorithmics [1606]. Algorithms [1568, 258, 534, 826, 92, 778, 1863, 1527, 300, 491, 33, 528, 240, 813, 864, 466, 1595, 1574, 315, 188, 773, 531, 757, 95, 786, 167, 769, 1530, 1863]. Algorithms- [258]. ALGOSENSORS [1863]. Aliasing [1710]. aligned [1010]. Alignment [90, 1109, 564]. Allocation [414, 185, 1170, 634, 193, 1157]. Alternate [810]. Alternating [504]. Alternations [1106]. Alternative [585, 83, 722, 1157]. Altruistic [182]. Alzheimer [569]. AMAS [1816]. AMAS-BT [1816]. Ambient [1168]. Ambiguity [907, 906]. Ambrosio [248]. Ameliorating [536]. American [1718]. AMG [1643]. AMGe [1643]. Amherst [1848]. Amino [559, 563].

Amoeba [257]. AMPLE [1851]. Amplifier [810]. Amyloidosis [717]. Analog [1141]. Analyses [1389, 840, 1004]. Analysing [1385]. Analysis [1143, 1665, 1336, 849, 720, 1061, 585, 366, 1708, 954, 1678, 692, 1570, 1483, 586, 1091, 942, 1186, 1520, 1540, 736, 778, 135, 794, 1491, 1709, 1445, 1250, 569, 1388, 909, 661, 281, 1141, 1494, 1134, 891, 1620, 1377, 785, 106, 487, 401, 997, 484, 486, 310, 1301, 271, 948, 1772, 1454, 762, 139, 844, 863, 1404, 1622, 1683, 77, 996, 85, 920, 1317, 1095, 780, 1123, 703, 714, 679, 804, 68, 691, 394, 1062, 511, 1346, 1342, 1352, 1730, 1086, 359, 1295, 1711, 1545, 578, 769, 399, 1419, 561, 562, 1485, 467, 363, 1366, 568, 497, 319, 943]. Analytic [1501]. Analytical [739, 1234]. Analyze [1271]. Analyzed [1589]. Analyzer [89]. Analyzing [125, 99, 92, 219, 883, 440, 112]. Anaphora [57]. Anaphoricity [542]. Anatomical [1357, 1295, 993]. Anatomy [1515, 996]. ANB [1818]. Ancient [539]. Android [1255, 1279, 878, 1451]. Angiogenic [220]. Angle [1022]. Angles [1050, 1023]. Anisotropic [245, 231]. Annotation [918, 217, 1210]. Annotation-Based [918]. Annotations [14, 694, 1110]. Anomaly [1293]. Anonymity [5]. Anonymization [1515]. Anonymous [942, 1748, 1057]. Answer [399]. Answering [451, 58, 1547]. Ant [1672, 549, 555]. Antennae [1571]. Anti [446]. Anti-Windup [446]. AntWars [758]. AnySURF [1259]. Ao [1168]. AOP [1116]. Aortic [836]. AOSE [1851, 1255]. AP2PC [1821]. API [213]. APIs [1239]. Apparent [1001]. Appearance [352, 134]. Appliances [1184]. Application [11, 258, 646, 266, 828, 1233, 1593, 932, 212, 882, 240, 887, 834, 1116, 819, 1753, 864, 1519, 1372, 839, 777, 501, 670, 1360, 673, 1633, 1314, 262, 703, 1192, 1311, 1225, 1545]. Applications [1218, 1389, 1336, 847, 1061, 1650, 827, 799, 1233, 1202, 1219, 1255, 1279, 1080, 12, 82, 1827, 1828, 1866, 1204, 1240, 1554, 947, 184, 1231, 295, 139, 1839, 34, 1237, 1673, 1458, 1224, 1226, 28, 1652, 1222, 683, 1273, 1328, 1062, 1842, 1651, 1220, 518, 419]. Applied [1861, 1345, 1670, 946, 836]. Applying [1130, 1455, 775, 399]. Approach [1766, 1759, 192, 585, 366, 1152, 296, 1074, 412, 313, 1480, 156, 233, 60, 1499, 1186, 1265, 1212, 1219, 986, 798, 822, 1460, 1546, 1077, 1252, 1613, 567, 123, 1494, 1213, 1611, 1189, 9, 1799, 785, 86, 1525, 1722, 1682, 1200, 459, 215, 1771, 573, 727, 811, 1326, 1215, 1335, 275, 1223, 78, 1126, 88, 679, 682, 867, 1082, 1356, 164, 1092, 1529, 96, 1469, 898, 173, 1193, 1509, 591, 1534, 638, 1325, 590, 1399, 249, 543]. Approaches [1490, 1018, 1294, 1169, 1339, 775, 112, 236]. Appropriate [1541]. Approximate [1042, 968, 1024, 1466]. Approximating [650, 663]. Approximation [1568, 1715, 1644, 795, 1599, 1648, 787, 390]. April [1839, 1685]. Arbitrary [1582]. Arc [642, 1037]. Architects [41]. Architectural [1114]. Architecture [1277, 1232, 1471, 445, 466, 1408, 26, 878, 1783, 899, 1463, 1775, 382, 1816, 858, 19, 1457, 1455]. Architectures [1792, 1176, 693, 1151]. Archive [1049, 763, 771]. Arcs [1036, 650]. Area [1023, 536, 1072]. ARIRANG [1622]. Arising [1360, 1693]. Arithmetic [1771]. ARMS [1851, 805]. Arrangement [1404]. Array [212, 1674]. Array-Oriented [212]. Arrhythmia [572]. Art [370, 73, 783]. Articulated [171, 281]. Artifact [974, 1412, 1428]. Artifact-Free [974, 1412]. Artificial [1669, 579, 416, 1827, 575, 524, 405, 314, 730]. Ascending [1172]. Ascending-Price [1172]. Asian [1718]. ASL [139]. ASLan [915]. ASM [328, 512]. ASP [112]. ASP-Based [112]. Aspect [1117, 701, 1220]. Aspect-Oriented [1117, 1220]. Aspects [1833, 1679, 877, 819, 1827, 1087, 740, 1841]. Aspiration [1175]. Assembly [585]. Assessing [1762]. Assessment [481, 548, 1736, 735, 220, 17, 19, 1187, 457, 454, 377, 378]. Assignment [834, 1182]. Assignments [1808]. Assist [842]. Assistance [1059]. Assisted [444, 121]. ASSL [1075]. Association [1481, 9, 1103, 1536, 517, 1528, 522, 518]. associative [599]. Assumption [150]. Asymmetric [434]. Asymptotic [1657]. Asynchronous [630, 383]. Athena [301]. Atlas [1284, 1359, 995]. Atlases [1855]. Atmosphere [1678]. Atomic [1702, 204]. Atrial [1364]. Attack [614, 1620, 1087, 612, 375]. Attacks [1140, 917, 1621, 1622, 1088, 1619, 96, 533]. Attention [505, 1393]. Attribute [566, 905, 100, 116, 115, 117, 1163]. Attribute-Based [116]. Attribute-TID [100]. Attributes [118, 100, 119]. Auction [1172]. Auctions [1263, 1494, 86]. Audio [1377]. Audio-Visual [1377]. Augmented [238, 877]. Augmenting [1074, 1255]. August [1818, 1826, 1827, 1828, 1866]. Austria [1838, 1818]. Authenticated [1138]. Authentication [1148, 872, 1784, 1146, 1633, 1786]. Authoring [1751]. Authority [440]. Authorization [734, 1777]. Auto [1274, 1793]. Auto-Adaptation [1274]. Auto-delegation [1793]. Autogenetic [821]. Automata [1766, 695, 324, 632, 31, 704, 750]. Automata-Based [750]. Automate [856]. Automated [1061, 1324, 1102, 1179, 716, 607, 1077, 926, 947, 1454, 981, 983, 858, 754, 999]. Automatic [1144, 1790, 163, 1511, 758, 90, 1348, 1364, 1742, 279, 780, 121, 1686, 359, 1343, 161, 814, 167, 372, 1456]. Automatically [1321]. Automating [1239]. Automation [713]. Automotive [1336]. Autonomic [1845, 1076, 1071, 1075]. Autonomous [1488, 1863, 1870, 873, 1478]. AUTOPIA [856]. AUTOSAR [1336]. Auxiliary [1626]. Availability [1322, 1793]. AVC [1408, 1392, 1409, 1427]. AVC/H.264 [1408]. Average [1594, 1595, 1575]. Averting [1175]. Avian [1192]. Avignon [1832]. Avoidance [1263]. Aware [1568, 1276, 1746, 191, 303, 49, 1077, 1184, 1237, 1215, 15, 1457, 620, 1073, 1785, 693, 175]. Awareness [1799, 1446]. Axis [641, 642, 1386]. Axon [1410]. azzurro [151].

B [1695, 99, 1697, 1699, 661, 1125, 921, 927, 98]. B-Spline [1699]. B-Splines [1695, 1697, 661]. Back [20, 50, 971, 1006, 1055, 1069, 1078, 1111, 1135, 1149, 1197, 1241, 1281, 1296, 1315, 1337, 1367, 1405, 1442, 1464, 1475, 1503, 1555, 1564, 1584, 1607, 1617, 1640, 101, 1733, 1757, 1778, 1802, 1813, 113, 142, 180, 195, 209, 224, 292, 307, 403, 499, 594, 603, 625, 639, 689, 711, 725, 807, 889, 913, 935, 949, 956]. Background [1440, 1376, 1319, 474]. Bag [1390]. Bag-of-Textons [1390].

Bagging [1552]. Bags [1293, 1288]. Balanced [318, 321]. Balanced-Sampling-Based [321]. Balancing [1754]. Balkan [311]. Balloons [1023]. Ban [1247]. Band [150, 342]. Band-Limited [342]. Bands [810]. Bandwidth [732]. Bang [1661]. Bank [789]. Bar [381]. Barcodes [562]. Bare [1081]. Bare-Bones [1081]. Barrier [1576]. Barycenter [262]. Base [1511, 753, 713]. Based [159, 1199, 258, 368, 1160, 1535, 645, 213, 800, 1144, 1284, 84, 1489, 928, 1148, 791, 918, 1205, 1746, 253, 370, 1708, 1650, 1490, 1227, 1196, 1293, 260, 313, 761, 75, 1741, 1124, 1480, 318, 992, 611, 554, 428, 1635, 344, 1181, 1100, 1483, 301, 1291, 1118, 700, 942, 347, 1177, 1435, 1283, 287, 736, 1355, 526, 458, 1294, 136, 471, 135, 357, 1212, 1559, 806, 129, 1235, 715, 1128, 1131, 908, 1709, 291, 822, 171, 341, 576, 336, 1369, 1250, 1195, 1370, 749, 1546, 340, 351, 172, 1288, 1106, 592, 451, 951, 349, 1211, 862, 1382, 1058, 1645, 631]. Based [1631, 567, 362, 662, 766, 455, 445, 73, 770, 984, 123, 1537, 489, 440, 538, 1142, 1329, 364, 1156, 439, 466, 832, 1210, 1271, 352, 1719, 1377, 61, 461, 1333, 1793, 1502, 56, 1272, 22, 1624, 470, 536, 1451, 574, 401, 452, 589, 1417, 1440, 1391, 321, 514, 469, 188, 320, 792, 767, 271, 1432, 468, 985, 1809, 1659, 1724, 670, 1115, 1117, 762, 1811, 1852, 337, 1326, 1374, 1166, 66, 14, 1484, 1421, 1180, 1518, 897, 126, 1414, 1253, 995, 1439, 994, 857, 1627, 1407, 74, 899, 1287, 1251, 1249, 193, 259, 920, 1317, 1412, 1104, 1463, 818, 1339, 524, 788]. Based [606, 1797, 376, 840, 62, 358, 678, 112, 1652, 802, 116, 46, 187, 1269, 1775, 1122, 1158, 789, 1278, 39, 1438, 79, 1171, 1479, 1498, 1092, 165, 298, 1433, 437, 1273, 1157, 1328, 1637, 431, 1099, 750, 1413, 1167, 96, 1371, 511, 1756, 415, 1346, 1469, 1201, 482, 898, 1352, 117, 312, 70, 1086, 1467, 1165, 1711, 1193, 383, 591, 331, 385, 396, 1492, 460, 581, 348, 559, 570, 525, 1534, 380, 325, 434, 638, 769, 1187, 1151, 1325, 169, 1810, 353, 354, 448, 1558, 400, 587, 1528, 571, 1651, 1806, 561, 562, 446, 522, 367, 1485, 588, 1632, 454, 377, 378]. Based [467, 363, 495, 329, 346, 493, 568, 314, 494, 497, 1427, 397, 533, 835, 1871]. Baseline [1395]. Bases [1570, 1696]. Basic [412, 1767, 1032]. Basis [1410, 851, 1614, 683]. Basket [785]. Batch [1665, 327, 740, 390]. Batched [1206]. Battery [766]. Bayesian [290, 1613, 1417, 998, 1550]. BDI [1253]. BE [23]. Beacon [444]. Become [1110]. Bee [416, 575, 524, 405]. Behavior [509, 1161, 1613, 1267, 1506, 1272, 139, 1512, 1192]. Behavioral [810, 922]. Behaviour [1483, 1805, 742, 684]. Belgium [1870]. Belief [964, 1192]. Beltrami [1372, 241, 237]. BEMD [347, 360]. Benchmark [136]. Benchmarking [498]. Bend [1019]. Benefit [34]. Benefits [1251]. Bernstein [654, 683]. Berth [414]. Beta [1695, 1699]. Beta-Function [1695, 1699]. Between [1812, 1247, 1696, 194, 1559, 487, 1449, 94, 450, 1541]. Beverage [1399]. Beyond [1085, 1606, 53, 787]. Bézier [643, 653, 654]. BFO [577]. Bhoomi [82]. Bi [146, 1134, 398]. Bi-affinity [146]. Bi-criteria [1134]. Bi-direction [398]. Bias [1360]. Biclustering [93, 1542]. Bidirectional [911]. Bifocal [1424]. Bifurcation [251]. Bilateral [146, 1412]. BiLingva [1124]. Bin [784]. Bin-Packing [784]. Binary [849, 1767, 489, 1517, 1772, 1414, 345, 531, 31, 456, 32, 250, 510]. Binding [218]. Bio [806, 1312, 1828]. Bio-Inspired [806, 1828]. Bio-physiological [1312]. Bioinformatic [223]. Bioinformatical [779]. Bioinformatics [93, 295, 298, 1823]. Biological [299, 1195, 106]. Biologically [1074]. Biologically-Inspired [1074]. Biology [1818]. Biomedical [1285]. Biometric [1635]. Biomimetic [463, 833, 730]. BioPax [216]. Bioprocess [1657, 1675, 740]. Biosensor [1635]. BIP [1759]. Bipartite [1098, 1027, 510]. BIRCH [1518]. BIRCH-Based [1518]. Bisimulation [1614]. Bits [1620]. Bitwise [1634]. Bitwise-Sharing [1634]. Bivariate [654]. Black [618, 892]. Black-Box [618, 892]. Blades [658]. Blast [777]. Bled [1834]. Blending [1695]. Blind [425, 1428, 568]. Blinding [611]. Block [1717, 1623, 607, 1137, 1621, 612, 79, 571]. Block-Preconditioners [1717]. Blocker [789]. Blocking [416]. Blocks [36, 1724]. Blog [1508]. Blogger [1508]. Blogger-Link-Topic [1508]. Blue [1680, 151, 1732]. Blur [239]. BNCOD [1815, 2, 3]. board [815]. Body [84, 1635]. Bold [1039]. Bond [829, 1730]. Bones [1081]. Book [1737]. Bookmarking [300, 1169]. Boolean [843]. Boosted [456]. Boosting [434]. Bootstrap [678]. Bootstrap-Based [678]. Born [1760]. Boronizing [501]. Both [491, 1343]. BOTTARI [1741]. Bottleneck [43]. Bottlenecks [30]. Bound [1035]. Boundaries [642, 1415]. Boundary [1688, 1645, 682]. Bounded [613, 632, 1266, 796, 1127]. Bounding [1143, 1142]. Bounds [1597, 1602, 1586, 1594, 1662]. Box [263, 618, 892]. Boxes [750]. BP [313]. Braced [1722]. Braga [1837]. Brain [1410, 569, 568, 835]. Branch [955]. Breast [566, 486]. Breathing [978]. Bright [144]. Brinkman [1693]. British [1815]. Broadband [22]. Broadcast [1579, 98, 476]. Broadcasting [1633]. Brownian [278]. Browsing [1451]. BT [1816]. Budapest [1821]. Buffer [15]. Bug [1774]. Bug-Finding [1774]. Building [794, 1631, 36, 477, 861, 1287, 713, 1676, 437, 167, 754]. Built [1727]. Bulgaria [1681, 1867, 1679]. Bundle [176, 1382, 282, 165]. Bundles [1021]. Bundling [1020]. BUPLE [1145]. Burden [37]. Bus [445]. Business [1177, 1116, 321, 1125]. Bytecode [703]. Byzantine [98].

C [1774, 896]. C# [1133]. Cable [1722]. Cable-Braced [1722]. Cadastral [163]. Cahn [1717]. Calculation [716]. Calculus [692, 291, 946, 597, 601, 598, 600, 599, 602, 748]. Calibration [1716, 772, 147, 825, 1422]. Call [694]. Calls [1453]. Camera [166, 147, 1083, 477, 279, 1404, 178]. Cameras [881, 1376, 124, 398]. Can [34, 28, 581, 1151]. Canada [1855, 1822, 1830, 1852, 1842]. Canada-France [1830]. Canaria [1835, 1836, 869]. Cancellation [948]. Cancer [566, 109, 770, 486, 841, 517, 581]. Cancer-Related [517]. Capabilities [1133, 810]. Capacitated [764]. Capacity [852, 1582, 1583]. Capsule [977]. Capture [832]. Capturing [1066]. Car [1661]. Car-Like [1661]. Cardiac [1362, 1365, 1358, 341, 1359, 1354, 1349, 1360, 1363, 1356, 1347, 1342, 1352, 1351, 1343, 1353]. Cards [608]. Care [1764, 1546, 1834]. Careflow [718]. Caregiver [472]. Carlo [1702, 1704]. Carriers [871]. CARTO [1339].

Carving [154]. Cascade [434]. Case [108, 1483, 1247, 930, 1736, 82, 1244, 1322, 1502, 1170, 1180, 1224, 717, 1786, 1638, 1684, 884, 903, 48]. CAT [498]. Catalytic [495]. Categorial [596, 597, 1829, 602]. Categorization [1100, 177]. Category [117]. cathartic [981]. Catheter [1352]. Catheter-Based [1352]. Catheterisation [1352]. Causal [1229, 668]. CBIR [1291]. CCC [322]. CCTS [1753]. CDMA [464]. CDS [1852]. CEE [1847]. CEE-SET [1847]. Cell [884, 330, 24, 23, 22]. Cells [1195, 552]. Cellular [750]. Center [36]. Centered [1859]. Central [1847]. Centralization [1153]. Centre [77]. Centric [299, 1671, 1067, 1114, 1125]. Centroid [781, 830]. cerevisiae [1665]. Certainty [1531]. Certified [741]. Cervical [1002]. Chain [816, 1320]. Chains [432]. Challenge [190, 1339, 1258, 1347, 1342]. Challenger [1053]. Challenges [306, 1780, 1061, 1855, 1334, 691, 1062]. Championship [327]. Change [1679, 1491, 583, 1117, 752, 967]. Changes [1519]. Changing [1123]. Channel [1141, 144]. Channels [615, 912]. Chaos [556, 802]. Chaos-Based [802]. Chaotic [393, 376, 506]. Characteristics [35, 45]. Characterization [1291]. Charging [766]. Check [1379]. Checking [797, 932, 930, 934, 945, 632, 1799, 697, 638, 1127]. Chemical [627, 387, 884]. Chemical-Reaction [387]. Chemotherapy [841]. China [1861, 1826, 1827, 1828, 1453]. Chinese [372, 1539]. Chip [49, 39]. Choice [795, 1161, 1266, 733]. Chordal [680]. Chosen [426]. Chromaticity [150]. Chromatography [564]. Chronic [1341]. Cine [1358, 1348]. Cipher [1623, 607, 1621]. Ciphers [1137, 612]. Ciphertext [426]. Ciphertexts [610]. Cipolla [1422]. Circuits [791, 850]. Circular [642, 1036, 650]. City [1491, 751, 167, 438]. Clamping [991]. Class [509, 1131, 300, 324, 1103, 667, 545, 1123, 912, 506, 390, 427, 337]. Classes [1523]. ClasSi [1523]. Classical [1656, 596]. Classification [1489, 1046, 299, 566, 366, 560, 458, 793, 1289, 803, 463, 1431, 300, 351, 417, 362, 7, 73, 427, 352, 63, 1308, 119, 1440, 486, 320, 337, 85, 259, 524, 13, 1312, 79, 1307, 1473, 95, 359, 572, 581, 578, 462, 317, 1533, 367, 467, 835]. Classifier [548, 429, 1417, 322]. Classifiers [772, 1310, 352, 976, 317]. Classifying [521]. Cleansing [974]. CLEFIA [1621]. Cli [713]. Cli-Knowme [713]. Client [1232]. Client-Side [1232]. Climate [1679]. Clinical [1284, 842, 716, 1283, 723, 1191, 1852, 717, 713, 1187, 1842]. Clinically [1357]. Clinically-Feasible [1357]. Clique [1602]. Clique-Width [1602]. Clock [551]. Clonal [351]. Cloning [1173, 47, 894]. Closed [645, 591]. Closed-Label [591]. Closeness [440]. Closure [1769]. Cloud [1068, 1060, 799, 1435, 1566, 1212, 1067, 1500]. Clouds [1210]. Cluster [614, 1520, 778, 1524, 1305, 1519, 1134, 1122, 1729, 1193, 1485]. Cluster-Based [1193]. Clustering [1535, 168, 585, 826, 1480, 60, 1614, 458, 1212, 1305, 1537, 1494, 64, 91, 1525, 315, 523, 1518, 1287, 62, 1509, 790, 553, 367, 314, 92]. Clusters [40, 1691, 25]. Co [1303, 1304, 1540, 395, 92, 160, 1414, 1609]. Co-clustering [92]. Co-evolving [1540]. Co-learning [1609]. Co-occurrence [1414].

Co-positive [395]. Co-segmentation [160]. Co-training [1303, 1304]. Coarse [1403]. Coarse-to-Fine [1403]. Code [1566, 1131, 947, 456]. Coded [1423]. Coder [1392]. Codes [533]. Coding [1381, 1437, 1392, 1771, 1439, 1407, 496, 1409]. Coefficients [1648, 361]. Coevolution [422]. Cofficients [853]. Cognitive [969, 1754]. Collaborative [1783, 1278]. Collaboratively [472]. Collateral [1573]. Collation [52, 1347]. Collections [179, 162]. Collision [1263]. Colloquium [1858, 1462]. Colon [974]. Colonography [974, 981, 976, 980, 975, 973]. Colonoscopy [982, 579, 279, 1374]. Colony [1672, 416, 575, 524, 549, 555, 405]. Color [266, 149, 147, 805, 1369, 146, 337, 1374, 151, 148, 977, 145, 474, 1407]. Color-Constant [145]. Colorectal [981, 976, 980, 581]. Colored [194]. Coloring [1604, 1577]. Colorization [1420]. Combination [322, 768]. Combinators [911]. Combined [885, 677, 492, 1311]. Combining [1048, 105, 1422, 523, 727, 116, 305, 1275, 1107]. Command [95]. Commerce [1512, 1273]. Commercial [1059]. Common [1610, 793]. Commons [1175]. Communication [1145, 608, 185, 733, 1638, 1743, 912, 398]. Communications [366, 1635, 1633]. Communicative [969]. Communities [1513]. Community [1049, 1028, 3, 1461, 1513]. Community-Driven [1049, 1461]. Commutative [611]. Compact [1776]. Compaction [1042]. Comparative [449, 67, 976, 1484, 94]. Compared [765]. Comparing [300, 8, 1064]. Comparison [566, 1018, 1313, 464, 908, 810, 1718, 1359, 218, 39, 1311, 768, 462, 450, 174]. Comparisons [1699]. Compartments [776]. Compensation [1376]. Compensator [788]. Competing [108]. Competition [1180]. Competitive [193]. Complement [1648]. Complementarities [1172]. Complete [986, 866, 1530]. Completion [1592, 258]. Complex [707, 643, 1174, 1593, 417, 1170, 326, 584, 430, 901, 1513]. Complexity [40, 1602, 1587, 1784, 1588, 8, 1636, 1226]. Compliance [720, 1247]. Component [707, 699, 1668, 1833, 709, 1118, 700, 135, 1051, 1329, 487, 923, 920, 693, 578, 638, 467, 363, 497]. Component-Based [1118, 700, 1329, 638]. Components [1199, 1838, 702, 710, 1071, 775, 1201, 448]. Composability [110]. Compose [120]. Composition [1746, 296, 938, 630, 295, 892, 921, 1492, 1415]. Composition/Decomposition [921]. Compositional [931, 924]. Compositions [697]. Comprehensive [445, 1318, 897, 780, 564, 1057]. Compressed [1516, 61, 562]. Compressible [817]. Compression [646, 1435, 1369, 670, 1425]. Compressive [874]. Compromised [1797]. Computation [641, 642, 1216, 1566, 1718, 30, 1634, 1864, 1683, 1259, 1775, 845, 1000, 1810, 353, 1528, 1601, 13]. Computational [613, 1855, 1845, 109, 1087, 1722, 1682, 1339, 1652, 1092, 1842]. Computations [1126, 1732]. Computed [999, 381]. Computer [1764, 1060, 718, 1824, 1680, 41, 719, 1869, 746, 1819, 1820, 975, 1071, 838, 1292, 1122, 1835, 1836, 1816]. Computer-Aided [975, 1292].

Computer-Interpretable [718, 719]. Computing [706, 1821, 1760, 1752, 799, 1696, 1566, 1845, 1688, 270, 779, 631, 1826, 1827, 1828, 721, 1461, 1867, 635, 539, 1323, 1449, 1335, 193, 1073, 1858, 46, 1690, 1151, 1558, 1452, 25]. Concept [1489, 824, 1285, 789, 298, 896, 591]. Concepts [718]. Conceptual [1208, 866, 1461, 1226, 70, 1750]. Concerns [1330, 895]. Conchoids [644]. Concrete [1722]. Concurrency [793, 1831, 1770]. Concurrent [912]. Conditioning [1382]. Conduction [1716]. Cone [1658]. Conference [1832, 1824, 1849, 1818, 1826, 1827, 1828, 1817, 1867, 1815, 1835, 1836, 1837, 1847]. Conferences [1846]. Configuration [1482]. Conflicts [1748, 1225]. Confluence [748]. Confluent [1009]. Conforming [1717]. Congruence [946]. Conjunction [1855, 1858, 1842, 372]. Connected [764, 443, 331, 503]. Connectionist [1311]. Connections [487]. Connectivity [569]. Consent [1066]. Consent-Capturing [1066]. Conservation [1072]. Consideration [327]. Considerations [414, 183, 473]. Consistency [961, 973, 354, 1366]. Consistent [1383]. Constant [867, 145]. Constrain [256]. Constrained [759, 1350, 1266, 1426, 771, 394, 385, 553]. Constraint [1152, 1658, 1595, 243, 1153, 1482, 704, 289, 434]. Constraints [1795, 782, 916, 471, 269, 1422, 9, 1301, 1230, 165]. Construct [1200]. Constructed [1682]. Constructing [1511, 1627]. Construction [737, 1639, 1115, 843, 493, 754]. Constructions [1035]. Constructive [939]. Consumer [480]. Consumption [1738]. Contact [1011, 1715, 1340, 1702]. Content [1284, 1472, 75, 1283, 749, 1628, 1852, 74, 1462]. Content-Based [1284, 1852]. Content-Distribution [1628]. Contest [1054]. Context [1746, 191, 303, 908, 1272, 1237, 1215, 1457, 1285, 892, 359, 752]. Context-Aware [1746, 191, 1237, 1215, 1457]. Context-Free [908]. Contexts [128]. Contextual [81, 805, 620]. Continua [1455]. Continuous [1578, 669, 134, 261, 133, 400, 249]. Contour [258, 1350, 1003]. Contour-Constrained [1350]. Contourlet [458, 354]. Contours [247, 257]. Contract [1157, 693]. Contract-Aware [693]. Contractible [672]. Contractility [1353]. Contracting [1590]. Contractions [1591]. Contracts [699]. Contrast [353]. Control [4, 1263, 1655, 1656, 817, 488, 656, 1658, 1067, 887, 885, 1423, 420, 621, 593, 622, 401, 1575, 816, 468, 734, 1794, 623, 620, 1084, 480, 1788, 867, 394, 1167, 1686, 1086, 858, 396, 392, 1632, 391, 1427, 397]. Controlled [1146, 857]. Controller [475, 932, 439, 381, 443, 525, 400, 446, 398]. Controlling [1800]. Controls [1644, 1660]. Convection [424]. Convergence [668, 1379, 1498]. Convergent [67, 659, 1611]. Conversation [692]. Conversion [1440]. Converters [788]. Convex [1045, 1030, 261, 236]. Convolution [263, 728, 846, 685]. Convolutions [27]. Cooperation [969, 855]. Cooperative [194, 576, 528, 1611, 767, 422]. Coordinated [48]. Coordination [1616, 831, 1266, 1254]. Copositive [394]. Copy [368, 169]. Copyright [1438]. Core [922, 703, 42, 466, 38]. Coreference [542]. Cores [1574, 44]. Corpus [90]. Correcting [849]. Correction [978, 244, 1360, 152]. correlate [1387]. Correlation [1561, 1425, 342, 1541]. Correlations [1454].

Correspondence [275, 274]. Correspondence-Less [275]. Correspondences [1080]. Cortex [253]. Cost [1276, 1579, 1140, 850, 888, 729, 1554, 871, 249]. Cost-Aware [1276]. Cost-Sensitive [1554]. Costs [722, 754]. Countries [1310]. Couple [679]. Coupled [887, 97]. Coupling [1131, 139, 261]. Coupon [1730]. Courses [87]. Covariance [338, 511]. Covariant [1365]. Cover [1605, 1606, 1606]. Coverage [1324, 930, 491, 452, 602, 1072, 955]. Covering [589]. Covering-Based [589]. Covers [843]. COW [1540]. CPN [751]. CPU [175]. Crack [348]. Cracks [348]. Create [1385, 747]. Creating [1458]. Creation [1174]. Credentials [1057]. Crete [1819, 1820]. Criteria [1324, 502, 723, 1546, 1134]. Criterion [388, 415]. Critic [428]. Critical [1318]. Crohn [983]. Cropping [1562]. Cross [1362, 1752, 1204]. Cross-Domain [1204]. Cross-Policy [1752]. Cross-Sections [1362]. Crossbar [1728]. Crossing [1022, 1034]. Crossings [1050, 1040]. Crossover [560, 531]. Crowd [1810]. Crowdsourcing [1217, 1215]. Cryptographic [940, 1141, 1063, 618]. Cryptography [1624, 939, 1139]. Cryptosystem [606, 1619]. Cryptosystems [1627]. CSP [1594]. CSPs [1153]. CSS [533]. CT [234, 974, 992, 986, 997, 981, 983, 976, 980, 995, 994, 975, 973, 1433, 1352]. CTL [1127]. CTP [1639]. Cubic [1005, 1032, 680]. CUDA [179, 174]. Cue [144]. Cultivation [1665]. Cultural [528, 1453, 139]. Culture [351, 1248]. Cumulative [736, 561]. Curl [660]. Curl-Free [660]. Current [677, 1850]. Curvature [1697, 663, 675, 267]. Curvatures [657]. Curve [645, 1435, 240, 663, 279, 1139]. Curve-Based [1435]. Curves [643, 644, 649, 650, 656, 1080, 664, 251, 684, 685, 561, 1832]. Customer [1789]. Customizing [917]. Customs [1244]. Cut [992, 1586, 995, 1356]. Cuts [979, 997, 996, 261]. Cutting [882]. Cutwidth [1605]. Cycle [627, 1597, 1596, 1353]. Cycles [526]. Cyclides [675]. Cyclone [1485]. Cyprus [1850]. Czech [1869].

D [27, 579, 1293, 1381, 176, 1355, 1623, 1294, 1358, 341, 1370, 1028, 281, 1213, 157, 160, 272, 978, 1440, 861, 1349, 1385, 279, 995, 1340, 867, 1400, 1433, 1352, 173, 1295, 1343, 578, 1560, 1409, 561]. D-Texture [1293]. Dai [1168]. Damage [1573]. Danish [1683]. Dark [220]. Dashboard [1744]. Data [11, 213, 52, 54, 1068, 1, 566, 585, 1208, 1490, 1227, 1860, 1861, 99, 1499, 1209, 1091, 1745, 736, 92, 772, 1488, 986, 212, 547, 908, 1736, 341, 491, 881, 1067, 76, 1526, 1077, 12, 6, 1859, 82, 1515, 498, 1800, 1228, 36, 1630, 1768, 1210, 785, 1308, 1459, 978, 244, 1815, 228, 1349, 1782, 1772, 727, 1214, 235, 1683, 77, 1073, 677, 1299, 1125, 186, 10, 1788, 1737, 1468, 214, 1378, 1312, 68, 1862, 1479, 1188, 1514, 1473, 1167, 1789, 1686, 1469, 1352, 19, 16, 769, 790, 1500, 1738, 1053, 1456, 521, 1298, 1691]. Data [1411, 1870, 1817, 1815]. Data- [1859]. Data-Aware [1077]. Data-Centric [1125]. Data-Driven [1499, 1214]. Data-Fidelity [244]. Data-Merging [566]. Database [2, 12, 3, 8, 483, 15, 17, 87, 13, 1466, 1342]. Database-Driven [12]. Databases [136, 104, 59, 97, 1815, 1518, 78, 1529]. Datacenters [37].

Dataflow [1132]. Dataset [1323, 62]. Datasets [777, 115, 161]. Dataspace [1471]. Dates [327, 389]. Dating [1507]. DBpedia [1736]. DC [788]. DCE [978]. DCT [1391]. De-noising [347]. Deadlines [783]. Deal [1099]. Deaths [568]. Deblurring [239, 1428]. December [1838]. Decentralized [1174, 190, 1726, 815]. Deception [139]. Decision [849, 1752, 1246, 1761, 1091, 1283, 1245, 631, 7, 1114, 1852, 13, 714, 323, 846, 544, 1409]. Decisions [1487, 714]. Declarative [713]. Decoder [1408]. Decoherence [1705]. Decomposition [1365, 660, 704, 422, 921]. Decompositions [1603]. Deconvolution [1428]. Dedicated [892]. Deductive [1829]. Deep [219]. Defect [775]. Defect-Prone [775]. Defense [1087]. Defined [1746, 1656, 430]. Defining [10]. Definite [1647, 291]. Definition [264, 1407]. Definitions [939]. Defocus [1398]. Deformable [283, 1348, 139, 990, 275]. Deformation [1362, 1345]. Deformations [669]. Degenerate [1730, 653]. Degradation [1766, 686]. Degraded [1421]. Degree [1591, 1594, 685]. Delaunay [1041]. Delay [759, 393, 1266, 485, 771, 392, 391]. Delay-Constrained [759, 771]. Delayed [1616, 1364]. Delayed-Enhancement [1364]. Delays [509, 310, 502, 394]. Delegation [1154, 1487, 1146, 1170, 1793]. Delivery [327]. Demand [321, 329, 754]. Demands [969]. Demo [1026]. Demodulation [1141]. Demons [987, 1344]. Demonstrations [1245]. DEMs [452]. Denoising [234, 243, 244, 1421]. Dense [1603, 178, 174]. Densities [1304, 737]. Density [421, 348]. Dependence [1570, 743]. Dependency [1081, 79, 1101]. Dependent [389, 530]. deployed [1635]. Deployment [799, 1801, 815, 1448]. Depth [1436, 1370, 138, 1402, 1440, 1398, 1371, 1409]. Depth-Reliability-Based [1371]. Derivation [1790, 1423]. Derivative [668]. Derivatives [1365]. Dermoscopic [1292]. Description [888, 1233, 355, 724]. Descriptive [875]. Descriptor [339, 1380]. Descriptors [73, 459, 259]. Design [475, 646, 1759, 436, 1227, 1181, 1186, 656, 1250, 451, 823, 621, 821, 1451, 1782, 1117, 198, 788, 824, 1674, 789, 1168, 896, 473, 525, 1731, 754]. Designing [902, 548, 1257, 1325]. Desktop [29, 1462]. Desktops [799]. Detail [1369, 1412]. Detail- [1412]. Details [1357]. Detect [888]. Detecting [1060, 917, 1106, 1184, 722, 137, 1225, 495]. Detection [258, 368, 475, 800, 168, 907, 253, 579, 850, 1490, 1561, 344, 1083, 1527, 803, 491, 1028, 619, 1629, 1229, 1262, 466, 72, 255, 1742, 1374, 981, 983, 976, 980, 975, 345, 854, 789, 1637, 584, 476, 836, 312, 250, 169, 438, 1513, 379]. Detector [376, 836, 434]. Deterioration [886]. Determination [1003]. Determined [268]. Developable [653]. Developed [429]. Developing [1218, 1219, 71, 1075, 1456]. Development [1124, 449, 1183, 700, 1177, 1217, 1329, 1323, 840, 1328, 98]. Developments [223]. Deviation [1271]. Devices [1140, 1559, 1058, 879, 1449, 1462, 875]. DFC [1132]. Diabetes [720]. Diaflux [719]. Diagnosis [1284, 1096, 986, 770, 486, 983, 1292, 581, 448, 835]. Diagnostics [906]. Diagram [631]. Diagram-Based [631]. Diagrams [849, 1009, 1051, 1213, 701, 748, 846]. Dialogue [1488]. Diathesis [1106].

Dictionaries [253, 1287]. Dictionary [1106]. Diffeomorphic [1362, 1345, 987, 282, 1366]. Difference [428, 1769]. Differences [1247, 1559, 1725]. Different [1696, 832, 85, 1684, 768, 462]. Differential [1669, 791, 1656, 1670, 756, 445, 389, 1093, 386, 326, 684, 385]. Diffusion [283, 1410, 739, 226, 737, 228, 231, 424]. Diffusion-Weighted [226]. Digital [475, 741, 1425, 438]. Dilemma [733]. Dimension [357]. Dimensional [784, 1710, 524, 62, 68, 564, 363, 827, 837, 1237]. Dimensionality [487, 1772, 578]. Dining [698]. Direct [818]. Directed [1029, 1038, 1044]. direction [398]. Directional [1571, 1416, 456]. Directions [795]. Disambiguation [1103]. Disaster [1258]. Disc [801, 984]. DISCO2 [564]. Discontinuity [1001]. Discounting [961]. Discover [1534, 1530]. Discovering [756, 100, 1512]. Discovery [306, 299, 1614, 301, 1571, 302, 1825, 583, 188, 295, 304, 305, 636, 382, 298, 167, 1539]. Discrete [847, 781, 1696, 1080, 388, 413, 409, 416, 766, 575, 502, 1725, 327, 682, 274]. Discrete-Time [502]. Discretization [1656]. Discretization-Optimization [1656]. Discretizations [1717, 1693]. Discriminant [586, 1301, 319]. Discriminative [253, 149, 1291]. DisCSP [1162]. Disease [563, 983]. Disease-Specific [563]. Diseased [997]. Diseases [835]. Disjoint [1592]. Disjoint-Paths [1592]. Disparity [1383, 863, 1379]. Dispatch [529]. Disperse [301]. Displays [1028]. Dispute [941]. Dissection [836]. Dissemination [1800]. Dissimilarity [352]. Dissimilarity-Based [352]. Distance [1143, 1410, 1291, 280, 1142, 550, 74, 117]. Distance-Bounding [1143]. Distances [265, 1387]. Distinct [327]. Distinctive [448]. Distinguisher [1623]. Distortion [1041, 274]. Distress [803]. Distributed [40, 698, 1795, 1669, 1481, 1152, 1190, 1490, 1277, 1154, 1474, 310, 1153, 734, 1073, 1084, 1473, 1157, 1086, 1075, 1193, 1690, 1500, 1577, 1625, 915]. Distributed/Mobile [1474]. Distributing [1174]. Distribution [982, 436, 1430, 1628, 1155, 406, 348, 442]. Distributions [1531]. Disturbed [1193]. Divergence [660, 1382, 1498]. Divergence-Based [1382]. Divergence-Free [660]. Diversity [1497, 1099]. DLP [1626]. DM [1622]. DM-Mode [1622]. DNA [104, 756, 218, 535, 561, 562]. DNA-Binding [218]. Do [1234, 1040, 1808, 535]. Do-It-Yourself [1234]. DOCM [1851]. Doctoral [1850, 1869, 1858, 1462]. Document [370, 1100, 1527, 1753, 91, 56, 1230]. Documents [368, 1751, 1753, 1547]. Dodgem [1661]. DoE [813]. Domain [1124, 449, 1437, 1234, 1487, 1392, 1613, 1511, 80, 1204, 1344, 848, 903, 374]. Domain-Specific [1234]. Domains [1720, 672]. Dominance [587, 588]. Dominant [722]. Dominating [1600, 1589, 1601]. Domination [1588]. Dots [273]. Down [30]. DP [479]. DPM [1870]. DPO [1126]. Dr. [906]. Drawing [651, 1018, 1038, 1054, 1050, 1032, 1039, 1843]. Drawings [1010, 1043, 1048, 1042, 1037, 1024, 1031]. Drawn [1029]. Drift [551]. Drifts [1489]. Drilling [316]. Drive [818]. Driven [1218, 306, 54, 648, 1049, 940, 296, 1499, 1219, 1221, 12, 1329, 1461, 1214, 1375, 1335, 32, 1328, 1413].

Driver [869]. Driverless [860]. Driving [860, 380]. Dropout [401]. DSA [1155]. DSC [475]. DSL [1116, 900]. DSLs [894]. DSR [627]. DSToolkit [1471]. Dual [974]. Dual-Energy [974]. Due [327, 1685]. Dundee [1815]. Dupin [675]. Duplicate [1527]. during [132]. DW [227]. DWT [1438]. Dynamic [1578, 800, 1066, 192, 1152, 1018, 153, 60, 1381, 1435, 1566, 702, 1572, 715, 1808, 1369, 414, 335, 813, 1603, 1459, 1200, 1153, 1261, 111, 31, 1093, 94, 286, 1122, 152, 32, 747, 750, 1062]. Dynamical [651, 1401]. Dynamically [1253, 1123]. Dynamics [648, 421, 1809, 1194, 1180, 1704, 743].

e-Book [1737]. E-commerce [1273]. e-Governance [82]. E-Graphs [693]. e-Health [842]. E-Learning [70]. e-Negotiation [1160]. e-WeddingThailand [1502]. EAMA [1816]. Ear [364]. Earliness [327]. Earliness/Tardiness [327]. Early [1322]. Earthquake [1722]. Easiness [1524]. East [1847]. Easy [794, 1451]. EB [622, 1020]. ECC [1148]. ECC-Based [1148]. ECCV [1819, 1820]. ECG [738, 572]. Echocardiographic [1366]. Echocardiography [1355]. Eco [1076]. Eco-systems [1076]. Ecological [110]. Economic [412, 1660, 529]. Ecosystem [22]. EDEVITALZH [842]. Edge [475, 1017, 1355, 1020, 1412, 1021, 251, 836, 250]. Edge-Preserving [1412]. Edges [1029, 137]. Editors [1323]. Educational [799, 1736]. EEG [463, 677, 1545, 568]. EEG/MEG [677]. Effect [1162, 1727, 1157]. Effective [130, 52, 1017, 1288, 409, 254, 619, 185, 91, 572, 1492, 169, 384, 317, 555, 405]. Effectively [765]. Effectiveness [1130, 183, 955]. Effects [772, 441, 717]. Efficacy [515]. Efficiency [40, 850, 37, 1720, 1439]. Efficient [1516, 1284, 1501, 1061, 75, 1437, 979, 105, 172, 6, 254, 1383, 1395, 1629, 97, 36, 9, 1634, 1768, 1210, 1525, 1574, 188, 1557, 1063, 1727, 1104, 148, 382, 845, 1529, 1469, 252, 237, 169, 521]. Effort [1049]. Egocentric [1393]. Egyptian [513]. eHealth [1059]. Eigenfunctions [1372]. Eigentemplate [1416]. Eigenvalue [1715]. Ein-Gedi [1824]. Eindhoven [1843]. Either [235]. Elastica [238, 240]. Elba [1840]. Elderly [1451]. Elections [606, 1806]. Elective [1156]. Electric [766, 865]. Electro [818]. Electroanatomic [1363]. Electrochemical [884]. Electromagnetic [326]. Electromechanics [1361]. Electron [1705]. Electronic [974, 99, 606, 1777]. Electronics [480]. Electrophysiologic [1341]. Electrophysiological [1357, 1340]. Electrophysiology [1339]. Element [1714, 1715, 1643, 1688, 1645, 661, 1693, 870]. Elements [1699, 1005, 1541]. ElGamal [606]. Eligibility [428, 723]. Elliptic [1644, 1648, 1721, 1725, 1139]. Embeddability [1045]. Embedded [1047, 371, 834, 493, 398]. Embedding [1043, 1046, 1013, 1012, 1204, 1041, 1290, 145, 1558]. Emergency [1678]. Emerging [183, 1165]. Emotion [1300, 1194]. Emotional [1510, 1312]. Empirical [736, 1531, 1164, 1840]. EMV [943]. Enable [1449, 788]. Enabling [1336, 1461, 13]. Encoding [797]. Encodings [768]. Encrypted [1207]. Encryption [611, 1138, 802, 610, 426, 1625]. End [1236, 1797].

End-User [1236, 1797]. Endocardial [1340]. Endogenous [207]. Endomicroscopy [1284]. Endoscopy [977]. Energy [1666, 190, 974, 1761, 318, 37, 240, 1631, 36, 1184, 1063, 1072, 1704, 175, 570, 342, 568, 379, 754, 437]. Energy-Aware [1184, 175]. Energy-Efficient [36, 1063]. Enforcement [621]. Enforcing [1277]. Engine [825, 373, 22]. Engineering [1238, 1845, 1790, 1244, 951, 1850, 819, 952, 1817, 1869, 1318, 634, 1840, 1063, 1335, 1320, 1254, 1837, 1847, 374]. Engines [1119, 658, 619]. English [541]. Enhanced [1255, 1482, 298, 1754, 977, 760]. Enhancement [350, 227, 226, 803, 1369, 1364, 1397]. Enhancements [1145]. Enhancing [1493, 1432, 184, 1273]. ENO [670]. ENO-Type [670]. Enriched [1481]. Ensemble [288, 352, 321, 507, 980, 1299, 309, 521]. Ensembles [1552]. Enterography [983]. Enterprise [707, 52, 1202, 1219, 1217, 1231, 1115]. Enterprise-Level [707]. Entities [1863]. Entries [235]. Entropy [736, 1392, 246, 546, 460]. Entropy-Scale [246]. Enumeration [765]. Environment [1669, 208, 192, 1274, 1660, 1449, 860, 187, 430, 1444]. Environmental [1267, 1722, 1686]. Environments [968, 157, 1373, 863, 193, 207, 1512, 1171, 1447, 814]. EP [1339]. EPC [624]. EPDiff [282]. Epicardial [1340]. Epipolar [576]. Episodes [1530]. Epsilon [1389]. Equation [1717, 1723, 1180, 1662, 1693, 1730, 424]. Equation-Based [1180]. Equations [1644, 1656, 441, 282]. Equilibria [627]. Equilibrium [868]. Equipment [1185]. Equivalence [930]. Equivalent [1400]. Equivariant [266]. Era [1752]. ERBS [1698]. Error [849, 850, 973]. Error-Correcting [849]. Errors [65, 793, 541, 111]. Eruption [1685]. Establishment [615, 624]. Estimates [1714, 1658]. Estimating [1454, 144, 998]. Estimation [645, 1410, 1365, 166, 1304, 284, 287, 1672, 288, 138, 349, 290, 538, 352, 461, 86, 111, 1421, 1375, 787, 804, 1356, 1379, 1371, 752, 406, 1429, 1550, 450, 140, 319]. Estimations [1667, 665]. Estimator [668]. Estoril [1821]. ESWC [1868]. Ethernet [493]. Euclidean [550]. Euclidean-Distance [550]. Eukaryotes [217]. EULAG [1684, 1732]. Euler [238, 172]. Eulerian [1683]. EUROCAST [1835, 1836]. European [1847]. Evacuation [1270, 1194]. Evader [1573]. Evaluate [1541]. Evaluating [125, 1029, 1268, 203, 1506, 1458, 1638]. Evaluation [1199, 1130, 718, 1102, 761, 1699, 1133, 1294, 1028, 451, 1403, 984, 1272, 1525, 880, 66, 871, 531, 818, 1122, 875, 1000, 167]. Event [99, 340, 766, 897, 1125, 921, 98, 1750, 921]. Event-B [99, 1125, 98, 921]. Events [1510]. Every [1039]. Everything [1444]. Evoking [1057]. Evolution [1669, 1074, 1227, 1118, 1670, 756, 279, 529, 389, 1093, 256, 386, 1705, 326, 282, 901, 385]. Evolutionary [534, 778, 1671, 763, 1095, 1082, 503, 769, 1528]. Evolvability [709]. Evolving [1493, 765, 567, 1792, 1278, 1514, 1540]. Exact [641, 642, 233, 1864, 268, 843]. Example [159, 1140, 212, 297, 218, 1810]. Example-Based [159]. Examples [1104]. Exchange [899]. Exchanges [311]. Exclusive [948]. Exclusive-OR [948]. Executable [926]. Execution [715, 1134, 31, 892]. Executions [1086]. Exhaustion [917]. Existential [1127]. Exogenous [207]. Expand [1286]. Expansion [1102]. Expenditure [570]. Experience [714, 433]. Experimentation [865]. Experiments [952, 1809, 894, 1101]. Expert [440, 447]. Explicit [1507, 45, 1153]. Exploiting [1487, 1639, 1312]. Exploration [1208, 1262, 473]. Explorations [1209]. EXPLORE [7]. Exploring [1140, 1322, 767, 1224, 1452]. Expo [1697]. Expo-Rational [1697]. Exponent [470]. Exponential [401, 485, 391]. Exposing [1]. exposure [1394]. Expression [130, 566, 1434, 1091, 129, 519, 107, 578]. Expressivity [1754]. Extended [412, 263, 829, 504]. Extending [731]. Extensible [1461, 1450, 638]. Extension [813]. Extracted [57]. Extracting [1430]. Extraction [655, 1420, 1235, 1511, 1630, 1470, 1414, 1547, 1104, 326, 360, 591]. Extremal [554]. Extremals [1661]. Extreme [592, 1287, 508]. Extrinsic [47].

Face [471, 805, 132, 455, 364, 483, 477, 573, 66, 345, 580, 456, 175, 365, 434, 1380]. Facial [130, 481, 118, 1434, 129, 461, 121]. Facilitate [1125]. Facility [1568, 1546, 764]. FACS [1833]. Factorization [92, 1514]. Factors [790]. Failure [526]. Fairing [229, 676]. Fairness [183]. Fake [1781]. False [1385, 1637]. Family [222, 1137]. Fast [258, 1294, 238, 240, 1403, 179, 1262, 91, 465, 271, 1771, 496, 988, 1419, 124, 1409, 1711]. Faster [1589, 571]. Fault [1140, 929, 312, 448]. Faults [491]. FAW [1394]. FE [1709, 1721]. Feasibility [1140]. Feasible [1357]. Feature [93, 357, 1178, 340, 1537, 1377, 1417, 334, 1117, 1454, 126, 1414, 1542, 1285, 358, 360, 79, 1387, 380, 769, 1380, 562, 993, 1300]. Feature-Fusion-Based [126]. Feature-Guided [993]. Features [130, 73, 905, 364, 470, 1105, 1385, 1394, 1259, 360, 267, 133, 339]. Featuring [530]. February [1835, 1836]. Fecal [974]. Fecal-Tagging [974]. Fed [1665, 740]. Fed-Batch [1665, 740]. Feed [1622]. Feed-Forward [1622]. Feedback [84, 155, 439, 64, 743, 506]. Feeds [302]. FEM [809, 1717, 1720, 1726, 1353]. FEM-Simulations [809]. Femtocell [1455]. Fermentation [108]. Ferret [368]. FGCM [985]. Fiber [1410, 226]. Fibers [1359, 1659]. Fidelity [244, 235]. Field [1578, 1404, 1727, 1638, 482, 870]. Field-Effect [1727]. Fields [291, 667, 1812]. File [40]. Filter [560, 347, 33, 288, 146, 1432, 121, 789, 1346]. Filter-Based [1432]. Filter-Wrapper [560]. Filtering [1355, 263, 1383, 465, 1421, 1783, 1397, 1412, 1637]. Filtering-Based [1421]. Filters [270, 230]. Finance [1719]. Financial [1484]. Find [1313]. Finding [1747, 300, 1603, 279, 1774]. Fine [1403]. Fingerprint [457]. Fingerprinting [464]. Fingerprints [350]. Finite [1714, 1715, 1699, 765, 1645, 661, 362, 1693, 1689, 846, 845, 504, 400]. Finite-Time [400]. Finland [1858]. Fins [1727]. FIPP [526]. Fire [1681, 466]. Firefighting [1587]. First [739, 904, 946, 583, 1809, 1853, 912]. First-Class [912]. First-Order [946]. First-Passage-Time [739]. Fishery [421]. FIST [988]. Fit [235]. Fitness [550, 762, 780]. Fitting [1410, 664, 688]. Five [381]. Five-Bar [381]. Fixed [1043, 1593]. Fixed-Parameter [1593]. Flammable [495]. Flash [15]. Flash-Aware [15]. Flat [1526]. Flavored [637]. Fleet [1268]. Flexible [1066, 1227, 411, 1232, 1431, 219, 1471, 387, 14, 1259, 1466, 1448, 406, 786, 405]. Flight [564]. Floating [1718]. Flood [1258]. Flow [99, 284, 285, 1350, 388, 288, 172, 413, 409, 290, 1383, 945, 1682, 410, 857, 389, 289, 804, 386, 1729, 1086, 408, 237, 384, 249]. Flow-Shop [410, 389, 384]. Flowshop [416, 415]. Flowtime [416, 415]. Flu [1192]. Fluid [287, 288, 812]. Flux [809]. FMCO [1838]. fMRI [1299]. Focused [131]. Folding [583]. Following [867]. Footprints [1431]. Force [1038, 1702, 482]. Force-Directed [1038]. Forecast [1678]. Forecasting [311, 61, 321, 507, 886, 309, 503]. Forensic [1311]. Forest [906, 129, 976, 1097]. Forest-Based [129]. Forests [1349]. Forever [1566]. Forex [441]. Forgeability [375]. Forgery [1783]. Form [1345, 1221, 662]. Formal [928, 616, 954, 942, 1762, 1845, 104, 953, 1071, 713, 98, 624, 943, 915, 1838, 1833]. Formalization [1333]. Formalize [1192]. Formalizing [903]. Format [931, 1526, 1753]. Formation [614, 1567, 1498, 330]. Formats [212]. Forms [545]. Formulation [1650, 1251]. Forum [1854]. Forward [1622, 1339]. Foundations [263, 1830]. Four [1032, 1777]. Fourier [1711, 853, 845, 1433, 1419]. FPEAP [504]. FPGA [1775]. FPGA-Based [1775]. FPGAs [850]. FPS [1830]. FPT [1602]. Fractal [357]. Frame [1676]. Framelet [260]. Framelet-Based [260]. Frames [1010, 946]. Framework [1218, 370, 1208, 1065, 75, 1291, 1250, 1178, 172, 1801, 1537, 1329, 294, 1318, 880, 1461, 727, 671, 701, 1182, 1020, 897, 1774, 1458, 1627, 998, 286, 476, 70, 16, 374, 1257]. France [1832, 1830, 1825, 1854, 1816]. Free [974, 1767, 1345, 908, 1709, 660, 158, 879, 1769, 978, 1384, 1162, 975, 1412, 1371]. Free-Breathing [978]. Free-Viewpoint [1371]. Freedom [960]. Frequencies [326, 516]. Frequency [1765, 888, 61, 787, 874, 448]. Frequent [1516, 97, 1529, 1492, 1530]. FREyA [1745]. Friends [876]. Frog [384]. Front [21, 51, 936, 950, 957, 958, 962, 966, 972, 1007, 1056, 1070, 1079, 1112, 1113, 1121, 1129, 1136, 1150, 1198, 1242, 1282, 1297, 1316, 1338, 1368, 1406, 1443, 102, 1465, 1476, 1477, 1486, 1495, 1496, 1504, 1505, 1521, 1522, 1532, 1538, 1543, 1548, 1549, 1556, 1565, 1585, 1608, 1618, 1641, 1642, 1646, 1649, 1653, 1663, 114, 1664, 1677, 1687, 1692, 1694, 1700, 1707, 1712, 1713, 1734, 1758, 1779, 1803, 1814, 143, 181, 196, 197, 199, 202, 206, 210, 225, 293, 308, 404, 500, 595, 626]. Front [640, 690, 712, 726, 808, 890, 914, 604, 431]. Frontier [1262]. Frontiers [1861]. Fuel [884]. Full [1622]. Full-ARIRANG [1622]. Fully [279, 1478, 858, 331]. Fun [1216]. Function [1695, 1614, 1699, 147, 738, 1369, 582, 538, 683, 557, 506]. Function-Based [1369]. Functional [905, 220, 241, 394]. Functionalities [731]. Functionality [1255, 44, 19]. Functioning [1454]. Functions [1410, 1765, 1695, 649, 266, 852, 1303, 1304, 1698, 395, 853, 556, 1142, 1385, 297, 1375, 871, 618, 1155, 843]. Fundus [801]. Furnace [777]. Fused [439]. Fusion [1570, 1527, 364, 1394, 126, 360, 133, 354, 317, 236]. Future [190, 2, 1320, 1478, 70]. Fuzzified [545]. Fuzzy [1667, 782, 548, 513, 1672, 432, 547, 546, 735, 443, 588, 391, 397]. Fuzzy-PID [397].

G2C [940]. GA [526, 530]. GA-Based [526]. Gabor [460, 363]. Gain [1582, 381]. Gait [364, 832, 342]. Galerkin [1730]. GAMA [1167]. Game [1667, 1615, 1808, 897, 378]. Game-Method [1667]. Games [1761, 1499, 1611, 1458, 1841]. GAP [1213]. Gas [564]. Gases [495]. Gathering [1499, 1193]. Gating [38]. Gauss [739]. Gauss-Diffusion [739]. Gaussian [263, 1421, 425, 1401]. GD [1843]. Gedi [1824]. Gel [584]. Gen2 [624]. GenComp [745]. Gender [1417]. Gene [566, 585, 560, 586, 1091, 567, 519, 107, 517, 578, 1680, 1732]. Gene/P [1680, 1732]. General [111, 1627, 1676]. Generalised [138]. Generalizations [682]. Generalized [1668, 488, 1773, 763, 593, 1620, 675, 683, 178, 590]. Generalizing [1025]. Generate [1212]. Generated [908, 684]. Generating [651, 851, 1736, 491, 1703, 57, 376, 1321]. Generation [1179, 715, 1077, 758, 1210]. Generative [1098]. Generator [29, 855, 88]. Generic [745, 120, 997, 929, 898]. Genes [535, 563]. Genetic [436, 333, 412, 491, 758, 427, 420, 830, 439, 1376, 531, 1094, 757, 358, 1674, 1675, 768, 760]. Genetic-Immune [439]. Genomes [1534]. Genomic [215]. GENote [217]. Genotype [212]. Geo [1211]. Geo-Referenced [1211]. Geodata [879]. Geodesic [265, 256]. Geographical [1167]. Geomaterials [1708]. Geometric [1025, 1035, 1005, 483]. Geometrical [1424]. Geometries [990]. Geometry [643, 652, 658, 671, 1020, 886, 178, 1560]. Geriatrics [842]. Germany [1863, 1864, 1839, 1853]. Gesture [479, 343, 480]. Get [268]. GF [851]. GHZ [532]. GHZ-Type [532]. Gliders [864]. Global [979, 1403, 574, 268, 542, 679, 1444, 1371]. Globally [659]. Globe [484]. Glowworm [556]. Goal [940, 201]. Goal-Driven [940]. Goals [944, 1249]. Good [1313, 1603, 955]. GOP [1437]. Gossip [1158]. Gossip-Based [1158]. Governance [1067, 82]. GPC [1858]. GPGPU [1701, 802]. GPS [747]. GPU [774, 761, 176, 171, 170, 172, 1688, 26, 1771, 1376, 173, 175, 169, 25, 177]. GPU-Based [761]. GPU-Clusters [25]. GPU-Powered [173]. GPUs [27, 1094, 231, 1732]. Gradient [258, 1413]. Gradient-Driven [1413]. Gradual [1489]. Grammar [1108, 905, 541, 598]. Grammars [908, 596, 597, 1829]. Gran [869, 1835, 1836]. Granularities [10]. granulation [590]. Graph [1048, 1049, 1042, 1018, 1227, 705, 1029, 992, 1038, 1098, 1054, 979, 1591, 1213, 56, 997, 995, 1039, 996, 1126, 829, 1356, 1401, 555, 1016, 1843]. Graph-Based [1227]. Graph-Cut [992]. Graphic [748]. Graphical [1699, 719, 638]. Graphics [1696]. Graphs [1011, 1043, 1015, 627, 1047, 1025, 1035, 1747, 1014, 1027, 1022, 1050, 1044, 1590, 1603, 1588, 162, 1574, 1032, 1739, 1030, 1033, 693, 1601]. GRASP [782]. Gravitation [332]. Gray [356]. Graz [1838]. Greece [1868, 1819, 1820]. Greedy [784, 764, 1089, 415, 571]. Grid [635, 443, 1858, 1731]. Grid-Connected [443]. Ground [346]. Group [5, 1593, 128, 1809, 1194, 286, 1441]. Group-Valued [286]. Group-Wised [1441]. Grouping [772]. Groups [1108, 846, 845]. Growth [1660]. Guarded [38]. Guessing [917]. Guided [1397, 1363, 993]. Guideline [717]. Guidelines [718, 719, 1191]. Guiding [1205]. Guimarães [1833]. Gwangju [1856, 1857]. GWASpi [212].

H.264 [1408, 1392, 1557, 1429, 1409, 1427]. H.264/AVC [1392, 1409, 1427]. H.264/SVC [1557, 1429]. Hadamard [848]. Hagenberg [1818]. Half [1398]. Half-Sweep [1398]. Hallucination [158]. Hallucination-Free [158]. Hand [135, 132, 343, 480, 134]. Handheld [178]. Handling [1489, 897]. Handshape [134]. HAoS [1775]. Haplotype [105]. Haplotyping [571]. Hardness [1042, 1598]. Hardware [806, 877, 779, 1775]. Harmanli [1681]. Harmonic [672]. Harmony [388, 413, 409]. Harnessing [1480]. Hash [618, 1776]. Hashing [483]. Hasse [1009]. HCRF [1307]. Head [132, 140]. Headway [867]. Health [842, 1546, 1453, 1454, 1455, 1834]. Health-Care [1834]. Healthcare [1544]. Healthgrids [1188]. Heart [1855, 1361, 1344, 840, 1341, 380]. Hearts [1359, 1339]. Heat [1716, 1372, 277]. Heavy [816]. Held [1855, 1850, 1858, 1865, 1842]. Helmholtz [1365, 660]. Hemodynamic [835]. Hepatic [1005]. Heraklion [1868, 1819, 1820]. Herb [1540]. Herb-Herb [1540]. Herbal [1547]. Herbs [1542, 1539]. Hermite [1695, 649, 1698, 1005]. Herniation [984]. Heston [1701]. Heteroassociative [423]. Heterogeneous [1199, 1615, 779, 1229, 321, 1473, 846, 48, 1711, 1690, 1691]. Heuristic [759, 200, 779, 766, 886, 775]. HeuristicLab [780]. Heuristics [1770, 1162, 783]. Hidden [66, 377]. Hiding [1559, 1862]. Hierarchical [931, 1148, 614, 1801, 631, 1423, 1390, 1213, 994, 276, 505, 1399, 367]. Hierarchy [1501, 248]. High [153, 313, 1679, 264, 779, 373, 466, 61, 1323, 1253, 1439, 1335, 1407, 886, 1093, 62, 743, 298, 1690, 379, 1691]. High-Level [1253]. High-Performance [373, 466, 1335, 1690]. High-Technology [743]. High-Throughput-Screening [1691]. Highly [1709, 49, 172]. HiL [1470]. HiLA [1220]. Hilliard [1717]. Hindi [90]. Hippocampus [569]. Histogram [233, 337, 345, 1387, 450]. Histology [1290]. History [370]. HMM [1307, 454]. Hoc [1863, 1580, 1072, 873]. Hodge [660]. Hodograph [656]. Holistic [120, 215]. Hologram [1425]. Holonic [540]. Home [1454, 1059]. Homeobox [218]. Homeokinetic [1306]. Homogeneous [228, 1473, 737]. Hood [1764]. hop [1581, 1580]. Horizontal [1395, 1682]. Hough [129]. HRTeam [1257]. HTCPNs [1122]. HTCPNs-Based [1122]. Human [84, 81, 253, 1216, 565, 1488, 335, 1267, 1506, 1430, 131, 1272, 1359, 713, 1257, 133, 1706, 1810, 1401, 278, 127, 1427, 1354]. Human-Machine [1488]. Human/Multi [1257]. Human/Multi-Robot [1257]. Humanoid [831, 729]. Humans [366, 1559]. Hummingbird [1138]. Hummingbird- [1138]. Hungary [1821]. HWTS [1858]. Hybrid [258, 1270, 560, 795, 554, 520, 526, 413, 420, 1502, 574, 1391, 1375, 529, 796, 1633, 389, 304, 1675, 323, 1729, 385, 384, 527, 314]. Hydraulic [818]. Hydrocarbons [888]. Hyper [290]. Hyper-parameters [290]. Hyperbolic [270, 1041]. Hypercomplex [1419]. Hyperelliptic [1080].

I/O [37]. IAPR [1853]. IBM [1680]. IC [826]. ICA [66, 1560, 361]. iCat [1256]. ICDEM [1817]. Iceland [1685]. ICIC [1826, 1827, 1828]. ICWE [1850]. Ideal [279]. Identification [1665, 982, 534, 813, 883, 832, 1417, 609, 542, 1675, 1551]. Identifying [517]. Identities [948]. Identity [611, 1624]. Identity-Based [1624]. Idle [415]. IEC [1630]. IEC61850 [445]. IFIP [1844, 1847]. iHOP [222]. II [1857, 1820, 1836, 519]. Illegal [1559]. Illumination [244]. Image [350, 1362, 266, 234, 242, 481, 166, 75, 1096, 347, 548, 1294, 357, 1358, 264, 232, 170, 1369, 417, 1288, 240, 1856, 1857, 254, 984, 1395, 1390, 239, 64, 666, 162, 243, 465, 1426, 1440, 334, 271, 1286, 670, 1309, 1421, 1375, 990, 1407, 74, 1397, 1285, 358, 802, 148, 1438, 360, 492, 79, 152, 282, 1433, 584, 1371, 356, 1290, 359, 977, 1415, 342, 457, 169, 354, 1428, 1419, 236, 687, 494, 1691, 1396, 1356, 1295]. Image-Driven [1375]. Image-Enhanced [977]. Imagery [463, 840]. Images [1002, 481, 153, 1293, 1561, 992, 280, 801, 987, 1370, 351, 146, 163, 741, 1423, 1348, 1001, 461, 80, 673, 995, 994, 973, 1292, 1003, 456, 1425, 1295, 1343, 353]. Imaging [1855, 269, 1398, 1360, 1357, 1341, 1842]. Imbalanced [1533]. Immune [351, 439, 98, 332, 314]. iModel [160]. Impact [1679, 991, 23, 1497, 1542, 1357]. Implanted [1635]. Implement [852]. Implementation [475, 1791, 1140, 850, 806, 1680, 172, 373, 343, 815, 1251, 1084, 1094, 1684, 383, 1731]. Implementations [937]. Implemented [850]. Implementing [1207, 26, 746]. Implications [627]. Implicit [1507, 665, 1287]. Importance [1182, 1020, 369]. Improve [423, 410, 186, 542, 1807, 581]. Improved [1410, 607, 592, 1392, 90, 1595, 1621, 469, 502, 1724, 776, 1689, 558, 339, 390]. Improvement [318, 444]. Improvements [843, 533]. Improves [218]. Improving [918, 986, 1515, 1552, 185, 1720, 1240, 154, 44, 1375, 783, 760]. Impulsive [1658, 391]. In-Vivo [1339]. Incident [438]. Incomplete [544]. Incompressible [1344]. Inconsistencies [1561, 1225]. Increasing [1591]. Incremental [1227, 428, 636, 530]. Indentifying [563]. Independence [617]. Independent [857, 1788, 1558, 361, 467]. Indeterminate [1291]. Index [1696, 29, 503, 450]. Indexing [1747, 1629, 483, 80, 1290, 1509]. India [1849, 1817, 539]. Indian [468, 71]. Indicators [716]. Indistinguishability [613]. Individual [1261]. Indoor [1373, 873]. Induced [1705]. Induction [930, 396]. Inductive [86, 1126]. Industrial [903]. Industries [316, 1744].

Inequalities [1647]. Inertial [125]. iNetSec [1844]. Infarction [1341]. Inference [1787, 105, 290, 519]. Inference-Proof [1787]. Inferring [514, 1180]. Influence [1010]. Influenza [1269, 1706]. Informatics [1540]. Information [52, 1846, 827, 1124, 1784, 1235, 805, 216, 866, 105, 2, 882, 749, 823, 945, 1523, 532, 215, 1470, 1740, 899, 1084, 1763, 1749, 1167, 1193, 145, 325, 1243, 474, 993, 1866]. Infrared [885, 334, 456, 379]. Infrastructure [859, 794, 1273]. Infrastructures [800]. Inherent [1799]. Inhomogeneous [1415]. Initialization [315]. Initiation [1327]. Initiative [1154, 1170]. Injection [1140, 817, 1781, 1770, 947, 929]. Ink [350]. Ink-on-Paper [350]. Innovation [53, 735]. Inpainting [258, 240, 228, 1413]. Input [509, 1626]. Insertions [1017]. Inspection [1453]. Inspired [1074, 806, 1845, 1076, 529, 332, 1828]. Instance [1489, 1524, 1309]. Instance-Window [1489]. Instruction [1061, 32]. Instruments [373]. Integer [652]. Integral [645, 1714, 1645, 1621, 345]. Integrated [739, 344, 1474, 49, 1546, 123, 1557]. Integrates [1167]. Integrating [905, 215, 1020, 1854, 1462]. Integration [299, 982, 954, 1202, 669, 216, 1116, 820, 1703, 59, 221, 1115, 1326, 744, 214, 1378, 637, 19, 16]. Intelligence [1174, 1752, 1493, 1827, 1168, 553]. Intelligent [859, 1501, 1846, 1491, 1826, 1827, 1828, 1189, 468, 1385, 438, 398, 201]. Intensity [1366]. Intensive [1073]. Intention [1171]. Intentions [204]. Inter [1002, 1407, 1429]. Inter-color [1407]. Inter-layer [1429]. Inter-subject [1002]. Interact [881]. Interacting [565]. Interaction [208, 1205, 1860, 1540, 1256, 1704, 558]. Interactions [194, 219, 1430, 514]. Interactive [1745, 451, 160, 343, 355, 1123, 988]. Interation [1860]. Interchange [931]. Interdependence [1380]. Interdiction [1573]. Interest [340, 1288, 160]. Interesting [300, 1528]. Interestingness [1541]. Interface [208, 695, 1233, 451, 218, 875, 559, 1203]. Interfaces [1058, 1215]. Interfacing [1279, 46]. Interference [1575]. Intermediate [280]. Intermediate-Level [280]. Intermolecular [1704]. International [1838, 1833, 1821, 1832, 1846, 1824, 1855, 1844, 1860, 1861, 1822, 1849, 1863, 1870, 1818, 1826, 1827, 1828, 1848, 1866, 1817, 1869, 1825, 1867, 1864, 1840, 1835, 1836, 1852, 1854, 1858, 1837, 1816, 1871, 1865, 1842, 1843]. Internet [1451, 1226, 1123, 1222]. Interologs [514]. Interoperability [1449]. Interpolant [645]. Interpolation [1695, 649, 1698, 1697, 673, 680, 683, 494]. Interpretable [718, 719]. Interrelations [994]. Intersection [937, 1636]. Intersections [1478]. Interval [1670, 432, 1588, 665, 1430, 484, 402, 325]. Intervention [992, 986, 441]. Interventions [722]. Interview [343]. Intra [1794]. Intra-role [1794]. Intraoperative [989]. Intrinsic [281, 47]. Introducing [1775, 1099, 1401]. Introduction [695, 1280]. Intrusion [1490, 1629]. Intuitionistic [1667, 1672, 588]. Invariant [1098, 149, 255, 997]. Inventing [1748]. Inver [407]. Inver-over [407]. Inverse [593, 1723, 973, 845, 424]. Inversion [549]. Inverters [443]. Investigating [1357]. Invited [139]. IOMMU [43]. iOS [1062]. IOTLB [43]. IPDT [883]. IPEC [1864]. iPhone [878]. IR [1559]. Iris [459].

Irreducible [851]. Irregular [1378]. ISAMAP [32]. ISCA [1816]. Island [1866, 1840]. Isogeometric [661]. Isolation [312]. Isometry [669]. Israel [1824]. Issue [1845]. Issuer [1147]. Issues [15]. Italian [151]. Italy [1840]. Item [9, 1497]. Item- [1497]. Items [1736, 97]. Itemset [1516]. Iterated [415]. Iterative [1437, 1307, 1413]. Iteratively [684]. ITMAS [1851].

Jacobi [668]. Jason [204]. Java [1120, 1132, 703]. JBI [1823]. JDBC [87]. Jeju [1866]. JMS [383]. Job [411, 1768, 387, 833, 406, 405]. Job-Shop [406, 405]. Join [294, 897]. Joint [234, 1846, 1304, 147, 1358, 1839]. jORCA [221]. JPEG [1423]. JPEG2000 [1771]. July [1818, 1817, 1815, 1854, 1834, 1837]. June [1832, 1846, 1824, 1844, 1850, 1848, 1867, 1815, 1816].

K2 [1088]. Kaizen [1271]. Kalman [288, 121]. Kendall [522]. Kendall- [522]. Kenzo [741]. Kernel [1597, 1305, 538, 1372, 1719, 277, 985, 282, 396, 317, 1601]. Kernel-Based [1719, 985]. Kernelization [1599, 1600]. Kernels [283, 1596, 1676]. Key [615, 1635, 1623, 1796, 1511, 1423, 1620, 1628, 1627, 612, 624, 426]. Key-Phrase [1511]. Keyword [778]. KFDA [365]. Kinds [34]. Kinetics [108]. kiosk [1450]. Kirchhoff [1723]. KLEIN [1137]. Knapsack [527]. KnightShift [37]. Know [41]. Knowledge [1735, 1183, 1487, 1859, 1613, 1511, 53, 583, 183, 468, 753, 1744, 1114, 1547, 714, 116, 713, 1749, 724, 1127, 374, 1834]. Knowledge-Centered [1859]. Knowme [713]. Known [1623]. Known-Key [1623]. Kolkata [1849]. KOM [216]. Korea [1856, 1857, 1866]. Korean [368]. Kozo [1008]. KPCA [364]. KR4HC [1834]. Krakow [1847]. Krasovskii [394].

Label [591, 249]. Labeled [1308, 1312]. Laboratory [839]. Lagrange [172]. Lagrangian [238]. Lambek [597, 601, 598, 600, 599, 602]. Laminar [1354]. Landau [1090]. Landmark [1420, 121, 167]. Landmarks [1261]. Landscape [762, 780]. Language [1124, 1233, 925, 1745, 1154, 136, 135, 1234, 132, 719, 922, 905, 923, 1829, 151, 116, 134, 892, 1837, 903, 898, 1165, 399, 915]. Languages [918, 1767, 902, 1128, 1769, 94, 207, 912]. Laparoscopic [837, 839]. Laplace [1372]. Large [1666, 1362, 585, 1747, 92, 1670, 1050, 6, 1859, 179, 189, 1768, 1811, 1518, 529, 1739, 115, 1728, 161, 141, 1867]. Large-Scale [1666, 585, 6, 1859, 189, 1811, 529, 115, 1728, 161, 1867]. LASER [1840, 861]. Latency [1737]. Lattice [1627, 591]. Lattice-Based [1627]. Lattices [1710]. Laxative [975]. Laxative-Free [975]. Layer [1145, 1632, 538, 1429]. Layered [1349, 1792]. Layout [1027, 1724, 893]. Lazy [1263, 427]. LBP [419]. LDA [66]. LEACH [318]. Leader [867]. Leaf [357]. League [327]. Leakage [809, 1620, 1628]. Leaping [384]. Learn [1613]. Learned [253, 1257]. Learnflows [628]. Learning [89, 283, 1610, 646, 200, 1615, 75, 1735, 118, 1291, 1265, 1616, 1524, 336, 592, 427, 324, 1611, 785, 86, 1554, 1432, 1612, 139, 1484, 17, 1737, 115, 1853, 1416, 1306, 1447, 95, 117, 70, 858, 1865, 570, 1545, 508, 433, 1553, 530, 1090, 1609]. Learning-Based [570]. Least [428, 268]. Least-Squares [428]. Lectures [1840]. Lednice [1869]. Left [1355, 1364, 1769, 1349, 1347]. Legacy [1131, 1222]. Legal [1247]. Lemmatisation [1108]. Length [488]. Lesion [1096]. Lesions [1291, 981, 976, 980]. Less [1709, 275]. Lessons [1195, 1257]. Leuven [1870]. Level [707, 81, 648, 280, 822, 1721, 486, 469, 811, 1253, 256, 298, 535, 250, 127, 1388]. Levels [1679, 1175, 438]. Leveraging [1613, 1243]. Libraries [703, 896]. Library [219]. Lifetime [1569, 790]. Ligation [571]. Light [859, 468, 1384, 1404, 151]. Light-Field [1404]. Lightweight [872, 1138, 1137, 1631, 1200, 1239, 1463]. Like [370, 1794, 1627, 1661]. Likelihood [1304, 1360, 511]. Likelihoods [171]. Limited [1376, 342]. LiMoSense [1572]. Line [1569, 1274, 576, 1231, 508]. Linear [641, 227, 1593, 395, 1670, 547, 1600, 349, 1213, 665, 1636, 1301, 710, 601, 268, 818, 1693, 682, 164, 394, 557, 494, 1601, 813]. Linear-Time [1600, 1601]. Lines [675, 924]. Link [1514, 1508]. Linked [1208, 1209, 1745, 1736, 1228, 1210, 1737, 1738]. Linking [1205, 1735, 909]. Linking-Based [1205]. Links [1029]. Linux [95]. Literature [370]. Live [1572, 178]. Liver [992, 1291, 991, 986, 979, 989, 997, 985, 996, 998]. LLE [365]. Load [11, 788, 1728, 1754]. Loading [44]. Local [1568, 1293, 284, 1698, 245, 987, 340, 1567, 244, 1414, 345, 1412, 328, 512, 1357, 679, 1259, 1157, 482, 516, 348, 339, 317, 247, 252]. Local-Affine [987]. Localisation [1431]. Localities [633]. Localization [81, 123, 131, 1574, 1261, 677, 873]. Locally [235]. Located [859]. Location [1568, 1741, 942, 805, 1546, 1266, 764, 876]. Location-Based [1741]. Locations [747]. Lock [743]. Lock-In [743]. Log [150, 1344, 1099]. Log-Based [1099]. Log-Chromaticity [150]. Log-Domain [1344]. Logarithm [1080]. Logic [613, 791, 852, 702, 513, 1808, 1163, 1116, 735, 597, 1829, 601, 1275, 724, 696, 963, 967]. Logical [970, 1077]. Logistic [1089, 686]. Logistics [1474]. Logitlinear [516]. Logs [67, 1506]. Lombardi [1038, 1037]. Lombardi-Style [1038]. Long [1388, 1517, 941, 515, 476]. Long-Range [1388]. Long-Term [941, 515]. Long-View [476]. Losing [1598]. Lossless [1392, 1407]. Lot [412, 413, 409, 386]. Lot-Streaming [413, 386]. Low [1140, 888, 421, 92, 264, 1771, 1041, 39, 535, 685, 353]. Low-Cost [888]. Low-Power [39]. Low-Rank [92]. Lower [1586]. LPSNR [1391]. LS [309]. LSSC [1867]. LTE [789]. Lucerne [1844]. Lucy [1428]. Lumbar [984]. Lung [1289]. Lyapunov [395, 394]. Lying [963].

M2M [1205]. MABS [1871]. MABSys [106]. Machine [89, 75, 1488, 592, 1109, 63, 337, 976, 1447, 95, 1064, 330, 570, 508, 1302, 435]. Machine-Part [330]. Machine-Readable [1064]. Machinery [448]. Machines [797, 765, 1797, 430, 448]. Macro [1270]. Macro-Micro [1270]. Madrid [1684]. Magallanes [221]. Magnet [400, 446]. Magnetic [809, 1364]. Magnification [494]. MAIDL [1233]. Main [525]. Maintenance [1185]. Make [1741]. Makespan [388]. Making [1487, 323]. Malayalam [89]. Malo [1816]. Malware [1061]. Mammograms [469]. Mammographic [724]. Mammography [486]. Manage [1789]. managed [1444]. Management [11, 54, 1666, 1472, 42, 715, 1128, 1474, 67, 1870, 37, 749, 1471, 1801, 731, 779, 1631, 1817, 53, 1630, 735, 297, 733, 15, 1125, 551, 48, 637, 190]. Managers [621]. Managing [1159]. MANETs [1082]. Maneuvers [1773]. Manifold [278, 1302]. Manifold-Regularized [1302]. Manifolds [487]. Manipulating [900]. Manipulation [1172]. Manipulator [381]. Manipulators [312]. Manufacturing [1744]. Manycores [28]. Map [1436, 1370, 340, 302, 1402, 861, 298, 1632]. Map-Layer-Based [1632]. Mapped [219]. Mapping [1362, 1696, 862, 23, 1224, 1125, 1339, 1363, 1340, 32]. Mappings [1654]. MapReduce [1739, 1500]. Maps [54, 153, 1313, 1520, 163, 1322, 1519, 154, 753, 1052, 672, 747]. March [1839]. Marching [258]. Margin [586, 63]. Margins [991]. Maritime [1244, 1807]. Marker [769]. Markerless [125]. Market [1246, 513, 785, 752, 309]. Marketing [742]. Marketplace [1750]. Marketplaces [1609]. Markets [190, 1493, 743, 1806]. Markov [1761, 432, 66, 377]. MARTE [1336]. MAS [192, 1246, 1274, 1277, 1482]. Mashup [1199, 1202, 1207, 1201]. Mashup-Providing [1207]. Mashups [1236, 1234, 1200, 1462, 1203]. Masquerader [95]. Mass [486, 469, 1659, 564]. Massive [1216]. Matching [1362, 1424, 1403, 1781, 275, 276, 1433, 274, 1441]. Materialized [60, 58]. Materials [1711]. Mathematical [1869]. Matheuristic [1666]. Matlab [786]. Matrices [103, 1688, 338]. Matrix [103, 548, 92, 1709, 291, 1582, 1724, 1662, 1380]. Matrix-Free [1709]. Matrix-Vector [1724]. Matroid [1570]. Matter [20, 50, 971, 1006, 1055, 1069, 1078, 1111, 1135, 1149, 1197, 1241, 1281, 1296, 1315, 1337, 1367, 1405, 1442, 1464, 1475, 1503, 1555, 1564, 1584, 1607, 1617, 1640, 101, 1733, 1757, 1778, 1802, 1813, 113, 142, 180, 195, 209, 224, 292, 307, 403, 499, 594, 603, 625, 639, 689, 711, 725, 807, 889, 913, 935, 949, 956, 21, 51, 936, 950, 957, 958, 962, 966, 972, 1007, 1056, 1070, 1079, 1112, 1113, 1121, 1129, 1136, 1150, 1198, 1242, 1282, 1297, 1316, 1338, 1368, 1406, 1443]. Matter [102, 1465, 1476, 1477, 1486, 1495, 1496, 1504, 1505, 1521, 1522, 1532, 1538, 1543, 1548, 1549, 1556, 1565, 1585, 1608, 1618, 1641, 1642, 1646, 1649, 1653, 1663, 114, 1664, 1677, 1687, 1692, 1694, 1700, 1707, 1712, 1713, 1734, 1758, 1779, 1803, 1814, 143, 181, 196, 197, 199, 202, 206, 210, 225, 293, 308, 404, 500, 595, 626, 640, 690, 712, 726, 808, 890, 914, 1040, 604, 220]. Max [249, 1594]. MAX-2-CSP [1594]. MAX-2-SAT [1594]. Max-Flow [249]. Maximal [63]. Maximization [1072]. Maximizing [1569]. Maximum [1304, 586, 9, 1604, 1360, 1725]. Maximum-Likelihood [1304]. May [1821, 1824, 1860, 1861, 1822, 1851, 1830, 1868, 1858, 1871, 1865]. MBR [526]. MC [464]. MCBR [1852]. MCBR-CDS [1852]. MCF52259 [493].

MDE4HPC [1335]. Mean [1761, 546, 334, 1812]. Mean-Entropy [546]. Mean-Payoff [1761]. Mean-Shift [334]. Meaning [355]. Means [245]. Measure [156, 1517, 259, 955, 1533, 1541, 442]. Measurement [809, 1456]. Measurements [1374, 998]. Measures [772, 305, 325]. Measuring [265, 1561, 1523]. Mechanics [490]. Mechanism [439]. Mechanisms [699, 1254, 1243]. Mechatronic [822, 823, 819, 820, 821]. Mechatronics [829]. Media [1655, 1449, 1729]. Medial [641, 642, 1000]. Medical [1293, 1283, 1294, 1288, 721, 1185, 1286, 840, 714, 1188, 1295, 1691, 1852]. Medicinal [71, 1547]. Medicine [1823, 1191, 1539]. MEEMD [361]. Meet [273]. MEG [677]. Melanoma [1096]. Members [1131]. Memetic [411, 1540, 1671, 550, 771]. MEMICS [1869]. Memories [423]. Memory [1764, 852, 336, 756, 41, 30, 34, 28]. Memristive [393]. Memristors [490]. Mendonca [1422]. Mental [1057]. Merging [566]. Mesh [1381, 655, 1378, 686, 1286]. Meshes [1726, 173]. Messages [1743]. Messaging [1115]. Messaging-Based [1115]. Meta [773, 328, 512, 1090]. Meta-learning [1090]. Meta-models [328, 512]. Meta-optimization [773]. Metabolism [728]. Metabolism-Repair [728]. Metacognitive [1460]. Metacompiling [910]. Metaheuristic [782, 773]. Metaheuristics [767]. Metamodel [621, 623, 901]. Metastases [515]. Method [1667, 1654, 851, 1102, 1017, 238, 1559, 423, 607, 1431, 340, 351, 1688, 1106, 1645, 362, 1589, 1395, 1390, 1364, 364, 100, 479, 1517, 1391, 1231, 1772, 1557, 523, 1518, 529, 694, 1104, 1797, 840, 1003, 854, 1689, 394, 1400, 1328, 1637, 330, 1711, 517, 460, 581, 250, 424, 354, 124, 1513, 474, 450, 521, 329, 494, 442, 1411]. Methodological [1223]. Methodologies [449, 875]. Methodology [797, 1181, 1479, 19]. Methods [1838, 918, 566, 1716, 1824, 799, 739, 1656, 654, 1300, 1845, 661, 498, 1718, 1869, 487, 315, 139, 1309, 787, 259, 87, 1071, 804, 214, 814]. Methylation [535]. Metric [1568, 1410, 1362, 1740, 1064, 1467]. Metric-Based [1467]. Metrics [641, 1614, 930, 979, 1131, 76, 203, 74]. Metrics-Based [1131]. Metro [1052]. MICCAI [1855, 1852, 1842]. Micro [1270, 1741, 1279, 1457]. Micro-agents [1279]. Micro-architecture [1457]. Micro-posts [1741]. Microarray [1308, 581]. Microarray-Based [581]. Microbial [213]. MicroRNA [519, 1534]. MicroRNAs [535]. Microscopy [1702]. Mid [1388]. Mid-level [1388]. Middle [1622, 127]. Middle-Level [127]. Middleware [1631, 1459]. Middleware-Based [1631]. Migrations [1520]. Mill [777]. Mind [1224]. Mine [9, 1529]. Miniatures [1294]. Minimal [1787, 843, 249]. Minimax [1302]. Minimization [1018, 797, 1019, 1384, 148, 267, 252]. Minimizer [268]. Minimizers [235]. Minimizing [409, 416, 1575, 327]. Minimum [1579, 759, 1591, 763, 1661, 771, 274]. Minimum-Cost [1579]. Mining [1516, 628, 1481, 1490, 1860, 1861, 69, 301, 1488, 67, 6, 82, 498, 97, 57, 1103, 1482, 18, 1536, 55, 1479, 1473, 1508, 558, 1492, 1500, 1553, 518]. Minority [1615]. MiRaE [519]. Missing [1004]. MITACS [1830]. Mitigating [43]. Mitigation [1258]. Mixed [1047, 652, 1154, 1170, 495]. Mixed-Initiative [1154, 1170]. Mixed-Integer [652]. Mixing [262]. Mixture [425, 1545]. MMW [492]. Mobile [1160, 1561, 1233, 877, 1863, 1474, 1255, 1279, 336, 1173, 881, 1058, 879, 1506, 878, 444, 734, 874, 1072, 187, 1462, 1544, 1512, 875, 473, 175]. Mobility [186, 912]. Mobyle [295]. MobyleNet [295]. Modal [630, 166, 550, 835]. Modality [1287]. Mode [1702, 593, 1622, 1409, 446]. Modechart [798]. Model [1218, 1578, 1270, 1701, 613, 1002, 35, 1227, 992, 1381, 1098, 932, 655, 1657, 238, 194, 513, 1358, 1221, 1581, 547, 822, 866, 1680, 104, 1804, 749, 6, 441, 1244, 934, 290, 945, 311, 1329, 708, 900, 632, 1271, 26, 1333, 536, 120, 923, 546, 767, 1809, 507, 1089, 1326, 66, 1683, 1335, 620, 920, 1317, 818, 1084, 358, 1340, 697, 1684, 892, 1447, 1328, 1685, 1508, 1165, 516, 402, 309, 425, 250, 426, 252, 638, 911, 1127, 1325, 1732, 686, 1577, 1632, 377, 378, 435, 217, 1166]. Model-Based [822, 1333, 1317, 818, 1325]. Model-Checking [932, 934, 945]. Model-Driven [1218, 1221, 1329, 1335, 1328]. Model-Preserving [708]. Modeled [751]. Modeling [1472, 628, 1650, 1678, 371, 902, 135, 1219, 715, 109, 823, 810, 1267, 393, 110, 1213, 1372, 160, 1272, 106, 941, 1361, 1117, 623, 1854, 994, 1071, 824, 1652, 68, 884, 1167, 511, 1756, 1728, 898, 1711, 383, 769, 278, 1651, 1806, 1220, 1226]. Modelled [1753]. Modelling [1667, 1174, 1855, 1755, 925, 817, 1808, 812, 1005, 157, 821, 816, 1248, 728, 742, 740, 1122, 929, 1754, 1165, 1187]. Models [108, 1716, 1855, 1181, 284, 1195, 926, 810, 820, 1831, 1109, 1630, 352, 22, 1432, 1659, 139, 1660, 1180, 899, 1251, 1223, 1339, 121, 788, 1675, 838, 1684, 927, 1321, 1311, 107, 1545, 1057, 1325, 1401, 587, 760, 328, 512]. Modern [372]. Modernization [1222]. Modes [607]. Modification [402]. Modifications [1274]. Modified [483, 248, 407, 1428]. Modular [455, 703, 858]. Modularization [895]. Module [741]. Modules [120, 558]. Molecular [1704]. MOLEN [46]. MOMA [328, 512]. Moment [1301, 1433]. Money [1216]. Monitor [147]. Monitor-Camera [147]. Monitoring [1578, 1572, 1799, 1454, 694, 620, 478, 752]. Monocular [800, 477, 1401]. Monotone [1043, 1034]. Montague [598]. Monte [1702, 1704]. Morlet [506]. Morphological [89, 423, 418, 88]. Morphosyntactic [1107]. Morphotronic [828, 829]. MOS [548]. Mosaic [152]. Mosaicing [1396]. MoScript [900]. Motifs [756]. Motion [125, 1365, 1263, 1345, 287, 1001, 832, 978, 1432, 1375, 998, 840, 286, 164, 1346, 1342, 1343, 1401, 1393, 278, 1366, 1411]. Motivated [418]. Motivation [1251]. Motivations [1809]. Motor [463, 396, 593]. Motors [400, 446]. Mouse [83]. Movement [84]. Movements [1445]. Moving [1436, 472, 346]. MPC [814]. MPEG [73, 1408]. MPEG-4 [1408]. MPEG-7 [73]. MPI [755]. MR [1358, 1348, 1001, 1433, 1343]. MRA [1797]. MRI [227, 1345, 226, 984, 978, 1349, 1360, 1339, 1363, 1340, 1347, 1346, 1351, 1353, 993]. MRI-Based [1339]. MRI-Guided [1363]. MTC [1455]. Muitlgranulation [543]. Multi [649, 436, 333, 42, 166, 560, 827, 1196, 1179, 1263, 1480, 1183, 837, 1822, 1091, 1265, 1616, 1849, 1566, 778, 1300, 1581, 1580, 1172, 576, 1178, 1546, 528, 351, 413, 659, 1382, 1163, 779, 1552, 945, 1792, 1267, 427, 158, 593, 203, 538, 439, 466, 1502, 550, 1185, 38, 184, 1612, 1394, 1117, 1261, 139, 1248, 1811, 1237, 1309, 337, 1166, 1404, 44, 577, 1439, 994, 674, 345, 198, 358, 453, 1378, 1258, 1356, 164, 282, 437, 431, 1004, 841, 1167, 1351, 1807, 161, 331, 1492, 1264, 408, 590, 433, 329, 405, 835, 1871]. Multi-Agent [1183, 1822, 1616, 1849, 1502, 1811, 198, 1196, 1179, 1480, 1178, 1267, 1185, 184, 1612, 1248, 1258, 1807, 433]. Multi-agent-Based [1492, 1871]. Multi-algorithm [1552]. Multi-attribute [1163]. Multi-class [427, 337]. Multi-core [42, 466, 38]. Multi-cores [44]. Multi-criteria-Based [1546]. Multi-cultural [139]. Multi-dimensional [827, 837, 1237]. Multi-energy [437]. Multi-exposure [1394]. Multi-feature [1300]. Multi-granulation [590]. Multi-hop [1581, 1580]. Multi-image [1356]. Multi-instance [1309]. Multi-layer [538]. Multi-layered [1792]. Multi-modal [166, 550, 835]. Multi-model [1166]. Multi-motor [593]. Multi-Objective [778, 331, 436, 413, 439, 577, 431, 841, 408, 329, 405]. Multi-Organ [994, 1004]. Multi-Paradigm [1117, 203]. Multi-Parent [560]. Multi-party [1566]. Multi-platform [779]. Multi-population [576, 528]. Multi-Region [1351]. Multi-Robot [1257, 1263, 1265, 1261, 1264]. Multi-scale [659, 674, 345, 1378, 282, 1167]. Multi-sensor [161]. Multi-spectral [351]. Multi-stage [164]. Multi-swarm [431]. Multi-talker [453]. Multi-Test [1091]. Multi-threaded [945]. Multi-unit [1172]. Multi-valued [649]. Multi-View [158, 1382, 1404, 1439]. Multiagent [1276, 1490, 1268, 1611, 1156, 1182]. Multiagent-Based [1490]. Multibody [1655]. Multiclass [63]. Multicore [23, 48]. Multidimensional [1214, 527]. Multidisciplinary [1841]. Multigranulation [589, 544, 587]. Multigrid [1709]. Multilevel [759, 1708, 1195, 697]. Multilinear [467]. Multimedia [1210, 1862, 1563, 25]. Multimodal [600, 602, 1285, 1342, 993]. Multiobjective [411, 755]. Multiparty [1634]. Multiplayer [1216]. Multiple [627, 1424, 1418, 1017, 1511, 946, 1588, 157, 334, 681, 1676, 610, 173, 1243, 124, 1298]. Multiple-Domain [1511]. Multiple-Interval [1588]. Multiplicative [291, 230]. Multiply [1724]. Multiprocessors [49, 39]. Multiresolution [646, 232]. Multiscale [1289, 288, 670]. Multithreading [1408]. Multivalued [727]. Multivariate [1696, 1454, 214]. Multiway [1586]. Multiwordnet [370]. Mumford [241, 250]. Music [757]. Musical [371]. Mutation [1133, 583]. Mutual [1636, 1786, 993]. Myanmar [435]. Mycobacterium [514]. Myocardial [1350, 1353]. Myocardium [1348]. Myoelectric [885]. MySQL [87].

Name [1026, 1105, 151]. Narrow [150]. Narrow-Band [150]. National [1815]. Natural [847, 1760, 1745, 1256, 1829, 326, 422, 1415]. Nature [1845, 414, 1076]. Nature-Inspired [1845, 1076]. Navigation [1260, 336, 1213, 1373, 1612]. Near [1527, 885, 1638, 456]. Near-Infrared [885, 456]. Nearest [429, 6]. Nearest-Neighbor [6]. Need [41]. Negative [791, 1119, 1290]. Negotiation [1160, 1159]. NEH [410]. Neighbor [1571, 429, 6]. Neighborcast [1579]. Neighborhood [784, 764, 316, 1673, 557]. NeP4B [192]. Nested [1015]. Net [1668, 629, 194, 634, 1157]. NetCDF [212]. Netherlands [1843]. Nets [1303, 630, 631, 1831, 633, 635, 601, 602, 796, 730]. Nets-Within-Nets [635]. Network [1569, 1844, 313, 1571, 715, 1798, 67, 336, 311, 593, 440, 1630, 189, 1799, 63, 1574, 501, 507, 1169, 811, 620, 747, 511, 331, 504, 309, 503, 790, 506, 1513, 454, 377, 378, 495, 510, 497]. Networked [1173, 1213, 401]. Networks [1568, 1389, 1270, 1669, 627, 436, 579, 1074, 827, 1493, 318, 1635, 526, 513, 1863, 1581, 1580, 1572, 509, 567, 770, 1159, 1459, 452, 1461, 502, 183, 1740, 220, 1510, 1633, 1162, 1073, 1071, 605, 1072, 1785, 1576, 1551, 312, 558, 505, 508, 1563]. Neural [1669, 579, 313, 1303, 513, 509, 336, 770, 311, 593, 501, 502, 507, 511, 312, 331, 504, 309, 503, 505, 506, 508, 495, 510, 497, 730]. Neurodegenerative [835]. Neuroelectric [677]. Neuroimage [1389]. Neurology [842]. Neuronal [736, 1659]. Neutral [484, 310, 392]. Newly [429]. Newly-Developed [429]. Newman [558]. Newton [1654]. Newtonian [817]. NGS [211]. NMF [471]. NNSC [360, 492]. NNTP [732]. No [481, 548, 388, 389, 415]. No-Idle [415]. No-Reference [481]. No-Wait [389]. Node [526, 1728, 1551]. Nodes [1747, 1748, 273]. Noise [1770, 1421, 1397, 230, 442]. noising [347]. Noisy [615, 862, 239]. Nominal [1108]. Non [1010, 366, 1717, 653, 1418, 227, 245, 217, 1702, 987, 737, 813, 1645, 247, 26, 978, 599, 981, 276, 286, 1155, 845, 1290, 274, 252, 354, 1625, 1396, 817]. Non-Abelian [845]. Non-adaptive [1155]. Non-aligned [1010]. Non-associative [599]. Non-cathartic [981]. Non-conforming [1717]. Non-contact [1702]. Non-degenerate [653]. Non-GPU [26]. Non-homogeneous [737]. Non-linear [227, 813]. Non-Local [245, 247, 252]. Non-model [217]. Non-negative [1290]. Non-Overlapped [1418]. Non-rigid [987, 978, 276, 286, 274, 1396]. Non-sampling [354]. Non-standard [1645]. Non-uniformly [1625]. Nonconforming [1714]. Nonconvex [235]. Nonessential [1184]. Noninteracting [887]. Noninvasive [1357]. Nonlinear [647, 547, 1369, 1719, 485, 507, 673, 1662, 231, 396, 392, 549]. Nonlocal [244, 229, 230]. Nonplanar [1042]. Nonsmooth [235]. Nonsymmetrical [883]. Nonverbal [139]. Norm [242]. Normal [659, 72, 667, 545, 674, 678]. Normalization [65, 1438]. Normals [686]. Normative [1613, 1280, 1125]. Norms [1277, 1247]. North [1485]. Nose [83]. Notation [1333]. Notion [617]. Novel [585, 333, 611, 371, 798, 270, 556, 7, 1142, 1364, 57, 470, 574, 507, 1740, 220, 327, 88, 854, 408, 1513, 1427]. November [1838, 1849, 1856, 1857, 1825]. NTRU [1627]. NTRU-Like [1627]. nu [63]. nu-Maximal [63]. Nucleotide [219]. Number [1696, 1044, 1634, 844, 1034]. Numbers [847]. Numerals [65]. Numeric [1818]. Numerical [739, 1483, 227, 1703, 1718, 1719]. Nyquist [787].

O [37]. Obfuscation [1566]. Object [149, 1133, 335, 1523, 8, 830, 160, 355, 337, 1425, 399]. Object-Oriented [1133, 399]. Objective [333, 778, 331, 1110, 436, 413, 439, 577, 431, 841, 408, 329, 405]. Objectives [1761, 235]. Objects [1838, 1120, 1418, 1436, 8, 665, 119, 338, 152]. Oblique [163]. Oblivious [1583]. Observation [1811]. Observations [520, 1157]. Observer [401, 400, 446]. Observer-Based [401]. Observer-Controller [400]. Obtaining [1482]. Occlusion [132, 580]. Occupancy [1771]. occurrence [1414]. Oceanic [269]. OCL [1221, 894]. October [1833, 1823, 1869, 1847]. Octree [1709]. Odd [1596]. Off [850, 44]. Off-Loading [44]. Offer [1151]. Offers [1123]. Offline [1018]. Older [1454]. Olfactory [582]. Oligonucleotide [516]. OLSR [614, 731]. Omics [214]. On-board [815]. On-Line [1274, 508]. One [421, 300, 664, 891, 479, 523]. One-Class [300]. One-Pass [479]. One-Stop-Shop [891]. Online [1507, 1570, 1494, 381, 1299, 605, 752]. Only [862, 559, 342]. Ontological [881]. Ontologies [297, 910, 742, 1110]. Ontology [213, 54, 1746, 1755, 449, 1211, 71, 1104, 298, 70]. Ontology-Based [1746, 1104]. OO [696]. OOP [786]. OPC [382]. Open [1010, 1049, 1061, 1274, 1277, 1736, 1461, 671, 1158, 1450, 1738, 1844]. OpenCL [1094]. OpenFlipper [671]. OpenMP [755]. Operating [871, 46]. Operation [607, 316]. Operations [1268, 1767, 1468, 96]. Operator [560, 531, 407]. Operators [646, 1647, 1645, 1725]. Opinion [72, 1498, 1110]. Opinionated [72]. Opinions [1742]. Opportunities [1452]. Optic [285, 801, 172, 290]. Optic-Flow [290]. Optical [284, 1350, 1383, 289, 804]. Optimal [1655, 1656, 979, 1023, 324, 1266, 1659, 1482, 1660, 1676, 348, 424, 319]. Optimisation [1152, 821, 762]. Optimising [228]. Optimization [1535, 774, 1668, 809, 436, 333, 1716, 1650, 652, 827, 781, 554, 1656, 778, 303, 1671, 264, 1672, 576, 592, 825, 556, 864, 439, 387, 574, 452, 550, 792, 1482, 1673, 577, 871, 1536, 542, 389, 1093, 757, 1652, 1704, 868, 437, 431, 732, 841, 552, 551, 330, 1351, 408, 434, 769, 549, 1651, 236, 332, 754, 773]. Optimization-Based [1652, 437, 1651]. Optimized [320, 996, 789]. Optimizer [331]. Optimizers [833]. Optimizing [1027, 786]. Option [1718]. Orchestrated [1203]. Orchestration [1795, 916]. Order [631, 946, 98, 1625]. Order-Preserving [1625]. Ordered [1021, 544, 325]. Ordering [1162]. Ordinal [437]. Ordinary [1656]. Organ [1002, 1001, 995, 994, 1004]. Organised [205]. organising [1158]. Organization [340, 1217, 179, 1072]. Organizational [1271]. Organizations [1248]. organized [185]. Organizing [1313, 1520, 1519, 162, 1567, 1073]. Organs [1004, 1000]. Orientation [982, 338, 398]. Oriented [1218, 706, 692, 1186, 1133, 212, 1792, 203, 721, 1176, 729, 1185, 1231, 1117, 1317, 1168, 1151, 399, 1533, 1220]. Original [977]. Orthogeodesic [1013]. Orthogonal [1042, 1696, 1031, 80, 358, 1560]. Orthotropic [1721]. Oscillation [627]. Oscillations [883]. Oscillatory [509, 1648]. Osmosis [270]. Other [1831]. Otsu [1400]. Oulu [1858]. Outline [357]. Output [651, 392]. Outsourcing [1217].

Over-Parameterized [289]. Overdetermined [425]. Overlapped [1418]. Overlapping [558]. Overlay [189]. Overloaded [1031]. Overview [1283, 1320]. OWL [753, 71, 910, 717, 744]. Ownership [1147, 1146]. Ozone [1679].

P [1680, 631, 1045, 1732]. P/T [631]. P2P [191, 876, 193, 187]. P2P-Based [193]. PACE [570]. Pacific [1856, 1857, 1485]. Package [1726, 1399]. Packet [347, 401]. Packing [784, 776]. Padé [402]. Page [1044, 1226]. Pages [1235, 1511, 1314]. Painting [73]. Pairing [1144]. PAKDD [1861]. Palindromes [281]. Palm [347, 496]. Palmas [869, 1835, 1836]. Palmprint [362, 482, 361, 363]. Pandemic [1269]. Paper [350]. Papers [1819, 1820, 1835, 1836, 1838, 1833, 1821, 1832, 1846, 1824, 1855, 1844, 1860, 1861, 1822, 1851, 1849, 1863, 1823, 1870, 1830, 1868, 1850, 1818, 1826, 1827, 1828, 1848, 1866, 1817, 1869, 1825, 1867, 1815, 1864, 1839, 1852, 1854, 1858, 1834, 1853, 1837, 1847, 1816, 1871, 1865, 1842, 1843]. Paphos [1850]. Papillomavirus [565]. Paprika [1323]. Parabola [536]. Parabolic [1730]. Paradigm [635, 939, 1117, 1063, 46, 203]. Parallel [828, 1680, 756, 90, 830, 1720, 767, 1726, 1683, 1269, 1689, 178, 1558, 397, 25, 1776]. Parallelization [33, 1468, 24]. Parallelizing [29]. Parameter [1665, 1593, 1377, 111, 773, 1095, 542, 757, 1675, 380, 769, 319]. Parameter-Based [380]. Parameterization [997, 672]. Parameterizations [1596]. Parameterized [1602, 1605, 1586, 1587, 1606, 1588, 1595, 1604, 289, 1864]. Parameters [1161, 871, 85, 1704, 290]. Parametric [1670, 632]. Parent [560]. Pareto [431]. Paris [1830, 1825]. Parity [1761]. Parse [906, 601]. Parser [602]. Parsimony [105, 760]. Parsing [908, 1081, 96, 1101]. Part [1667, 1650, 1856, 1857, 1819, 1820, 1835, 1836, 1652, 248, 330, 1127, 1651]. Partial [1153, 728, 39, 201]. Partially [1029, 1308, 1309, 503, 1853]. Participatory [842, 1181]. Particle [1535, 781, 761, 576, 33, 592, 550, 1432, 529, 1536, 431, 551, 331, 408]. Partition [571]. Partitioning [1120, 91]. Partitions [249]. Partner [1256]. party [1566]. Pass [69, 479]. Passage [739]. Passengers [871]. Passing [1153]. Passive [1144, 1145, 429]. Patch [245, 1417, 1413, 237]. Patch-Based [1417, 1413]. Patch-Space [237]. Patches [653, 708]. Patent [1619]. Path [1597, 1604, 444, 279, 1739, 743, 1637, 1264]. Paths [1592, 491, 1590, 1052]. Pathway [216, 581]. Pathway-Based [581]. Pathways [109, 112]. Patient [1190, 715, 1005]. Patient-Specific [1005]. Patient-Tailored [715]. Patients [720, 997]. Pattern [69, 520, 869, 67, 463, 451, 1213, 844, 18, 55, 1469, 898, 1075, 317]. Pattern-Based [451, 1469]. Pattern-Recognition [1075]. Patterns [1183, 1164, 1517, 723, 1414, 345, 68, 1512, 1529, 1492, 517]. Pavement [348]. Payment [99, 1777]. Payoff [1761]. PC [1558]. PC-Based [1558]. PCA [76, 66]. PCB [316]. PDE [271, 494]. PDE-Based [271]. PDEs [270, 1719, 492]. Peak [564]. Pedestrian [1426, 1809, 1612]. Pedestrians [1270]. Pedigree [105]. Peer [1821, 1804, 189, 63, 184, 1628, 17, 193, 186, 1785, 1467]. Peer-to-Peer [1821, 189, 63, 184, 1628, 193, 186, 1785, 1467]. Penalties [1266, 327]. Penalty [268]. Penrose [748]. People [1445, 751, 1451, 1454, 124]. Perception [1559, 1370, 207, 967]. Perceptually [418]. Perfect [1023]. Performance [707, 849, 1133, 423, 1680, 45, 779, 1552, 8, 373, 23, 466, 735, 66, 976, 1323, 44, 1335, 85, 1317, 531, 1123, 1689, 462, 1690]. Period [412]. Periodic [1529, 688]. Periodic-Frequent [1529]. Permanent [400, 446]. Permutation [1515, 415]. Permutations [1703]. Personal [1238, 1511, 1789]. Personalisation [1340]. Personalization [718]. Personalized [842, 1823, 1544]. Perspective [629, 540, 1028, 634, 468, 70]. Perspectives [110]. Persuasive [1458]. Perturbation [1662, 424]. Perturbations [392]. Pervasive [695, 866, 1858, 1452]. Pessimistic [589]. Petri [1831, 629, 194, 630, 633, 634, 796]. Petrov [1730]. PGCM [985]. Phantom [1352]. Phase [1782, 1729, 342]. Phase-Only [342]. Phenomenon [1706]. Philosophers [698]. Phishing [619, 96]. Phone [881, 1506]. Phoneme [1377, 703]. Phones [1544]. Phonetic [1781]. Phonon [1705]. Phonon-Induced [1705]. Photo [179]. Photometric [159, 147, 277, 1384]. Photos [253]. Phrase [1511]. Phrases [435]. Physical [1145, 371, 1454, 462, 1577]. Physically [1142]. physiological [1312]. Pi [946, 443, 446]. Pi-Calculus [946]. PID [439, 397]. Piecewise [641]. Pigmented [1292]. Pinning [1023, 30]. Pipeline [35]. Pipelined [1500]. Pipelines [705]. Pixels [1420]. PKI [731]. Placements [462]. Plaintexts [1625]. Planar [1592, 1011, 1010, 645, 1015, 1046, 1035, 1014, 1037, 1044, 657, 1600, 664, 1030, 1033, 1601]. Planarity [1047, 1022]. Planarization [1017, 555]. Plane [644, 1012, 1041]. Planning [991, 1546, 1773, 864, 444, 1115, 1264]. Plans [715]. Plant [883, 71, 1534, 467]. Platform [842, 1268, 1272, 878, 876, 1166, 1363, 865, 1450, 1167, 175, 1558, 779]. Platforms [774, 1207, 874, 46, 48, 638]. Platoons [867]. Player [476]. Players [758]. Playing [1097]. Please [706]. PLSR [507]. Plucked [373]. Plucked-String [373]. Plugin [744]. PMG [1328]. PMG-Pro [1328]. Pocket [271, 1473]. Poincaré [1647]. Point [168, 1015, 1435, 1013, 1014, 1491, 1045, 788]. Point-of-Load [788]. Point-Set [1013]. Point-Sets [1014]. Pointer [1709]. Pointer-Less [1709]. Points [1672, 340, 1288, 679, 1441, 399]. Points-to [399]. Poland [1846, 1847]. Polar [1419]. Policies [616, 622, 734, 623, 1064]. Policy [1752, 412, 1065, 1265, 1748, 1245, 1790, 1244, 1267, 621, 1251, 867, 1754]. Policy-Reuse [1265]. Polish [1108, 1106, 1105, 1107, 1101]. Political [1742, 1743, 1806]. Pollution [1679, 1680, 1685]. Poly [1037]. Poly-arc [1037]. Pólya [743]. Polyakov [266, 148]. Polygons [656]. Polymorphic [46]. Polynomial [1698, 1355, 1596, 667, 867]. Polynomials [643, 851, 80, 358, 683]. Polyp [982, 975]. Polyps [579]. Polytopes [1030]. Popular [619]. Population [554, 1450, 576, 528]. Population-Based [554].

Populations [1615]. Porcine [1341]. Porous [1729]. Portal [991]. Portfolio [333, 546]. Portfolios [1789]. Portugal [1833, 1821, 1837]. Pose [166, 138, 349, 140]. Positioning [444]. Positions [1050]. Positive [1647, 291, 743, 394, 395]. Positives [1385]. Possibilities [832]. posts [1741]. Posture [133]. Potential [1679, 1589, 1704, 356]. Potts [252]. Power [1568, 40, 647, 42, 412, 1480, 1493, 810, 766, 38, 39, 884]. Power-Aware [1568]. Power-Efficiency [40]. Power-of-Two [412]. Powered [173]. Powerful [617]. Powers [1583]. PPR [309]. PR [1804]. PR-1 [1804]. Practical [652, 176, 464, 727, 1619, 1399]. Practice [1849, 1830]. Practices [84]. Pragmatic [1252]. Pragmatics [1743]. Praise [965]. Pre [1635, 13]. Pre-computation [13]. Pre-deployed [1635]. Preceding [344]. Precise [801]. Precision [504, 379]. Preconditioners [1717, 1721]. Preconditioning [1693]. Precursors [1534]. Predator [421]. Predicted [220]. Predicting [583, 515, 775, 516]. Prediction [1669, 1765, 565, 513, 822, 1407, 1514, 511, 559, 1429, 557, 1393, 1806]. Predictions [1493]. Predictive [842, 42, 488]. Predictor [1411]. Preference [1163]. Preferences [1507, 200, 1553]. Preferentially [535]. Prefetching [1509]. Prefix [1769]. Prefix-Free [1769]. Preimage [1622]. Preliminary [1101]. Prescriptions [1539]. Presence [1755, 1523]. PreSense [1755]. PRESENT [1796]. Preservation [1515, 893]. Preserving [669, 1369, 1526, 1001, 708, 1461, 1052, 1412, 605, 977, 1625]. Pressure [525]. Pretzel [824]. Pretzel-Shaped [824]. Preventing [614, 1559]. Prevention [1445, 947]. Preventive [842]. Prey [421]. Price [1172, 503]. Pricing [1494, 1730]. PRIMA [1849]. Principal [135, 487, 578, 363, 497]. Principle [447, 314]. Principles [1255, 1725, 1849]. Print [347]. Prior [1426, 235, 256, 249]. Priority [1332, 414]. Priors [261]. Privacy [1780, 1, 1065, 942, 617, 1870, 1515, 1848, 1799, 1461, 620, 1783, 605, 1638, 1785, 1273, 1064]. Privacy-Aware [620, 1785]. Privacy-Enhancing [1273]. Privacy-Preserving [1461]. Privacy-Protecting [1783]. Private [1781, 1636, 609, 876]. Pro [1328]. Probabilistic [569, 1793, 1169, 1311, 497, 783]. Probability [1302]. Probe [447]. Problem [1568, 1715, 1666, 1174, 759, 412, 781, 784, 526, 388, 882, 414, 887, 413, 763, 721, 633, 575, 387, 634, 1723, 762, 1661, 755, 771, 386, 1258, 682, 1082, 415, 330, 768, 406, 385, 407, 424, 384, 527, 530, 510, 555, 329, 405, 1601, 783]. Problem-Dependent [530]. Problems [1048, 1061, 1597, 560, 1602, 1065, 1844, 1656, 1616, 779, 1567, 1648, 1721, 410, 529, 524, 274, 408, 1533, 390]. Procedural [713]. Procedure [111]. Procedures [784, 1682, 1352]. Proceedings [1856, 1857]. Process [1501, 715, 1790, 445, 1125, 636, 424, 1401, 760, 1550, 1456]. Processes [739, 1761, 1177, 817, 737, 733, 743, 1749, 1187]. Processing [652, 1696, 738, 1119, 1459, 792, 671, 1739, 148, 1466, 1469, 1419]. Processor [37, 23, 1674]. Processors [42]. Produced [847]. Product [1696, 294, 1231, 924]. Product-Line [1231]. Production [824, 1738]. Products [103]. Profiles [246, 1134, 1440, 578]. Profiling [1784].

Prognostic [1280]. Program [919, 909, 856]. Programming [1236, 35, 705, 1603, 758, 427, 905, 708, 26, 897, 207, 46, 682, 1275, 399, 760, 1822]. Programs [1132, 1164, 945, 203, 1774, 696]. Progress [621, 622, 623, 620, 624]. Progression [1794]. Progressive [97]. Progressively [1656]. Project [82, 713, 1188, 385]. Projection [320, 504]. Projective [168]. ProLD [1228]. ProMAS [1822]. Promoter [535]. Promotion [699]. Prone [775, 973]. Proof [1787, 1058, 601, 602]. Proofs [939]. Propagate [1228]. Propagation [1192, 557, 1243]. Proper [1105]. Properties [645, 1714, 1074, 1769, 1710, 183, 1727, 740, 1075, 559]. Property [791, 1547, 618]. Proportional [1011]. Proposal [1223]. Proposed [1619]. Prospect [1188]. Prostate [1003, 993]. Prostheses [885]. Protecting [1783]. Protection [445, 1438]. Protégé [744]. Protein [219, 583, 514, 220, 584, 558, 559, 557]. Protein-Nucleotide [219]. Protein-Protein [514, 558]. Proteins [565, 218]. Protocol [1144, 937, 1148, 614, 1146, 1161, 917, 944, 1142, 948, 1360, 1786, 1157, 732, 624, 943, 1085]. Protocols [1143, 940, 1205, 938, 617, 1634, 1786, 1777]. Prototype [217]. Provenance [1068, 1748]. Providing [5, 1207, 1468, 637]. Provisioning [12]. Proxemic [1809]. Proximity [1024]. Prudent [1791]. Pruning [1545]. Pseudorandom [855]. PSIVT [1856, 1857]. PSL [1853]. PSO [452]. Psychographic [742]. Psychological [447]. PTZ [1376]. Public [866, 1627, 1638, 426]. Public-Key [1627]. Publication [1515]. Publishing [1751]. Pullbacks [1126]. Punjabi [90]. Purposes [1665]. Pursuing [1452]. Putting [1472]. Pyicos [219]. Pythagorean [656, 667]. Pythagorean-Hodograph [656].

QoS [296, 303, 693]. QoS-Aware [303]. QSDC [533]. Qt [878]. Quadratic [645, 1715, 1721]. Quadrature [273, 1346]. Qualitative [1245]. Quality [306, 481, 548, 716, 49, 76, 1523, 61, 1525, 1240, 1374, 1224, 1249, 259, 1223, 1686, 19, 1429, 16, 457]. Quality-Driven [306]. Quality-of-Service [49]. Quantification [1345, 999, 86]. Quantitative [1018, 984, 691]. Quantities [1751]. Quantization [73]. Quantum [532, 529, 490, 1763]. Quantum-Inspired [529]. Quasi [645, 680, 1033, 568]. Quasi-Interpolant [645]. Quaternion [349]. Queries [1211, 58, 1286, 1468, 1099]. Query [1119, 1739, 1466, 1099, 1469]. Querying [1745, 900, 1466]. Question [451, 1547]. Questionnaire [1772]. Quick [170]. Quincunx [670]. Quintic [656]. Quotient [1769].

R2RIF [744]. RAC [1048]. Radar [313, 876, 786]. Radial [1389, 1027]. Radiation [234]. Radio [888, 442]. Radio-Frequency [888]. Radius [1352]. Railway [886, 870]. Rainfall [507]. Random [560, 1747, 1634, 1620, 1402, 1426, 976, 1287, 1509, 1097]. Randomized [784, 764, 1562, 248]. Randomly [502]. Range [1120, 153, 1369, 1388, 862, 14, 1376]. Range-Limited [1376]. Range-Only [862]. Ranges [1569]. Rank [92]. Ranking [1523, 440]. Rapid [1323]. Rare [9]. Rascal [909]. Raster [10]. Rate [583, 668, 380, 1427]. Rating [1497]. Rating-Diversity [1497]. Ratings [1783]. Ratio [561]. Rational [1697, 738, 657, 675]. Rationale [1319]. Ray [1382, 1352]. Rayleigh [549]. RBAC [1794]. RBF [495]. RDF [1747, 1739, 1469]. RDFa [1753, 14]. RDFPath [1739]. Re [1226]. Re-modeling [1226]. Reachable [795, 1658]. Reaction [627, 387]. Reactions [1510]. Readable [1064]. Readers [1737]. Reads [219, 211]. Ready [1257]. Real [706, 344, 1499, 933, 1580, 798, 1408, 1331, 1554, 1510, 1376, 345, 1320, 1363, 453, 854, 1258, 1379, 584, 175, 1127, 1558, 1456, 140, 1396]. Real-Time [344, 1499, 933, 1580, 798, 1408, 1331, 1376, 345, 1320, 1363, 453, 1379, 175, 1127, 1558, 1456, 140, 1396]. Real-World [1554, 1510, 1258]. Reality [877]. Realization [303, 1117, 728]. Realizing [1030]. Really [170]. Rearranged [455]. Reasoning [1741, 1245, 1280, 717, 304, 1275, 201]. Receptors [582]. Recognition [130, 81, 313, 1434, 1096, 520, 136, 135, 357, 1300, 129, 463, 340, 455, 123, 1390, 364, 1377, 479, 343, 459, 1105, 573, 355, 66, 126, 1287, 480, 453, 134, 854, 496, 580, 456, 1171, 430, 482, 133, 117, 1075, 175, 365, 460, 342, 1441, 380, 361, 339, 1380, 1399, 127, 372, 363, 497]. Recommendation [749, 1178, 523, 1169, 1749, 1201]. Recommender [1535, 1497, 187, 1750]. Reconciling [108, 1748]. Reconfiguration [1265, 1122, 693]. Reconfigurations [702]. Reconnaissance [346]. Reconstructed [1425]. Reconstructing [901]. Reconstruction [176, 232, 853, 658, 240, 461, 477, 678, 492, 267, 167]. Recording [1559]. Records [721]. Recovering [1436]. Rectangle [1010]. Rectangle-of-Influence [1010]. Rectification [1395]. Recurrent [509, 612]. Recursive [1724, 530]. RED [1825]. Redistribution [279]. Reduce [722, 1797, 1737, 578]. Reduced [1623, 1621]. Reduced-Round [1621]. Reducing [1131, 541]. Reduct [62]. Reduction [432, 1134, 487, 1772, 1397, 524, 323, 402, 1411, 754]. Refactoring [893]. Reference [481, 548, 953, 1455, 899]. Referenced [1211]. Referral [183]. Refinement [630, 670, 788, 1307]. Refining [1656, 1071]. Refusals [1787]. Region [982, 805, 162, 1374, 535, 584, 1351]. Regional [1353]. Regions [1330, 137]. Registration [234, 1002, 1358, 986, 987, 163, 1348, 1001, 989, 990, 973, 282, 1433, 1346, 1352, 161, 993, 1396]. Regression [1355, 547, 1494, 777, 1089, 396, 570, 760]. Regular [1516, 1769]. Regular-Frequent [1516]. Regularities [1539]. Regularity [1506]. Regularization [266, 291, 254, 244, 235, 286, 1371, 319]. Regularized [268, 1302, 319]. Regularizing [1432]. Regulating [396]. Regulation [859, 1797, 519, 400]. Regulatory [567]. Reinforced [1722]. Reinforcement [1610, 200, 1265, 1616, 336, 1611, 1612, 1306, 858, 433]. Related [612, 473, 517]. Related-Key [612]. Relatedness [116]. Relation [1599, 1229, 1104]. Relational [548, 8, 78]. Relations [946]. Relationship [582, 823]. Relationships [127]. Relative [165]. Relaxed [1656]. Relay [883]. Release [389]. Relevance [156, 64]. Reliability [537, 822, 1371]. Reliable [786]. Relocation [1131]. Remarks [845]. Remediation [1682].

Remote [351, 1453, 398]. Removing [159, 948, 230]. Rendering [671]. Reorganization [1276]. Repair [728]. Replication [952]. Report [1054, 583]. Reports [1637]. Repositories [301, 900]. Repository [1114]. Represent [194, 1165]. Representation [648, 705, 1235, 1526, 631, 1768, 320, 670, 531, 1542, 1788, 1834, 846, 724, 1000, 1513, 127]. Representation-Independent [1788]. Representations [1011, 1434, 837, 117, 687]. Represented [649]. Republic [1869]. Requests [1798]. Requirements [1790, 1322, 1792, 1182, 28, 1125, 1225]. Research [449, 2, 335, 1805, 713, 495, 329]. Resection [991]. Reservoir [1493]. Residual [1392, 1384, 1429]. Residues [559]. Resilience [1576]. Resilient [1628]. Resistance [791]. Resistive [1655]. Resolution [313, 1679, 57, 461, 465, 941, 1404, 542]. Resolved [1371]. Resonance [1364, 326]. Resonating [812]. Resource [306, 299, 1499, 917, 779, 185, 188, 634, 193, 1749, 48, 385, 1825]. Resource-Constrained [385]. Resources [300, 834]. Respect [641, 810]. Respiratory [998]. Response [147, 1722]. REST [213]. RESTful [1205, 1206]. Restoration [242, 576]. Restrained [424]. Restructuring [59]. Retail [785]. Retargeting [156, 155]. Retinal [801]. Retinas [806]. Retrieval [283, 52, 1284, 1501, 1293, 75, 1283, 1294, 1289, 1288, 418, 80, 64, 1286, 1740, 1852, 74, 1285, 358, 359, 1295, 374]. Retrievals [370]. Retrieving [128]. Retrospect [1188]. Returns [546]. Reusable [1209]. Reuse [1183, 1265, 1255, 1252, 1231]. Reuse-Oriented [1231]. Reusing [898]. Revalidation [1230]. Revenue [86]. Reverberated [453]. Reverse [819, 424]. Reversible [848]. Review [1481, 1804, 878]. Revised [1838, 1833, 1821, 1832, 1846, 1824, 1855, 1844, 1860, 1861, 1822, 1851, 1849, 1863, 1823, 1870, 1830, 1868, 1850, 1818, 1826, 1827, 1828, 1848, 1866, 1817, 1869, 1819, 1820, 1825, 1867, 1815, 1864, 1840, 1839, 1835, 1836, 1852, 1854, 1858, 1834, 1853, 1837, 1847, 1816, 1871, 1865, 1842, 1843]. Revision [964]. Revisited [1587]. Rewriting [1126]. RF [810]. RFID [1144, 1780, 1148, 872, 1145, 1147, 1146, 1142, 792, 609, 1139, 1786, 624]. RFID. [1848]. RFIDs [1141]. RFIDSec [1848]. Rich [1105, 1226, 259, 1222]. Richardson [1428]. Ridge [396]. Ridgelet [362]. Right [1022]. Rights [1246, 960, 1058]. rigid [987, 978, 276, 286, 274, 1396]. Rigorous [1759]. Rim [1856, 1857]. Ringing [1428]. Risk [1678, 1445, 1793, 1187]. Risk-Based [1793]. RLSRunner [909]. Road [1270, 709, 803, 1390, 871, 854, 747]. Roads [344]. Robot [1260, 336, 830, 1262, 1257, 473, 312, 1263, 1265, 1261, 1264]. Robotic [1567, 472]. Robotics [397]. Robots [828, 1173, 729]. Robust [1152, 135, 1647, 1172, 290, 484, 1557, 1673, 116, 1438, 580, 312, 525, 250, 342, 1545, 1380]. Robustness [1721, 186]. Rocchio [367]. Roche [585]. Role [1246, 1128, 1176, 898, 127, 1794]. Role-Based [1128]. Roles [191]. Root [1026]. Rooted [759, 771]. Rotating [448]. ROTIV [1147]. Rough [589, 1257, 544, 590, 587, 588, 543]. Rough-and-Ready [1257]. Roughness [325]. Round [1623, 1621]. Round-Reduced [1623]. Route [1773, 868].

Routh [402]. Routine [1187]. Routing [318, 1581, 1639, 879, 776, 1021, 768, 329]. RPG [1794]. RPG-Like [1794]. RSS [302]. Rule [200, 1629, 1536, 744, 1092, 591, 530]. Rule-Based [1092]. Rule-Set [530]. Rules [4, 1746, 1481, 1116, 9, 1103, 1528, 522, 518]. Runtime [708]. RWA [755].

S [1665, 750]. S-Boxes [750]. S3E [1858]. SA [1643]. SA- [1643]. Saarbrücken [1863, 1864, 1839]. Sabotage [1083]. Safe [1599]. Safety [699, 991]. Sailing [221]. Saint [1816]. Saint-Malo [1816]. Salesman [407, 783]. Saliency [153, 1393]. Sample [1305]. Sample-to-Cluster [1305]. Sampling [1402, 272, 1710, 321, 354]. SANETs [837]. SAT [1594]. Satisfaction [1595, 201]. Saturation [629, 510]. Sausages [824]. Saving [437]. SB [216]. SBDO [1152]. SCADE [1318]. Scala [911]. Scalability [1429]. Scalable [1120, 464, 1146, 1709, 708]. Scale [1666, 585, 1098, 285, 1670, 6, 1859, 246, 189, 470, 255, 1811, 529, 1162, 1093, 137, 115, 251, 356, 1728, 161, 687, 659, 674, 345, 1378, 282, 1167, 1824, 1867]. Scale-Free [1162]. Scale-Invariant [255]. Scale-Up [687]. Scaled [865]. Scaling [907]. Scanner [861]. Scannerless [907]. Scar [1364, 1341]. Scenario [453, 1312]. Scenarios [1484]. Scene [1422, 120, 119]. Schedule [1796]. Schedules [1768, 612]. Scheduling [1190, 412, 411, 1332, 1581, 388, 413, 779, 766, 1189, 1156, 387, 316, 410, 762, 389, 386, 833, 437, 415, 406, 385, 408, 384, 390, 405]. Schema [1227, 59]. Scheme [258, 611, 464, 1423, 489, 1794, 1628, 523, 1375, 606, 874, 1438, 79, 178, 312, 1086, 1533]. Schemes [647, 227, 270, 663, 681, 375, 533]. Schools [1840]. Schur [1648]. Science [1845, 1869]. Scientific [301, 212, 1867, 1323, 28]. SCOOP [953]. Scopes [842]. Screening [1691]. Scripting [94, 912]. SDL [1854, 1332, 1333, 1331, 1854, 1319]. SDL-2010 [1319]. Seam [154]. Seamless [152]. Search [370, 784, 388, 413, 409, 951, 619, 36, 86, 764, 316, 29, 1089, 1673, 1376, 268, 1278, 1467, 385, 141]. Searching [489, 1785, 1528]. SEC [622]. Secant [1654]. Second [1855, 1283, 1817, 1852]. Secret [615]. Sectional [664]. Sections [1362]. Secure [1144, 937, 1635, 938, 608, 1207, 945, 1785, 1562, 426, 1560]. Securing [1068, 1145, 1798]. Security [1143, 1795, 616, 1844, 916, 1027, 1870, 1830, 1790, 944, 1801, 1631, 1792, 1848, 1866, 1630, 1620, 1391, 1815, 939, 1839, 1763, 1786, 1638, 1082, 1862, 1448, 511, 1807, 454, 377, 378, 915, 1846]. Segment [168, 1388, 251]. Segmental [996]. Segmentation [258, 1002, 260, 992, 655, 1358, 1350, 979, 801, 170, 132, 1211, 1388, 254, 246, 123, 489, 1348, 1390, 1364, 470, 1426, 255, 997, 486, 469, 985, 1349, 1309, 995, 994, 996, 259, 757, 1609, 286, 256, 1356, 988, 1347, 1004, 1351, 348, 160]. Segmented [341]. Segregation [1808]. Selected [1833, 1821, 1832, 1846, 1824, 1855, 1844, 1860, 1861, 1822, 1851, 1849, 1863, 1823, 1870, 1830, 1868, 1850, 1818, 1826, 1827, 1828, 1848, 1866, 1817, 1869, 1819, 1820, 1825, 1867, 1815, 1864, 1839, 1835, 1836, 1852, 1858, 1834, 1853, 1837, 1847, 1816, 1871, 1865, 769, 1842, 1843]. Selecting [1541]. Selection [566, 560, 60, 586, 93, 1178, 351, 6, 1244, 1267, 1537, 58, 546, 1089, 995, 78, 1254, 1637, 769]. Selective [44]. Selectivity [450]. Self [1074, 190, 797, 1313, 1520, 1300, 340, 1422, 1519, 185, 1567, 381, 465, 1073, 1072, 1158, 1444, 1387, 506]. Self- [1300]. Self-calibration [1422]. Self-Checking [797]. Self-correlate [1387]. Self-feedback [506]. Self-managed [1444]. Self-management [190]. Self-organising [1158]. Self-Organization [340]. Self-organized [185]. Self-Organizing [1520, 1519, 1567, 1073]. Self-star [1074]. Self-trained [465]. Semantic [1238, 52, 370, 1751, 1209, 1868, 1106, 451, 302, 1753, 1210, 188, 1470, 1740, 701, 14, 1449, 116, 1275, 298, 1290, 1295, 374, 1755]. Semantically [1754]. Semantics [11, 904, 1128, 1200, 1239, 1278, 1829]. Semi [1304, 118, 1305, 1403, 1310, 189, 999, 1301, 1299, 1371, 557, 377]. Semi-Global [1403, 1371]. Semi-structured [189]. Semi-Supervised [1305, 118, 1310, 1301, 1299, 557]. Semi-unsupervised [1304]. Semilinear [1644]. Sense [1741, 1103]. Sensing [1569, 1755, 351, 874]. Sensitive [1554]. Sensitivity [1665, 1701, 567, 1683, 1095]. Sensitivity-Based [567]. Sensomotory [831]. Sensor [1568, 1578, 1074, 827, 1755, 318, 1635, 1570, 1571, 877, 1863, 1572, 862, 1159, 1459, 452, 1576, 570, 462, 790, 495, 161]. Sensors [1260, 859, 84, 125, 540, 812, 873]. Sensory [1559]. Sentences [541]. Separation [1330, 666, 425, 696]. September [1855, 1863, 1870, 1819, 1820, 1864, 1852, 1853, 1842, 1843]. Sequence [213, 1077, 100, 389, 1307]. Sequence-Dependent [389]. Sequences [1345, 129, 217, 1703, 514, 672, 694, 121, 95, 1346, 1530, 561, 562, 1366]. Sequencing [219]. Sequential [850, 797, 18, 557]. Series [1669, 651, 61, 1517, 1518]. Server [1026, 44]. Servers [77]. Service [1218, 1276, 699, 706, 1746, 296, 705, 692, 1186, 1327, 49, 1792, 1176, 1200, 1185, 1317, 1463, 697, 1468, 305, 298, 1328, 1756, 1750, 1492, 1151, 1452]. Service-Based [1463, 1328]. Service-Oriented [1185, 1317, 1151]. Service-Oriented-Architectures [1792]. Services [1795, 1791, 1174, 718, 1202, 942, 303, 222, 866, 1753, 221, 53, 1206, 1185, 1059, 1633, 1224, 304, 742, 382, 383]. Servicing [1173]. Session [1327]. Sessions [705]. SET [1847, 809, 648, 1654, 1013, 194, 1300, 341, 1600, 1310, 1589, 1045, 1636, 589, 486, 469, 1105, 323, 256, 431, 250, 590, 399, 588, 530, 1601]. Set-Up [809]. Set-Valued [1654]. SETOP [1870]. Sets [1015, 1362, 795, 92, 1014, 1658, 1045, 243, 544, 587, 543]. Setting [38]. Settings [1261, 757]. Setup [389]. SFIM [1380]. Shadow [1684]. Shadows [144]. Shah [241, 250]. Shallow [1108]. Shape [283, 644, 669, 657, 281, 418, 1426, 997, 486, 1360, 355, 121, 256, 134, 261, 1000, 274, 1441, 278]. Shape-Appearance [134]. Shape-Prior [1426]. Shaped [824]. Shapes [245, 281, 275, 276, 286]. SHARE [1188]. Shared [756, 921]. Sharing [1634, 183, 1782, 876, 182, 433]. Sharpness [1397, 353]. Shearlets [666]. Shedding [11]. Sheets [430]. Shenzhen [1861]. Shift [170, 334]. Shifting [37]. Shop [411, 388, 413, 409, 891, 387, 410, 389, 386, 406, 408, 384, 405].

Short [336, 72]. Shortest [1052]. Shortest-Paths [1052]. Shot [489]. Should [1764, 268]. Shrinkage [492]. Shuffled [384]. Side [1232, 1141]. Side-Channel [1141]. SIFT [489, 459, 1385]. Sign [136, 135, 132, 1310, 134, 141]. Signal [475, 792, 1545]. Signaling [109, 1331]. Signals [646, 885, 857, 874]. Signature [941, 1086, 375]. Signatures [277, 1797]. Signcryption [1624]. Significance [211]. Significant [747, 1551]. SIIS [1846]. Silicon [806, 1727]. Similar [1747, 80]. Similarities [211]. Similarity [1293, 1212, 1106, 1523, 1517, 305, 218, 1064, 1467, 561, 562]. Similarity-Based [1106]. Simple [619, 1689]. Simpler [1600]. Simplex [385]. Simplified [1088]. Simply [1015]. Simply-Nested [1015]. SimSpectrum [1212]. Simulating [1244, 1612, 1805]. Simulation [1270, 1196, 739, 1179, 1735, 1268, 991, 471, 1808, 1250, 1804, 934, 820, 766, 1267, 1681, 839, 1272, 816, 792, 762, 1194, 811, 1166, 198, 788, 884, 1685, 1167, 1192, 1728, 1729, 1807, 1871, 870, 505, 1810]. Simulation-Aided [198]. Simulation-Based [762, 788]. Simulations [809, 1679, 837, 1720, 1811, 1727, 1704, 1498]. Simulator [1702, 825, 1484, 1269, 786]. Simultaneous [1048, 441, 244]. Single [166, 69, 778, 56, 583, 465, 137, 687, 495]. Single-Objective [778]. Singular [1661]. Sink [1314]. SINR [1581]. Sites [1223, 516]. Situation [1194, 1446, 511, 454, 377, 378]. Situation-Awareness [1446]. Size [1579, 1437]. Sizes [784]. Skew [294]. Skewed [521]. Skin [1096, 805, 1292]. Skinless [824]. SLAM [1373]. SLE [1837]. Sliding [1001, 990, 446]. Slope [523, 1378]. Slopes [1032]. Slovenia [1834]. Small [27, 1015, 1431]. Smart [1284, 1130, 540, 715, 1445, 1460, 1446, 1447, 1448, 438, 608]. SmartCard [606]. SmartCard-Based [606]. Smoking [1247]. Smoothed [1643]. Smoothing [1372, 1412]. Snoop [39]. Snoop-Based [39]. SOA [1065]. Soccer [476]. Social [1238, 1752, 1247, 960, 1808, 1250, 300, 1753, 440, 1169, 1510, 959, 605, 1498, 1551, 1841]. Societies [1158]. Socio [1249]. Socio-Technical [1249]. Soft [323, 1082]. Software [1174, 1833, 954, 919, 692, 205, 1118, 700, 960, 223, 749, 951, 1217, 952, 891, 1318, 634, 1840, 1191, 297, 1854, 959, 1320, 1463, 775, 924, 691, 1321, 1847, 1841, 1837]. Soil [538]. Solid [501]. Solids [641]. Solution [763, 1217, 1719, 268, 771]. Solutions [1424, 1593, 1334, 1445, 1115, 1544, 898]. Solver [1709]. Solvers [1708, 1647]. Solving [784, 1616, 779, 1153, 1626, 1082, 330, 424, 384]. Some [1679, 819]. SOS [903]. Sound [371, 373]. Soundness [948]. Source [671, 425]. Sources [1229, 1384, 116]. South [1856, 1857]. Sozopol [1867]. Space [168, 650, 827, 1824, 285, 150, 264, 340, 122, 239, 251, 1433, 237]. Space-Time [340]. Space-Variant [239]. Spaces [654, 1460, 664, 1446, 1285, 1448]. Spain [1823, 1835, 1836, 223]. Spanish [1823]. Spanning [759, 763, 771]. Sparse [1362, 1644, 1434, 586, 1593, 1616, 1358, 477, 320, 1724, 787, 998, 874, 1416, 496, 356, 117, 687]. Sparse-Representations [687]. Sparseness [471]. Spatial [656, 1392, 665, 228, 994, 1171, 1429, 1380]. Spatialized [1193]. Spatially [245]. Spatio [1358, 1383, 1349, 126, 10, 1264, 339, 127].

Spatio-Temporal [126, 10, 1264, 127, 1358, 1383, 1349, 339]. Spatiotemporal [130]. SPEA2 [525]. Speaker [497]. Special [395, 1845, 834, 667, 545, 375]. Specialization [919]. Specialized [560]. Species [1431]. Specific [1124, 1234, 1005, 903, 563]. Specification [918, 616, 233, 1154, 922, 622, 915]. Specifications [940, 917, 1253]. Specified [1050]. Specify [1165]. Spectral [1643, 1614, 1212, 351]. Spectrometry [564]. Spectrum [379]. Specular [1384]. Specular-Free [1384]. Speech [1377, 453]. Speed [1270, 886, 396, 400, 446]. Speed-Regulating [396]. Speedboat [825]. Speeding [802]. Spherical [157, 63]. Spiking [505]. Spin [1727]. Spirals [657]. Spit [445]. Splicing [585]. Spline [645, 1699, 656, 676, 680]. Splines [1695, 654, 1697, 661, 688]. Splitting [654, 532]. Sponsored [86]. Spontaneus [1870]. Spots [584]. Spreading [1706]. SQP [529]. Squares [428, 268]. SR [627]. SSVM [1824]. Stability [401, 485, 484, 310, 502, 394]. Stabilization [1657, 391]. Stable [627, 654]. Stacks [1358]. STACOM [1855]. STACOM’11 [1339]. stage [164]. Stand [706]. Standard [426, 1645]. Standardizing [1334]. Standby [537]. STANSE [1774]. star [1074]. Start [516]. Starting [1672]. State [797, 765, 783]. State-of-the-Art [783]. States [532, 1312]. Static [794, 1580, 813, 631, 157, 755]. Station [1321]. Stationary [1723]. Statistical [646, 1002, 1561, 232, 269, 1109, 1359, 211, 1637, 783, 435, 1855]. Statistically [361]. Statistics [341]. Statistics-Based [341]. Status [1320]. Stealthy [1061]. Steam [525]. Steganalysis [79]. Steganography [1562, 1563]. Stem [552]. Step [488, 1519]. Steps [709]. Stereo [159, 806, 171, 269, 1403, 1388, 1382, 158, 1395, 157, 1384, 174]. Stereo-Based [171]. Stereo-Vision [806]. Stereophotogrammetry [1386]. Stereoscopic [1370, 1028, 289]. Stippling [273]. Stochastic [1701, 648, 534, 284, 414, 934, 1611, 502, 330, 107, 378, 329]. Stock [1493, 513, 882, 311, 309, 503]. Stool [1374]. Stop [891]. Storage [613, 1411]. Store [1120, 714]. Story [1219]. Strategic [436, 1172]. Strategies [43, 1781, 1567, 964, 965]. Strategy [1160, 1499, 232, 341, 1801, 1403, 418, 1157, 24, 473, 571, 397]. Stream [11, 1741, 1119]. Streaming [1580, 413, 409, 386, 1563]. Streamlining [1226]. Streams [521]. Strength [1796]. Strengthen [855]. Stress [1018, 380]. Stress-Minimization [1018]. Strike [1718]. String [373]. Strong [1725]. Structural [1596, 740, 298, 448]. Structure [627, 582, 1526, 538, 1372, 536, 1385, 164, 482, 1401, 1354]. Structure-Function [582]. Structured [705, 1526, 63, 189]. Structures [260, 1005, 1519, 1213, 1722, 601, 1165, 730]. Studies [139, 1652, 372]. Study [108, 1701, 869, 930, 67, 1680, 82, 1244, 741, 1164, 445, 1170, 1484, 1180, 857, 717, 94, 1684, 868, 903, 1347, 1706, 1352, 236, 346, 397]. Studying [1300]. Style [1038]. Stylistics [1092]. Sub [1140, 134, 1295]. Sub-image [1295]. Sub-threshold [1140]. Sub-units [134]. Subband [458]. Subclass [1611]. Subcortical [1372]. Subcutaneous [999]. Subdivision [647, 662, 663, 665, 681]. Subgraph [1602, 510]. subject [1002]. Subpixel [836]. Subset [78]. Subsidy [1267]. Subspace [573]. Subspaces [1610]. Substances [1678]. Substantiating [1249]. Subsurface [1682]. Subtraction [474]. Successful [1501, 1452]. Succinct [315]. Suffix [1767]. Suffix-Free [1767]. Sugeno [391]. Suggestions [1099]. Sugiyama [1008]. Suite [1134, 943, 295]. Sulci [253]. Summaries [57]. Summarization [56]. Summer [1840]. Super [461, 465, 1404, 1371]. Super-Resolution [461, 465, 1404]. Super-Resolved [1371]. Superpage [45]. Superpixel [1288]. Superpixel-Based [1288]. Supervised [366, 579, 1303, 1524, 1305, 429, 785, 255, 1309, 1853, 118, 1310, 1301, 1299, 557]. Supine [973]. Support [1199, 649, 366, 1752, 1246, 205, 1209, 1096, 1283, 1014, 1488, 1245, 45, 49, 302, 770, 97, 9, 63, 1185, 1852, 337, 976, 28, 430, 448]. Supporting [1227, 1146, 1446, 1547]. Suppression [1783, 1428]. SURF [149, 126]. SURF-Based [126]. Surface [234, 803, 664, 272, 229, 679, 1258, 267, 688]. Surfaces [1832, 667, 674, 1000]. Surgery [1189, 1156]. Surgical [838]. Surveillance [1083]. Survey [55, 1319, 419]. Survivable [526]. Survival [1216]. SVC [1408, 1557, 1429]. SVD [826, 1391]. SVM [566, 362, 1676, 1534]. SVM-Based [1534]. SVR [321, 309]. Swarm [1535, 781, 761, 1566, 576, 592, 556, 550, 529, 815, 1536, 551, 331, 408, 553, 431]. Swarms [781]. Sweep [1398]. Swine [1339]. Switch [1728]. Switched [395, 485]. Switzerland [1844]. SWRL [71]. Symbolic [777]. SyMGiza [1109]. Symmetric [1647]. Symmetries [281]. Symmetrized [1109]. Symposium [1838, 1863, 1823, 1850, 1856, 1857, 1864, 1843]. Symptoms [1542, 1541]. Symptoms-Herbs [1542]. Synchronous [593, 400, 446]. Syndrome [1541]. Syntactic [1108, 520, 597]. Syntax [1829]. Synthesis [791, 371, 1570, 932, 373, 633, 920, 1379, 848, 1371]. Synthesizer [371]. Synthesizing [1321]. SysML [819, 820]. System [1535, 40, 84, 1759, 1746, 1752, 1490, 888, 534, 745, 1096, 1133, 1762, 1658, 822, 67, 803, 749, 741, 779, 1631, 770, 593, 393, 1630, 57, 1502, 64, 479, 1318, 343, 1451, 447, 816, 1470, 734, 811, 268, 1854, 871, 1547, 480, 187, 868, 743, 580, 437, 95, 98, 478, 482, 1754, 70, 175, 544, 396, 392, 570, 141, 462, 1731, 1558, 1456, 140, 1427, 1777]. System-Level [822]. Systematic [1240, 714, 457]. Systemic [1775]. Systems [11, 707, 1276, 928, 1766, 651, 1846, 191, 692, 1124, 99, 1183, 795, 1655, 1822, 700, 1593, 933, 1849, 395, 537, 1670, 1083, 1863, 1671, 798, 822, 738, 300, 2, 1859, 834, 934, 813, 819, 820, 8, 1629, 30, 189, 821, 1497, 1502, 106, 401, 485, 484, 310, 184, 1835, 1836, 1182, 111, 15, 871, 198, 1249, 87, 1317, 186, 1084, 1693, 1123, 13, 829, 46, 1786, 789, 884, 394, 846, 624, 1075, 1467, 1193, 402, 325, 1409, 391, 915, 1076, 1845]. Systems-Theoretic [700]. Systolic [774].

T [631, 688]. T-Splines [688]. Table [1776]. Tables [714]. Tag [1212, 1169, 1139, 39]. Tagged [1345, 1346, 1343]. Taggers [1107]. Tagging [974]. Tags [609, 182]. Tailored [715]. Taipei [1860, 1851, 1871, 1865]. Taiwan [1860, 1851, 1871, 1865]. Takagi [391]. Talk [139]. talker [453].

Tamil [88]. Tamper [1561]. Tangential [279]. Tanks [887]. Tardiness [327]. Target [800, 313, 1432, 535, 346]. Targets [220, 326]. Task [1183, 1154, 715, 834, 23, 1170, 1182, 1157, 1201]. Task-Based [1201]. TASTE [1320]. Taxi [1268]. TaxiSim [1268]. Taxonomy [951]. TC [1847]. TC3 [1853]. TCM [1540, 1542]. TDFFD [1345]. Teaching [17, 87]. Team [1265, 1257]. Teams [473]. Teamwork [970]. Technical [1249]. Technique [436, 520, 539, 492]. Techniques [75, 93, 303, 1762, 67, 951, 839, 410, 85, 838, 1847]. Technologies [1130, 715, 1454, 1455]. Technology [932, 1851, 882, 1856, 1857, 743]. Telco [1202]. Temper [777]. Template [1416]. Temporal [1472, 428, 1434, 1345, 702, 1119, 126, 10, 1264, 127, 1366, 1411, 1358, 1383, 1349, 339]. Temporally [1383]. Ten [629]. Tensor [1696, 1514, 482]. Tensor-Product [1696]. Term [336, 941, 515]. Termination [703]. Terms [946]. Test [1091, 930, 908, 491, 12, 1134, 1325]. Testing [1324, 1047, 1132, 954, 1593, 933, 1133, 1770, 1326, 1104, 955, 1325, 568]. Tests [783]. Tetrahedral [1726]. Text [301, 57, 47, 367]. Text-Mining [57]. Texton [417]. Textons [1390]. Texts [1105]. Textual [1333]. Texture [1293, 458, 1289, 417, 246, 470, 262, 474]. Textured [173]. TGI [1020]. Their [1235, 1658, 810, 1519]. Them [1472]. Theorem [265, 272]. Theoretic [700, 844]. Theoretical [263, 1404]. Theories [1827, 946, 419]. Theory [828, 821, 602, 1835, 1836, 1090, 376, 829, 1192, 356, 1839]. Therapy [234, 713]. Thermal [884]. Thesaurus [370]. Third [1824, 1825, 1842]. threaded [945]. Threatening [1798]. Three [1195, 1710, 438, 561]. Three-Dimensional [1710]. Threshold [606, 610, 348, 1140]. Thresholding [1385, 1400, 178]. Throttling [41]. throughout [1353]. Throughput [1728, 1691]. TID [100]. Tidsets [1516]. Tight [1602]. Tightening [930]. Time [1144, 1669, 782, 739, 344, 176, 1499, 933, 961, 1358, 1580, 509, 798, 737, 340, 1600, 409, 1519, 393, 1408, 61, 1331, 1517, 502, 1376, 1518, 345, 1320, 1661, 1363, 453, 854, 867, 1379, 394, 564, 175, 1127, 1558, 400, 508, 1456, 140, 391, 1396, 1601]. Time-Based [1144]. Time-Delay [393]. Time-of-Flight [564]. Time-Varying [509, 502, 394, 508]. Timed [928, 1766, 934, 632]. Times [389]. Timing [1336]. Tip [83]. Tiruchirappalli [1817]. Tissue [999, 1344]. ToF [234]. Together [1472, 221]. Token [897]. Token-Based [897]. Tomography [999]. Tonal [228]. Tone [152]. Tones [977]. Tool [718, 1246, 919, 1027, 217, 219, 254, 1320, 870]. Tool-Chain [1320]. Tools [1199, 646, 299, 794, 920]. Top [1516]. Top- [1516]. Topic [54, 753, 1508]. Topics [1819, 1820]. Topological [258]. Topology [49, 1575, 1020]. Topology-Aware [49]. TOR [1085]. Toronto [1855, 1822, 1852, 1842]. Torque [381]. Torremolinos [1823]. Torsion [1697]. Tortorelli [248]. TOSCA [1839]. Total [409, 416, 254, 98, 415]. Toulouse [1854]. Tourism [1483, 735]. Tourist [1553]. Toxic [1678]. Trace [428, 31]. Tracing [1061]. Tracker [1416]. Tracking [648, 125, 1418, 149, 1355, 136, 135, 1350, 171, 83, 132, 335, 1388, 122, 338, 334, 1432, 1344, 1261, 121, 480, 134, 1343, 392, 786, 1401, 124, 140, 346]. Tracks [886, 1485]. Tractability [1593]. Tractography [569]. Trading [850]. Trading-Off [850]. Traditional [1539]. Traffic [859, 1196, 1491, 1310, 468, 856, 1166, 857, 1797, 478, 1728, 438]. Tragedy [1175]. Train [1321]. trained [465]. Training [1300, 1310, 1104, 838, 508, 1533, 1303, 1304]. Trajectories [747]. Trajectory [864, 767]. Trajectory-Based [767]. Transactional [296, 1330, 34, 1529]. Transactional-QoS [296]. Transactions [1787, 1206, 1845, 1859, 1831, 1862]. Transcription [516]. Transcriptome [585]. Transfection [519]. Transfer [1610, 1147, 1146, 1369, 751, 116]. Transform [5, 347, 1437, 80, 363, 1711]. Transform-Wiener [347]. Transformation [745, 362, 891, 482, 380, 1560]. Transformations [944, 892, 911, 893]. Transformers [1449]. Transforming [753]. Transforms [847, 1696, 659, 844, 674, 845]. Transistors [1727]. Transit [1807]. Transitivity [1366]. Translated [541]. Translation [1109, 31, 927, 32, 435]. Transmission [1381]. Transmural [1341]. Transparent [25]. Transport [800, 1650, 866, 751, 1727, 1652, 1651]. Transportation [1659, 1638]. Transportation-Based [1659]. Transversal [1596]. Travel [868]. Traveling [407, 783]. Treatment [720, 715, 722, 841]. Tree [759, 69, 979, 1526, 1639, 763, 7, 123, 1494, 13, 771, 248]. Tree-Based [123]. Tree-Structured [1526]. Trees [1091, 1013, 1012, 1590, 664, 1087, 256, 1041]. Trend [513]. Trends [1850, 1819, 1820]. Trial [723, 1191, 1545]. Triangle [686]. Triangular [653, 1699]. Triangulated [641]. Triangulations [1036, 1695]. Trifocal [165]. Trigonometric [649, 1698]. Trinomials [851]. Trojan [376]. Tropical [1485]. True [892]. TrueNyms [1085]. Trust [1128]. Trustworthy [99, 187]. TS [1689]. TSP [575]. TTCN [1325]. TTCN-3 [1325]. TTP [606]. Tuberculosis [514]. Tubular [260, 256, 688]. Tumor [979, 989, 997, 985, 996, 578, 769]. Tumors [988]. TUMS [1756]. Tunable [542]. Tuning [439, 381, 1095, 814]. Tuples [1119]. Turbine [658]. Tutorial [951, 1840]. Tutorials [1850]. TV [1450, 24]. TV-kiosk [1450]. Tweets [1742]. Twice [1553]. Twice-Learning [1553]. Twin [1606]. Twin-Cover [1606]. TwitAg [1178]. Twitter [1756, 1178, 1743]. Twitter-Based [1756]. Two [1143, 847, 412, 781, 784, 421, 1519, 1718, 1721, 410, 1684, 1729, 564, 167, 375, 363, 398]. Two-Dimensional [784, 564, 363]. Two-Level [1721]. Two-Phase [1729]. Two-Step [1519]. Type [1714, 432, 146, 1718, 532, 546, 670, 683]. Type- [432, 546]. Typed [912]. Types [937, 1126]. Typing [213]. Typographic [65].

U [863]. UAV [1170, 346]. Ubiquitous [1461, 1449, 1458]. UHF [1139]. UI [1323]. UK [1815, 3]. Ulm [1853]. Ultra [1407]. Ultrasound [1003]. UML [928, 1221, 926, 1333, 701, 927]. UML-Based [928]. Unambiguous [1039]. Unbiased [1525]. Uncalibrated [79]. Uncapacitated [1568]. Uncertain [1657]. Uncertainty [1263, 933, 18, 1356, 325]. Unclonable [1142].

Uncorrelated [305]. Undergraduate [87]. Underground [442]. Underlying [1574]. Underspecified [1099]. Understanding [1247, 823, 120, 355, 1249, 743, 1444]. Underwater [864]. Unfinished [217]. Unfoldings [796]. Unified [1213, 1271, 382]. Uniform [793, 895, 1728]. Uniformization [265]. uniformly [1625]. Unify [895]. Union [937]. Unison [1097]. unit [1172]. Units [1751, 1696, 134]. Universal [746, 151, 713]. Universities [1313]. Unknown [1310, 1384]. Unlabeled [1312, 1298]. Unmanned [1258]. Unnatural [422]. Unstructured [1467]. Unsupervised [1303, 1537, 470, 1304]. Untagged [1343]. Untangling [1035]. Unwanted [1797]. Up-Sampling [1402]. Update [1787, 967]. Updates [1299]. upon [1531, 1370, 583]. Upper [1035, 1594]. Upward [1046, 1047, 1044, 1045]. Urban [344, 1195, 863, 1166, 860, 478]. USA [1848]. Usability [918, 1453, 880]. Usable [1058]. Usage [1535, 1506, 734, 1084, 1788, 1544, 732, 563]. Usages [372]. Use [27, 318, 794, 1322, 1240, 1239, 1224, 1674, 1539]. Used [872]. User [1236, 1507, 299, 1746, 1755, 1671, 1067, 451, 1497, 72, 1214, 1253, 1797, 868, 1512, 1099, 430, 1756, 1754, 875, 1509, 1203, 25]. User- [1497]. User-Centric [299, 1671]. User-Defined [1746, 430]. User-Driven [1214]. Users [742]. Uses [17]. Using [89, 4, 11, 475, 1535, 859, 1516, 646, 1362, 1501, 1701, 809, 125, 1174, 1365, 1120, 35, 1190, 1490, 1424, 75, 1741, 826, 99, 992, 5, 411, 1614, 1313, 1745, 1265, 548, 1698, 1697, 1108, 1520, 65, 930, 702, 1219, 513, 1350, 979, 1762, 1289, 987, 336, 463, 491, 1119, 1790, 881, 749, 417, 76, 138, 270, 409, 416, 569, 1322, 1005, 1703, 427, 420, 179, 1348, 895, 830, 1395, 1390, 364, 1494, 664, 439, 665, 466, 157, 9, 57, 821, 352, 63, 479, 666, 22, 1402, 343, 1574, 1517, 532, 338, 1417, 1200, 486, 861, 1301, 1153, 985, 1286]. Using [1772, 1115, 507, 1089, 1482, 71, 1182, 111, 297, 1375, 672, 1376, 1449, 577, 694, 787, 1335, 85, 345, 74, 1536, 717, 515, 87, 780, 1084, 1339, 121, 606, 1285, 112, 1340, 1737, 802, 742, 289, 755, 182, 1082, 39, 360, 496, 456, 1425, 1514, 903, 1447, 1478, 747, 95, 1562, 98, 1192, 1371, 693, 1789, 551, 482, 1290, 1807, 1343, 516, 365, 1353, 309, 250, 342, 1441, 380, 578, 783, 1203, 505, 1732, 1393, 1513, 1409, 563, 474, 687, 1550, 1366, 174, 1396]. Utility [1, 1308]. Utilization [77].

v [217, 1859, 1831, 863]. Vaadin [1204]. Vaccine [515]. Valence [1106]. Validating [76, 1182]. Validation [585, 616, 1179, 1221, 1364, 941, 1810]. Validity [1524]. Valley [982]. Valuable [1512]. Value [682]. Valued [1654, 286, 649]. Values [1089]. Valuing [1]. VANETs [872, 855]. Vanishing [168]. Variability [925, 902, 1334, 1354, 380]. Variable [784, 1594, 764, 316, 1673, 1162, 415, 543]. Variable-Ordering [1162]. Variables [1696]. Variance [1412, 62, 353]. Variance-Based [1412]. Variant [239]. Variants [1264]. Variation [215]. Variational [1824, 233, 269, 172, 243]. Variations [254]. Various [74]. Varying [1579, 509, 8, 502, 394, 508]. Varying-Size [1579]. Vector [73, 593, 63, 667, 1724, 337, 976, 430, 448]. Vectors [562]. Vehicle [344, 766, 420, 762, 865, 854, 768, 329]. Vehicles [860, 776, 1258, 1478, 858]. Vehicular [867]. Vein [991, 357]. Ventricle [1355, 1349]. Ventricular [1340, 1347]. Verbal [366]. Verbs [1106]. Verging [1386]. Verifiability [426]. Verifiable [1324]. Verification [707, 1791, 928, 1766, 918, 919, 1147, 1762, 798, 607, 109, 926, 622, 1840, 924, 927, 624, 354]. Verified [665, 1125]. Verifying [1071]. Versa [1151]. Versioning [1230]. Versions [1088]. Vertex [1605, 1606]. Vertices [1050]. Vessel [996, 1352]. Vessel-Radius [1352]. via [265, 1610, 919, 739, 1570, 853, 779, 1383, 1372, 751, 1426, 997, 244, 322, 44, 996, 1783, 519, 557, 1500, 1560, 553, 339, 391]. Vibration [420]. Vice [1151]. Video [81, 1436, 118, 1381, 1437, 136, 135, 1580, 129, 464, 1856, 1857, 489, 466, 131, 1439, 121, 1401, 1393, 1558, 1409, 346]. Video-Based [136, 135]. Videos [1559, 1374, 476]. ViennaCL [1690]. Vietnam [1196]. View [1787, 827, 60, 700, 970, 158, 477, 1379, 476, 1750, 1382, 1404, 1439]. Viewpoint [1379, 1371]. Viewport [1051]. Views [1418, 58, 173, 1298]. VII [1862]. Virtual [579, 799, 839, 746, 279, 1379, 1171]. Visceral [999]. Visibility [154]. Visible [456]. Vision [800, 1824, 344, 280, 806, 831, 830, 1819, 1820, 120]. Vision-Based [800, 344]. Visual [1124, 1235, 1289, 1028, 1288, 270, 1377, 505, 1393, 1427, 177]. Visualization [974, 341, 891, 1517, 1214, 1363, 214, 1479, 1016]. Visualize [1026, 1053]. Visualizing [1027, 1519]. Visually [80]. Vivo [1339]. VNS [782]. Vocabulary [141]. Volatility [1701]. Volcano [1685]. Volume [341, 676, 1689]. Volumetric [658, 1361, 231]. Vote [617]. Vote-Independence [617]. Voting [617, 256]. Voxel [1708, 1720]. Voxels [1710]. VRP [782]. vs [1336, 30, 422]. Vulnerability [1762]. VURTIGO [1363].

W [670, 532]. W-Type [532]. w.r.t [1594]. W3C [1203]. Wait [388, 389]. Walk [1402]. Walks [1426]. Wall [840]. Walsh [853, 848]. Warehouse [76]. Warfare [447]. Warsaw [1846]. Wasserstein [262]. Water [1246]. Water-Rights [1246]. Watermark [1438, 1558]. Watermarking [1391, 1557, 78, 1560]. Watershed [469]. Wavelet [5, 347, 1696, 287, 506, 363, 494]. Wavelet-Based [287]. Wavelets [660, 666]. Waves [269, 549]. Way [1745, 1240, 1053]. WCB [360]. WDM [526]. Weak [1260, 269, 1725, 612, 1387]. Weakly [261]. Wearable [874, 462]. Weaving [701]. Web [1868, 1850, 1498, 1238, 1218, 1236, 1535, 1795, 1791, 299, 1205, 296, 1755, 1209, 1219, 303, 1221, 1235, 217, 222, 1511, 221, 1206, 72, 1240, 947, 1239, 1231, 1237, 1215, 1224, 1223, 1314, 304, 55, 94, 1222, 742, 1468, 1462, 1275, 1444, 1225, 1509, 383, 1492, 1220]. Website [1472]. WeddingThailand [1502]. WEED [1816]. Wei [1786]. Weight [548, 97]. Weighted [647, 69, 1304, 226, 1647, 288, 661, 1536, 358, 369, 317, 433]. Weighted-Power [647]. Weighting [1385, 1352]. Weights [1305]. Wellbeing [1450]. Western [1485]. Wetlands [1682]. Weyl [107]. WG [1844]. Wheelchair [472]. While [1]. Whither [2, 3]. Wide [602]. Wide-Coverage [602]. Widgets [1203]. Width [1602, 1598]. Wiener [347].

Wikipedia [1100]. Wikipedia-Based [1100]. Will [706]. Willmore [682]. Window [1489]. Windows [782]. Windup [446]. Wine [108]. WIOSCA [1816]. Wireless [84, 1074, 827, 318, 1863, 1581, 1580, 1582, 1159, 452, 790, 398]. Wiring [1207]. WISA [1866]. Wised [1441]. WISP [1139]. Within [635, 1747]. without [40, 1635, 1153]. WLAN [873]. WO [1619]. WO/2009/066313 [1619]. Wood [460]. Word [90, 1109, 1103]. Wordnet [1102]. Words [1289, 1288, 72]. Work [621, 622, 623, 620, 624, 1353]. Workflow [303, 895, 1799, 1191, 637]. Workflows [301, 620]. Working [192, 593]. Workloads [35]. Workshop [1833, 1821, 1855, 1844, 1860, 1822, 1283, 1870, 1830, 1848, 1866, 1869, 1825, 1839, 1852, 1834, 1853, 1871, 1865, 1842]. Workshops [1861, 1851, 1868, 1850, 1819, 1820, 1858, 1816]. World [1554, 1510, 1763, 1258]. Worms [1060]. Woz [1312]. Wrapper [560, 1540]. WSN [826, 1801, 733, 473]. WSN-Aided [473]. WSNs [811, 1637]. Wyner [1437].

X [1470, 1352]. X-ray [1352]. XHTML [1753]. XII [1871]. XML [59, 1230]. XV [1845].

Years [629]. Yield [884]. Yourself [1234].

Zero [1438, 1730]. Zero-Coupon [1730]. Zero-Watermark [1438]. Zhengzhou [1826, 1827, 1828]. ZigBee [1631]. Ziv [1437].

References

Barker:2012:VPW

[1] Ken Barker. “valuing” privacy while exposing data utility. Lecture Notes in Computer Science, 6121:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

      accesspage/chapter/10.1007/978-3-642-25704-9_1.

Gray:2012:WBF

[2] Alasdair J. G. Gray. Whither BNCOD? The future of database and information systems research. Lecture Notes in Computer Science, 6121: 3–6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

      tronic). URL http://link.springer.com/chapter/10.1007/978-3-
      642-25704-9_2/.

Jeffery:2012:WBU

[3] Keith G. Jeffery. Whither BNCOD and the UK database community. Lecture Notes in Computer Science, 6121:7–11, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-25704-9_3/.

Adaikkalavan:2012:ACU

[4] Raman Adaikkalavan and Sharma Chakravarthy. Access control using active rules. Lecture Notes in Computer Science, 6121:12–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25704-9_4/.

Chertov:2012:PGA

[5] Oleg Chertov and Dan Tavrov. Providing group anonymity using wavelet transform. Lecture Notes in Computer Science, 6121:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25704-9_5/.

Hamerly:2012:EMS

[6] Greg Hamerly and Greg Speegle. Efficient model selection for large-scale nearest-neighbor data mining. Lecture Notes in Computer Science, 6121: 37–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25704-9_6/.

Islam:2012:END

[7] Md Zahidul Islam. EXPLORE: a novel decision tree classification algorithm. Lecture Notes in Computer Science, 6121:55–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25704-9_7/.

Kalantari:2012:CPO

[8] Reza Kalantari and Christopher H. Bryant. Comparing the performance of object and object relational database systems on objects of varying complexity. Lecture Notes in Computer Science, 6121:72–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25704-9_8/.

Kiran:2012:EAM

[9] R. Uday Kiran and Polepalli Krishna Reddy. An efficient approach to mine rare association rules using maximum items’ support constraints. Lecture Notes in Computer Science, 6121:84–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25704-9_9/.

Pozzani:2012:DST

[10] Gabriele Pozzani and Esteban Zimányi. Defining spatio-temporal granularities for raster data. Lecture Notes in Computer Science, 6121:96– 107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25704-9_10/.

Adaikkalavan:2012:LSD

[11] Raman Adaikkalavan. Load shedding in data stream management systems using application semantics. Lecture Notes in Computer Science, 6121:108–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25704-9_11/.

Haller:2012:TDP

[12] Klaus Haller. Test data provisioning for database-driven applications. Lecture Notes in Computer Science, 6121:113–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25704-9_12/.

Rehman:2012:EDT

[13] Nafees Ur Rehman and Marc H. Scholl. Enabling decision tree classification in database systems through pre-computation. Lecture Notes in Computer Science, 6121:118–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25704-9_13/.

Navarro-Galindo:2012:FRS

[14] José Luis Navarro-Galindo and José Samos Jiménez. Flexible range semantic annotations based on RDFa. Lecture Notes in Computer Science, 6121:122–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25704-9_14/.

Ou:2012:IFA

[15] Yi Ou and Theo Härder. Issues of flash-aware buffer management for database systems. Lecture Notes in Computer Science, 6121:127–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25704-9_15/.

Wang:2012:QFD

[16] Jianing Wang. A quality framework for data integration. Lecture Notes in Computer Science, 6121:131–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25704-9_16/.

Paterson:2012:UPA

[17] James Paterson, John N. Wilson, and Petra Leimich. Uses of peer assessment in database teaching and learning. Lecture Notes in Computer Science, 6121:135–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25704-9_17/.

Muzammal:2012:USP

[18] Muhammad Muzammal and Rajeev Raman. Uncertainty in sequential pattern mining. Lecture Notes in Computer Science, 6121:147–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25704-9_18/.

Wang:2012:DIM

[19] Jianing Wang. A data integration methodology and architecture with quality assessment functionality. Lecture Notes in Computer Science, 6121:151–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25704-9_19/.

Anonymous:2012:BMa

[20] Anonymous. Back matter. Lecture Notes in Computer Science, 6121: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-25704-9/1.

Anonymous:2012:FMa

[21] Anonymous. Front matter. Lecture Notes in Computer Science, 6121: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-25704-9/1.

Lange:2012:AAB

[22] Michael Lange and Tony Field. Accelerating agent-based ecosystem models using the Cell Broadband Engine. Lecture Notes in Computer Science, 6161:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24322-6_1/.

Keller:2012:PIT

[23] Jörg Keller and Ana Lucia Varbanescu. Performance impact of task mapping on the Cell BE multicore processor. Lecture Notes in Computer Science, 6161:13–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24322-6_2/.

Takayama:2012:PSC

[24] Motohiro Takayama and Ryuji Sakai. Parallelization strategy for CELL TV. Lecture Notes in Computer Science, 6161:24–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-24322-6_3/.

vanWerkhoven:2012:TUT

[25] Ben van Werkhoven, Jason Maassen, and Frank J. Seinstra. Towards user transparent parallel multimedia computing on GPU-clusters. Lecture Notes in Computer Science, 6161:28–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-24322-6_4/.

Kofsky:2012:IGP

[26] Stephen M. Kofsky, Daniel R. Johnson, John A. Stratton, Wen mei W. Hwu, and Sanjay J. Patel. Implementing a GPU programming model on a non-GPU accelerator architecture. Lecture Notes in Computer Science, 6161:40–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24322-6_5/.

AlUmairy:2012:USC

[27] Shams A. H. Al Umairy, Alexander S. van Amesfoort, and Irwan D. Setija. On the use of small 2D convolutions on GPUs. Lecture Notes in Computer Science, 6161:52–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-24322-6_6/.

Pavlovic:2012:CMS

[28] Milan Pavlovic, Yoav Etsion, and Alex Ramirez. Can manycores support the memory requirements of scientific applications? Lecture Notes in Computer Science, 6161:65–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-24322-6_7/.

Meder:2012:PIG

[29] David J. Meder and Walter F. Tichy. Parallelizing an index generator for desktop search. Lecture Notes in Computer Science, 6161:77– 85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24322-6_8/.

Kim:2012:CVM

[30] Martha A. Kim and Stephen A. Edwards. Computation vs. memory systems: Pinning down accelerator bottlenecks. Lecture Notes in Computer Science, 6161:86–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24322-6_9/.

Porto:2012:TEA

[31] João Porto, Guido Araujo, Edson Borin, and Youfeng Wu. Trace execution automata in dynamic binary translation. Lecture Notes in Computer Science, 6161:99–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24322-6_10/.

Souza:2012:IIM

[32] Maxwell Souza, Daniel Nicácio, and Guido Araújo. ISAMAP: Instruction mapping driven by dynamic binary translation. Lecture Notes in Computer Science, 6161:117–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-24322-6_11/.

Goodrum:2012:PPF

[33] Matthew A. Goodrum, Michael J. Trotter, Alla Aksel, Scott T. Acton, and Kevin Skadron. Parallelization of particle filter algorithms. Lecture Notes in Computer Science, 6161:139–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-24322-6_12/.

Moir:2012:WKA

[34] Mark Moir and Dan Nussbaum. What kinds of applications can benefit from transactional memory? Lecture Notes in Computer Science, 6161:150–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24322-6_13/.

Bienia:2012:CWU

[35] Christian Bienia and Kai Li. Characteristics of workloads using the pipeline programming model. Lecture Notes in Computer Science, 6161:161–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24322-6_14/.

Keys:2012:SEE

[36] Laura Keys, Suzanne Rivoire, and John D. Davis. The search for energyefficient building blocks for the data center. Lecture Notes in Computer Science, 6161:172–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24322-6_15/.

Ghosh:2012:KSB

[37] Sabyasachi Ghosh, Mark Redekopp, and Murali Annavaram. Knight-Shift: Shifting the I/O burden in datacenters to management processor for energy efficiency. Lecture Notes in Computer Science, 6161:183– 197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24322-6_16/.

Madan:2012:GPG

[38] Niti Madan, Alper Buyuktosunoglu, Pradip Bose, and Murali Annavaram. Guarded power gating in a multi-core setting. Lecture Notes in Computer Science, 6161:198–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-24322-6_17/.

Shafiee:2012:UPT

[39] Ali Shafiee, Narges Shahidi, and Amirali Baniasadi. Using partial tag comparison in low-power snoop-based chip multiprocessors. Lecture Notes in Computer Science, 6161:211–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-24322-6_18/.

Amur:2012:APE

[40] Hrishikesh Amur and Karsten Schwan. Achieving power-efficiency in clusters without distributed file system complexity. Lecture Notes in Computer Science, 6161:222–232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-24322-6_19/.

Hanson:2012:WCA

[41] Heather Hanson and Karthick Rajamani. What computer architects need to know about memory throttling. Lecture Notes in Computer Science, 6161:233–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24322-6_20/.

Bircher:2012:PPM

[42] William Lloyd Bircher and Lizy John. Predictive power management for multi-core processors. Lecture Notes in Computer Science, 6161:243–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24322-6_21.

Amit:2012:ISM

[43] Nadav Amit, Muli Ben-Yehuda, and Ben-Ami Yassour. IOMMU: Strategies for mitigating the IOTLB bottleneck. Lecture Notes in Computer Science, 6161:256–274, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-24322-6_22.

Nellans:2012:ISP

[44] David Nellans, Kshitij Sudan, Erik Brunvand, and Rajeev Balasubramonian. Improving server performance on multi-cores via selective off-loading of OS functionality. Lecture Notes in Computer Science, 6161:275–292, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24322-6_23.

Gorman:2012:PCE

[45] Mel Gorman and Patrick Healy. Performance characteristics of explicit superpage support. Lecture Notes in Computer Science, 6161:293–310, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24322-6_24.

Sabeghi:2012:IOS

[46] Mojtaba Sabeghi and Koen Bertels. Interfacing operating systems and polymorphic computing platforms based on the MOLEN programming paradigm. Lecture Notes in Computer Science, 6161:311–323, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24322-6_25.

Kleanthous:2012:EIT

[47] Marios Kleanthous, Yiannakis Sazeides, and Marios D. Dikaiakos. Extrinsic and intrinsic text cloning. Lecture Notes in Computer Science, 6161:324–340, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24322-6_26.

Tembey:2012:CCR

[48] Priyanka Tembey, Ada Gavrilovska, and Karsten Schwan. A case for coordinated resource management in heterogeneous multicore platforms. Lecture Notes in Computer Science, 6161:341–356, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24322-6_27.

Grot:2012:TAQ

[49] Boris Grot, Stephen W. Keckler, and Onur Mutlu. Topology-aware quality-of-service support in highly integrated chip multiprocessors. Lecture Notes in Computer Science, 6161:357–375, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24322-6_28.

Anonymous:2012:BMb

[50] Anonymous. Back matter. Lecture Notes in Computer Science, 6161: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-24322-6/1.

Anonymous:2012:FMb

[51] Anonymous. Front matter. Lecture Notes in Computer Science, 6161: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-24322-6/1.

Anantaram:2012:SCE

[52] C. Anantaram and Shailly Goyal. Semantic collation of enterprise data for effective information retrieval. Lecture Notes in Computer Science, 6411:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27872-3_1/.

Kawtrakul:2012:BKM

[53] Asanee Kawtrakul. Beyond knowledge management: Knowledge services innovation. Lecture Notes in Computer Science, 6411:9–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27872-3_2/.

Andres:2012:ODD

[54] Frederic Andres and Rajkumar Kannan. Ontology driven data management with topic maps. Lecture Notes in Computer Science, 6411:16– 23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27872-3_3/.

Rajimol:2012:WAP

[55] A. Rajimol and G. Raju. Web access pattern mining — a survey. Lecture Notes in Computer Science, 6411:24–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27872-3_4/.

Kumaresh:2012:GBS

[56] Nandhini Kumaresh and Balasundaram Sadhu Ramakrishnan. Graph based single document summarization. Lecture Notes in Computer Science, 6411:32–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27872-3_5/.

Kishore:2012:NTM

[57] Ayyalu Hariharan Nandhu Kishore and Mohan Saravanan. A novel textmining system for generating abstract from extracted summaries using anaphora resolution. Lecture Notes in Computer Science, 6411:36–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27872-3_6/.

Kumar:2012:MVS

[58] T. V. Vijay Kumar and Mohammad Haider. Materialized views selection for answering queries. Lecture Notes in Computer Science, 6411:44– 51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27872-3_7/.

Kanchana:2012:IXD

[59] R. Kanchana, Aishwarya Rajagopal, S. Kaavya, and R. Bakiyalakshmi. Integration of XML databases by schema restructuring. Lecture Notes in Computer Science, 6411:52–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27872-3_8/.

Chaudhari:2012:DMV

[60] Manoj S. Chaudhari and Chandrashekhar Dhote. Dynamic materialized view selection algorithm: a clustering approach. Lecture Notes in Computer Science, 6411:57–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27872-3_9/.

Korczak:2012:QFB

[61] Jerzy Korczak and Krzysztof Drelczuk. Quality of forecasting based on compressed high frequency time series. Lecture Notes in Computer Science, 6411:67–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27872-3_10/.

Rajput:2012:RVB

[62] Dharmveer Singh Rajput, P. K. Singh, and M. Bhattacharya. Reduct and variance based clustering of high dimensional dataset. Lecture Notes in Computer Science, 6411:75–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27872-3_11/.

Kumar:2012:PPN

[63] Santosh Kumar, Sukumar Nandi, and Santosh Biswas. Peer-to-peer network classification using nu-maximal margin spherical structured multiclass support vector machine. Lecture Notes in Computer Science, 6411: 80–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27872-3_12/.

Kumar:2012:AIR

[64] K. Susheel Kumar, Pradeep Kumar Saroj, and R. C. Tripathi. An adaptive image retrieval system with relevance feedback and clustering. Lecture Notes in Computer Science, 6411:85–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27872-3_13/.

Deshmukh:2012:UNT

[65] Sachin N. Deshmukh, Suresh C. Mehrotra, and Hardeep Singh. Using the normalization for typographic errors in numerals. Lecture Notes in Computer Science, 6411:93–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27872-3_14/.

Nallammal:2012:PEF

[66] N. Nallammal and V. Radha. Performance evaluation of face recognition based on PCA, LDA, ICA and hidden Markov model. Lecture Notes in Computer Science, 6411:96–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27872-3_15/.

Gangopadhyay:2012:CSP

[67] Bodhisattwa Gangopadhyay, Artur Arsenio, and Claudia Antunes. Comparative study of pattern mining techniques for network management system logs for convergent network. Lecture Notes in Computer Science, 6411:101–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27872-3_16/.

Schneider:2012:APD

[68] Stephan Schneider and Dirk Frosch-Wilke. Analysis patterns in dimensional data modeling. Lecture Notes in Computer Science, 6411:109– 116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27872-3_17/.

Castelino:2012:MSP

[69] Olivia Castelino, Preetham Kumar, and Srivatsa Maddodi. Mining single pass weighted pattern tree. Lecture Notes in Computer Science, 6411:117–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27872-3_18/.

Vanjulavalli:2012:OBC

[70] N. Vanjulavalli. Ontology based conceptual framework of E-learning system — the future perspective. Lecture Notes in Computer Science, 6411:125–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27872-3_19/.

Mohan:2012:DIM

[71] A. Raja Mohan and G. Arumugam. Developing Indian medicinal plant ontology using OWL and SWRL. Lecture Notes in Computer Science, 6411:131–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27872-3_20/.

Kumar:2012:DWU

[72] K. M. Anil Kumar and Suresha. Detection of Web users’ opinion from normal and short opinionated words. Lecture Notes in Computer Science, 6411:139–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27872-3_21.

Ivanova:2012:FAP

[73] Krassimira Ivanova, Peter Stanchev, Evgeniya Velikova, Koen Vanhoof, and Benoit Depaire. Features for art painting classification based on vector quantization of MPEG-7 descriptors. Lecture Notes in Computer Science, 6411:146–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27872-3_22.

Patil:2012:CBI

[74] Sanjay Patil and Sanjay Talbar. Content based image retrieval using various distance metrics. Lecture Notes in Computer Science, 6411:154–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27872-3_23.

Celia:2012:ECB

[75] B. Celia and I. Felci Rajam. An efficient content based image retrieval framework using machine learning techniques. Lecture Notes in Computer Science, 6411:162–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27872-3_24.

Gupta:2012:VDW

[76] Rolly Gupta and Anjana Gosain. Validating data warehouse quality metrics using PCA. Lecture Notes in Computer Science, 6411:170–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27872-3_25.

Palshikar:2012:UAS

[77] Girish Keshav Palshikar, Amrit Lal Ahuja, and Harrick M. Vin. Utilization analysis of servers in a data centre. Lecture Notes in Computer Science, 6411:173–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27872-3_26.

Rao:2012:SSA

[78] Burepalli V. S. Rao and Munaga V. N. K. Prasad. Subset selection approach for watermarking relational databases. Lecture Notes in Computer Science, 6411:181–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27872-3_27.

Shankar:2012:BDF

[79] Deepa D. Shankar, T. Gireeshkumar, K. Praveen, R. Jithin, and Ashji S. Raj. Block dependency feature based classification scheme for uncalibrated image steganalysis. Lecture Notes in Computer Science, 6411: 189–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27872-3_28.

Krishnamoorthy:2012:IRV

[80] R. Krishnamoorthy and J. Kalpana. Indexing and retrieval of visually similar images in the orthogonal polynomials transform domain. Lecture Notes in Computer Science, 6411:196–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27872-3_29.

Arunothayam:2012:HAR

[81] Mahendiran Arunothayam, Baskaran Ramachandran, and Dhavachelvan Ponnurangam. Human action recognition and localization in video at contextual level. Lecture Notes in Computer Science, 6411:204–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27872-3_30.

Hanumanthappa:2012:ADM

[82] M. Hanumanthappa, B. R. Prakash, and Manish Kumar. Applications of data mining in e-governance: a case study of Bhoomi project. Lecture Notes in Computer Science, 6411:208–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27872-3_31.

Gireeshkumar:2012:TNT

[83] T. Gireeshkumar, K. J. Poornaselvan, Sattviksharma, Gulshankumar, and R. Sreevathsan. Tracking of nose tip: an alternative for mouse. Lecture Notes in Computer Science, 6411:219–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27872-3_32.

Aravind:2012:WSB

[84] Alex Aravind and Viswanathan Manickam. A wireless sensors based feedback system for human body movement practices. Lecture Notes in Computer Science, 6411:226–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27872-3_33.

Panchal:2012:PAC

[85] Gaurang Panchal, Amit Ganatra, Y. P. Kosta, and Devyani Panchal. Performance analysis of classification techniques using different parameters. Lecture Notes in Computer Science, 6411:234–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27872-3_34.

Kumari:2012:REQ

[86] Madhu Kumari and Kamal K. Bharadwaj. Revenue estimation and quantification in sponsored search auctions: an inductive learning approach. Lecture Notes in Computer Science, 6411:242–244, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27872-3_35.

Pereira:2012:UMJ

[87] Anil L. Pereira, Mehdi Raoufi, and Jerrod C. Frost. Using MySQL and JDBC in new teaching methods for undergraduate database systems courses. Lecture Notes in Computer Science, 6411:245–248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27872-3_36.

Rekha:2012:NAM

[88] R. U. Rekha, M. Anand Kumar, V. Dhanalakshmi, K. P. Soman, and S. Rajendran. A novel approach to morphological generator for Tamil. Lecture Notes in Computer Science, 6411:249–251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27872-3_37.

Abeera:2012:MAM

[89] V. P. Abeera, S. Aparna, R. U. Rekha, M. Anand Kumar, V. Dhanalakshmi, and K. P. Soman. Morphological analyzer for Malayalam using machine learning. Lecture Notes in Computer Science, 6411:252–254, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27872-3_38.

Jindal:2012:IAA

[90] Karuna Jindal and Vishal Goyal. Improved algorithm for automatic word alignment for Hindi–Punjabi parallel corpus. Lecture Notes in Computer Science, 6411:255–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27872-3_39.

Kumar:2012:FEP

[91] Rajeev Kumar, Alok Ranjan, and Joydip Dhar. A fast and effective partitioning algorithm for document clustering. Lecture Notes in Computer Science, 6411:264–271, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27872-3_40.

Donavalli:2012:LRM

[92] Archana Donavalli, Manjeet Rege, Xumin Liu, and Kourosh Jafari-Khouzani. Low-rank matrix factorization and co-clustering algorithms for analyzing large data sets. Lecture Notes in Computer Science, 6411: 272–279, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27872-3_41.

Desai:2012:BFS

[93] Bhavik Desai, Pankaj Andhale, Manjeet Rege, and Qi Yu. Biclustering and feature selection techniques in bioinformatics. Lecture Notes in Computer Science, 6411:280–287, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27872-3_42.

Ranjan:2012:CSB

[94] Alok Ranjan, Rajeev Kumar, and Joydip Dhar. A comparative study between dynamic Web scripting languages. Lecture Notes in Computer Science, 6411:288–295, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27872-3_43.

Subbulakshmi:2012:MCS

[95] T. Subbulakshmi, S. Mercy Shalinie, and A. Ramamoorthi. Masquerader classification system with Linux command sequences using machine learning algorithms. Lecture Notes in Computer Science, 6411:296–302, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27872-3_44.

Tak:2012:POB

[96] Gaurav Kumar Tak and Shashikala Tapaswi. Parsing operations based approach towards phishing attacks. Lecture Notes in Computer Science, 6411:303–308, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27872-3_45.

Keshavamurthy:2012:EMF

[97] B. N. Keshavamurthy, Mitesh Sharma, and Durga Toshniwal. Efficient mining of frequent items coupled with weight and /or support over progressive databases. Lecture Notes in Computer Science, 6411:309–316, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27872-3_46.

Suryavanshi:2012:FDB

[98] Raghuraj Suryavanshi and Divakar Yadav. Formal development of Byzantine immune total order broadcast system using event-B. Lecture Notes in Computer Science, 6411:317–324, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27872-3_47.

Chandra:2012:ADF

[99] Girish Chandra and Divakar Yadav. Analyzing data flow in trustworthy electronic payment systems using event-B. Lecture Notes in Computer Science, 6411:325–332, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27872-3_48.

Kumar:2012:ATM

[100] Preetham Kumar and V. S. Ananthanarayana. Attribute-TID method for discovering sequence of attributes. Lecture Notes in Computer Science, 6411:333–340, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27872-3_49.

Anonymous:2012:BMc

[101] Anonymous. Back matter. Lecture Notes in Computer Science, 6411: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27872-3/1.

Anonymous:2012:FMc

[102] Anonymous. Front matter. Lecture Notes in Computer Science, 6411: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27872-3/1.

Banaji:2012:MPM

[103] Murad Banaji. P_{0}-matrix products of matrices. Lecture Notes in Computer Science, 6479:1–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28067-2_1/.

Gillis:2012:FMD

[104] Joris J. M. Gillis and Jan Van den Bussche. A formal model for databases in DNA. Lecture Notes in Computer Science, 6479:18–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28067-2_2/.

Graca:2012:EAH

[105] Ana Graça, Inês Lynce, and João Marques-Silva. Efficient and accurate haplotype inference by combining parsimony and pedigree information. Lecture Notes in Computer Science, 6479:38–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28067-2_3/.

Lemaire:2012:MMA

[106] François Lemaire and Asli Ürgüplü. MABSys: Modeling and analysis of biological systems. Lecture Notes in Computer Science, 6479:57– 75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28067-2_4/.

Vidal:2012:MSG

[107] Samuel Vidal, Michel Petitot, and François Boulier. Models of stochastic gene expression and Weyl algebra. Lecture Notes in Computer Science, 6479:76–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28067-2_5/.

Assar:2012:RCM

[108] Rodrigo Assar, Felipe A. Vargas, and David J. Sherman. Reconciling competing models: a case study of wine fermentation kinetics. Lecture Notes in Computer Science, 6479:98–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28067-2_6/.

Gong:2012:CMV

[109] Haijun Gong, Paolo Zuliani, and Anvesh Komuravelli. Computational modeling and verification of signaling pathways in cancer. Lecture Notes in Computer Science, 6479:117–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28067-2_7/.

Kahramanogullari:2012:CPE

[110] Ozan Kahramanoğulları, Ferenc Jordán, and Corrado Priami. Composability: Perspectives in ecological modeling. Lecture Notes in Computer Science, 6479:136–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28067-2_8/.

Nakatsui:2012:GPA

[111] Masahiko Nakatsui, Alexandre Sedoglavic, and François Lemaire. A general procedure for accurate parameter estimation in dynamic systems using new estimation errors. Lecture Notes in Computer Science, 6479:149– 166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28067-2_9/.

Ray:2012:APU

[112] Oliver Ray, Takehide Soh, and Katsumi Inoue. Analyzing pathways using ASP-based approaches. Lecture Notes in Computer Science, 6479:167– 183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28067-2_10/.

Anonymous:2012:BMd

[113] Anonymous. Back matter. Lecture Notes in Computer Science, 6479: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28067-2/1.

Anonymous:2012:FMd

[114] Anonymous. Front matter. Lecture Notes in Computer Science, 6479: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28067-2/1.

Russakovsky:2012:ALL

[115] Olga Russakovsky and Li Fei-Fei. Attribute learning in large-scale datasets. Lecture Notes in Computer Science, 6553:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-35749-7_1/.

Rohrbach:2012:CLS

[116] Marcus Rohrbach, Michael Stark, György Szarvas, and Bernt Schiele. Combining language sources and robust semantic relatedness for attribute-based knowledge transfer. Lecture Notes in Computer Science, 6553:15–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35749-7_2/.

Tsagkatakis:2012:SRD

[117] Grigorios Tsagkatakis and Andreas Savakis. Sparse representations and distance learning for attribute based category recognition. Lecture Notes in Computer Science, 6553:29–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35749-7_3/.

Cherniavsky:2012:SSL

[118] Neva Cherniavsky, Ivan Laptev, Josef Sivic, and Andrew Zisserman. Semi-supervised learning of facial attributes in video. Lecture Notes in Computer Science, 6553:43–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35749-7_4/.

Li:2012:OAS

[119] Li-Jia Li, Hao Su, Yongwhan Lim, and Li Fei-Fei. Objects as attributes for scene classification. Lecture Notes in Computer Science, 6553:57– 69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35749-7_5/.

Li:2012:GMC

[120] Congcong Li, Adarsh Kowdle, Ashutosh Saxena, and Tsuhan Chen. A generic model to compose vision modules for holistic scene understanding. Lecture Notes in Computer Science, 6553:70–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-35749-7_6/.

Prabhu:2012:AFL

[121] Utsav Prabhu, Keshav Seshadri, and Marios Savvides. Automatic facial landmark tracking in video sequences using Kalman filter assisted active shape models. Lecture Notes in Computer Science, 6553:86–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35749-7_7/.

Herzog:2012:TAS

[122] Dennis L. Herzog and Volker Krüger. Tracking in action space. Lecture Notes in Computer Science, 6553:100–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35749-7_8/.

Jiang:2012:TBA

[123] Zhuolin Jiang, Zhe Lin, and Larry S. Davis. A tree-based approach to integrated action localization, recognition and segmentation. Lecture Notes in Computer Science, 6553:114–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35749-7_9/.

Yildiz:2012:FMT

[124] Alparslan Yildiz and Yusuf Sinan Akgul. A fast method for tracking people with multiple cameras. Lecture Notes in Computer Science, 6553:128– 138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35749-7_10/.

Baak:2012:AEM

[125] Andreas Baak, Thomas Helten, Meinard Müller, Gerard Pons-Moll, and Bodo Rosenhahn. Analyzing and evaluating markerless motion tracking using inertial sensors. Lecture Notes in Computer Science, 6553:139– 152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35749-7_11/.

Noguchi:2012:SBS

[126] Akitsugu Noguchi and Keiji Yanai. A SURF-based spatio-temporal feature for feature-fusion-based action recognition. Lecture Notes in Computer Science, 6553:153–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35749-7_12/.

Yuan:2012:MLR

[127] Fei Yuan, Véronique Prinet, and Junsong Yuan. Middle-level representation for human activities recognition: The role of spatio-temporal relationships. Lecture Notes in Computer Science, 6553:168–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35749-7_13/.

Lan:2012:RAG

[128] Tian Lan, Yang Wang, Greg Mori, and Stephen N. Robinovitch. Retrieving actions in group contexts. Lecture Notes in Computer Science, 6553:181–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35749-7_14/.

Fanelli:2012:HFB

[129] Gabriele Fanelli, Angela Yao, Pierre-Luc Noel, Juergen Gall, and Luc Van Gool. Hough forest-based facial expression recognition from video sequences. Lecture Notes in Computer Science, 6553:195–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35749-7_15/.

Akakin:2012:SFE

[130] Hatice Çınar Akakın and Bülent Sankur. Spatiotemporal features for effective facial expression recognition. Lecture Notes in Computer Science, 6553:207–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35749-7_16/.

Klaser:2012:HFA

[131] Alexander Kläser, Marcin Marszałek, Cordelia Schmid, and Andrew Zisserman. Human focused action localization in video. Lecture Notes in Computer Science, 6553:219–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35749-7_17/.

Gonzalez:2012:HTH

[132] Matilde Gonzalez, Christophe Collet, and Rémi Dubot. Head tracking and hand segmentation during hand over face occlusion in sign language. Lecture Notes in Computer Science, 6553:234–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-35749-7_18/.

Tran:2012:FHP

[133] Khai Tran, Ioannis A. Kakadiaris, and Shishir K. Shah. Fusion of human posture features for continuous action recognition. Lecture Notes in Computer Science, 6553:244–257, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35749-7_19/.

Roussos:2012:HTA

[134] Anastasios Roussos, Stavros Theodorakis, Vassilis Pitsikalis, and Petros Maragos. Hand tracking and affine shape-appearance handshape subunits in continuous sign language recognition. Lecture Notes in Computer Science, 6553:258–272, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35749-7_20/.

Du:2012:HMT

[135] Wei Du and Justus Piater. Hand modeling and tracking for videobased sign language recognition by robust principal component analysis. Lecture Notes in Computer Science, 6553:273–285, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-35749-7_21.

Dreuw:2012:TBD

[136] Philippe Dreuw, Jens Forster, and Hermann Ney. Tracking benchmark databases for video-based sign language recognition. Lecture Notes in Computer Science, 6553:286–297, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-35749-7_22.

Rapantzikos:2012:DRS

[137] Konstantinos Rapantzikos, Yannis Avrithis, and Stefanos Kollias. Detecting regions from single scale edges. Lecture Notes in Computer Science, 6553:298–311, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-35749-7_23.

Hadfield:2012:GPE

[138] Simon Hadfield and Richard Bowden. Generalised pose estimation using depth. Lecture Notes in Computer Science, 6553:312–325, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-35749-7_24.

Metaxas:2012:ITC

[139] Dimitris Metaxas. Invited talk: Coupling deformable models and learning methods for nonverbal behavior analysis: Applications to deception, multi-cultural studies and ASL. Lecture Notes in Computer Science, 6553:326–328, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-35749-7_25.

Zhang:2012:RTS

[140] Zengyin Zhang, Minyoung Kim, Fernando de la Torre, and Wende Zhang. A real-time system for head tracking and pose estimation. Lecture Notes in Computer Science, 6553:329–341, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-35749-7_26.

Wang:2012:SLV

[141] Haijing Wang, Alexandra Stefan, Sajjad Moradi, Vassilis Athitsos, and Carol Neidle. A system for large vocabulary sign search. Lecture Notes in Computer Science, 6553:342–353, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-35749-7_27.

Anonymous:2012:BMe

[142] Anonymous. Back matter. Lecture Notes in Computer Science, 6553: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-35749-7/1.

Anonymous:2012:FMe

[143] Anonymous. Front matter. Lecture Notes in Computer Science, 6553: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35749-7/1.

Panagopoulos:2012:ESB

[144] Alexandros Panagopoulos, Chaohui Wang, Dimitris Samaras, and Nikos Paragios. Estimating shadows with the bright channel cue. Lecture Notes in Computer Science, 6554:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35740-4_1/.

Wang:2012:CCI

[145] Fan Wang and Roberto Manduchi. Color-constant information embedding. Lecture Notes in Computer Science, 6554:13–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-35740-4_2/.

Gupta:2012:BAF

[146] Mithun Das Gupta and Jing Xiao. Bi-affinity filter: a bilateral type filter for color images. Lecture Notes in Computer Science, 6554:27– 40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35740-4_3/.

Elbrandt:2012:PCC

[147] Tobias Elbrandt and Jörn Ostermann. Photometric color calibration of the joint monitor-camera response function. Lecture Notes in Computer Science, 6554:41–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35740-4_4/.

Rosman:2012:PAM

[148] Guy Rosman, Xue-Cheng Tai, Lorina Dascal, and Ron Kimmel. Polyakov action minimization for efficient color image processing. Lecture Notes in Computer Science, 6554:50–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35740-4_5/.

Chu:2012:CIS

[149] Dung Manh Chu and Arnold W. M. Smeulders. Color invariant SURF in discriminative object tracking. Lecture Notes in Computer Science, 6554: 62–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35740-4_6/.

Eibenberger:2012:NBA

[150] Eva Eibenberger and Elli Angelopoulou. The narrow-band assumption in log-chromaticity space. Lecture Notes in Computer Science, 6554:76– 89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35740-4_7/.

Paggetti:2012:LBA

[151] Giulia Paggetti and Gloria Menegaz. Is light blue (azzurro) color name universal in the Italian language? Lecture Notes in Computer Science, 6554:90–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35740-4_8/.

Shin:2012:TCD

[152] Yong-Ho Shin, Min-Gyu Park, Young-Sun Jeon, Young-Su Moon, Shi-Hwa Lee, and Kuk-Jin Yoon. Tone correction with dynamic objects for seamless image mosaic. Lecture Notes in Computer Science, 6554:104– 117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35740-4_9/.

Bremond:2012:SMH

[153] Roland Brémond, Josselin Petit, and Jean-Philippe Tarel. Saliency maps of high dynamic range images. Lecture Notes in Computer Science, 6554:118–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35740-4_10/.

Mansfield:2012:VMI

[154] Alex Mansfield, Peter Gehler, Luc Van Gool, and Carsten Rother. Visibility maps for improving seam carving. Lecture Notes in Computer Science, 6554:131–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35740-4_11/.

Kav-Venaki:2012:FR

[155] Eitam Kav-Venaki and Shmuel Peleg. Feedback retargeting. Lecture Notes in Computer Science, 6554:145–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35740-4_12/.

Chamaret:2012:HMR

[156] Christel Chamaret, Olivier Le Meur, Philippe Guillotel, and Jean-Claude Chevet. How to measure the relevance of a retargeting approach? Lecture Notes in Computer Science, 6554:156–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35740-4_13/.

Kim:2012:MSE

[157] Hansung Kim and Adrian Hilton. 3D modelling of static environments using multiple spherical stereo. Lecture Notes in Computer Science, 6554:169–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35740-4_14/.

Jancosek:2012:HFM

[158] Michal Jancosek and Tomas Pajdla. Hallucination-free multi-view stereo. Lecture Notes in Computer Science, 6554:184–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-35740-4_15/.

Ackermann:2012:REE

[159] Jens Ackermann, Martin Ritz, André Stork, and Michael Goesele. Removing the example from example-based photometric stereo. Lecture Notes in Computer Science, 6554:197–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35740-4_16/.

Kowdle:2012:IIC

[160] Adarsh Kowdle, Dhruv Batra, Wen-Chao Chen, and Tsuhan Chen. iModel: Interactive co-segmentation for object of interest 3D modeling. Lecture Notes in Computer Science, 6554:211–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-35740-4_17/.

Wang:2012:ARL

[161] Quan Wang and Suya You. Automatic registration of large-scale multisensor datasets. Lecture Notes in Computer Science, 6554:225–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35740-4_18/.

Ladikos:2012:RGO

[162] Alexander Ladikos, Edmond Boyer, Nassir Navab, and Slobodan Ilic. Region graphs for organizing image collections. Lecture Notes in Computer Science, 6554:239–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35740-4_19/.

Habbecke:2012:ARO

[163] Martin Habbecke and Leif Kobbelt. Automatic registration of oblique aerial images with cadastral maps. Lecture Notes in Computer Science, 6554:253–266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35740-4_20/.

Sinha:2012:MSL

[164] Sudipta N. Sinha, Drew Steedly, and Richard Szeliski. A multi-stage linear approach to structure from motion. Lecture Notes in Computer Science, 6554:267–281, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-35740-4_21.

Steffen:2012:RBA

[165] Richard Steffen, Jan-Michael Frahm, and Wolfgang Förstner. Relative bundle adjustment based on trifocal constraints. Lecture Notes in Computer Science, 6554:282–295, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-35740-4_22.

Bodensteiner:2012:ASI

[166] Christoph Bodensteiner, Marcus Hebel, and Michael Arens. Accurate single image multi-modal camera pose estimation. Lecture Notes in Computer Science, 6554:296–309, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-35740-4_23.

Weyand:2012:ETA

[167] Tobias Weyand, Jan Hosang, and Bastian Leibe. An evaluation of two automatic landmark building discovery algorithms for city reconstruction. Lecture Notes in Computer Science, 6554:310–323, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-35740-4_24.

Andalo:2012:VPD

[168] Fernanda A. Andaló, Gabriel Taubin, and Siome Goldenstein. Vanishing point detection by segment clustering on the projective space. Lecture Notes in Computer Science, 6554:324–337, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-35740-4_25.

Xie:2012:EEI

[169] Hongtao Xie, Ke Gao, Yongdong Zhang, Jintao Li, Yizhi Liu, and Huamin Ren. Effective and efficient image copy detection based on GPU. Lecture Notes in Computer Science, 6554:338–349, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-35740-4_26.

Fulkerson:2012:RQS

[170] Brian Fulkerson and Stefano Soatto. Really quick shift: Image segmentation on a GPU. Lecture Notes in Computer Science, 6554:350–358, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
35740-4_27.

Friborg:2012:GAL

[171] Rune Møllegaard Friborg, Søren Hauberg, and Kenny Erleben. GPU accelerated likelihoods for stereo-based articulated tracking. Lecture Notes in Computer Science, 6554:359–371, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-35740-4_28.

Gwosdek:2012:HEG

[172] Pascal Gwosdek, Henning Zimmer, Sven Grewenig, Andrés Bruhn, and Joachim Weickert. A highly efficient GPU implementation for variational optic flow based on the Euler–Lagrange framework. Lecture Notes in Computer Science, 6554:372–383, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-35740-4_29.

Tzevanidis:2012:MVT

[173] K. Tzevanidis, X. Zabulis, T. Sarmis, P. Koutlemanis, N. Kyriazis, and A. Argyros. From multiple views to textured 3D meshes: a GPUpowered approach. Lecture Notes in Computer Science, 6554:384–397, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
35740-4_30.

Zhu:2012:CDS

[174] Ke Zhu, Matthias Butenuth, and Pablo d’Angelo. Comparison of dense stereo using CUDA. Lecture Notes in Computer Science, 6554:398–410, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
35740-4_31.

Wang:2012:EAR

[175] Yi-Chu Wang, Bryan Donyanavard, and Kwang-Ting (Tim) Cheng. Energy-aware real-time face recognition system on mobile CPU–GPU platform. Lecture Notes in Computer Science, 6554:411–422, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
35740-4_32.

Choudhary:2012:PTB

[176] Siddharth Choudhary, Shubham Gupta, and P. J. Narayanan. Practical time bundle adjustment for 3D reconstruction on the GPU. Lecture Notes in Computer Science, 6554:423–435, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-35740-4_33.

vandeSande:2012:AVC

[177] Koen E. A. van de Sande, Theo Gevers, and Cees G. M. Snoek. Accelerating visual categorization with the GPU. Lecture Notes in Computer Science, 6554:436–449, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-35740-4_34.

Stuhmer:2012:PGT

[178] Jan Stühmer, Stefan Gumhold, and Daniel Cremers. Parallel generalized thresholding scheme for live dense geometry from a handheld camera. Lecture Notes in Computer Science, 6554:450–462, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-35740-4_35.

Johnson:2012:FOL

[179] Tim Johnson, Pierre Fite-Georgel, Rahul Raguram, and Jan-Michael Frahm. Fast organization of large photo collections using CUDA. Lecture Notes in Computer Science, 6554:463–476, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-35740-4_36.

Anonymous:2012:BMf

[180] Anonymous. Back matter. Lecture Notes in Computer Science, 6554: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-35740-4/1.

Anonymous:2012:FMf

[181] Anonymous. Front matter. Lecture Notes in Computer Science, 6554: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35740-4/1.

Savarimuthu:2012:ASU

[182] Sharmila Savarimuthu, Maryam Purvis, and Martin K. Purvis. Altruistic sharing using tags. Lecture Notes in Computer Science, 6573:1– 12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31809-2_1/.

Manavalan:2012:EPK

[183] Priyadarshini Manavalan and Munindar P. Singh. Emerging properties of knowledge sharing referral networks: Considerations of effectiveness and fairness. Lecture Notes in Computer Science, 6573:13–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31809-2_2/.

Mari:2012:EPP

[184] Marco Mari, Agostino Poggi, Michele Tomaiuolo, and Paola Turci. Enhancing peer-to-peer applications with multi-agent systems. Lecture Notes in Computer Science, 6573:24–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31809-2_3/.

Kafali:2012:ISO

[185] Özgür Kafalı and Pınar Yolum. Improving self-organized resource allocation with effective communication. Lecture Notes in Computer Science, 6573:35–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31809-2_4/.

Pommier:2012:DMP

[186] Hugo Pommier and François Bourdon. Data mobility in peer-to-peer systems to improve robustness. Lecture Notes in Computer Science, 6573: 47–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31809-2_5/.

Sahli:2012:TAB

[187] Nabil Sahli, Gabriele Lenzini, and Henk Eertink. Trustworthy agentbased recommender system in a mobile P2P environment. Lecture Notes in Computer Science, 6573:59–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31809-2_6/.

Lopes:2012:EAA

[188] António Luís Lopes and Luís Miguel Botelho. Efficient algorithms for agent-based semantic resource discovery. Lecture Notes in Computer Science, 6573:71–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31809-2_7/.

Kimura:2012:SSO

[189] Kousaku Kimura, Satoshi Amamiya, Tsunenori Mine, and Makoto Amamiya. A semi-structured overlay network for large-scale peer-to-peer systems. Lecture Notes in Computer Science, 6573:83–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31809-2_8/.

Brazier:2012:FEM

[190] Frances Brazier, Elth Ogston, and Martijn Warnier. The future of energy markets and the challenge of decentralized self-management. Lecture Notes in Computer Science, 6573:95–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31809-2_9/.

Cabri:2012:ARC

[191] Giacomo Cabri. Agent roles for context-aware P2P systems. Lecture Notes in Computer Science, 6573:104–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31809-2_10/.

Bergamaschi:2012:WDE

[192] Sonia Bergamaschi, Francesco Guerra, Federica Mandreoli, and Maurizio Vincini. Working in a dynamic environment: The NeP4B approach as a MAS. Lecture Notes in Computer Science, 6573:115–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31809-2_11/.

Peleg:2012:APP

[193] Yoni Peleg and Jeffrey S. Rosenschein. Agents and peer-to-peer computing: Towards P2P-based resource allocation in competitive environments. Lecture Notes in Computer Science, 6573:129–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31809-2_12/.

Ebadi:2012:CPN

[194] Toktam Ebadi, Maryam Purvis, and Martin K. Purvis. A colored Petri net model to represent the interactions between a set of cooperative agents. Lecture Notes in Computer Science, 6573:141–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31809-2_13/.

Anonymous:2012:BMg

[195] Anonymous. Back matter. Lecture Notes in Computer Science, 6573: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31809-2/1.

Anonymous:2012:FMg

[196] Anonymous. Front matter. Lecture Notes in Computer Science, 6573: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31809-2/1.

Anonymous:2012:FMh

[197] Anonymous. Front matter. Lecture Notes in Computer Science, 6599: 1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28939-2/1/1.

Pechoucek:2012:TSA

[198] Michal Pěchouček, Michal Jakob, and Peter Novák. Towards simulationaided design of multi-agent systems. Lecture Notes in Computer Science, 6599:3–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28939-2_1/.

Anonymous:2012:FMi

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28939-2/2/1.

Broekens:2012:RLH

[200] Joost Broekens, Koen Hindriks, and Pascal Wiggers. Reinforcement learning as heuristic for action-rule preferences. Lecture Notes in Computer Science, 6599:25–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28939-2_2/.

vanRiemsdijk:2012:TRP

[201] M. Birna van Riemsdijk and Neil Yorke-Smith. Towards reasoning with partial goal satisfaction in intelligent agents. Lecture Notes in Computer Science, 6599:41–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28939-2_3/.

Anonymous:2012:FMj

[202] Anonymous. Front matter. Lecture Notes in Computer Science, 6599: 61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28939-2/3/1.

Jordan:2012:EAO

[203] Howell R. Jordan and Rem Collier. Evaluating agent-oriented programs: Towards multi-paradigm metrics. Lecture Notes in Computer Science, 6599:63–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28939-2_4/.

Kiss:2012:AIJ

[204] Daniel Kiss, Neil Madden, and Brian Logan. Atomic intentions in Jason^{+}. Lecture Notes in Computer Science, 6599:79–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28939-2_5/.

Carr:2012:SSO

[205] Hugo Carr, Alexander Artikis, and Jeremy Pitt. Software support for organised adaptation. Lecture Notes in Computer Science, 6599:96–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28939-2_6/.

Anonymous:2012:FMk

[206] Anonymous. Front matter. Lecture Notes in Computer Science, 6599: 117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28939-2/4/1.

Ricci:2012:APA

[207] Alessandro Ricci, Andrea Santi, and Michele Piunti. Action and perception in agent programming languages: From exogenous to endogenous environments. Lecture Notes in Computer Science, 6599:119–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28939-2_7/.

Behrens:2012:IAE

[208] Tristan Behrens, Koen V. Hindriks, Rafael H. Bordini, Lars Braubach, and Mehdi Dastani. An interface for agent-environment interaction. Lecture Notes in Computer Science, 6599:139–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28939-2_8/.

Anonymous:2012:BMh

[209] Anonymous. Back matter. Lecture Notes in Computer Science, 6599: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28939-2/1.

Anonymous:2012:FMl

[210] Anonymous. Front matter. Lecture Notes in Computer Science, 6599: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28939-2/1.

Munoz-Merida:2012:SSN

[211] Antonio Muñoz-Mérida and Javier Ríos. Statistical significance for NGS reads similarities. Lecture Notes in Computer Science, 6620:1– 7, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28062-7_1/.

Fernandez:2012:AAO

[212] Fernando Muñiz Fernandez and Angel Carreño Torres. Application of array-oriented scientific data formats (NetCDF) to genotype data, GWASpi as an example. Lecture Notes in Computer Science, 6620:8– 20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28062-7_2/.

Almeida:2012:ORA

[213] João Almeida, João Tiple, and Mário Ramirez. An ontology and a REST API for sequence based microbial typing data. Lecture Notes in Computer Science, 6620:21–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28062-7_3/.

Sanchez:2012:MMI

[214] Alex Sánchez and José Fernández-Real. Multivariate methods for the integration and visualization of omics data. Lecture Notes in Computer Science, 6620:29–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28062-7_4/.

Lopes:2012:HAI

[215] Pedro Lopes and José Luís Oliveira. A holistic approach for integrating genomic variation information. Lecture Notes in Computer Science, 6620: 42–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28062-7_5/.

Garcia-Godoy:2012:SKI

[216] María Jesús García-Godoy and Ismael Navas-Delgado. SB–KOM: Integration of pathway information with BioPax. Lecture Notes in Computer Science, 6620:50–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28062-7_6/.

Fernandez-Pozo:2012:GVW

[217] Noé Fernández-Pozo and Darío Guerrero-Fernández. GENote v. β: a Web tool prototype for annotation of unfinished sequences in nonmodel eukaryotes. Lecture Notes in Computer Science, 6620:66–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28062-7_7/.

Sebastian:2012:ISI

[218] Álvaro Sebastián and Carlos P. Cantalapiedra. Interface similarity improves comparison of DNA-binding proteins: The homeobox example. Lecture Notes in Computer Science, 6620:72–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28062-7_8/.

Gonzalez-Vallinas:2012:PFT

[219] Juan González-Vallinas and Sonja Althammer. Pyicos: a flexible tool library for analyzing protein-nucleotide interactions with mapped reads from deep sequencing. Lecture Notes in Computer Science, 6620:83– 88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28062-7_9/.

Morilla:2012:NAF

[220] Ian Morilla, Miguel A. Medina, and Juan A. G. Ranea. Novel angiogenic functional targets predicted through “dark matter” assessment in protein networks. Lecture Notes in Computer Science, 6620:89–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28062-7_10/.

Karlsson:2012:JMS

[221] Johan Karlsson and Oswaldo Trelles. jORCA and Magallanes sailing together towards integration of Web services. Lecture Notes in Computer Science, 6620:94–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28062-7_11/.

Fernandez:2012:IWS

[222] José M. Fernández and Robert Hoffmann. iHOP Web services family. Lecture Notes in Computer Science, 6620:102–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28062-7_12/.

Fernandez:2012:BSD

[223] José M. Fernández and Alfonso Valencia. Bioinformatic software developments in Spain. Lecture Notes in Computer Science, 6620:108– 120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28062-7_13/.

Anonymous:2012:BMi

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28062-7/1.

Anonymous:2012:FMm

[225] Anonymous. Front matter. Lecture Notes in Computer Science, 6620: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28062-7/1.

Duits:2012:FED

[226] Remco Duits, Tom C. J. Dela Haije, Arpan Ghosh, Eric Creusen, and Anna Vilanova. Fiber enhancement in diffusion-weighted MRI. Lecture Notes in Computer Science, 6667:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-24785-9_1/.

Creusen:2012:NSL

[227] Eric J. Creusen, Remco Duits, and Tom C. J. Dela Haije. Numerical schemes for linear and non-linear enhancement of DW–MRI. Lecture Notes in Computer Science, 6667:14–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-24785-9_2/.

Mainberger:2012:OST

[228] Markus Mainberger, Sebastian Hoffmann, Joachim Weickert, and Ching Hoo Tang. Optimising spatial and tonal data for homogeneous diffusion inpainting. Lecture Notes in Computer Science, 6667:26–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24785-9_3/.

Morigi:2012:NSF

[229] Serena Morigi, Marco Rucci, and Fiorella Sgallari. Nonlocal surface fairing. Lecture Notes in Computer Science, 6667:38–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-24785-9_4/.

Teuber:2012:NFR

[230] Tanja Teuber and Annika Lang. Nonlocal filters for removing multiplicative noise. Lecture Notes in Computer Science, 6667:50–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24785-9_5/.

Schwarzkopf:2012:VNA

[231] Andreas Schwarzkopf, Thomas Kalbe, Chandrajit Bajaj, Arjan Kuijper, and Michael Goesele. Volumetric nonlinear anisotropic diffusion on GPUs. Lecture Notes in Computer Science, 6667:62–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-24785-9_6/.

Frick:2012:SMS

[232] Klaus Frick and Philipp Marnitz. A statistical multiresolution strategy for image reconstruction. Lecture Notes in Computer Science, 6667:74– 85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24785-9_7/.

Chan:2012:VAE

[233] Raymond Chan, Mila Nikolova, and You-Wei Wen. A variational approach for exact histogram specification. Lecture Notes in Computer Science, 6667:86–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24785-9_8/.

Bauer:2012:JTI

[234] Sebastian Bauer, Benjamin Berkels, Joachim Hornegger, and Martin Rumpf. Joint ToF image denoising and registration with a CT surface in radiation therapy. Lecture Notes in Computer Science, 6667:98– 109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24785-9_9/.

Nikolova:2012:EFD

[235] Mila Nikolova. Either fit to data entries or locally to prior: The minimizers of objectives with nonsmooth nonconvex data fidelity and regularization. Lecture Notes in Computer Science, 6667:110–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-24785-9_10/.

Yuan:2012:SCO

[236] Jing Yuan, Juan Shi, Xue-Cheng Tai, and Yuri Boykov. A study on convex optimization approaches to image fusion. Lecture Notes in Computer Science, 6667:122–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24785-9_11/.

Wetzler:2012:EBF

[237] Aaron Wetzler and Ron Kimmel. Efficient Beltrami flow in patchspace. Lecture Notes in Computer Science, 6667:134–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-24785-9_12/.

Duan:2012:FAL

[238] Yuping Duan, Yu Wang, Xue-Cheng Tai, and Jooyoung Hahn. A fast augmented Lagrangian method for Euler’s elastica model. Lecture Notes in Computer Science, 6667:144–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-24785-9_13/.

Klapp:2012:DSV

[239] Iftach Klapp, Nir Sochen, and David Mendlovic. Deblurring spacevariant blur by adding noisy image. Lecture Notes in Computer Science, 6667:157–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24785-9_14/.

Hahn:2012:FAE

[240] Jooyoung Hahn, Ginmo J. Chung, Yu Wang, and Xue-Cheng Tai. Fast algorithms for p-elastica energy with the application to image inpainting and curve reconstruction. Lecture Notes in Computer Science, 6667:169– 182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24785-9_15/.

Sochen:2012:BMS

[241] Nir Sochen and Leah Bar. The Beltrami–Mumford–Shah functional. Lecture Notes in Computer Science, 6667:183–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-24785-9_16/.

Bertaccini:2012:ANA

[242] Daniele Bertaccini, Raymond H. Chan, Serena Morigi, and Fiorella Sgallari. An adaptive norm algorithm for image restoration. Lecture Notes in Computer Science, 6667:194–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-24785-9_17/.

Lenzen:2012:VID

[243] Frank Lenzen, Florian Becker, Jan Lellmann, Stefania Petra, and Christoph Schnörr. Variational image denoising with adaptive constraint sets. Lecture Notes in Computer Science, 6667:206–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-24785-9_18/.

Liu:2012:SDI

[244] Jun Liu, Xue cheng Tai, Haiyang Huang, and Zhongdan Huan. Simultaneous denoising and illumination correction via local data-fidelity and nonlocal regularization. Lecture Notes in Computer Science, 6667:218– 230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24785-9_19/.

Deledalle:2012:ANL

[245] Charles-Alban Deledalle, Vincent Duval, and Joseph Salmon. Anisotropic non-local means with spatially adaptive patch shapes. Lecture Notes in Computer Science, 6667:231–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-24785-9_20/.

Hong:2012:ESP

[246] Byung-Woo Hong, Kangyu Ni, and Stefano Soatto. Entropy-scale profiles for texture segmentation. Lecture Notes in Computer Science, 6667:243–254, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24785-9_21.

Jung:2012:NLA

[247] Miyoun Jung, Gabriel Peyré, and Laurent D. Cohen. Non-local active contours. Lecture Notes in Computer Science, 6667:255–266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24785-9_22.

Tari:2012:MAT

[248] Sibel Tari and Murat Genctav. From a modified Ambrosio–Tortorelli to a randomized part hierarchy tree. Lecture Notes in Computer Science, 6667:267–278, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24785-9_23.

Yuan:2012:CMF

[249] Jing Yuan, Egil Bae, Yuri Boykov, and Xue-Cheng Tai. A continuous max-flow approach to minimal partitions with label cost prior. Lecture Notes in Computer Science, 6667:279–290, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24785-9_24.

Wang:2012:RED

[250] Li-Lian Wang, Yuying Shi, and Xue-Cheng Tai. Robust edge detection using Mumford–Shah model and binary level set method. Lecture Notes in Computer Science, 6667:291–301, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24785-9_25.

Sakai:2012:BSE

[251] Tomoya Sakai, Haruhiko Nishiguchi, Hayato Itoh, and Atsushi Imiya. Bifurcation of segment edge curves in scale space. Lecture Notes in Computer Science, 6667:302–313, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24785-9_26.

Werlberger:2012:EMN

[252] Manuel Werlberger, Markus Unger, Thomas Pock, and Horst Bischof. Efficient minimization of the non-local Potts model. Lecture Notes in Computer Science, 6667:314–325, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24785-9_27.

Berkels:2012:SDP

[253] Benjamin Berkels, Marc Kotowski, Martin Rumpf, and Carlo Schaller. Sulci detection in photos of the human cortex based on learned discriminative dictionaries. Lecture Notes in Computer Science, 6667:326–337, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24785-9_28.

Hochbaum:2012:EET

[254] Dorit S. Hochbaum. An efficient and effective tool for image segmentation, total variations and regularization. Lecture Notes in Computer Science, 6667:338–349, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-24785-9_29.

Li:2012:SSI

[255] Yan Li, David M. J. Tax, and Marco Loog. Supervised scale-invariant segmentation (and detection). Lecture Notes in Computer Science, 6667: 350–361, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24785-9_30.

Rouchdy:2012:GVS

[256] Youssef Rouchdy and Laurent D. Cohen. A geodesic voting shape prior to constrain the level set evolution for the segmentation of tubular trees. Lecture Notes in Computer Science, 6667:362–373, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24785-9_31.

Welk:2012:AAC

[257] Martin Welk. Amoeba active contours. Lecture Notes in Computer Science, 6667:374–385, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-24785-9_32.

Ahipo:2012:HSC

[258] Y. Ahipo, D. Auroux, L. D. Cohen, and M. Masmoudi. A hybrid scheme for contour detection and completion based on topological gradient and fast marching algorithms- application to inpainting and segmentation. Lecture Notes in Computer Science, 6667:386–397, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24785-9_33.

Peles:2012:SQM

[259] David Peles and Michael Lindenbaum. A segmentation quality measure based on rich descriptors and classification methods. Lecture Notes in Computer Science, 6667:398–410, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24785-9_34.

Cai:2012:FBA

[260] Xiaohao Cai, Raymond H. Chan, Serena Morigi, and Fiorella Sgallari. Framelet-based algorithm for segmentation of tubular structures. Lecture Notes in Computer Science, 6667:411–422, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24785-9_35.

Schmitzer:2012:WCC

[261] Bernhard Schmitzer and Christoph Schnörr. Weakly convex coupling continuous cuts and shape priors. Lecture Notes in Computer Science, 6667:423–434, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24785-9_36.

Rabin:2012:WBA

[262] Julien Rabin, Gabriel Peyré, Julie Delon, and Marc Bernot. Wasserstein barycenter and its application to texture mixing. Lecture Notes in Computer Science, 6667:435–446, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24785-9_37.

Gwosdek:2012:TFG

[263] Pascal Gwosdek, Sven Grewenig, Andrés Bruhn, and Joachim Weickert. Theoretical foundations of Gaussian convolution by extended box filtering. Lecture Notes in Computer Science, 6667:447–458, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24785-9_38.

Feigin:2012:HDI

[264] Micha Feigin, Dan Feldman, and Nir Sochen. From high definition image to low space optimization. Lecture Notes in Computer Science, 6667:459–470, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24785-9_39.

Aflalo:2012:MGD

[265] Yonathan Aflalo and Ron Kimmel. Measuring geodesic distances via the uniformization theorem. Lecture Notes in Computer Science, 6667: 471–482, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24785-9_40.

Batard:2012:PAE

[266] Thomas Batard and Nir Sochen. Polyakov action on (ρ, G)-equivariant functions application to color image regularization. Lecture Notes in Computer Science, 6667:483–494, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24785-9_41.

Shi:2012:CMS

[267] Juan Shi, Min Wan, Xue-Cheng Tai, and Desheng Wang. Curvature minimization for surface reconstruction with features. Lecture Notes in Computer Science, 6667:495–507, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24785-9_42.

Nikolova:2012:SWS

[268] Mila Nikolova. Should we search for a global minimizer of least squares regularized with an ℓ_{0} penalty to get the exact solution of an under determined linear system? Lecture Notes in Computer Science, 6667:508–519, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24785-9_43.

Gallego:2012:WSC

[269] Guillermo Gallego, Anthony Yezzi, Francesco Fedele, and Alvise Benetazzo. Weak statistical constraints for variational stereo imaging of oceanic waves. Lecture Notes in Computer Science, 6667:520–531, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24785-9_44.

Hagenburg:2012:NSH

[270] Kai Hagenburg, Michael Breuß, Joachim Weickert, and Oliver Vogel. Novel schemes for hyperbolic PDEs using osmosis filters from visual computing. Lecture Notes in Computer Science, 6667:532–543, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24785-9_45.

Luxenburger:2012:FPB

[271] Andreas Luxenburger, Henning Zimmer, Pascal Gwosdek, and Joachim Weickert. Fast PDE-based image analysis in your pocket. Lecture Notes in Computer Science, 6667:544–555, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24785-9_46.

Lee:2012:STS

[272] Deokwoo Lee and Hamid Krim. A sampling theorem for a 2D surface. Lecture Notes in Computer Science, 6667:556–567, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24785-9_47.

Graf:2012:QNM

[273] Manuel Gräf, Daniel Potts, and Gabriele Steidl. Quadrature nodes meet stippling dots. Lecture Notes in Computer Science, 6667:568–579, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24785-9_48.

Wang:2012:DMD

[274] Chaohui Wang, Michael M. Bronstein, Alexander M. Bronstein, and Nikos Paragios. Discrete minimum distortion correspondence problems for non-rigid shape matching. Lecture Notes in Computer Science, 6667:580–591, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24785-9_49.

Pokrass:2012:CLA

[275] Jonathan Pokrass, Alexander M. Bronstein, and Michael M. Bronstein. A correspondence-less approach to matching of deformable shapes. Lecture Notes in Computer Science, 6667:592–603, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24785-9_50.

Raviv:2012:HMN

[276] Dan Raviv, Anastasia Dubrovina, and Ron Kimmel. Hierarchical matching of non-rigid shapes. Lecture Notes in Computer Science, 6667: 604–615, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24785-9_51.

Kovnatsky:2012:PHK

[277] Artiom Kovnatsky, Michael M. Bronstein, Alexander M. Bronstein, and Ron Kimmel. Photometric heat kernel signatures. Lecture Notes in Computer Science, 6667:616–627, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24785-9_52.

Yi:2012:HAM

[278] Sheng Yi, Hamid Krim, and Larry K. Norris. Human activity modeling as Brownian motion on shape manifold. Lecture Notes in Computer Science, 6667:628–639, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24785-9_53.

Mikula:2012:CEA

[279] Karol Mikula and Jozef Urbán. 3D curve evolution algorithm with tangential redistribution for a fully automatic finding of an ideal camera path in virtual colonoscopy. Lecture Notes in Computer Science, 6667:640–652, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24785-9_54.

Dimitrov:2012:DII

[280] Pavel Dimitrov, Matthew Lawlor, and Steven W. Zucker. Distance images and intermediate-level vision. Lecture Notes in Computer Science, 6667:653–664, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24785-9_55.

Hooda:2012:SPA

[281] Amit Hooda, Michael M. Bronstein, Alexander M. Bronstein, and Radu P. Horaud. Shape palindromes: Analysis of intrinsic symmetries in 2D articulated shapes. Lecture Notes in Computer Science, 6667:665–676, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24785-9_56.

Sommer:2012:KBE

[282] Stefan Sommer, François Lauze, Mads Nielsen, and Xavier Pennec. Kernel bundle EPDiff: Evolution equations for multi-scale diffeomorphic image registration. Lecture Notes in Computer Science, 6667:677–688, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24785-9_57.

Aflalo:2012:DSR

[283] Yonathan Aflalo, Alexander M. Bronstein, Michael M. Bronstein, and Ron Kimmel. Deformable shape retrieval by learning diffusion kernels. Lecture Notes in Computer Science, 6667:689–700, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24785-9_58.

Corpetti:2012:SML

[284] Thomas Corpetti and Etienne Mémin. Stochastic models for local optical flow estimation. Lecture Notes in Computer Science, 6667:701–712, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24785-9_59.

Demetz:2012:OFS

[285] Oliver Demetz, Joachim Weickert, Andrés Bruhn, and Henning Zimmer. Optic flow scale space. Lecture Notes in Computer Science, 6667:713–724, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24785-9_60.

Rosman:2012:GVR

[286] Guy Rosman, Michael M. Bronstein, Alexander M. Bronstein, Alon Wolf, and Ron Kimmel. Group-valued regularization framework for motion segmentation of dynamic non-rigid shapes. Lecture Notes in Computer Science, 6667:725–736, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24785-9_61.

Derian:2012:WBF

[287] Pierre Dérian, Patrick Héas, Cédric Herzet, and Étienne Mémin. Wavelet-based fluid motion estimation. Lecture Notes in Computer Science, 6667:737–748, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-24785-9_62.

Gorthi:2012:MWE

[288] Sai Gorthi, Sébastien Beyou, Thomas Corpetti, and Etienne Mémin. Multiscale weighted ensemble Kalman filter for fluid flow estimation. Lecture Notes in Computer Science, 6667:749–760, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24785-9_63.

Rosman:2012:POF

[289] Guy Rosman, Shachar Shem-Tov, David Bitton, Tal Nir, Gilad Adiv, Ron Kimmel, and Arie Feuer. Over-parameterized optical flow using a stereoscopic constraint. Lecture Notes in Computer Science, 6667: 761–772, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24785-9_64.

Heas:2012:ROF

[290] P. Héas, C. Herzet, and E. Mémin. Robust optic-flow estimation with Bayesian inference of model and hyper-parameters. Lecture Notes in Computer Science, 6667:773–785, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24785-9_65.

Florack:2012:RPD

[291] Luc Florack. Regularization of positive definite matrix fields based on multiplicative calculus. Lecture Notes in Computer Science, 6667: 786–796, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24785-9_66.

Anonymous:2012:BMj

[292] Anonymous. Back matter. Lecture Notes in Computer Science, 6667: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-24785-9/1.

Anonymous:2012:FMn

[293] Anonymous. Front matter. Lecture Notes in Computer Science, 6667: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-24785-9/1.

Kyritsis:2012:NFJ

[294] Victor Kyritsis, Paraskevas V. Lekeas, Dora Souliou, and Foto Afrati. A new framework for join product skew. Lecture Notes in Computer Science, 6799:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27392-6_1/.

Menager:2012:BAD

[295] Hervé Ménager, Vivek Gopalan, Bertrand Néron, Sandrine Larroudé, and Julien Maupetit. Bioinformatics applications discovery and composition with the Mobyle Suite and MobyleNet. Lecture Notes in Computer Science, 6799:11–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27392-6_2/.

Blanco:2012:TQD

[296] Eduardo Blanco, Yudith Cardinale, María-Esther Vidal, Joyce El Haddad, and Maude Manouvrier. A transactional-QoS driven approach for Web service composition. Lecture Notes in Computer Science, 6799:23– 42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27392-6_3/.

Neveu:2012:UOS

[297] Pascal Neveu, Caroline Domerg, Juliette Fabre, Vincent Négre, and Emilie Gennari. Using ontologies of software: Example of R functions management. Lecture Notes in Computer Science, 6799:43–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27392-6_4/.

Strauser:2012:SMS

[298] Edouard Strauser, Mikaël Naveau, Hervé Ménager, Julien Maupetit, and Zoé Lacroix. Semantic map for structural bioinformatics: Enhanced service discovery based on high level concept ontology. Lecture Notes in Computer Science, 6799:57–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27392-6_5/.

Bazzi:2012:UCC

[299] Rida A. Bazzi, Jeffrey M. Kiefer, and Zoé Lacroix. A user-centric classification of tools for biological resource discovery and integration on the Web. Lecture Notes in Computer Science, 6799:71–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27392-6_6/.

Godoy:2012:COC

[300] Daniela Godoy. Comparing one-class classification algorithms for finding interesting resources in social bookmarking systems. Lecture Notes in Computer Science, 6799:88–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27392-6_7/.

Costa:2012:ATM

[301] Flavio Costa, Daniel de Oliveira, Eduardo Ogasawara, and Alexandre A. B. Lima. Athena: Text mining based discovery of scientific workflows in disperse repositories. Lecture Notes in Computer Science, 6799:104– 121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27392-6_8/.

Hochard:2012:SMR

[302] Gaïane Hochard, Zoé Lacroix, Jordi Creus, and Bernd Amann. A semantic map of RSS feeds to support discovery. Lecture Notes in Computer Science, 6799:122–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27392-6_9/.

ElHaddad:2012:OTQ

[303] Joyce El Haddad. Optimization techniques for QoS-aware workflow realization in Web services context. Lecture Notes in Computer Science, 6799:134–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27392-6_10/.

Quafafou:2012:HRW

[304] Mohamed Quafafou, Omar Boucelma, Yacine Sam, and Zahi Jarir. Hybrid reasoning for Web services discovery. Lecture Notes in Computer Science, 6799:150–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27392-6_11/.

Sanchez-Vilas:2012:CUS

[305] Fernando Sánchez-Vilas, Manuel Lama, Juan C. Vidal, and Eduardo Sánchez. Combining uncorrelated similarity measures for service discovery. Lecture Notes in Computer Science, 6799:160–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27392-6_12/.

Amann:2012:CQD

[306] Bernd Amann, Laure Berti-Equille, Zoé Lacroix, and María-Esther Vidal. Challenges of quality-driven resource discovery. Lecture Notes in Computer Science, 6799:181–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27392-6_13/.

Anonymous:2012:BMk

[307] Anonymous. Back matter. Lecture Notes in Computer Science, 6799: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27392-6/1.

Anonymous:2012:FMo

[308] Anonymous. Front matter. Lecture Notes in Computer Science, 6799: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27392-6/1.

Wang:2012:NNE

[309] Lingzhi Wang and Jiansheng Wu. Neural network ensemble model using PPR and LS–SVR for stock market forecasting. Lecture Notes in Computer Science, 6838:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-24728-6_1/.

Liu:2012:SAN

[310] Duyu Liu and Qinzhen Huang. Stability analysis of neutral systems with distributed delays. Lecture Notes in Computer Science, 6838:9– 16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24728-6_2/.

Janeski:2012:NNM

[311] Miroslav Janeski and Slobodan Kalajdziski. Neural network model for forecasting Balkan stock exchanges. Lecture Notes in Computer Science, 6838:17–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24728-6_3/.

Van:2012:RFD

[312] Mien Van, Hee-Jun Kang, and Young-Shick Ro. A robust fault detection and isolation scheme for robot manipulators based on neural networks. Lecture Notes in Computer Science, 6838:25–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-24728-6_4/.

Cao:2012:AHR

[313] Wei Cao, Hui Zhou, Zhimin Zhou, and Zuowei Fu. An approach for high resolution radar target recognition based on BP neural network. Lecture Notes in Computer Science, 6838:33–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-24728-6_5/.

Zhou:2012:HCA

[314] Yan Zhou and Zhifeng Hu. Hybrid clustering algorithm based on the artificial immune principle. Lecture Notes in Computer Science, 6838:40– 46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24728-6_6/.

Liang:2012:SIM

[315] Xueru Liang, Shangkun Ren, and Lei Yang. Succinct initialization methods for clustering algorithms. Lecture Notes in Computer Science, 6838: 47–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24728-6_7/.

Liang:2012:VNS

[316] Yun-Chia Liang and Chia-Yun Tien. Variable neighborhood search for drilling operation scheduling in PCB industries. Lecture Notes in Computer Science, 6838:55–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-24728-6_8/.

Yang:2012:AWF

[317] Shixin Yang, Wangmeng Zuo, Lei Liu, Yanlai Li, and David Zhang. Adaptive weighted fusion of local kernel classifiers for effective pattern classification. Lecture Notes in Computer Science, 6838:63–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24728-6_9/.

Chen:2012:ILR

[318] Jie Chen. Improvement of LEACH routing algorithm based on use of balanced energy in wireless sensor networks. Lecture Notes in Computer Science, 6838:71–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24728-6_10/.

Zhu:2012:ORP

[319] Lin Zhu. Optimal regularization parameter estimation for regularized discriminant analysis. Lecture Notes in Computer Science, 6838:77–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24728-6_11/.

Lu:2012:OPS

[320] Can-Yi Lu. Optimized projection for sparse representation based classification. Lecture Notes in Computer Science, 6838:83–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-24728-6_12/.

Liu:2012:BSB

[321] Yue Liu, Wang Wei, Kang Wang, Zhenjiang Liao, and Jun jun Gao. Balanced-sampling-based heterogeneous SVR ensemble for business demand forecasting. Lecture Notes in Computer Science, 6838:91–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24728-6_13/.

Lu:2012:CCC

[322] Can-Yi Lu. CCC: Classifier combination via classifier. Lecture Notes in Computer Science, 6838:100–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-24728-6_14/.

Rose:2012:HRS

[323] Ahmad Nazari Mohd Rose, Mohd Isa Awang, Hasni Hassan, and Aznida Hayati Zakaria. Hybrid reduction in soft set decision making. Lecture Notes in Computer Science, 6838:108–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-24728-6_15/.

Jiang:2012:NCO

[324] Wen Jiang. A new class of ϵ-optimal learning automata. Lecture Notes in Computer Science, 6838:116–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-24728-6_16/.

Wang:2012:UMR

[325] Jie Wang. Uncertainty measures of roughness based on interval ordered information systems. Lecture Notes in Computer Science, 6838:122–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24728-6_17/.

Secmen:2012:DEA

[326] Mustafa Secmen and M. Fatih Tasgetiren. A differential evolution algorithm for the extraction of complex natural resonance frequencies of electromagnetic targets. Lecture Notes in Computer Science, 6838:131– 138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24728-6_18/.

Pourali:2012:NDL

[327] Zahra Pourali and Majid Aminnayeri. A novel discrete league championship algorithm for minimizing earliness/tardiness penalties with distinct due dates and batch delivery consideration. Lecture Notes in Computer Science, 6838:139–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-24728-6_19/.

Pilat:2012:LMMa

[328] Martin Pilát and Roman Neruda. Local meta-models for ASM–MOMA. Lecture Notes in Computer Science, 6838:147–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-24728-6_20/.

Zhao:2012:RVR

[329] Yanwei Zhao, Chuan Li, Jing ling Zhang, Xingqiu Ren, and Wei Ren. Research on vehicle routing problem with stochastic demand based on multi-objective method. Lecture Notes in Computer Science, 6838: 153–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24728-6_21.

Trinh:2012:SOM

[330] Khoa Trinh, Jacques Ferland, and Tien Dinh. A stochastic optimization method for solving the machine-part cell formation problem. Lecture Notes in Computer Science, 6838:162–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24728-6_22.

Wang:2012:FCM

[331] Zenghui Wang and Yanxia Sun. Fully connected multi-objective particle swarm optimizer based on neural network. Lecture Notes in Computer Science, 6838:170–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_23.

Zhang:2012:IGI

[332] Yu Zhang, Lihua Wu, Ying Zhang, and Jianxin Wang. Immune gravitation inspired optimization algorithm. Lecture Notes in Computer Science, 6838:178–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24728-6_24.

Bevilacqua:2012:NMO

[333] Vitoantonio Bevilacqua, Vincenzo Pacelli, and Stefano Saladino. A novel multi objective genetic algorithm for the portfolio optimization. Lecture Notes in Computer Science, 6838:186–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24728-6_25.

Liu:2012:TMF

[334] Ruiming Liu and Miao Yang. Tracking multiple feature in infrared image with mean-shift. Lecture Notes in Computer Science, 6838:194–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24728-6_26.

He:2012:RDH

[335] Yongjian He, Qiong Wu, Shoupeng Feng, Rongkun Zhou, Yonghua Xing, and Fei Wang. Research on dynamic human object tracking algorithm. Lecture Notes in Computer Science, 6838:202–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24728-6_27.

Gavrilov:2012:MRN

[336] Andrey V. Gavrilov and Artem Lenskiy. Mobile robot navigation using reinforcement learning based on neural network with short term memory. Lecture Notes in Computer Science, 6838:210–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24728-6_28.

Mumcu:2012:HBC

[337] Tarik Veli Mumcu, Ibrahim Aliskan, Kayhan Gulez, and Gurkan Tuna. Histogram based color object classification by multi-class support vector machine. Lecture Notes in Computer Science, 6838:218–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24728-6_29.

Li:2012:TOU

[338] Peihua Li and Qi Sun. Tracking objects using orientation covariance matrices. Lecture Notes in Computer Science, 6838:226–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24728-6_30.

Yang:2012:ARI

[339] Kai Yang, Ji-Xiang Du, and Chuan-Min Zhai. Action recognition via an improved local descriptor for spatio-temporal features. Lecture Notes in Computer Science, 6838:234–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_31.

Guo:2012:ERB

[340] Yi-Lan Guo, Ji-Xiang Du, and Chuan-Min Zhai. Event recognition based on a local space-time interest points and self-organization feature map method. Lecture Notes in Computer Science, 6838:242–249, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24728-6_32.

Gai:2012:SSB

[341] Changqing Gai, Kuanquan Wang, Lei Zhang, and Wangmeng Zuo. Strategy of statistics-based visualization for segmented 3D cardiac volume data set. Lecture Notes in Computer Science, 6838:250–256, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24728-6_33.

Wang:2012:RGR

[342] Min Wang, Wei Jia, Huanglin Zeng, and Xiao-Feng Wang. Robust gait recognition using gait energy image and band-limited phase-only correlation. Lecture Notes in Computer Science, 6838:257–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24728-6_34.

Lee:2012:III

[343] Yang Weon Lee. Implementation of interactive interview system using hand gesture recognition. Lecture Notes in Computer Science, 6838: 264–269, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24728-6_35.

Chong:2012:IRT

[344] Yanwen Chong, Wu Chen, Zhilin Li, William H. K. Lam, and Qingquan Li. Integrated real-time vision-based preceding vehicle detection in urban roads. Lecture Notes in Computer Science, 6838:270–275, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24728-6_36.

Paris:2012:RTF

[345] Sébastien Paris, Hervé Glotin, and Zhong-Qiu Zhao. Real-time face detection using integral histogram of multi-scale local binary patterns. Lecture Notes in Computer Science, 6838:276–281, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24728-6_37.

Zhao:2012:SUV

[346] Wen-Bo Zhao, Wei Chen, Guang-Zheng Zheng, Ke-Ming Huang, Kong-Jin Zhao, and Yu-Ge Li. Study on UAV video reconnaissance based adaptively tracking algorithm for the ground moving target. Lecture Notes in Computer Science, 6838:282–289, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_38.

Dai:2012:PPI

[347] Gui-Ping Dai. Palm print image de-noising based on BEMD and wavelet packet transform-Wiener filter. Lecture Notes in Computer Science, 6838:290–297, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24728-6_39.

Wang:2012:PCS

[348] Shengchun Wang and Wensheng Tang. Pavement crack segmentation algorithm based on local optimal threshold of cracks density distribution. Lecture Notes in Computer Science, 6838:298–302, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24728-6_40.

He:2012:LPE

[349] Yongjian He, Caigui Jiang, Chengwei Hu, Jingmin Xin, Qiong Wu, and Fei Wang. Linear pose estimation algorithm based on quaternion. Lecture Notes in Computer Science, 6838:303–310, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24728-6_41.

Agrawal:2012:IEA

[350] Raghav Agrawal, Badrinath Srinivas, and Phalguni Gupta. Image enhancement algorithm for ink-on-paper fingerprints. Lecture Notes in Computer Science, 6838:311–318, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_42.

Guo:2012:MSR

[351] Yi-Nan Guo, Dawei Xiao, Shuguo Zhang, and Jian Cheng. Multi-spectral remote sensing images classification method based on adaptive immune clonal selection culture algorithm. Lecture Notes in Computer Science, 6838:319–326, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24728-6_43.

Kohli:2012:AEU

[352] Sharad Kohli, Surya Prakash, and Phalguni Gupta. Age estimation using active appearance models and ensemble of classifiers with dissimilaritybased classification. Lecture Notes in Computer Science, 6838:327–334, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24728-6_44.

Xu:2012:AVB

[353] Xin Xu, Yinglin Wang, Jinshan Tang, Xiaolong Zhang, and Xiaoming Liu. Adaptive variance based sharpness computation for low contrast images. Lecture Notes in Computer Science, 6838:335–341, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24728-6_45.

Xuan:2012:NSC

[354] Shi-Bin Xuan, Gao-Li Sang, Bo Zhao, and Zeng-Guo Zheng. Nonsampling contourlet based “consistency verification” method of image fusion. Lecture Notes in Computer Science, 6838:342–349, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24728-6_46.

Mori:2012:UMS

[355] Satoshi Mori, Yoshinori Kobayashi, and Yoshinori Kuno. Understanding the meaning of shape description for interactive object recognition. Lecture Notes in Computer Science, 6838:350–356, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24728-6_47.

Tang:2012:GSP

[356] Wen-Sheng Tang, Shao-Hua Jiang, and Shu-Lin Wang. Gray scale potential theory of sparse image. Lecture Notes in Computer Science, 6838:357–363, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24728-6_48.

Du:2012:RLI

[357] Ji-Xiang Du, Chuan-Min Zhai, and Qing-Ping Wang. Recognition of leaf image based on outline and vein fractal dimension feature. Lecture Notes in Computer Science, 6838:364–369, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_49.

Ramasamy:2012:WMF

[358] Krishnamoorthi Ramasamy and Sathiya Devi Shanmugam. Weighted multi feature based image retrieval with orthogonal polynomials model and genetic algorithm. Lecture Notes in Computer Science, 6838: 370–376, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24728-6_50.

Vavilin:2012:ACA

[359] Andrey Vavilin, Kang-Hyun Jo, Moon-Ho Jeong, Jong-Eun Ha, and Dong-Joong Kang. Automatic context analysis for image classification and retrieval. Lecture Notes in Computer Science, 6838:377–382, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24728-6_51.

Shang:2012:IFE

[360] Li Shang and Jie Chen. Image feature extraction using the fusion features of BEMD and WCB–NNSC. Lecture Notes in Computer Science, 6838:383–390, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24728-6_52.

Yan:2012:PRM

[361] Ting-Qin Yan, Shu-Fen Liu, and Chang-Xiong Zhou. Palmprint recognition with MEEMD and statistically independent coefficients ICA. Lecture Notes in Computer Science, 6838:391–397, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24728-6_53.

Huai:2012:PCM

[362] Wen-Jun Huai and Li Shang. A palmprint classification method based on finite ridgelet transformation and SVM. Lecture Notes in Computer Science, 6838:398–404, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_54.

Zhang:2012:PRB

[363] Yu Zhang, Mei-Xing Qi, and Li Shang. Palmprint recognition based on two-dimensional Gabor wavelet transform and two-dimensional principal component analysis. Lecture Notes in Computer Science, 6838:405–411, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24728-6_55.

Katiyar:2012:RBF

[364] Rohit Katiyar and Vinay Kumar Pathak. Recognition based on fusion of gait, ear and face features using KPCA method. Lecture Notes in Computer Science, 6838:412–419, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_56.

Wang:2012:FRU

[365] Guoqiang Wang and Guoqiang Ding. Face recognition using KFDA– LLE. Lecture Notes in Computer Science, 6838:420–425, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24728-6_57.

Bevilacqua:2012:SAS

[366] Vitoantonio Bevilacqua, Marco Suma, Dario D’Ambruoso, and Giovanni Mandolino. A supervised approach to support the analysis and the classification of non verbal humans communications. Lecture Notes in Computer Science, 6838:426–431, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_58.

Zeng:2012:TCA

[367] Anping Zeng and Yongping Huang. A text classification algorithm based on Rocchio and hierarchical clustering. Lecture Notes in Computer Science, 6838:432–439, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_59.

Ahn:2012:KDC

[368] Byung Ryul Ahn, Won gyum Kim, Won Young Yu, and Moon-Hyun Kim. Korean documents copy detection based on Ferret. Lecture Notes in Computer Science, 6838:440–447, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_60.

Ren:2012:IWA

[369] Shangkun Ren, Yuexian Hou, Peng Zhang, and Xueru Liang. Importance weighted AdaRank. Lecture Notes in Computer Science, 6838:448–455, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24728-6_61.

Bevilacqua:2012:SSF

[370] Vitoantonio Bevilacqua, Vito Santarcangelo, Alberto Magarelli, and Annalisa Bianco. A semantic search framework for document retrievals (literature, art and history) based on thesaurus multiwordnet like. Lecture Notes in Computer Science, 6838:456–463, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24728-6_62.

Cho:2012:NMS

[371] Sangjin Cho, Myeongsu Kang, and Uipil Chong. A novel musical synthesizer embedded with physical modeling sound synthesis algorithm. Lecture Notes in Computer Science, 6838:464–471, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24728-6_63.

Zan:2012:SAR

[372] Hongying Zan, Lijuan Zhou, and Kunli Zhang. Studies on the automatic recognition of modern Chinese conjunction usages. Lecture Notes in Computer Science, 6838:472–479, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_64.

Kang:2012:IHP

[373] Myeongsu Kang, Jiwon Choi, Yongmin Kim, Cheol-Hong Kim, and Jong-Myon Kim. Implementation of high-performance sound synthesis engine for plucked-string instruments. Lecture Notes in Computer Science, 6838:480–487, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24728-6_65.

Zhang:2012:SRF

[374] Xutang Zhang, Xiaofeng Chen, Xin Hou, and Ting Zhuang. A semantic retrieval framework for engineering domain knowledge. Lecture Notes in Computer Science, 6838:488–493, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_66.

Zhang:2012:FAT

[375] Jianhong Zhang, Yuanbo Cui, and Xi Wu. Forgeability attack of two special signature schemes. Lecture Notes in Computer Science, 6838: 494–501, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24728-6_67.

Qin:2012:TDG

[376] Jie Qin, Qun Si, Huijuan Yan, and Fuliang Yan. A Trojan detector generating algorithm based on chaotic theory. Lecture Notes in Computer Science, 6838:502–508, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_68.

Zhang:2012:NSSa

[377] Boyun Zhang, Zhigang Chen, Xiai Yan, Shulin Wang, and Qiang Fan. Network security situation assessment based on hidden Semi–Markov model. Lecture Notes in Computer Science, 6838:509–516, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24728-6_69.

Zhang:2012:NSSb

[378] Boyun Zhang, Zhigang Chen, Wensheng Tang, Qiang Fan, Xiai Yan, and Shulin Wang. Network security situation assessment based on stochastic game model. Lecture Notes in Computer Science, 6838:517–525, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24728-6_70.

Zhu:2012:HPD

[379] Xueguang Zhu. High precision detection of infrared energy spectrum. Lecture Notes in Computer Science, 6838:526–531, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24728-6_71.

Wang:2012:UHR

[380] Jeen-Shing Wang, Che-Wei Lin, and Ya-Ting C. Yang. Using heart rate variability parameter-based feature transformation algorithm for driving stress recognition. Lecture Notes in Computer Science, 6838:532–537, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24728-6_72.

Le:2012:OSG

[381] Tien Dung Le, Hee-Jun Kang, and Young-Soo Suh. An online self gain tuning computed torque controller for a five-bar manipulator. Lecture Notes in Computer Science, 6838:538–543, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24728-6_73.

Son:2012:TED

[382] Mai Son and Myeong-Jae Yi. Towards an efficient discovery services in OPC unified architecture. Lecture Notes in Computer Science, 6838: 544–552, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24728-6_74.

Wang:2012:AMI

[383] Jie Wang and Liang Tong. Asynchronous modeling and implementation of Web services based on JMS. Lecture Notes in Computer Science, 6838:553–559, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24728-6_75.

Xu:2012:ESF

[384] Ye Xu, Ling Wang, Gang Zhou, and Shengyao Wang. An effective shuffled frog leaping algorithm for solving hybrid flow-shop scheduling problem. Lecture Notes in Computer Science, 6838:560–567, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24728-6_76.

Wang:2012:HAB

[385] Ling Wang, Ye Xu, and Chen Fang. A hybrid algorithm based on simplex search and differential evolution for resource-constrained project scheduling problem. Lecture Notes in Computer Science, 6838:568–575, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24728-6_77.

Sang:2012:DEA

[386] Hongyan Sang, Liang Gao, and Xinyu Li. A differential evolution algorithm for lot-streaming flow shop scheduling problem. Lecture Notes in Computer Science, 6838:576–583, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_78.

Li:2012:FJS

[387] Junqing Li, Yuanzhen Li, Huaqing Yang, Kaizhou Gao, Yuting Wang, and Tao Sun. Flexible job shop scheduling problem by chemicalreaction optimization algorithm. Lecture Notes in Computer Science, 6838:584–591, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24728-6_79.

Gao:2012:DHS

[388] Kaizhou Gao, Shengxian Xie, Hua Jiang, and Junqing Li. Discrete harmony search algorithm for the no wait flow shop scheduling problem with makespan criterion. Lecture Notes in Computer Science, 6838:592–599, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24728-6_80.

Qian:2012:HDE

[389] Bin Qian, Hua-Bin Zhou, Rong Hu, and Feng-Hong Xiang. Hybrid differential evolution optimization for no-wait flow-shop scheduling with sequence-dependent setup times and release dates. Lecture Notes in Computer Science, 6838:600–611, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_81.

Zhang:2012:IAA

[390] Jianwei Zhang, Baowei Zhang, Zengyu Cai, and Zhaoyang Li. An improved approximation algorithm for a class of batch scheduling problems. Lecture Notes in Computer Science, 6838:612–618, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24728-6_82.

Zhong:2012:EST

[391] Qishui Zhong and Shungang Xu. Exponential stabilization for Takagi– Sugeno fuzzy systems with time delay via impulsive control. Lecture Notes in Computer Science, 6838:619–625, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24728-6_83.

Wang:2012:OTC

[392] Cuihong Wang and Huijuan Cao. H_{∞} output tracking control for neutral delay system with nonlinear perturbations. Lecture Notes in Computer Science, 6838:626–633, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_84.

Jin:2012:CMT

[393] Ju Jin, Yongbin Yu, Yijing Liu, Xiaorong Pu, and Xiaofeng Liao. Chaotic modeling of time-delay memristive system. Lecture Notes in Computer Science, 6838:634–641, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_85.

Song:2012:SAC

[394] Yulin Song. Stability analysis and constrained control of positive systems with time-varying delays: a linear copositive Lyapunov–Krasovskii functional method. Lecture Notes in Computer Science, 6838:642–649, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24728-6_86.

Ding:2012:LCP

[395] Xiuyong Ding, Lan Shu, and Changcheng Xiang. On linear co-positive Lyapunov functions for a special of switched linear positive systems. Lecture Notes in Computer Science, 6838:650–657, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24728-6_87.

Wang:2012:IMS

[396] Feng Wang, Jinlin Ding, and Zhifeng Hu. Induction motor speedregulating control system based on nonlinear kernel ridge regression. Lecture Notes in Computer Science, 6838:658–663, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24728-6_88.

Zhu:2012:PRC

[397] Caihong Zhu and Hongtao Zhang. Parallel robotics control strategy study based on fuzzy-PID. Lecture Notes in Computer Science, 6838:664–669, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24728-6_89.

Zhu:2012:ERC

[398] Xueguang Zhu. Embedded remote controller with two cameras intelligent orientation and bi-direction wireless communication. Lecture Notes in Computer Science, 6838:670–675, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_90.

Yang:2012:AAS

[399] Bo Yang, Mingyi Zhang, and Ying Zhang. Applying answer set programming to points-to analysis of object-oriented language. Lecture Notes in Computer Science, 6838:676–685, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_91.

Yan:2012:CFT

[400] Yan Yan, Shuanghe Yu, Zhenqiang Yang, and Jialu Du. Continuous finite-time observer-controller based speed regulation of permanent magnet synchronous motors. Lecture Notes in Computer Science, 6838: 686–693, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24728-6_92.

Li:2012:OBE

[401] Xue Li, Jia min Weng, Dajun Du, and Haoliang Bai. Observerbased exponential stability analysis for networked control systems with packet dropout. Lecture Notes in Computer Science, 6838:694–700, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24728-6_93.

Wang:2012:MAR

[402] Zhi zhen Wang, Li Li, and Xiao fang Wang. Modification algorithm on Routh–Padé model reduction of interval systems. Lecture Notes in Computer Science, 6838:701–704, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24728-6_94.

Anonymous:2012:BMl

[403] Anonymous. Back matter. Lecture Notes in Computer Science, 6838: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-24728-6/1.

Anonymous:2012:FMp

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-24728-6/1.

Zhou:2012:EAB

[405] Gang Zhou, Ling Wang, Ye Xu, and Shengyao Wang. An effective artificial bee colony algorithm for multi-objective flexible job-shop scheduling problem. Lecture Notes in Computer Science, 6839:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25944-9_1/.

Wang:2012:EDA

[406] Shengyao Wang, Ling Wang, Gang Zhou, and Ye Xu. An estimation of distribution algorithm for the flexible job-shop scheduling problem. Lecture Notes in Computer Science, 6839:9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-25944-9_2/.

Wang:2012:MIO

[407] Yuting Wang, Jian Sun, Junqing Li, and Kaizhou Gao. A modified inver-over operator for the traveling salesman problem. Lecture Notes in Computer Science, 6839:17–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25944-9_3/.

Wang:2012:NMO

[408] Wanliang Wang, Lili Chen, Jing Jie, Yanwei Zhao, and Jing Zhang. A novel multi-objective particle swarm optimization algorithm for flow shop scheduling problems. Lecture Notes in Computer Science, 6839:24– 31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25944-9_4/.

Han:2012:MTFa

[409] Hong-Yan Han. Minimizing the total flow time for lot streaming flow shop using an effective discrete harmony search algorithm. Lecture Notes in Computer Science, 6839:32–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25944-9_5/.

Liu:2012:TTI

[410] Gengcheng Liu, Shiji Song, and Cheng Wu. Two techniques to improve the NEH algorithm for flow-shop scheduling problems. Lecture Notes in Computer Science, 6839:41–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25944-9_6/.

Chiang:2012:FJS

[411] Tsung-Che Chiang and Hsiao-Jou Lin. Flexible job shop scheduling using a multiobjective memetic algorithm. Lecture Notes in Computer Science, 6839:49–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25944-9_7/.

Bulut:2012:GAE

[412] Onder Bulut, M. Fatih Tasgetiren, and M. Murat Fadiloglu. A genetic algorithm for the economic lot scheduling problem under extended basic period approach and power-of-two policy. Lecture Notes in Computer Science, 6839:57–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25944-9_8/.

Han:2012:MOH

[413] Hong-Yan Han. A multi-objective hybrid discrete harmony search algorithm for lot-streaming flow shop scheduling problem. Lecture Notes in Computer Science, 6839:66–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25944-9_9/.

Guldogan:2012:DBA

[414] Evrim Ursavas Guldogan, Onder Bulut, and M. Fatih Tasgetiren. A dynamic berth allocation problem with priority considerations under stochastic nature. Lecture Notes in Computer Science, 6839:74–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25944-9_10/.

Tasgetiren:2012:BVI

[415] M. Fatih Tasgetiren, Quan-Ke Pan, Ling Wang, and Angela H.-L. Chen. A DE based variable iterated greedy algorithm for the no-idle permutation flowshop scheduling problem with total flowtime criterion. Lecture Notes in Computer Science, 6839:83–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-25944-9_11/.

Han:2012:MTFb

[416] Yu-Yan Han, Jun-Hua Duan, Yu-Jie Yang, Min Zhang, and Bao Yun. Minimizing the total flowtime flowshop with blocking using a discrete artificial bee colony. Lecture Notes in Computer Science, 6839:91–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25944-9_12/.

Guo:2012:TIC

[417] Zhenhua Guo, Qin Li, Lin Zhang, Jane You, Wenhuang Liu, and Jinghua Wang. Texture image classification using complex texton. Lecture Notes in Computer Science, 6839:98–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25944-9_13/.

Hu:2012:PMM

[418] Rong-Xiang Hu. A perceptually motivated morphological strategy for shape retrieval. Lecture Notes in Computer Science, 6839:105–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25944-9_14/.

Zhao:2012:TAL

[419] Yang Zhao. Theories and applications of LBP: a survey. Lecture Notes in Computer Science, 6839:112–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25944-9_15/.

Jabeen:2012:VCV

[420] Syeda Darakhshan Jabeen and Rathindra Nath Mukherjee. Vibration control of a vehicle using hybrid genetic algorithm. Lecture Notes in Computer Science, 6839:121–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25944-9_16/.

Das:2012:DTP

[421] T. Das, Rathindra Nath Mukherjee, and K. S. Chaudhuri. Dynamics of a two prey one predator fishery with low predator density. Lecture Notes in Computer Science, 6839:129–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25944-9_17/.

Shi:2012:NVU

[422] Min Shi. Natural vs. unnatural decomposition in cooperative coevolution. Lecture Notes in Computer Science, 6839:138–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25944-9_18/.

Feng:2012:MIP

[423] Naiqin Feng, Yushan Zhang, Lianhui Ao, and Shuangxi Wang. A method to improve performance of heteroassociative morphological memories. Lecture Notes in Computer Science, 6839:148–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25944-9_19/.

Wang:2012:ROP

[424] Bo Wang, Guang an Zou, and Peng Zhao. A restrained optimal perturbation method for solving the inverse problem in reverse process of convection diffusion equation. Lecture Notes in Computer Science, 6839:154– 161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25944-9_20/.

Wang:2012:OBS

[425] Yujia Wang and Yunfeng Xue. Overdetermined blind source separation by Gaussian mixture model. Lecture Notes in Computer Science, 6839:162–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25944-9_21.

Weng:2012:NCC

[426] Zhiwei Weng, Jian Weng, Kai He, and Yingkai Li. New chosen ciphertext secure public key encryption in the standard model with public verifiability. Lecture Notes in Computer Science, 6839:170–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_22.

Jabeen:2012:LLM

[427] Hajira Jabeen and Abdul Rauf Baig. Lazy learning for multi-class classification using genetic programming. Lecture Notes in Computer Science, 6839:177–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25944-9_23.

Cheng:2012:ACA

[428] Yuhu Cheng, Huanting Feng, and Xuesong Wang. Actor-critic algorithm based on incremental least-squares temporal difference with eligibility trace. Lecture Notes in Computer Science, 6839:183–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_24.

Feng:2012:APN

[429] KaiYan Feng, JunHui Gao, KaiRui Feng, Lei Liu, and YiXue Li. Active and passive nearest neighbor algorithm: a newly-developed supervised classifier. Lecture Notes in Computer Science, 6839:189–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25944-9_25.

Tang:2012:SVM

[430] Wen sheng Tang, Sheng chun Wang, and He long Xiao. Support vector machines for user-defined sheets recognition in complex environment. Lecture Notes in Computer Science, 6839:197–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_26.

Sun:2012:NMS

[431] Yanxia Sun, Barend Jacobus van Wyk, and Zenghui Wang. A new multiswarm multi-objective particle swarm optimization based on Pareto front set. Lecture Notes in Computer Science, 6839:203–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_27.

Figueroa-Garcia:2012:ITF

[432] Juan C. Figueroa-García, Dusko Kalenatic, and Cesar Amilcar Lopez. Interval type-2 fuzzy Markov chains: Type reduction. Lecture Notes in Computer Science, 6839:211–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_28.

Yu:2012:MAR

[433] Lasheng Yu and Issahaku Abdulai. A multi-agent reinforcement learning with weighted experience sharing. Lecture Notes in Computer Science, 6839:219–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25944-9_29.

Wen:2012:ACO

[434] Jia-Bao Wen and Yue-Shan Xiong. Asymmetric constraint optimization based adaptive boosting for cascade face detector. Lecture Notes in Computer Science, 6839:226–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_30.

Zin:2012:TMM

[435] Thet Thet Zin, Khin Mar Soe, and Ni Lar Thein. Translation model of Myanmar phrases for statistical machine translation. Lecture Notes in Computer Science, 6839:235–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_31.

Bevilacqua:2012:MOG

[436] Vitoantonio Bevilacqua, Mariagrazia Dotoli, Marco Falagario, and Fabio Sciancalepore. A multi-objective genetic optimization technique for the strategic design of distribution networks. Lecture Notes in Computer Science, 6839:243–250, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_32.

Su:2012:OOB

[437] Zhong-Hua Su, Qing-Shan Jia, and Chen Song. Ordinal optimizationbased multi-energy system scheduling for building energy saving. Lecture Notes in Computer Science, 6839:251–259, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25944-9_33.

Yan:2012:TLI

[438] Hongyan Yan, Xiaojuan Zhang, and Hongxia Xu. Three levels intelligent incident detection algorithm of smart traffic in the digital city. Lecture Notes in Computer Science, 6839:260–266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25944-9_34.

Khoie:2012:PCT

[439] Maryam Khoie, Karim Salahshoor, Ehsan Nouri, and Ali Khaki Sedigh. PID controller tuning using multi-objective optimization based on fused genetic-immune algorithm and immune feedback mechanism. Lecture Notes in Computer Science, 6839:267–276, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25944-9_35.

Jin:2012:BAC

[440] Ling Jin, Jae Yeol Yoon, Young Hee Kim, and Ung Mo Kim. Based on analyzing closeness and authority for ranking expert in social network. Lecture Notes in Computer Science, 6839:277–283, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_36.

Han:2012:EFI

[441] Feng Han and Chi Xie. The effects of Forex intervention: a simultaneous equations model. Lecture Notes in Computer Science, 6839:284–291, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25944-9_37.

Zi-jian:2012:NMU

[442] Tian Zi-jian, Hou Yan, and Zhang XiangYang. A new method of underground radio noise distribution measure. Lecture Notes in Computer Science, 6839:292–299, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_38.

Nguyen:2012:FPC

[443] Ngoc-Tung Nguyen and Hong-Hee Lee. Fuzzy PI controller for gridconnected inverters. Lecture Notes in Computer Science, 6839:300–308, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25944-9_39.

Li:2012:IPP

[444] Jirui Li and Kai Yang. Improvement of path planning in mobile beacon assisted positioning. Lecture Notes in Computer Science, 6839:309–316, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25944-9_40.

Islam:2012:CSI

[445] Mojaharul Islam and Hong-Hee Lee. A comprehensive study on IEC61850 process bus architecture and spit bus based differential protection. Lecture Notes in Computer Science, 6839:317–324, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_41.

Yu:2012:SMO

[446] Shuanghe Yu, Zhenqiang Yang, Jialu Du, and Jingcong Ma. Sliding mode observer based anti-windup PI speed controller for permanent magnet synchronous motors. Lecture Notes in Computer Science, 6839:325–332, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25944-9_42.

Li:2012:PPE

[447] Shouqi Li, Fangcheng Long, and Yongchang Wang. Probe into principle of expert system in psychological warfare. Lecture Notes in Computer Science, 6839:333–340, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_43.

Xue:2012:SFD

[448] Hongtao Xue, Huaqing Wang, Liuyang Song, and Peng Chen. Structural fault diagnosis of rotating machinery based on distinctive frequency components and support vector machines. Lecture Notes in Computer Science, 6839:341–348, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_44.

Changrui:2012:CRM

[449] Yu Changrui and Luo Yan. Comparative research on methodologies for domain ontology development. Lecture Notes in Computer Science, 6839:349–356, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25944-9_45.

Zhang:2012:CBH

[450] Weiqi Zhang and Kunlong Zhang. The comparison between histogram method and index method in selectivity estimation. Lecture Notes in Computer Science, 6839:357–362, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_46.

Hao:2012:SPB

[451] Tianyong Hao, Wenyin Liu, and Chunshen Zhu. Semantic pattern-based user interactive question answering: User interface design and evaluation. Lecture Notes in Computer Science, 6839:363–370, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_47.

Li:2012:PBW

[452] Wenli Li. PSO based wireless sensor networks coverage optimization on DEMs. Lecture Notes in Computer Science, 6839:371–378, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_48.

Rotili:2012:RTS

[453] Rudy Rotili, Emanuele Principi, Stefano Squartini, and Björn Schuller. Real-time speech recognition in a multi-talker reverberated acoustic scenario. Lecture Notes in Computer Science, 6839:379–386, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_49.

Zhang:2012:NSS

[454] Boyun Zhang, Zhigang Chen, Shulin Wang, Xiai Yan, Dingxing Zhang, and Qiang Fan. Network security situation assessment based on HMM. Lecture Notes in Computer Science, 6839:387–394, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_50.

Huxidan:2012:FRB

[455] Huxidan, Wanquan Liu, and Chong Lu. Face recognition based on rearranged modular 2DPCA. Lecture Notes in Computer Science, 6839: 395–403, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25944-9_51.

Shen:2012:FRV

[456] Linlin Shen, Jinwen He, Shipei Wu, and Songhao Zheng. Face recognition from visible and near-infrared images using boosted directional binary code. Lecture Notes in Computer Science, 6839:404–411, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_52.

Wu:2012:SAF

[457] Min Wu, A. Yong, Tong Zhao, and Tiande Guo. A systematic algorithm for fingerprint image quality assessment. Lecture Notes in Computer Science, 6839:412–420, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_53.

Dong:2012:TCB

[458] Yongsheng Dong and Jinwen Ma. Texture classification based on contourlet subband clustering. Lecture Notes in Computer Science, 6839: 421–426, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25944-9_54.

Liu:2012:IRA

[459] Xiaomin Liu and Peihua Li. An iris recognition approach with SIFT descriptors. Lecture Notes in Computer Science, 6839:427–434, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25944-9_55.

Wang:2012:NWR

[460] Hang jun Wang, Heng nian Qi, and Xiao-Feng Wang. A new wood recognition method based on Gabor entropy. Lecture Notes in Computer Science, 6839:435–440, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_56.

Kou:2012:AEF

[461] Jie Kou, Ji-Xiang Du, and Chuan-Min Zhai. Age estimation of facial images based on a super-resolution reconstruction algorithm. Lecture Notes in Computer Science, 6839:441–446, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_57.

Wang:2012:WPA

[462] Jeen-Shing Wang, Fang-Chen Chuang, and Ya-Ting C. Yang. A wearable physical activity sensor system: Its classification algorithm and performance comparison of different sensor placements. Lecture Notes in Computer Science, 6839:447–454, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_58.

Ge:2012:TAC

[463] Yanbin Ge and Yan Wu. Towards adaptive classification of motor imagery EEG using biomimetic pattern recognition. Lecture Notes in Computer Science, 6839:455–460, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_59.

Feng:2012:CSA

[464] Liu Feng and Seong Whan Kim. Comparison of scalable ACC and MC–CDMA for practical video fingerprinting scheme. Lecture Notes in Computer Science, 6839:461–468, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_60.

Li:2012:FSI

[465] Dalong Li and Steven Simske. Fast single image super-resolution by self-trained filtering. Lecture Notes in Computer Science, 6839:469–475, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25944-9_61.

Kim:2012:HPV

[466] Yongmin Kim, Myeongsu Kang, and Jong-Myon Kim. High-performance video based fire detection algorithms using a multi-core architecture. Lecture Notes in Computer Science, 6839:476–483, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25944-9_62.

Zhang:2012:PCB

[467] Shan-Wen Zhang, Min-Rong Zhao, and Xiao-Feng Wang. Plant classification based on multilinear independent component analysis. Lecture Notes in Computer Science, 6839:484–490, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_63.

Mandava:2012:KBA

[468] V. Mandava, P. Nimmagadda, T. R. Korrapati, and K. R. Anne. Knowledge based agent for intelligent traffic light control — an Indian perspective. Lecture Notes in Computer Science, 6839:491–501, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_64.

Liu:2012:MSM

[469] Jun Liu, Xiaoming Liu, Jianxun Chen, and J. Tang. Mass segmentation in mammograms based on improved level set and watershed algorithm. Lecture Notes in Computer Science, 6839:502–508, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_65.

Lenskiy:2012:UTS

[470] Artem Lenskiy. Unsupervised texture segmentation algorithm based on novel scale exponent features. Lecture Notes in Computer Science, 6839:509–515, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25944-9_66.

Du:2012:FAS

[471] Ji-Xiang Du, Chuan-Min Zhai, and Yong-Qing Ye. Face aging simulation based on NMF algorithm with sparseness constraints. Lecture Notes in Computer Science, 6839:516–522, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_67.

Kobayashi:2012:RWM

[472] Yoshinori Kobayashi, Yuki Kinpara, Erii Takano, Yoshinori Kuno, and Keiichi Yamazaki. Robotic wheelchair moving with caregiver collaboratively. Lecture Notes in Computer Science, 6839:523–532, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_68.

Tuna:2012:ESR

[473] Gurkan Tuna, Kayhan Gulez, Vehbi Cagri Gungor, and Tarik Veli Mumcu. Exploration strategy related design considerations of WSNaided mobile robot exploration teams. Lecture Notes in Computer Science, 6839:533–540, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-25944-9_69.

Yuan:2012:NBS

[474] Guo-Wu Yuan, Yun Gao, Dan Xu, and Mu-Rong Jiang. A new background subtraction method using texture and color information. Lecture Notes in Computer Science, 6839:541–548, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25944-9_70.

Ajaz:2012:DIE

[475] Sabooh Ajaz, Prashan Premaratne, and Malin Premaratne. Design and implementation of edge detection algorithm using digital signal controller (DSC). Lecture Notes in Computer Science, 6839:549–556, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_71.

Tran:2012:LVP

[476] Quang Tran, An Tran, Tien Ba Dinh, and Duc Duong. Long-view player detection framework algorithm in broadcast soccer videos. Lecture Notes in Computer Science, 6839:557–564, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_72.

Le:2012:BFR

[477] My-Ha Le and Kang-Hyun Jo. Building face reconstruction from sparse view of monocular camera. Lecture Notes in Computer Science, 6839: 565–572, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25944-9_73.

Tang:2012:UTM

[478] Nam Tang, Cuong Do, Tien Ba Dinh, and Thang Ba Dinh. Urban traffic monitoring system. Lecture Notes in Computer Science, 6839:573–580, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25944-9_74.

Kuremoto:2012:GRS

[479] Takashi Kuremoto, Yasuhiro Kinoshita, Liang bing Feng, and Shun Watanabe. A gesture recognition system using one-pass DP method. Lecture Notes in Computer Science, 6839:581–587, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25944-9_75.

Premaratne:2012:HGT

[480] Prashan Premaratne, Sabooh Ajaz, and Malin Premaratne. Hand gesture tracking and recognition system for control of consumer electronics. Lecture Notes in Computer Science, 6839:588–593, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_76.

Bhattacharjee:2012:NRI

[481] Debalina Bhattacharjee, Surya Prakash, and Phalguni Gupta. Noreference image quality assessment for facial images. Lecture Notes in Computer Science, 6839:594–601, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_77.

Tiwari:2012:PBR

[482] Kamlesh Tiwari, Devendra Kumar Arya, and Phalguni Gupta. Palmprint based recognition system using local structure tensor and force field transformation. Lecture Notes in Computer Science, 6839:602–607, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25944-9_78.

Kaushik:2012:MGH

[483] Vandana Dixit Kaushik, Amit K. Gupta, Umarani Jayaraman, and Phalguni Gupta. Modified geometric hashing for face database indexing. Lecture Notes in Computer Science, 6839:608–613, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25944-9_79.

Liu:2012:GRS

[484] Duyu Liu and Xin Gao. Globe robust stability analysis for interval neutral systems. Lecture Notes in Computer Science, 6839:614–621, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25944-9_80.

Liu:2012:ESN

[485] Xiu Liu, Shouming Zhong, and Changcheng Xiang. Exponential stability of nonlinear switched delay systems. Lecture Notes in Computer Science, 6839:622–629, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25944-9_81.

Liu:2012:MCL

[486] Xiaoming Liu, Xin Xu, Jun Liu, and J. Tang. Mass classification with level set segmentation and shape analysis for breast cancer diagnosis using mammography. Lecture Notes in Computer Science, 6839:630–637, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25944-9_82.

Li:2012:CBP

[487] Bo Li and Jin Liu. The connections between principal component analysis and dimensionality reduction methods of manifolds. Lecture Notes in Computer Science, 6839:638–643, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_83.

Dong:2012:SLA

[488] Wenyong Dong and Jin Liu. Step length adaptation by generalized predictive control. Lecture Notes in Computer Science, 6839:644–649, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25944-9_84.

Jiang:2012:VSS

[489] Xinghao Jiang, Tanfeng Sun, Jin Liu, Wensheng Zhang, and Juan Chao. An video shot segmentation scheme based on adaptive binary searching and SIFT. Lecture Notes in Computer Science, 6839:650–655, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25944-9_85.

Prevenslik:2012:MQM

[490] Thomas Prevenslik. Memristors by quantum mechanics. Lecture Notes in Computer Science, 6839:656–663, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_86.

Gong:2012:GTD

[491] Dun wei Gong and Yan Zhang. Generating test data for both paths coverage and faults detection using genetic algorithms. Lecture Notes in Computer Science, 6839:664–671, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_87.

Shang:2012:MIR

[492] Li Shang and Pin gang Su. MMW image reconstruction combined NNSC shrinkage technique and PDEs algorithm. Lecture Notes in Computer Science, 6839:672–679, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_88.

Zheng:2012:CEE

[493] Hong-Jing Zheng and Na Tun. Construction of embedded Ethernet based on MCF52259. Lecture Notes in Computer Science, 6839:680–686, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25944-9_89.

Zhou:2012:IMM

[494] Changxiong Zhou, Chunmei Lu, Yubo Tian, and Chuanlin Zhou. Image magnification method based on linear interpolation and wavelet and PDE. Lecture Notes in Computer Science, 6839:687–693, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25944-9_90.

Zhang:2012:RDM

[495] Yu Zhang. Research of detecting mixed flammable gases with a single catalytic sensor based on RBF neural network. Lecture Notes in Computer Science, 6839:694–700, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_91.

Shang:2012:PRU

[496] Li Shang, Ming Cui, and Jie Chen. Palm recognition using fast sparse coding algorithm. Lecture Notes in Computer Science, 6839:701–707, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25944-9_92.

Zhou:2012:SRB

[497] Yan Zhou and Li Shang. Speaker recognition based on principal component analysis and probabilistic neural network. Lecture Notes in Computer Science, 6839:708–715, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25944-9_93.

Ince:2012:BDM

[498] Ibrahim Furkan Ince, Adem Karahoca, and Dilek Karahoca. Benchmarking data mining methods in CAT. Lecture Notes in Computer Science, 6839:716–726, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25944-9_94.

Anonymous:2012:BMm

[499] Anonymous. Back matter. Lecture Notes in Computer Science, 6839: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-25944-9/1.

Anonymous:2012:FMq

[500] Anonymous. Front matter. Lecture Notes in Computer Science, 6839: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-25944-9/1.

Liu:2012:ANN

[501] YuXi Liu and ZhiFeng Zhang. Application of neural network on solid boronizing. Lecture Notes in Computer Science, 6840:1–7, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-24553-4_1/.

Luo:2012:ISC

[502] Mengzhuo Luo and Shouming Zhong. Improved stability criteria for discrete-time stochastic neural networks with randomly time-varying delays. Lecture Notes in Computer Science, 6840:8–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-24553-4_2/.

Wang:2012:PCN

[503] Didi Wang, Pei-Chann Chang, Jheng-Long Wu, and Changle Zhou. A partially connected neural evolutionary network for stock price index forecasting. Lecture Notes in Computer Science, 6840:14–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24553-4_3/.

Wang:2012:FPE

[504] Yanfei Wang and Jingen Wang. Finite precision extended alternating projection neural network (FPEAP). Lecture Notes in Computer Science, 6840:20–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24553-4_4/.

Wu:2012:SVA

[505] QingXiang Wu, T. Martin McGinnity, Liam Maguire, Rongtai Cai, and Meigui Chen. Simulation of visual attention using hierarchical spiking neural networks. Lecture Notes in Computer Science, 6840:26–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24553-4_5/.

Xu:2012:CCN

[506] Yqoqun Xu and Xueling Yang. A class of chaotic neural network with Morlet wavelet function self-feedback. Lecture Notes in Computer Science, 6840:32–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24553-4_6/.

Meng:2012:NNN

[507] Chun Meng and Jiansheng Wu. A novel nonlinear neural network ensemble model using K–PLSR for rainfall forecasting. Lecture Notes in Computer Science, 6840:41–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-24553-4_7/.

Ye:2012:LEL

[508] Yibin Ye, Stefano Squartini, and Francesco Piazza. On-line extreme learning machine for training time-varying neural networks. Lecture Notes in Computer Science, 6840:49–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-24553-4_8/.

Feng:2012:OBC

[509] Chunhua Feng and Zhenkun Huang. Oscillatory behavior for a class of recurrent neural networks with time-varying input and delays. Lecture Notes in Computer Science, 6840:55–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-24553-4_9/.

Zhang:2012:SBN

[510] Cui Zhang, Li-Qing Zhao, and Rong-Long Wang. A saturation binary neural network for bipartite subgraph problem. Lecture Notes in Computer Science, 6840:64–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-24553-4_10/.

Tang:2012:MAN

[511] Chenghua Tang, Xin Wang, Reixia Zhang, and Yi Xie. Modeling and analysis of network security situation prediction based on covariance likelihood neural. Lecture Notes in Computer Science, 6840:71– 78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24553-4_11/.

Pilat:2012:LMMb

[512] Martin Pilát and Roman Neruda. Local meta-models for ASM– MOMA. Lecture Notes in Computer Science, 6840:79–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-24553-4_12/.

ElAal:2012:SMT

[513] Maha Mahmoud Abd ElAal, Gamal Selim, and Waleed Fakhr. Stock market trend prediction model for the Egyptian stock market using neural networks and fuzzy logic. Lecture Notes in Computer Science, 6840: 85–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24553-4_13/.

Liu:2012:IPP

[514] Zhi-Ping Liu, Jiguang Wang, Yu-Qing Qiu, Ross K. K. Leung, and Xiang-Sun Zhang. Inferring protein-protein interactions based on sequences and interologs in mycobacterium tuberculosis. Lecture Notes in Computer Science, 6840:91–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24553-4_14/.

Pennisi:2012:PLT

[515] Marzio Pennisi, Dario Motta, Alessandro Cincotti, and Francesco Pappalardo. Predicting long-term vaccine efficacy against metastases using agents. Lecture Notes in Computer Science, 6840:97–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-24553-4_15/.

Wang:2012:APT

[516] Jia Wang, Chuang Ma, Dao Zhou, Libin Zhang, and Yanhong Zhou. Accurately predicting transcription start sites using logitlinear model and local oligonucleotide frequencies. Lecture Notes in Computer Science, 6840:107–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24553-4_16/.

Wang:2012:NMI

[517] Hong-Qiang Wang, Xin-Ping Xie, and Ding Li. A new method for identifying cancer-related gene association patterns. Lecture Notes in Computer Science, 6840:115–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-24553-4_17/.

Zhang:2012:NMA

[518] Sheng-Li Zhang. A new mining algorithm of association rules and applications. Lecture Notes in Computer Science, 6840:123–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-24553-4_18/.

Taguchi:2012:MIG

[519] Y h. Taguchi and Jun Yasuda. MiRaE: Inference of gene expression regulation via MicroRNA Transfection II. Lecture Notes in Computer Science, 6840:129–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24553-4_19/.

DiLecce:2012:SPR

[520] Vincenzo Di Lecce and Marco Calabrese. Syntactic pattern recognition from observations: a hybrid technique. Lecture Notes in Computer Science, 6840:136–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-24553-4_20/.

Zhang:2012:EEM

[521] Juan Zhang, Xuegang Hu, Yuhong Zhang, and Peipei Li. An efficient ensemble method for classifying skewed data streams. Lecture Notes in Computer Science, 6840:144–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_21.

Zeng:2012:ARA

[522] Anping Zeng and Yongping Huang. An association rules algorithm based on Kendall-τ. Lecture Notes in Computer Science, 6840:152–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24553-4_22.

Mi:2012:RAC

[523] Zhenzhen Mi and Congfu Xu. A recommendation algorithm combining clustering method and slope one scheme. Lecture Notes in Computer Science, 6840:160–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_23.

Prasartvit:2012:DRB

[524] Thananan Prasartvit, Boonserm Kaewkamnerdpong, and Tiranee Achalakul. Dimensional reduction based on artificial bee colony for classification problems. Lecture Notes in Computer Science, 6840:168–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24553-4_24.

Wang:2012:RCD

[525] Shuan Wang, Dapeng Hua, Zhiguo Zhang, Ming Li, Ke Yao, and Zhanyou Wen. Robust controller design for main steam pressure based on SPEA2. Lecture Notes in Computer Science, 6840:176–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24553-4_25.

Din:2012:GBH

[526] Der-Rong Din. GA-based hybrid algorithm for MBR problem of FIPP p-cycles for node failure on survivable WDM networks. Lecture Notes in Computer Science, 6840:183–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_26.

Zhang:2012:NHA

[527] Xiaoxia Zhang, Zhe Liu, and Qiuying Bai. A new hybrid algorithm for the multidimensional knapsack problem. Lecture Notes in Computer Science, 6840:191–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_27.

Guo:2012:MPC

[528] Yi nan Guo, Dandan Liu, and Jian Cheng. Multi-population cooperative cultural algorithms. Lecture Notes in Computer Science, 6840:199–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24553-4_28.

Niu:2012:HQI

[529] Qun Niu, Zhuo Zhou, and Tingting Zeng. A hybrid quantum-inspired particle swarm evolution algorithm and SQP method for large-scale economic dispatch problems. Lecture Notes in Computer Science, 6840: 207–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_29.

Zhang:2012:RIL

[530] Haofan Zhang, Lei Fang, and Sheng-Uei Guan. Recursive and incremental learning GA featuring problem-dependent rule-set. Lecture Notes in Computer Science, 6840:215–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_30.

Picek:2012:ECO

[531] Stjepan Picek, Marin Golub, and Domagoj Jakobovic. Evaluation of crossover operator performance in genetic algorithms with binary representation. Lecture Notes in Computer Science, 6840:223–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24553-4_31.

Li:2012:QIS

[532] Lvzhou Li and Daowen Qiu. Quantum information splitting using GHZtype and W-type states. Lecture Notes in Computer Science, 6840: 231–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_32.

Zou:2012:AIQ

[533] Xiangfu Zou and Daowen Qiu. Attacks and improvements of QSDC schemes based on CSS codes. Lecture Notes in Computer Science, 6840:239–246, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_33.

Cao:2012:SSI

[534] Yi Cao, Yuehui Chen, and Yaou Zhao. Stochastic system identification by evolutionary algorithms. Lecture Notes in Computer Science, 6840:247–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_34.

Su:2012:DMP

[535] Zhixi Su, Junfeng Xia, and Zhongming Zhao. Do MicroRNAs preferentially target the genes with low DNA methylation level at the promoter region? Lecture Notes in Computer Science, 6840:253–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24553-4_35.

Li:2012:AMB

[536] Cuifeng Li, Jianbo Ye, and Fatai Zheng. Ameliorating GM(1, 1) model based on the structure of the area under parabola. Lecture Notes in Computer Science, 6840:259–266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_36.

Distefano:2012:RSS

[537] Salvatore Distefano. Reliability of standby systems. Lecture Notes in Computer Science, 6840:267–275, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_37.

Kang:2012:KFB

[538] Min-Jae Kang, Chang-Jin Boo, and Ho-Chan Kim. A kernel function based estimation algorithm for multi-layer soil structure. Lecture Notes in Computer Science, 6840:276–281, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_38.

Mehera:2012:CTA

[539] Chitralekha Mehera. Computing technique in ancient India. Lecture Notes in Computer Science, 6840:282–289, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24553-4_39.

DiLecce:2012:SSH

[540] Vincenzo Di Lecce and Marco Calabrese. Smart sensors: a holonic perspective. Lecture Notes in Computer Science, 6840:290–298, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24553-4_40.

Lin:2012:RGE

[541] Nay Yee Lin, Khin Mar Soe, and Ni Lar Thein. Reducing grammar errors for translated English sentences. Lecture Notes in Computer Science, 6840:299–306, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_41.

Qi:2012:ICR

[542] Shuhan Qi, Xuan Wang, and Xinxin Li. Improve coreference resolution with parameter tunable anaphoricity identification and global optimization. Lecture Notes in Computer Science, 6840:307–314, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24553-4_42.

Zhang:2012:VMR

[543] Ming Zhang, Zhenmin Tang, Weiyan Xu, and Xibei yang. A variable muitlgranulation rough sets approach. Lecture Notes in Computer Science, 6840:315–322, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-24553-4_43.

Wang:2012:IMR

[544] Li juan Wang, Xi bei Yang, Jing yu Yang, and Chen Wu. Incomplete multigranulation rough sets in incomplete ordered decision system. Lecture Notes in Computer Science, 6840:323–330, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24553-4_44.

Morales:2012:SCF

[545] Omar Salazar Morales and José Jairo Soriano Méndez. A special class of fuzzified normal forms. Lecture Notes in Computer Science, 6840: 331–344, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_45.

Liu:2012:MEM

[546] Ying Liu and Yanju Chen. Mean-entropy model for portfolio selection with type-2 fuzzy returns. Lecture Notes in Computer Science, 6840:345–352, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_46.

Figueroa-Garcia:2012:LRM

[547] Juan Carlos Figueroa-García and Jesus Rodriguez-Lopez. A linear regression model for nonlinear fuzzy data. Lecture Notes in Computer Science, 6840:353–360, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_47.

De:2012:NRI

[548] Indrajit De and Jaya Sil. No reference image quality assessment by designing fuzzy relational classifier using MOS weight matrix. Lecture Notes in Computer Science, 6840:361–369, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_48.

Xu:2012:ACO

[549] Jiangqiao Xu and Xianhai Song. Ant colony optimization for nonlinear inversion of Rayleigh waves. Lecture Notes in Computer Science, 6840:370–377, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_49.

Liang:2012:MFE

[550] J. J. Liang, Bo Yang Qu, Song Tao Ma, and Ponnuthurai Nagaratnam Suganthan. Memetic fitness Euclidean-distance particle swarm optimization for multi-modal optimization. Lecture Notes in Computer Science, 6840:378–385, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_50.

Tekchandani:2012:CDM

[551] Prakash Tekchandani and Aditya Trivedi. Clock drift management using particle swarm optimization. Lecture Notes in Computer Science, 6840:386–393, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_51.

Taherdangkoo:2012:SCO

[552] Mohammad Taherdangkoo, Mehran Yazdi, and Mohammad Hadi Bagheri. Stem cells optimization algorithm. Lecture Notes in Computer Science, 6840:394–403, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_52.

Xu:2012:CCS

[553] Xiaohua Xu, Zhoujin Pan, Ping He, and Ling Chen. Constrained clustering via swarm intelligence. Lecture Notes in Computer Science, 6840:404–409, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_53.

Chen:2012:PBH

[554] Yu Chen, Kai Zhang, and Xiufen Zou. A population-based hybrid extremal optimization algorithm. Lecture Notes in Computer Science, 6840:410–417, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_54.

Zhao:2012:EAC

[555] Li-Qing Zhao, Cui Zhang, and Rong-Long Wang. An effective ant colony algorithm for graph planarization problem. Lecture Notes in Computer Science, 6840:418–425, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_55.

Huang:2012:NCG

[556] Kai Huang and Yong quan Zhou. A novel chaos glowworm swarm optimization algorithm for optimization functions. Lecture Notes in Computer Science, 6840:426–434, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_56.

Wang:2012:SSP

[557] Jingyan Wang, Yongping Li, Ying Zhang, and Jianhua He. Semisupervised protein function prediction via sequential linear neighborhood propagation. Lecture Notes in Computer Science, 6840:435–441, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24553-4_57.

Wang:2012:INA

[558] Xuesong Wang, Lijing Li, and Yuhu Cheng. An improved Newman algorithm for mining overlapping modules from protein-protein interaction networks. Lecture Notes in Computer Science, 6840:442–447, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24553-4_58.

Wang:2012:PIR

[559] Bing Wang, Peng Chen, and Jun Zhang. Protein interface residues prediction based on amino acid properties only. Lecture Notes in Computer Science, 6840:448–452, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_59.

Bonilla-Huerta:2012:HFW

[560] Edmundo Bonilla-Huerta, Béatrice Duval, José C. Hernández Hernández, and Jin-Kao Hao. Hybrid filter-wrapper with a specialized random multi-parent crossover operator for gene selection and classification problems. Lecture Notes in Computer Science, 6840:453–461, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24553-4_60.

Yu:2012:SADa

[561] Hong-Jie Yu. Similarity analysis of DNA sequences based on three 2-D cumulative ratio curves. Lecture Notes in Computer Science, 6840: 462–469, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_61.

Yu:2012:SADb

[562] Hong-Jie Yu. Similarity analysis of DNA barcodes sequences based on compressed feature vectors. Lecture Notes in Computer Science, 6840:470–477, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_62.

Yuan:2012:IDG

[563] Fang Yuan, Jing Li, and Lun Li. Indentifying disease genes using disease-specific amino acid usage. Lecture Notes in Computer Science, 6840:478–485, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_63.

Wang:2012:DCP

[564] Bing Wang, Aiqin Fang, Xue Shi, Seong Ho Kim, and Xiang Zhang. DISCO2: a comprehensive peak alignment algorithm for two-dimensional gas chromatography time-of-flight mass spectrometry. Lecture Notes in Computer Science, 6840:486–491, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_64.

Cui:2012:PHP

[565] Guangyu Cui, Chao Fang, and Kyungsook Han. Prediction of human proteins interacting with human papillomavirus proteins. Lecture Notes in Computer Science, 6840:492–497, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_65.

Bevilacqua:2012:CDM

[566] Vitoantonio Bevilacqua, Paolo Pannarale, Mirko Abbrescia, and Claudia Cava. Comparison of data-merging methods with SVM attribute selection and classification in breast cancer gene expression. Lecture Notes in Computer Science, 6840:498–507, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_66.

Hsiao:2012:EGR

[567] Yu-Ting Hsiao and Wei-Po Lee. Evolving gene regulatory networks: a sensitivity-based approach. Lecture Notes in Computer Science, 6840: 508–513, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_67.

Zhou:2012:BTQ

[568] Wei Zhou, Gang Liu, Qiwei Shi, Shilei Cui, Yina Zhou, Huili Zhu, and Rubin Wang. Blind testing of quasi brain deaths based on analysis of EEG energy. Lecture Notes in Computer Science, 6840:514–520, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24553-4_68.

Hasan:2012:CAH

[569] Md. Kamrul Hasan, Wook Lee, Byungkyu Park, and Kyungsook Han. Connectivity analysis of hippocampus in Alzheimer’s brain using probabilistic tractography. Lecture Notes in Computer Science, 6840:521–528, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24553-4_69.

Wang:2012:PSS

[570] Jeen-Shing Wang, Che-Wei Lin, Ya-Ting C. Yang, Tzu-Ping Kao, and Wei-Hsin Wang. A PACE sensor system with machine learning-based energy expenditure regression algorithm. Lecture Notes in Computer Science, 6840:529–536, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-24553-4_70.

Yao:2012:FHA

[571] Xiaohui Yao, Yun Xu, and Jiaoyun Yang. A faster haplotyping algorithm based on block partition, and greedy ligation strategy. Lecture Notes in Computer Science, 6840:537–544, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_71.

Wang:2012:EEA

[572] Jeen-Shing Wang, Wei-Chun Chiang, Ya-Ting C. Yang, and Yu-Liang Hsu. An effective ECG arrhythmia classification algorithm. Lecture Notes in Computer Science, 6840:545–550, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_72.

Mi:2012:NSA

[573] Jian-Xun Mi. A new subspace approach for face recognition. Lecture Notes in Computer Science, 6840:551–557, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24553-4_73.

Li:2012:NAB

[574] Li Li, Fangmin Yao, Lijing Tan, Ben Niu, and Jun Xu. A novel DE– ABC-based hybrid algorithm for global optimization. Lecture Notes in Computer Science, 6840:558–565, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_74.

Li:2012:DAB

[575] Li Li, Yurong Cheng, Lijing Tan, and Ben Niu. A discrete artificial bee colony algorithm for TSP problem. Lecture Notes in Computer Science, 6840:566–573, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_75.

Gao:2012:REL

[576] Hongwei Gao, Xiaofeng Liu, Jinguo Liu, Fuguo Chen, and Ben Niu. Restoration of epipolar line based on multi-population cooperative particle swarm optimization. Lecture Notes in Computer Science, 6840: 574–581, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_76.

Niu:2012:MOO

[577] Ben Niu, Hong Wang, Lijing Tan, and Jun Xu. Multi-objective optimization using BFO algorithm. Lecture Notes in Computer Science, 6840:582–587, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_77.

Wang:2012:UPC

[578] Shu-Lin Wang, Min Li, and Hongqiang Wang. Using 2D principal component analysis to reduce dimensionality of gene expression profiles for tumor classification. Lecture Notes in Computer Science, 6840:588–595, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24553-4_78.

Bevilacqua:2012:VCP

[579] Vitoantonio Bevilacqua, Domenico De Fano, Silvia Giannini, and Giuseppe Mastronardi. 3D virtual colonoscopy for polyps detection by supervised artificial neural networks. Lecture Notes in Computer Science, 6840:596–603, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_79.

Sharma:2012:FRS

[580] Mohit Sharma, Surya Prakash, and Phalguni Gupta. Face recognition system robust to occlusion. Lecture Notes in Computer Science, 6840:604–609, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_80.

Wang:2012:PBC

[581] Hong-Qiang Wang, Xin-Ping Xie, and Chun-Hou Zheng. A pathwaybased classification method that can improve microarray-based colorectal cancer diagnosis. Lecture Notes in Computer Science, 6840:610–617, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24553-4_81.

Gromiha:2012:SFR

[582] M. Michael Gromiha, R. Sowdhamini, and K. Fukui. Structure-function relationship in olfactory receptors. Lecture Notes in Computer Science, 6840:618–623, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_82.

Lai:2012:FRK

[583] Lien-Fu Lai, Chao-Chin Wu, and Liang-Tsung Huang. First report of knowledge discovery in predicting protein folding rate change upon single mutation. Lecture Notes in Computer Science, 6840:624–631, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24553-4_83.

Sun:2012:DPS

[584] Cheng li Sun, Yong Xu, Jie Jia, and Yu He. Detection of protein spots from complex region on real gel image. Lecture Notes in Computer Science, 6840:632–640, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_84.

Bevilacqua:2012:NAC

[585] Vitoantonio Bevilacqua, Fabio Stroppa, Stefano Saladino, and Ernesto Picardi. A novel approach to clustering and assembly of large-scale Roche 454 transcriptome data for gene validation and alternative splicing analysis. Lecture Notes in Computer Science, 6840:641–648, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24553-4_85.

Cui:2012:SMM

[586] Yan Cui, Jian Yang, and Chun-Hou Zheng. Sparse maximum margin discriminant analysis for gene selection. Lecture Notes in Computer Science, 6840:649–656, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_86.

Yang:2012:MDB

[587] Xibei Yang. The models of dominance–based multigranulation rough sets. Lecture Notes in Computer Science, 6840:657–664, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-24553-4_87.

Zhang:2012:IFD

[588] Yanqin Zhang and Xibei Yang. An intuitionistic fuzzy dominance– based rough set. Lecture Notes in Computer Science, 6840:665–672, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24553-4_88.

Lin:2012:CBP

[589] Guoping Lin and Jinjin Li. A covering-based pessimistic multigranulation rough set. Lecture Notes in Computer Science, 6840:673–680, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
24553-4_89.

Xu:2012:GMG

[590] Weihua Xu, Xiantao Zhang, and Qiaorong Wang. A generalized multigranulation rough set approach. Lecture Notes in Computer Science, 6840:681–689, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_90.

Wang:2012:CLC

[591] Junhong Wang, Jiye Liang, and Yuhua Qian. Closed-label concept lattice based rule extraction approach. Lecture Notes in Computer Science, 6840:690–698, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-24553-4_91.

Han:2012:IEL

[592] Fei Han, Hai-Fen Yao, and Qing-Hua Ling. An improved extreme learning machine based on particle swarm optimization. Lecture Notes in Computer Science, 6840:699–704, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-24553-4_92.

Jin-lin:2012:NNG

[593] Ding Jin-lin and Wang Feng. Neural network generalized inverse of multimotor synchronous system working on vector control mode. Lecture Notes in Computer Science, 6840:705–710, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-24553-4_93.

Anonymous:2012:BMn

[594] Anonymous. Back matter. Lecture Notes in Computer Science, 6840: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-24553-4/1.

Anonymous:2012:FMr

[595] Anonymous. Front matter. Lecture Notes in Computer Science, 6840: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-24553-4/1.

Moot:2012:CCG

[596] Richard Moot and Christian Retoré. Classical categorial grammars: AB grammars. Lecture Notes in Computer Science, 6850:1–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31555-8_1/.

Moot:2012:LCGa

[597] Richard Moot and Christian Retoré. A logic for categorial grammars: Lambek’s syntactic calculus. Lecture Notes in Computer Science, 6850: 23–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31555-8_2/;   http://link.springer.com/content/pdf/10.1007/
978-3-642-31555-8_2.

Moot:2012:LCM

[598] Richard Moot and Christian Retoré. Lambek calculus and Montague grammar. Lecture Notes in Computer Science, 6850:65–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31555-8_3/.

Moot:2012:NAL

[599] Richard Moot and Christian Retoré. The non-associative Lambek calculus. Lecture Notes in Computer Science, 6850:101–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31555-8_4/.

Moot:2012:MLC

[600] Richard Moot and Christian Retoré. The multimodal Lambek calculus. Lecture Notes in Computer Science, 6850:149–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31555-8_5/.

Moot:2012:LCL

[601] Richard Moot and Christian Retoré. Lambek calculus and linear logic: Proof nets as parse structures. Lecture Notes in Computer Science, 6850:193–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31555-8_6/.

Moot:2012:PNM

[602] Richard Moot and Christian Retoré. Proof nets for the multimodal Lambek calculus: From theory to a wide-coverage categorial parser. Lecture Notes in Computer Science, 6850:239–297, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31555-8_7/.

Anonymous:2012:BMo

[603] Anonymous. Back matter. Lecture Notes in Computer Science, 6850: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31555-8/1.

Moot:2012:FM

[604] Richard Moot and Christian Retoré. Front matter. Lecture Notes in Computer Science, 6850:??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/bfm:978-3-642-31555-8/1.

Raji:2012:PPO

[605] Fatemeh Raji, Ali Miri, and Mohammad Davarpanah Jazi. Preserving privacy in online social networks. Lecture Notes in Computer Science, 6888:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27901-0_1/.

Pujol-Ahullo:2012:TSB

[606] Jordi Pujol-Ahulló, Roger Jardí-Cedó, Jordi Castellà-Roca, and Oriol Farràs. TTP SmartCard-based ElGamal cryptosystem using threshold scheme for electronic elections. Lecture Notes in Computer Science, 6888: 14–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27901-0_2/.

Gagne:2012:AVB

[607] Martin Gagné, Pascal Lafourcade, Yassine Lakhnech, and Reihaneh Safavi-Naini. Automated verification of block cipher modes of operation, an improved method. Lecture Notes in Computer Science, 6888:23– 31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27901-0_3/.

Dragoni:2012:SSC

[608] Nicola Dragoni, Eduardo Lostal, Davide Papini, and Javier Fabra. SC^{2}: Secure communication over Smart Cards. Lecture Notes in Computer Science, 6888:32–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27901-0_4/.

Malek:2012:PIR

[609] Behzad Malek and Ali Miri. Private identification of RFID tags. Lecture Notes in Computer Science, 6888:49–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27901-0_5/.

Stanek:2012:TEM

[610] Martin Stanek. Threshold encryption into multiple ciphertexts. Lecture Notes in Computer Science, 6888:62–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27901-0_6/.

Chen:2012:NCB

[611] Yu Chen, Song Luo, Jianbin Hu, and Zhong Chen. A novel commutative blinding identity based encryption scheme. Lecture Notes in Computer Science, 6888:73–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27901-0_7/.

Pudovkina:2012:RKA

[612] Marina Pudovkina. A related-key attack on block ciphers with weak recurrent key schedules. Lecture Notes in Computer Science, 6888:90– 101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27901-0_8/.

Barthe:2012:CIL

[613] Gilles Barthe, Mathilde Duclos, and Yassine Lakhnech. A computational indistinguishability logic for the bounded storage model. Lecture Notes in Computer Science, 6888:102–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27901-0_9/.

Cervera:2012:PCF

[614] Gimer Cervera, Michel Barbeau, Joaquin Garcia-Alfaro, and Evangelos Kranakis. Preventing the cluster formation attack against the hierarchical OLSR protocol. Lecture Notes in Computer Science, 6888:118– 131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27901-0_10/.

Ahmadi:2012:SKE

[615] Hadi Ahmadi and Reihaneh Safavi-Naini. Secret key establishment over noisy channels. Lecture Notes in Computer Science, 6888:132– 147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27901-0_11/.

Bourdier:2012:FSV

[616] Tony Bourdier, Horatiu Cirstea, Mathieu Jaume, and Hélène Kirchner. Formal specification and validation of security policies. Lecture Notes in Computer Science, 6888:148–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27901-0_12/.

Dreier:2012:VIP

[617] Jannik Dreier, Pascal Lafourcade, and Yassine Lakhnech. Vote-Independence: a powerful privacy notion for voting protocols. Lecture Notes in Computer Science, 6888:164–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27901-0_13/.

Rjasko:2012:BBP

[618] Michal Rjaško. Black-box property of cryptographic hash functions. Lecture Notes in Computer Science, 6888:181–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27901-0_14/.

Huh:2012:PDP

[619] Jun Ho Huh and Hyoungshick Kim. Phishing detection with popular search engines: Simple and effective. Lecture Notes in Computer Science, 6888:194–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27901-0_15/.

Papagiannakopoulou:2012:CPA

[620] Eugenia I. Papagiannakopoulou, Maria N. Koukovini, and Georgios V. Lioudakis. A contextual privacy-aware access control model for network monitoring workflows: Work in progress. Lecture Notes in Computer Science, 6888:208–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27901-0_16/.

Jiague:2012:MDA

[621] Michel Embe Jiague, Marc Frappier, Frédéric Gervais, Régine Laleau, and Richard St-Denis. A metamodel for the design of access-control policy enforcement managers: Work in progress. Lecture Notes in Computer Science, 6888:218–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27901-0_17/.

Konopacki:2012:SVA

[622] Pierre Konopacki, Hakim Belhaouari, Marc Frappier, and Régine Laleau. Specification and verification of access control policies in EB^{3} SEC: Work in progress. Lecture Notes in Computer Science, 6888:227–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27901-0_18/.

Milhau:2012:MMA

[623] Jérémy Milhau, Marc Frappier, and Régine Laleau. A metamodel of the B modeling of access-control policies: Work in progress. Lecture Notes in Computer Science, 6888:234–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27901-0_19/.

Tounsi:2012:FVK

[624] Wiem Tounsi, Nora Cuppens-Boulahia, Frédéric Cuppens, and Joaquin Garcia-Alfaro. Formal verification of a key establishment protocol for EPC gen2 RFID systems: Work in progress. Lecture Notes in Computer Science, 6888:242–251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27901-0_20/.

Anonymous:2012:BMp

[625] Anonymous. Back matter. Lecture Notes in Computer Science, 6888: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27901-0/1.

Anonymous:2012:FMs

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27901-0/1.

Banaji:2012:CSS

[627] Murad Banaji. Cycle structure in SR and DSR graphs: Implications for multiple equilibria and stable oscillation in chemical reaction networks. Lecture Notes in Computer Science, 6900:1–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29072-5_1/.

Bergenthum:2012:MML

[628] Robin Bergenthum, Jörg Desel, and Andreas Harrer. Modeling and mining of learnflows. Lecture Notes in Computer Science, 6900:22–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29072-5_2/.

Ciardo:2012:TYS

[629] Gianfranco Ciardo, Yang Zhao, and Xiaoqing Jin. Ten years of saturation: a Petri net perspective. Lecture Notes in Computer Science, 6900: 51–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29072-5_3/.

Elhog-Benzina:2012:RAC

[630] Dorsaf Elhog-Benzina and Serge Haddad. Refinement and asynchronous composition of modal Petri nets. Lecture Notes in Computer Science, 6900:96–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29072-5_4/.

Hong:2012:CHS

[631] Silien Hong and Fabrice Kordon. Computing a hierarchical static order for decision diagram-based representation from P/T nets. Lecture Notes in Computer Science, 6900:121–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29072-5_5/.

Knapik:2012:BMC

[632] Michał Knapik and Wojciech Penczek. Bounded model checking for parametric timed automata. Lecture Notes in Computer Science, 6900:141– 159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29072-5_6/.

Koutny:2012:SPP

[633] Maciej Koutny and Marta Pietkiewicz-Koutny. Synthesis problem for Petri nets with localities. Lecture Notes in Computer Science, 6900:160– 180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29072-5_7/.

Lopez-Grao:2012:PNP

[634] Juan-Pablo López-Grao and José-Manue Colom. A Petri net perspective on the resource allocation problem in software engineering. Lecture Notes in Computer Science, 6900:181–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29072-5_8/.

Mascheroni:2012:NWN

[635] Marco Mascheroni and Fabio Farina. Nets-within-nets paradigm and grid computing. Lecture Notes in Computer Science, 6900:201–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29072-5_9/.

Sole:2012:IPD

[636] Marc Solé and Josep Carmona. Incremental process discovery. Lecture Notes in Computer Science, 6900:221–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29072-5_10/.

Wagner:2012:PAF

[637] Thomas Wagner, José Quenum, and Daniel Moldt. Providing an agent flavored integration for workflow management. Lecture Notes in Computer Science, 6900:243–264, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29072-5_11/.

Westergaard:2012:GAC

[638] Michael Westergaard and Lars Michael Kristensen. A graphical approach to component-based and extensible model checking platforms. Lecture Notes in Computer Science, 6900:265–291, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29072-5_12/.

Anonymous:2012:BMq

[639] Anonymous. Back matter. Lecture Notes in Computer Science, 6900: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29072-5/1.

Anonymous:2012:FMt

[640] Anonymous. Front matter. Lecture Notes in Computer Science, 6900: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29072-5/1.

Aichholzer:2012:EMAa

[641] Oswin Aichholzer, Wolfgang Aigner, Franz Aurenhammer, and Bert Jüttler. Exact medial axis computation for triangulated solids with respect to piecewise linear metrics. Lecture Notes in Computer Science, 6920:1–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27413-8_1/.

Aichholzer:2012:EMAb

[642] Oswin Aichholzer, Wolfgang Aigner, Thomas Hackl, and Nicola Wolpert. Exact medial axis computation for circular arc boundaries. Lecture Notes in Computer Science, 6920:28–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27413-8_2/.

Ait-Haddou:2012:CBC

[643] Rachid Ait-Haddou and Taishin Nomura. Complex Bézier curves and the geometry of polynomials. Lecture Notes in Computer Science, 6920:43– 65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27413-8_3/.

Alcazar:2012:SCP

[644] Juan Gerardo Alcázar. The shape of conchoids to plane algebraic curves. Lecture Notes in Computer Science, 6920:66–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27413-8_4/.

Allouch:2012:EIP

[645] C. Allouch, P. Sablonnière, and D. Sbibih. Estimation of integral properties of a planar closed curve based on a quadratic spline quasi-interpolant. Lecture Notes in Computer Science, 6920:80–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27413-8_5/.

Arandiga:2012:DMO

[646] Francesc Aràndiga, Albert Cohen, and Dionisio F. Yáñez. Design of multiresolution operators using statistical learning tools: Application to compression of signals. Lecture Notes in Computer Science, 6920:94– 108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27413-8_6/.

Arandiga:2012:WPN

[647] Francesc Aràndiga, Rosa Donat, and Maria Santágueda. Weightedpower_{p} nonlinear subdivision schemes. Lecture Notes in Computer Science, 6920:109–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27413-8_7/.

Avenel:2012:TLS

[648] Christophe Avenel, Etienne Mémin, and Patrick Pérez. Tracking level set representation driven by a stochastic dynamics. Lecture Notes in Computer Science, 6920:130–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27413-8_8/.

Bastl:2012:HIC

[649] Bohumír Bastl, Miroslav Lávička, and Zbyněk Šír. G^{2} Hermite interpolation with curves represented by multi-valued trigonometric support functions. Lecture Notes in Computer Science, 6920:142–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27413-8_9/.

Bela:2012:AAS

[650] Szilvia Béla and Bert Jüttler. Approximating algebraic space curves by circular arcs. Lecture Notes in Computer Science, 6920:157–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27413-8_10/.

Benmakrouha:2012:GSD

[651] Farida Benmakrouha, Christiane Hespel, and Edouard Monnier. Generating series for drawing the output of dynamical systems. Lecture Notes in Computer Science, 6920:178–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27413-8_11/.

Bommes:2012:PMI

[652] David Bommes, Henrik Zimmer, and Leif Kobbelt. Practical mixedinteger optimization for geometry processing. Lecture Notes in Computer Science, 6920:193–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27413-8_12/.

Canton:2012:NDD

[653] Alicia Cantón and Leonardo Fernández-Jambrina. Non-degenerate developable triangular Bézier patches. Lecture Notes in Computer Science, 6920:207–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27413-8_13/.

Davydov:2012:SSB

[654] Oleg Davydov and Abid Saeed. Stable splitting of bivariate splines spaces by Bernstein–Bézier methods. Lecture Notes in Computer Science, 6920:220–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27413-8_14/.

Digne:2012:MSM

[655] Julie Digne, Jean-Michel Morel, and Charyar Mehdi-Souzani. Mesh segmentation and model extraction. Lecture Notes in Computer Science, 6920:236–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27413-8_15/.

Farouki:2012:DSP

[656] Rida T. Farouki, Carla Manni, Francesca Pelosi, and Maria Lucia Sampoli. Design of C^{2} spatial Pythagorean-hodograph quintic spline curves by control polygons. Lecture Notes in Computer Science, 6920:253–269, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27413-8_16/.

Georgiev:2012:SCP

[657] Georgi H. Georgiev. Shape curvatures of planar rational spirals. Lecture Notes in Computer Science, 6920:270–279, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27413-8_17/.

Grossmann:2012:VGR

[658] David Großmann and Bert Jüttler. Volumetric geometry reconstruction of turbine blades for aircraft engines. Lecture Notes in Computer Science, 6920:280–295, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27413-8_18/.

Harizanov:2012:GCA

[659] Stanislav Harizanov. Globally convergent adaptive normal multiscale transforms. Lecture Notes in Computer Science, 6920:296–310, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27413-8_19/.

Harouna:2012:HHD

[660] Souleymane Kadri Harouna and Valérie Perrier. Helmholtz–Hodge decomposition on [0, 1]^{d} by divergence-free and curl-free wavelets. Lecture Notes in Computer Science, 6920:311–329, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27413-8_20/.

Hollig:2012:FEA

[661] Klaus Höllig, Jörg Hörner, and Axel Hoffacker. Finite element analysis with B-splines: Weighted and isogeometric methods. Lecture Notes in Computer Science, 6920:330–350, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27413-8_21.

Huang:2012:BFS

[662] Jinghao Huang and Peter Schröder. \sqrt{3}-based 1-form subdivision. Lecture Notes in Computer Science, 6920:351–368, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27413-8_22.

Karciauskas:2012:CAC

[663] Kȩstutis Karčiauskas and Jörg Peters. Curvature of approximating curve subdivision schemes. Lecture Notes in Computer Science, 6920:369–381, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27413-8_23.

Kergosien:2012:FSO

[664] Yannick L. Kergosien. Fitting a surface to one of its sectional planar curves using adaptive trees in spaces of curves. Lecture Notes in Computer Science, 6920:382–401, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27413-8_24.

Kiel:2012:VSS

[665] Stefan Kiel. Verified spatial subdivision of implicit objects using implicit linear interval estimations. Lecture Notes in Computer Science, 6920:402–415, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27413-8_25.

Kutyniok:2012:ISU

[666] Gitta Kutyniok and Wang-Q Lim. Image separation using wavelets and shearlets. Lecture Notes in Computer Science, 6920:416–430, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27413-8_26.

Lavicka:2012:SCP

[667] Miroslav Lávička and Jan Vršek. On a special class of polynomial surfaces with Pythagorean normal vector fields. Lecture Notes in Computer Science, 6920:431–444, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27413-8_27.

Liu:2012:CRC

[668] Da yan Liu, Olivier Gibaru, and Wilfrid Perruquetti. Convergence rate of the causal Jacobi derivative estimator. Lecture Notes in Computer Science, 6920:445–455, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27413-8_28.

Esturo:2012:CDI

[669] Janick Martinez Esturo, Christian Rössl, and Holger Theisel. Continuous deformations by isometry preserving shape integration. Lecture Notes in Computer Science, 6920:456–472, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27413-8_29.

Matei:2012:WET

[670] Basarab Mateï, Sylvain Meignen, and Anastasia Zakharova. On a (W)ENO-type multiscale representation based on quincunx refinement: Application to image compression. Lecture Notes in Computer Science, 6920:473–487, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27413-8_30.

Mobius:2012:OOS

[671] Jan Möbius and Leif Kobbelt. OpenFlipper: an open source geometry processing and rendering framework. Lecture Notes in Computer Science, 6920:488–500, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27413-8_31.

Nguyen:2012:PCD

[672] Thien Nguyen and Bert Jüttler. Parameterization of contractible domains using sequences of harmonic maps. Lecture Notes in Computer Science, 6920:501–514, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-27413-8_32.

Nyiri:2012:NIA

[673] E. Nyiri, O. Gibaru, and Ph. Auquiert. Nonlinear L_{1} C^{1} interpolation: Application to images. Lecture Notes in Computer Science, 6920:515–526, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27413-8_33.

Oswald:2012:NMS

[674] Peter Oswald. Normal multi-scale transforms for surfaces. Lecture Notes in Computer Science, 6920:527–542, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27413-8_34.

Peternell:2012:GDC

[675] Martin Peternell. Generalized Dupin cyclides with rational lines of curvature. Lecture Notes in Computer Science, 6920:543–552, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27413-8_35.

Pettersen:2012:SVF

[676] Kjell Fredrik Pettersen and Vibeke Skytt. Spline volume fairing. Lecture Notes in Computer Science, 6920:553–561, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27413-8_36.

Pitolli:2012:NCL

[677] Francesca Pitolli. Neuroelectric current localization from combined EEG/MEG data. Lecture Notes in Computer Science, 6920:562–574, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27413-8_37.

Ramli:2012:BBN

[678] Ahmad Ramli and Ioannis Ivrissimtzis. Bootstrap-based normal reconstruction. Lecture Notes in Computer Science, 6920:575–585, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27413-8_38.

Rossl:2012:CPL

[679] Christian Rössl and Holger Theisel. Couple points — a local approach to global surface analysis. Lecture Notes in Computer Science, 6920: 586–602, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27413-8_39.

Sablonniere:2012:CCS

[680] P. Sablonnière, D. Sbibih, and M. Tahrichi. Chordal cubic spline quasi interpolation. Lecture Notes in Computer Science, 6920:603–611, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27413-8_40.

Sauer:2012:MSS

[681] Tomas Sauer. Multiple subdivision schemes. Lecture Notes in Computer Science, 6920:612–628, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27413-8_41.

Schoenemann:2012:LPA

[682] Thomas Schoenemann, Simon Masnou, and Daniel Cremers. On a linear programming approach to the discrete Willmore boundary value problem and generalizations. Lecture Notes in Computer Science, 6920:629–646, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27413-8_42.

Simsek:2012:IFG

[683] Yilmaz Simsek. Interpolation function of generalized q-Bernstein-type basis polynomials and applications. Lecture Notes in Computer Science, 6920:647–662, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27413-8_43.

Sokolov:2012:DBI

[684] Dmitry Sokolov, Christian Gentil, and Hicham Bensoudane. Differential behaviour of iteratively generated curves. Lecture Notes in Computer Science, 6920:663–680, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27413-8_44.

Vrsek:2012:ACL

[685] Jan Vršek and Miroslav Lávička. Algebraic curves of low convolution degree. Lecture Notes in Computer Science, 6920:681–696, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27413-8_45.

Yang:2012:LMD

[686] Ying Yang and Ioannis Ivrissimtzis. A logistic model for the degradation of triangle mesh normals. Lecture Notes in Computer Science, 6920:697–710, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27413-8_46.

Zeyde:2012:SIS

[687] Roman Zeyde, Michael Elad, and Matan Protter. On single image scaleup using sparse-representations. Lecture Notes in Computer Science, 6920:711–730, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27413-8_47.

Zheng:2012:PST

[688] Jianmin Zheng and Yimin Wang. Periodic T-splines and tubular surface fitting. Lecture Notes in Computer Science, 6920:731–746, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27413-8_48.

Anonymous:2012:BMr

[689] Anonymous. Back matter. Lecture Notes in Computer Science, 6920: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27413-8/1.

Anonymous:2012:FMu

[690] Anonymous. Front matter. Lecture Notes in Computer Science, 6920: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27413-8/1.

Seshia:2012:QAS

[691] Sanjit A. Seshia. Quantitative analysis of software: Challenges and recent advances. Lecture Notes in Computer Science, 6921:1–5, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27269-1_1/.

Caires:2012:ASO

[692] Luís Caires and Hugo Torres Vieira. Analysis of service oriented software systems with the conversation calculus. Lecture Notes in Computer Science, 6921:6–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27269-1_2/.

Tamura:2012:QCA

[693] Gabriel Tamura, Rubby Casallas, and Anthony Cleve. QoS contractaware reconfiguration of component architectures using E-graphs. Lecture Notes in Computer Science, 6921:34–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27269-1_3/.

Nobakht:2012:MMC

[694] B. Nobakht, M. M. Bonsangue, and F. S. de Boer. Monitoring method call sequences using annotations. Lecture Notes in Computer Science, 6921:53–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27269-1_4/.

Calder:2012:IPI

[695] M. Calder, P. Gray, A. Miller, and C. Unsworth. An introduction to pervasive interface automata. Lecture Notes in Computer Science, 6921:

71–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27269-1_5/.

Yijing:2012:SLO

[696] Liu Yijing and Qiu Zongyan. A separation logic for OO programs. Lecture Notes in Computer Science, 6921:88–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27269-1_6/.

Rossi:2012:MCA

[697] Sabina Rossi. Model checking adaptive multilevel service compositions. Lecture Notes in Computer Science, 6921:106–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27269-1_7/.

Andova:2012:DAD

[698] S. Andova, L. P. J. Groenewegen, and E. P. de Vink. Distributed adaption of dining philosophers. Lecture Notes in Computer Science, 6921:125–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27269-1_8/.

Andre:2012:CSP

[699] Pascal André, Gilles Ardourel, and Mohamed Messabihi. Component service promotion: Contracts, mechanisms and safety. Lecture Notes in Computer Science, 6921:145–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27269-1_9/.

Cote:2012:STV

[700] Daniel Côté, Michel Embe Jiague, and Richard St-Denis. Systemstheoretic view of component-based software development. Lecture Notes in Computer Science, 6921:163–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27269-1_10/.

Mouheb:2012:AWU

[701] Djedjiga Mouheb, Dima Alhadidi, and Mariam Nouh. Aspect weaving in UML activity diagrams: a semantic and algorithmic framework. Lecture Notes in Computer Science, 6921:182–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27269-1_11/.

Dormoy:2012:UTL

[702] Julien Dormoy, Olga Kouchnarenko, and Arnaud Lanoix. Using temporal logic for dynamic reconfigurations of components. Lecture Notes in Computer Science, 6921:200–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27269-1_12/.

Ramirez-Deantes:2012:MTA

[703] D. Ramírez-Deantes, J. Correas, and G. Puebla. Modular termination analysis of Java bytecode and its application to phoneME core libraries. Lecture Notes in Computer Science, 6921:218–236, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27269-1_13/.

Pourvatan:2012:DCA

[704] Bahman Pourvatan, Marjan Sirjani, and Farhad Arbab. Decomposition of constraint automata. Lecture Notes in Computer Science, 6921:237– 258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27269-1_14/.

Bruni:2012:GRS

[705] Roberto Bruni, Zhiming Liu, and Liang Zhao. Graph representation of sessions and pipelines for structured service programming. Lecture Notes in Computer Science, 6921:259–276, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27269-1_15/.

Arbab:2012:WRS

[706] Farhad Arbab. Will the real service oriented computing please stand up? Lecture Notes in Computer Science, 6921:277–285, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27269-1_16/.

Adam:2012:PVC

[707] Ludwig Adam. Performance verification in complex enterprise-level component systems. Lecture Notes in Computer Science, 6921:286–289, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27269-1_17/.

Kirsch:2012:RPT

[708] Christoph M. Kirsch and Luís Lopes. Runtime programming through model-preserving, scalable runtime patches. Lecture Notes in Computer Science, 6921:290–294, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27269-1_18/.

Bravetti:2012:SRC

[709] Mario Bravetti, Cinzia Di Giusto, and Jorge A. Pérez. Steps on the road to component evolvability. Lecture Notes in Computer Science, 6921:295– 299, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27269-1_19/.

Macedo:2012:TLA

[710] Hugo Daniel Macedo and José Nuno Oliveira. Towards linear algebras of components. Lecture Notes in Computer Science, 6921:300–303, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27269-1_20/.

Anonymous:2012:BMs

[711] Anonymous. Back matter. Lecture Notes in Computer Science, 6921: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27269-1/1.

Anonymous:2012:FMv

[712] Anonymous. Front matter. Lecture Notes in Computer Science, 6921: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27269-1/1.

Shahar:2012:HCK

[713] Yuval Shahar. The “Human Cli-Knowme” project: Building a universal, formal, procedural and declarative clinical knowledge base, for the automation of therapy and research. Lecture Notes in Computer Science, 6924:1–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27697-2_1/.

Riano:2012:SAM

[714] David Riaño. A systematic analysis of medical decisions: How to store knowledge and experience in decision tables. Lecture Notes in Computer Science, 6924:23–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27697-2_2/.

Fdez-Olivares:2012:TNB

[715] Juan Fdez-Olivares, Inmaculada Sánchez-Garzón, Arturo González-Ferrer, and Juan A. Cózar. Task network based modeling, dynamic generation and adaptive execution of patient-tailored treatment plans based on smart process management technologies. Lecture Notes in Computer Science, 6924:37–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27697-2_3/.

Dentler:2012:TAC

[716] Kathrin Dentler, Annette ten Teije, Ronald Cornet, and Nicolette de Keizer. Towards the automated calculation of clinical quality indicators. Lecture Notes in Computer Science, 6924:51–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27697-2_4/.

Peleg:2012:REC

[717] Mor Peleg, Samson W. Tu, Giorgio Leonardi, Silvana Quaglini, and Paola Russo. Reasoning with effects of clinical guideline actions using OWL: AL amyloidosis as a case study. Lecture Notes in Computer Science, 6924:65–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27697-2_5/.

Bonacin:2012:CPS

[718] Rodrigo Bonacin, Cédric Pruski, and Marcos Da Silveira. Careflow personalization services: Concepts and tool for the evaluation of computerinterpretable guidelines. Lecture Notes in Computer Science, 6924:80– 93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27697-2_6/.

Hatko:2012:DGL

[719] Reinhard Hatko, Joachim Baumeister, Volker Belli, and Frank Puppe. Diaflux: a graphical language for computer-interpretable guidelines. Lecture Notes in Computer Science, 6924:94–107, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27697-2_7/.

Bahati:2012:ATC

[720] Raphael Bahati, Stacey Guy, and Femida Gwadry-Sridhar. Analysis of treatment compliance of patients with diabetes. Lecture Notes in Computer Science, 6924:108–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27697-2_8/.

Juarez:2012:CPO

[721] Jose M. Juarez, Manuel Campos, Antonio Gomariz, and Antonio Morales. Computing problem oriented medical records. Lecture Notes in Computer Science, 6924:117–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27697-2_9/.

Lopez-Vallverdu:2012:DDA

[722] Joan Albert López-Vallverdú, David Riaño, and Antoni Collado. Detecting dominant alternative interventions to reduce treatment costs. Lecture Notes in Computer Science, 6924:131–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27697-2_10/.

Milian:2012:PCT

[723] Krystyna Milian, Annette ten Teije, Anca Bucur, and Frank van Harmelen. Patterns of clinical trial eligibility criteria. Lecture Notes in Computer Science, 6924:145–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27697-2_11/.

Taylor:2012:MKR

[724] Paul Taylor and Igor Toujilov. Mammographic knowledge representation in description logic. Lecture Notes in Computer Science, 6924:158– 169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27697-2_12/.

Anonymous:2012:BMt

[725] Anonymous. Back matter. Lecture Notes in Computer Science, 6924: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27697-2/1.

Anonymous:2012:FMw

[726] Anonymous. Front matter. Lecture Notes in Computer Science, 6924: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27697-2/1.

Miro-Julia:2012:FCM

[727] Margaret Miró-Julià. A framework for combining multivalued data: a practical approach. Lecture Notes in Computer Science, 6927:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27549-4_1/.

Pichler:2012:MMR

[728] Franz Pichler. On modelling metabolism-repair by convolution and partial realization. Lecture Notes in Computer Science, 6927:9–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27549-4_2/.

Kopacek:2012:COH

[729] P. Kopacek. Cost oriented humanoid robots. Lecture Notes in Computer Science, 6927:18–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27549-4_3/.

deBlasio:2012:NBN

[730] Gabriel de Blasio, Arminda Moreno-Díaz, Roberto Moreno-Díaz Jr., and Roberto Moreno-Díaz. New biomimetic neural structures for artificial neural nets. Lecture Notes in Computer Science, 6927:25–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27549-4_4/.

Hernandez-Goya:2012:EOF

[731] C. Hernández-Goya, P. Caballero-Gil, J. Molina-Gil, and C. Caballero-Gil. Extending OLSR functionalities to PKI management. Lecture Notes in Computer Science, 6927:32–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27549-4_5/.

Surmacz:2012:BUO

[732] Tomasz Surmacz. Bandwidth usage optimization for NNTP protocol. Lecture Notes in Computer Science, 6927:40–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27549-4_6/.

Nikodem:2012:DCM

[733] Jan Nikodem. The dilemma of choice in management of communication processes in WSN. Lecture Notes in Computer Science, 6927:48– 55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27549-4_7/.

Martinelli:2012:DAS

[734] Fabio Martinelli and Paolo Mori. A distributed authorization system with mobile usage control policies. Lecture Notes in Computer Science, 6927:56–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27549-4_8/.

Lozada:2012:FLP

[735] Dayana Lozada, Jose Manuel Castillo, Alberto Salguero, and Francisco Araque. Fuzzy logic for the performance assessment of the innovation management in tourism. Lecture Notes in Computer Science, 6927:64– 71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27549-4_9/.

DiCrescenzo:2012:NDA

[736] Antonio Di Crescenzo and Maria Longobardi. Neuronal data analysis based on the empirical cumulative entropy. Lecture Notes in Computer Science, 6927:72–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27549-4_10/.

Giorno:2012:CDT

[737] Virginia Giorno, Amelia G. Nobile, and Luigi M. Ricciardi. On the construction of densities for time non-homogeneous diffusion processes. Lecture Notes in Computer Science, 6927:80–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27549-4_11/.

Fridli:2012:RFS

[738] Sándor Fridli, Levente Lócsi, and Ferenc Schipp. Rational function systems in ECG processing. Lecture Notes in Computer Science, 6927:88– 95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27549-4_12/.

Buonocore:2012:FPT

[739] Aniello Buonocore, Luigia Caputo, and Enrica Pirozzi. First-passage-time for Gauss-diffusion processes via integrated analytical, simulation and numerical methods. Lecture Notes in Computer Science, 6927:96– 104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27549-4_13/.

Roman:2012:MAS

[740] Monica Roman. Modelling aspects and structural properties of a fedbatch bioprocess. Lecture Notes in Computer Science, 6927:105–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27549-4_14/.

Heras:2012:CMS

[741] Jónathan Heras, Vico Pascual, and Julio Rubio. A certified module to study digital images with the Kenzo system. Lecture Notes in Computer Science, 6927:113–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27549-4_15/.

Rodriguez:2012:MPB

[742] Abraham Rodríguez Rodríguez, Nicolás Iglesias García, and José María Quinteiro-González. Modelling the psychographic behaviour of users using ontologies in Web marketing services. Lecture Notes in Computer Science, 6927:121–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27549-4_16/.

Schwaninger:2012:USD

[743] Markus Schwaninger and Christoph Mandl. Understanding the system dynamics of high-technology markets: Pólya processes with positive feedback, path dependence and lock-in. Lecture Notes in Computer Science, 6927:129–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27549-4_17/.

Pomarolli:2012:RRI

[744] Andreas Pomarolli, Stefan Anderlik, and Josef Küng. R2RIF — rule integration plugin for Protégé OWL. Lecture Notes in Computer Science, 6927:137–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27549-4_18/.

Chen:2012:GGT

[745] Qiao Chen and Heinz Dobler. GenComp — a generic transformation system. Lecture Notes in Computer Science, 6927:145–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27549-4_19/.

Krebs:2012:IUV

[746] Nico Krebs, Lothar Schmitz, and Uwe M. Borghoff. Implementing the universal virtual computer. Lecture Notes in Computer Science, 6927:153–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27549-4_20/.

Stumptner:2012:UGT

[747] Reinhard Stumptner, Bernhard Freudenthaler, Jürgen Hönigl, Karl Rehrl, and Josef Küng. Using GPS trajectories to create a dynamic network of significant locations as an abstraction of road maps. Lecture Notes in Computer Science, 6927:161–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27549-4_21.

Nistal:2012:CGC

[748] J. L. Freire Nistal, A. Blanco Ferro, J. M. Molinelli Barba, and E. Freire Brañas. On the confluence of the graphic calculus with Penrose diagrams (I). Lecture Notes in Computer Science, 6927:169–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27549-4_22.

Grimon:2012:SRI

[749] Francisca Grimón, Marylin Giugni, Joaquín Fernández, and Joseph Monguet. System for recommendation of information based on a management content model using software agents. Lecture Notes in Computer Science, 6927:177–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_23.

Szaban:2012:DCA

[750] Miroslaw Szaban and Franciszek Seredynski. Dynamic cellular automatabased S-boxes. Lecture Notes in Computer Science, 6927:184–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_24.

Kolar:2012:PTC

[751] Dušan Kolář and Šárka Květoňová. People transfer in city transport modeled via CPN. Lecture Notes in Computer Science, 6927:192–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_25.

Walchhofer:2012:ACE

[752] Norbert Walchhofer, Karl Anton Froeschl, and Kurt Hornik. Adaptive change estimation in the context of online market monitoring. Lecture Notes in Computer Science, 6927:200–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27549-4_26.

Matousek:2012:TKB

[753] Kamil Matoušek, Petr Křemen, Josef Küng, Reinhard Stumptner, and Stefan Anderlik. On transforming a knowledge base from topic maps to OWL. Lecture Notes in Computer Science, 6927:208–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27549-4_27.

Zwettler:2012:ABC

[754] Gerald Zwettler, Paul Track, Florian Waschaurek, Richard Woschitz, and Elmar Hagmann. Automated building construction design optimization for reduction of construction costs and energy demand. Lecture Notes in Computer Science, 6927:216–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_28.

Rubio-Largo:2012:UMO

[755] Álvaro Rubio-Largo, Miguel A. Vega-Rodríguez, and Juan A. Gómez-Pulido. Using a multiobjective OpenMP+MPI DE for the static RWA problem. Lecture Notes in Computer Science, 6927:224–231, 2012.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_29.

Gonzalez-Alvarez:2012:DDM

[756] David L. González-Álvarez, Miguel A. Vega-Rodríguez, and Juan A. Gómez-Pulido. Discovering DNA motifs with a parallel shared memory differential evolution. Lecture Notes in Computer Science, 6927:232–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_30.

Rafael:2012:OPS

[757] Brigitte Rafael, Stefan Oertl, Michael Affenzeller, and Stefan Wagner. Optimization of parameter settings for genetic algorithms in music segmentation. Lecture Notes in Computer Science, 6927:240–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_31.

Infuhr:2012:AGA

[758] Johannes Inführ and Günther R. Raidl. Automatic generation of 2-AntWars players with genetic programming. Lecture Notes in Computer Science, 6927:248–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_32.

Berlakovich:2012:MHR

[759] Martin Berlakovich, Mario Ruthmair, and Günther R. Raidl. A multilevel heuristic for the rooted delay-constrained minimum spanning tree problem. Lecture Notes in Computer Science, 6927:256–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_33.

Zavoianu:2012:IPR

[760] Alexandru-Ciprian Zăvoianu, Gabriel Kronberger, Michael Kommenda, and Daniela Zaharie. Improving the parsimony of regression models for an enhanced genetic programming process. Lecture Notes in Computer Science, 6927:264–271, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_34.

Cardenas-Montes:2012:GBE

[761] Miguel Cárdenas-Montes, Miguel A. Vega-Rodríguez, and Juan José Rodríguez-Vázquez. GPU-based evaluation to accelerate particle swarm algorithm. Lecture Notes in Computer Science, 6927:272–279, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_35.

Merkuryeva:2012:SBF

[762] Galina Merkuryeva and Vitaly Bolshakov. Simulation-based fitness landscape analysis and optimisation for vehicle scheduling problem. Lecture Notes in Computer Science, 6927:280–286, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27549-4_36.

Hu:2012:EAS

[763] Bin Hu and Günther R. Raidl. An evolutionary algorithm with solution archive for the generalized minimum spanning tree problem. Lecture Notes in Computer Science, 6927:287–294, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27549-4_37.

Leitner:2012:VNG

[764] Markus Leitner and Günther R. Raidl. Variable neighborhood and greedy randomized adaptive search for capacitated connected facility location. Lecture Notes in Computer Science, 6927:295–302, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27549-4_38.

Ediger:2012:EEF

[765] Patrick Ediger, Rolf Hoffmann, and Sylvia Grüner. Effectively evolving finite state machines compared to enumeration. Lecture Notes in Computer Science, 6927:303–310, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_39.

Hutterer:2012:HPS

[766] Stephan Hutterer, Michael Affenzeller, and Franz Auinger. Heuristic power scheduling of electric vehicle battery charging based on discrete event simulation. Lecture Notes in Computer Science, 6927:311–318, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_40.

Luque:2012:EAP

[767] Gabriel Luque, Francisco Luna, Enrique Alba, and Sergio Nesmachnow. Exploring the accuracy of a parallel cooperative model for trajectorybased metaheuristics. Lecture Notes in Computer Science, 6927:319–326, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_41.

Vonolfen:2012:CCD

[768] Stefan Vonolfen, Andreas Beham, Michael Affenzeller, Stefan Wagner, and Andreas Mayr. Combination and comparison of different genetic encodings for the vehicle routing problem. Lecture Notes in Computer Science, 6927:327–334, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_42.

Winkler:2012:ASE

[769] Stephan M. Winkler, Michael Affenzeller, Gabriel Kronberger, and Michael Kommenda. Analysis of selected evolutionary algorithms in feature selection and parameter optimization for data based tumor marker modeling. Lecture Notes in Computer Science, 6927:335–342, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_43.

Jacak:2012:NNB

[770] Witold Jacak and Karin Pröll. Neural networks based system for cancer diagnosis support. Lecture Notes in Computer Science, 6927:343–350, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_44.

Ruthmair:2012:MAS

[771] Mario Ruthmair and Günther R. Raidl. A memetic algorithm and a solution archive for the rooted delay-constrained minimum spanning tree problem. Lecture Notes in Computer Science, 6927:351–358, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_45.

Dreiseitl:2012:EDG

[772] Stephan Dreiseitl and Melanie Osl. Effects of data grouping on calibration measures of classifiers. Lecture Notes in Computer Science, 6927:359–366, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27549-4_46.

Neumuller:2012:PMO

[773] Christoph Neumüller, Stefan Wagner, Gabriel Kronberger, and Michael Affenzeller. Parameter meta-optimization of metaheuristic optimization algorithms. Lecture Notes in Computer Science, 6927:367–374, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_47.

Alba:2012:SOG

[774] Enrique Alba and Pablo Vidal. Systolic optimization on GPU platforms. Lecture Notes in Computer Science, 6927:375–383, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27549-4_48.

Ramler:2012:AHA

[775] Rudolf Ramler and Thomas Natschläger. Applying heuristic approaches for predicting defect-prone software components. Lecture Notes in Computer Science, 6927:384–391, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_49.

Pirkwieser:2012:IPR

[776] Sandro Pirkwieser, Günther R. Raidl, and Jens Gottlieb. Improved packing and routing of vehicles with compartments. Lecture Notes in Computer Science, 6927:392–399, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_50.

Kommenda:2012:ASR

[777] Michael Kommenda, Gabriel Kronberger, Christoph Feilmayr, and Leonhard Schickmair. Application of symbolic regression on blast furnace and temper mill datasets. Lecture Notes in Computer Science, 6927:400–407, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_51.

Dorfer:2012:ASO

[778] Viktoria Dorfer, Stephan M. Winkler, Thomas Kern, Gerald Petz, and Patrizia Faschang. Analysis of single-objective and multi-objective evolutionary algorithms in keyword cluster optimization. Lecture Notes in Computer Science, 6927:408–415, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_52.

Holzlwimmer:2012:HSR

[779] Andreas Hölzlwimmer, Hannes Brandstätter-Müller, Bahram Parsapour, and Gerald Lirk. A heuristic scheduling and resource management system for solving bioinformatical problems via high performance computing on heterogeneous multi-platform hardware. Lecture Notes in Computer Science, 6927:416–423, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_53.

Pitzer:2012:CAF

[780] Erik Pitzer, Michael Affenzeller, Andreas Beham, and Stefan Wagner. Comprehensive and automatic fitness landscape analysis using HeuristicLab. Lecture Notes in Computer Science, 6927:424–431, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27549-4_54.

Campos-Rodriguez:2012:PSO

[781] Clara Campos-Rodríguez, José A. Moreno-Pérez, and Dolores R. Santos-Peñate. Particle swarm optimization with two swarms for the discrete (r|p)-centroid problem. Lecture Notes in Computer Science, 6927: 432–439, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27549-4_55.

Brito:2012:AGV

[782] J. Brito, F. J. Martínez, José A. Moreno-Pérez, and J. L. Verdegay. ACO–GRASP–VNS metaheuristic for VRP with fuzzy Windows time constraints. Lecture Notes in Computer Science, 6927:440–447, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_56.

Weyland:2012:UST

[783] Dennis Weyland, Roberto Montemanni, and Luca Maria Gambardella. Using statistical tests for improving state-of-the-art heuristics for the Probabilistic Traveling Salesman Problem with deadlines. Lecture Notes in Computer Science, 6927:448–455, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_57.

Chwatal:2012:STD

[784] Andreas M. Chwatal and Sandro Pirkwieser. Solving the two-dimensional bin-packing problem with variable bin sizes by greedy randomized adaptive search procedures and variable neighborhood search. Lecture Notes in Computer Science, 6927:456–463, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_58.

Kronberger:2012:MBA

[785] Gabriel Kronberger and Michael Affenzeller. Market basket analysis of retail data: Supervised learning approach. Lecture Notes in Computer Science, 6927:464–471, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_59.

Weiss:2012:FRR

[786] Andreas Weiss. A flexible and reliable radar simulator in Matlab OOP for optimizing tracking algorithms. Lecture Notes in Computer Science, 6927:472–476, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27549-4_60.

Onic:2012:FEB

[787] Alexander Onic and Mario Huemer. Frequency estimation beyond Nyquist using sparse approximation methods. Lecture Notes in Computer Science, 6927:477–484, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_61.

Priewasser:2012:RSM

[788] Robert Priewasser, Matteo Agostinelli, and Stefano Marsili. Refinement of simulation models for point-of-load DC–DC converters to enable accurate simulation-based compensator design. Lecture Notes in Computer Science, 6927:485–490, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_62.

Schlechter:2012:OFD

[789] Thomas Schlechter. Optimized filter design for a filter bank based blocker detection concept for LTE systems. Lecture Notes in Computer Science, 6927:491–496, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27549-4_63.

Wojciechowski:2012:CDA

[790] Bartosz Wojciechowski, Maciej Nikodem, and Tomasz Surmacz. Clustering and data aggregation as factors of wireless sensor network lifetime. Lecture Notes in Computer Science, 6927:497–504, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27549-4_64.

Bawiec:2012:SLC

[791] Marek A. Bawiec, Bartosz Wojciechowski, Maciej Nikodem, and Janusz Biernat. Synthesis of logic circuits based on negative differential resistance property. Lecture Notes in Computer Science, 6927:505–512, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_65.

Lunglmayr:2012:SBO

[792] Michael Lunglmayr and Mario Huemer. Simulation based optimization of signal processing for RFID. Lecture Notes in Computer Science, 6927:513–518, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27549-4_66.

Fiedor:2012:UCC

[793] Jan Fiedor, Bohuslav Křena, Zdeněk Letko, and Tomáš Vojnar. A uniform classification of common concurrency errors. Lecture Notes in Computer Science, 6927:519–526, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_67.

Dudka:2012:EUI

[794] Kamil Dudka, Petr Peringer, and Tomáš Vojnar. An easy to use infrastructure for building static analysis tools. Lecture Notes in Computer Science, 6927:527–534, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_68.

Chen:2012:CDA

[795] Xin Chen and Erika Ábrahám. Choice of directions for the approximation of reachable sets for hybrid systems. Lecture Notes in Computer Science, 6927:535–542, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27549-4_69.

Novosad:2012:UBH

[796] Petr Novosad and Milan Češka. Unfoldings of bounded hybrid Petri nets. Lecture Notes in Computer Science, 6927:543–550, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27549-4_70.

Brzozowska:2012:SEM

[797] Agata Brzozowska, Jerzy Greblicki, and Jerzy Kotowski. State encoding and minimization methodology for self-checking sequential machines. Lecture Notes in Computer Science, 6927:551–558, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27549-4_71.

Fiedor:2012:NAM

[798] Jan Fiedor, Marek Gach, and Milan Češka. A novel approach to modechart verification of real-time systems. Lecture Notes in Computer Science, 6927:559–567, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-27549-4_72.

Brzozowska:2012:CCE

[799] Agata Brzozowska, Jerzy Greblicki, and Jerzy Kotowski. Cloud computing in educational applications methods of virtual desktops deployment. Lecture Notes in Computer Science, 6927:568–575, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27549-4_73.

Alvarez:2012:MVB

[800] S. Álvarez, M. A. Sotelo, D. F. Llorca, R. Quintero, and O. Marcos. Monocular vision-based target detection on dynamic transport infrastructures. Lecture Notes in Computer Science, 6927:576–583, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_74.

Fraga:2012:PSO

[801] A. Fraga, N. Barreira, M. Ortega, M. G. Penedo, and M. J. Carreira. Precise segmentation of the optic disc in retinal fundus images. Lecture Notes in Computer Science, 6927:584–591, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_75.

Rodriguez-Vazquez:2012:SCB

[802] Juan José Rodríguez-Vázquez, Sixto Romero-Sánchez, and Miguel Cárdenas-Montes. Speeding up a chaos-based image encryption algorithm using GPGPU. Lecture Notes in Computer Science, 6927:592–599, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_76.

Gavilan:2012:SCR

[803] M. Gavilán, D. Balcones, M. A. Sotelo, D. F. Llorca, O. Marcos, C. Fernández, and I. García. Surface classification for road distress detection system enhancement. Lecture Notes in Computer Science, 6927:600–607, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27549-4_77.

Sanchez:2012:ARA

[804] Javier Sánchez. Analysis of recent advances in optical flow estimation methods. Lecture Notes in Computer Science, 6927:608–615, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27549-4_78.

Fernandez:2012:CSC

[805] A. Fernandez, M. Ortega, B. Cancela, and M. G. Penedo. Contextual and skin color region information for face and arms location. Lecture Notes in Computer Science, 6927:616–623, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27549-4_79.

Eibensteiner:2012:SVA

[806] Florian Eibensteiner, Jürgen Kogler, Christoph Sulzbachner, and Josef Scharinger. Stereo-vision algorithm based on bio-inspired silicon retinas for implementation in hardware. Lecture Notes in Computer Science, 6927:624–631, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27549-4_80.

Anonymous:2012:BMu

[807] Anonymous. Back matter. Lecture Notes in Computer Science, 6927: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27549-4/1.

Anonymous:2012:FMx

[808] Anonymous. Front matter. Lecture Notes in Computer Science, 6927: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27549-4/1.

Atzlesberger:2012:OMF

[809] Johannes Atzlesberger and Bernhard G. Zagar. Optimization of a magnetic flux leakage measurement set-up using FEM-simulations. Lecture Notes in Computer Science, 6928:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27579-1_1/.

Hoflehner:2012:CRP

[810] Markus Hoflehner and Andreas Springer. Comparison of RF power amplifier behavioral models with respect to their modeling capabilities in adjacent and alternate bands. Lecture Notes in Computer Science, 6928: 9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27579-1_2/.

Mostl:2012:NSL

[811] Georg Möstl, Richard Hagelauer, Gerhard Müller, and Andreas Springer. A network and system level approach towards an accurate simulation of WSNs. Lecture Notes in Computer Science, 6928:17–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27579-1_3/.

Heinisch:2012:MRF

[812] Martin Heinisch, Erwin K. Reichel, and Bernhard Jakoby. On the modelling of resonating fluid sensors. Lecture Notes in Computer Science, 6928:25–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27579-1_4/.

Hirsch:2012:ESN

[813] Markus Hirsch and Thomas E. Passenbrunner. Extension of static nonlinear DoE identification algorithms to dynamic systems. Lecture Notes in Computer Science, 6928:33–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27579-1_5/.

Waschl:2012:ATM

[814] Harald Waschl, Daniel Alberer, and Luigi del Re. Automatic tuning methods for MPC environments. Lecture Notes in Computer Science, 6928:41–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27579-1_6/.

Passenbrunner:2012:BID

[815] Thomas E. Passenbrunner and Luigi del Re. On-board implementation of a decentralized algorithm for deployment of a swarm. Lecture Notes in Computer Science, 6928:49–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27579-1_7/.

Ludwig:2012:MSC

[816] P. Ludwig, K. Rieger, and K. Schlacher. Modelling, simulation and control of a heavy chain system. Lecture Notes in Computer Science, 6928: 57–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27579-1_8/.

Daxberger:2012:MCC

[817] H. Daxberger, K. Rieger, and K. Schlacher. On modelling and control of compressible Non–Newtonian injection processes. Lecture Notes in Computer Science, 6928:65–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27579-1_9/.

Poltschak:2012:MBE

[818] Florian Poltschak, Peter Hehenberger, Babak Farrokhzad, Wolfgang Amrhein, and Klaus Zeman. Model-based evaluation of a linear electro hydraulic direct drive. Lecture Notes in Computer Science, 6928:73– 80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27579-1_10/.

Hochwallner:2012:SAS

[819] Martin Hochwallner, Matthias Hörl, Stefan Dierneder, and Rudolf Scheidl. Some aspects of SysML application in the reverse engineering of mechatronic systems. Lecture Notes in Computer Science, 6928: 81–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27579-1_11/.

Horl:2012:ISS

[820] Matthias Hörl, Martin Hochwallner, Stefan Dierneder, and Rudolf Scheidl. Integration of SysML and simulation models for mechatronic systems. Lecture Notes in Computer Science, 6928:89–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27579-1_12/.

Kittel:2012:MOM

[821] Konstantin Kittel, Peter Hehenberger, Sándor Vajna, and Klaus Zeman. Modelling and optimisation of mechatronic systems using the autogenetic design theory. Lecture Notes in Computer Science, 6928:97–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27579-1_13/.

Follmer:2012:MBA

[822] Martin Follmer, Peter Hehenberger, and Klaus Zeman. Model-based approach for the reliability prediction of mechatronic systems on the system-level. Lecture Notes in Computer Science, 6928:105–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27579-1_14/.

Hehenberger:2012:URI

[823] Peter Hehenberger, Alexander Egyed, and Klaus Zeman. Understanding the relationship of information in mechatronic design modeling. Lecture Notes in Computer Science, 6928:113–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27579-1_15/.

Punz:2012:MDP

[824] Stefan Punz, Peter Hehenberger, and Martin Follmer. Modeling and design of a production concept for skinless pretzel-shaped sausages. Lecture Notes in Computer Science, 6928:121–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27579-1_16/.

Hirsch:2012:OSS

[825] Markus Hirsch, Thomas Schwarzgruber, Michael Aschaber, and Herbert Pöllhuber. Optimization of a speedboat simulator for engine calibration. Lecture Notes in Computer Science, 6928:129–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27579-1_17/.

Chaczko:2012:WCU

[826] Zenon Chaczko. WSN clustering using IC–SVD algorithms. Lecture Notes in Computer Science, 6928:137–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27579-1_18/.

Braun:2012:MDI

[827] Robin Braun and Zenon Chaczko. Multi-dimensional information space view of wireless sensor networks with optimization applications. Lecture Notes in Computer Science, 6928:146–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27579-1_19/.

Chaczko:2012:AMT

[828] Zenon Chaczko and Germano Resconi. Application of morphotronic theory to parallel robots. Lecture Notes in Computer Science, 6928:153– 160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27579-1_20/.

Resconi:2012:MBG

[829] Germano Resconi and Zenon Chaczko. Mechatronics and the bond graph theory extended by the morphotronic systems. Lecture Notes in Computer Science, 6928:161–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_21.

Kale:2012:PRV

[830] Anup Kale, Zenon Chaczko, and Imre Rudas. Parallel robot vision using genetic algorithm and object centroid. Lecture Notes in Computer Science, 6928:170–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_22.

Hoefer:2012:TSC

[831] Gerhard Hoefer and Manfred Mauerkirchner. Towards sensomotory coordination of vision and action in humanoid agents. Lecture Notes in Computer Science, 6928:179–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_23.

Klempous:2012:DPG

[832] Ryszard Klempous. The different possibilities for gait identification based on motion capture. Lecture Notes in Computer Science, 6928:187–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27579-1_24.

Smutnicki:2012:BOJ

[833] Czesław Smutnicki. Biomimetic optimizers for job scheduling. Lecture Notes in Computer Science, 6928:195–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27579-1_25.

Handzlik:2012:TRA

[834] Adam Handzlik, Andrzej Jablonski, Ryszard Klempous, and Agnieszka Skotarczyk. Task and resources assignment in special application embedded systems. Lecture Notes in Computer Science, 6928:203–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27579-1_26.

Zwettler:2012:DND

[835] Gerald Zwettler, Robert Pichler, and Werner Backfrieder. Diagnosis of neurodegenerative diseases based on multi-modal hemodynamic classification of the brain. Lecture Notes in Computer Science, 6928:209–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27579-1_27.

Trujillo-Pino:2012:SED

[836] A. Trujillo-Pino, K. Krissian, D. Santana-Cedrés, and J. Esclarín-Monreal. A subpixel edge detector applied to aortic dissection detection. Lecture Notes in Computer Science, 6928:217–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27579-1_28.

Chiu:2012:MDR

[837] Christopher Chiu and Zenon Chaczko. Multi-dimensional representations of laparoscopic simulations for SANETs. Lecture Notes in Computer Science, 6928:225–232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_29.

Rozenblit:2012:MTC

[838] Jerzy W. Rozenblit. Models and techniques for computer aided surgical training. Lecture Notes in Computer Science, 6928:233–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27579-1_30.

Klempous:2012:AST

[839] Ryszard Klempous, Jan Nikodem, and Andrzej Wytyczak-Partyka. Application of simulation techniques in a virtual laparoscopic laboratory. Lecture Notes in Computer Science, 6928:242–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27579-1_31.

Quatember:2012:DAM

[840] Bernhard Quatember, Martin Mayr, Wolfgang Recheis, and Stefanos Demertzis. Development of an accurate method for motion analyses of the heart wall based on medical imagery. Lecture Notes in Computer Science, 6928:248–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27579-1_32.

Szlachcic:2012:MOO

[841] Ewa Szlachcic, Pawel Porombka, and Jerzy Kotowski. Multi-objective optimization of cancer chemotherapy treatment. Lecture Notes in Computer Science, 6928:256–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_33.

Araujo:2012:EPP

[842] Carmen Paz Suárez Araujo, Miguel Ángel Pérez del Pino, and Patricio García Báez. EDEVITALZH: Predictive, preventive, participatory and personalized e-health platform to assist in the geriatrics and neurology clinical scopes. Lecture Notes in Computer Science, 6928:264–271, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27579-1_34.

Steinbach:2012:ICE

[843] Bernd Steinbach and Christian Posthoff. Improvements of the construction of exact minimal covers of Boolean functions. Lecture Notes in Computer Science, 6928:272–279, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_35.

Moraga:2012:PAU

[844] Claudio Moraga. Pattern analysis under number theoretic transforms. Lecture Notes in Computer Science, 6928:280–287, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27579-1_36.

Stankovic:2012:REC

[845] Radomir S. Stanković, Jaakko T. Astola, Claudio Moraga, and Stanislav Stanković. Remarks on efficient computation of the inverse Fourier transforms on finite non-Abelian groups. Lecture Notes in Computer Science, 6928:288–295, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27579-1_37.

Stankovic:2012:RCS

[846] Stanislav Stanković, Radomir S. Stanković, Jaakko T. Astola, and Claudio Moraga. Representation of convolution systems on finite groups by heterogeneous decision diagrams. Lecture Notes in Computer Science, 6928:296–303, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27579-1_38.

Atreas:2012:DTP

[847] Nikolaos Atreas and Costas Karanikas. Discrete transforms produced from two natural numbers and applications. Lecture Notes in Computer Science, 6928:304–310, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_39.

Stankovic:2012:RSW

[848] Milena Stanković and Suzana Stojković. Reversible synthesis in the Walsh Hadamard domain. Lecture Notes in Computer Science, 6928: 311–318, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27579-1_40.

Astola:2012:PAE

[849] Helena Astola, Stanislav Stanković, and Jaakko T. Astola. Performance analysis of error-correcting binary decision diagrams. Lecture Notes in Computer Science, 6928:319–326, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_41.

Borowik:2012:TED

[850] Grzegorz Borowik and Andrzej Kraśniewski. Trading-off error detection efficiency with implementation cost for sequential circuits implemented with FPGAs. Lecture Notes in Computer Science, 6928:327–334, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27579-1_42.

Borowik:2012:MGI

[851] Grzegorz Borowik and Andrzej Paszkiewicz. Method of generating irreducible polynomials over GF(3) on the basis of trinomials. Lecture Notes in Computer Science, 6928:335–342, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_43.

Borowik:2012:MCI

[852] Grzegorz Borowik, Tadeusz Łuba, and Paweł Tomaszewicz. On memory capacity to implement logic functions. Lecture Notes in Computer Science, 6928:343–350, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_44.

Gat:2012:RFW

[853] György Gát. Reconstruction of functions via Walsh–Fourier cofficients. Lecture Notes in Computer Science, 6928:351–358, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27579-1_45.

Sanchez:2012:RTV

[854] Adrián Peñate Sánchez, Alexis Quesada-Arencibia, and Carlos M. Travieso González. Real time vehicle recognition: a novel method for road detection. Lecture Notes in Computer Science, 6928:359–364, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27579-1_46.

Molina-Gil:2012:PGS

[855] J. Molina-Gil, P. Caballero-Gil, A. Fúster-Sabater, and C. Caballero-Gil. Pseudorandom generator to strengthen cooperation in VANETs. Lecture Notes in Computer Science, 6928:365–373, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27579-1_47.

Milanes:2012:APA

[856] Vicente Milanés, Enrique Onieva, Joshué Pérez, Jorge Villagrá, and Jorge Godoy. AUTOPIA program advances: How to automate the traffic? Lecture Notes in Computer Science, 6928:374–381, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27579-1_48.

Onieva:2012:STF

[857] Enrique Onieva, Vicente Milanés, Joshué Pérez, Javier Alonso, and Teresa de Pedro. Study of traffic flow controlled with independent agentbased traffic signals. Lecture Notes in Computer Science, 6928:382–389, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27579-1_49.

Villagra:2012:RLM

[858] Jorge Villagrá, Vicente Milanés, Joshué Pérez, Jorge Godoy, and Enrique Onieva. A reinforcement learning modular control architecture for fully automated vehicles. Lecture Notes in Computer Science, 6928:390–397, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27579-1_50.

Alonso:2012:TLI

[859] Javier Alonso, Jorge Godoy, Roberto Sanz, Enrique Onieva, and Vicente Milanés. Traffic light intelligent regulation using infrastructure located sensors. Lecture Notes in Computer Science, 6928:398–403, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27579-1_51.

Perez:2012:DDV

[860] Joshué Pérez, Jorge Villagrá, Enrique Onieva, Vicente Milanés, and Teresa de Pedro. Driving by driverless vehicles in urban environment. Lecture Notes in Computer Science, 6928:404–411, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27579-1_52.

Llamazares:2012:MBU

[861] Á. Llamazares, E. J. Molinos, M. Ocaña, L. M. Bergasa, N. Hernández, and F. Herranz. 3D map building using a 2D laser scanner. Lecture Notes in Computer Science, 6928:412–419, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_53.

Herranz:2012:MBN

[862] F. Herranz, M. Ocaña, L. M. Bergasa, N. Hernández, A. Llamazares, and C. Fernández. Mapping based on a noisy range-only sensor. Lecture Notes in Computer Science, 6928:420–425, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_54.

Musleh:2012:UVD

[863] Basam Musleh, Arturo de la Escalera, and José María Armingol. U– V disparity analysis in urban environments. Lecture Notes in Computer Science, 6928:426–432, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_55.

Isern-Gonzalez:2012:AOA

[864] José Isern-González, Daniel Hernández-Sosa, and Enrique Fernández-Perdomo. Application of optimization algorithms to trajectory planning for underwater gliders. Lecture Notes in Computer Science, 6928: 433–440, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27579-1_56.

Sanchez-Medina:2012:ESV

[865] Javier J. Sanchez-Medina, Moises Diaz-Cabrera, and Manuel J. Galan-Moreno. Electric scaled vehicle as ITS experimentation platform. Lecture Notes in Computer Science, 6928:441–448, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27579-1_57.

Garcia-Rodriguez:2012:CCM

[866] Carmelo R. García-Rodríguez, Ricardo Pérez-García, and Gabino Padrón-Morales. A complete conceptual model for pervasive information services for public transport. Lecture Notes in Computer Science, 6928:449–456, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27579-1_58.

Sebek:2012:CTH

[867] Michael Šebek and Zdeněk Hurák. Constant time headway control policy in leader following vehicular platoons: 2-D polynomial approach. Lecture Notes in Computer Science, 6928:457–464, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27579-1_59.

Sanchez-Medina:2012:UES

[868] Javier J. Sanchez-Medina, Moises Diaz-Cabrera, and Manuel J. Galan-Moreno. User equilibrium study of AETROS travel route optimization system. Lecture Notes in Computer Science, 6928:465–472, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27579-1_60.

Diaz-Cabrera:2012:DPS

[869] Moises Diaz-Cabrera, Javier J. Sanchez-Medina, Idaira Perez-Armas, and Elisa Medina-Machin. Driver pattern study of Las Palmas de Gran Canaria. Lecture Notes in Computer Science, 6928:473–480, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27579-1_61.

Villeta:2012:RFE

[870] Miguel Villeta, José G. Zato Recellado, José E. Naranjo, and Lourdes Cecilia. Railway field element simulation tool. Lecture Notes in Computer Science, 6928:481–485, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_62.

Padron-Morales:2012:EPF

[871] Gabino Padrón-Morales, Carmelo R. García-Rodríguez, and Francisco Alayón-Hernández. Evaluation of parameters of functions of cost optimization system for operating systems for carriers of passengers by road. Lecture Notes in Computer Science, 6928:486–492, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27579-1_63.

Caballero-Gil:2012:LAR

[872] C. Caballero-Gil, P. Caballero-Gil, A. Peinado-Domínguez, and J. Molina-Gil. Lightweight authentication for RFID used in VANETs. Lecture Notes in Computer Science, 6928:493–500, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27579-1_64.

Schmitzberger:2012:AWS

[873] Heinrich Schmitzberger. Autonomous WLAN sensors for ad hoc indoor localization. Lecture Notes in Computer Science, 6928:501–509, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27579-1_65.

Ribeiro:2012:CSS

[874] Stephan da Costa Ribeiro, Martin Kleinsteuber, Andreas Möller, and Matthias Kranz. A compressive sensing scheme of frequency sparse signals for mobile and wearable platforms. Lecture Notes in Computer Science, 6928:510–518, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_66.

Tschernuth:2012:EDU

[875] Michael Tschernuth, Michael Lettner, and Rene Mayrhofer. Evaluation of descriptive user interface methodologies for mobile devices. Lecture Notes in Computer Science, 6928:519–526, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27579-1_67.

Mayrhofer:2012:FRT

[876] Rene Mayrhofer, Clemens Holzmann, and Romana Koprivec. Friends radar: Towards a private P2P location sharing platform. Lecture Notes in Computer Science, 6928:527–535, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_68.

Erifiu:2012:HSA

[877] Alexander Erifiu and Gerald Ostermayer. Hardware sensor aspects in mobile augmented reality. Lecture Notes in Computer Science, 6928: 536–543, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27579-1_69.

Lettner:2012:MPA

[878] Michael Lettner, Michael Tschernuth, and Rene Mayrhofer. Mobile platform architecture review: Android, iPhone, Qt. Lecture Notes in Computer Science, 6928:544–551, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_70.

Kastner:2012:RFG

[879] Karl-Heinz Kastner and Gerald Ostermayer. Routing with free geodata on mobile devices. Lecture Notes in Computer Science, 6928:552–559, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27579-1_71.

Lettner:2012:UEF

[880] Florian Lettner and Clemens Holzmann. Usability evaluation framework. Lecture Notes in Computer Science, 6928:560–567, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27579-1_72.

Grillo:2012:UMP

[881] Pierluigi Grillo, Silvia Likavec, and Ilaria Lombardi. Using mobile phone cameras to interact with ontological data. Lecture Notes in Computer Science, 6928:568–576, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_73.

Greblicki:2012:CSP

[882] Jerzy Greblicki, Jerzy Kotowski, and Ewa Szlachcic. The cutting stock problem recent application in information technology. Lecture Notes in Computer Science, 6928:577–584, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_74.

Huba:2012:RII

[883] Mikuláš Huba and Peter Ťapák. Relay identification of IPDT plant by analyzing nonsymmetrical oscillations. Lecture Notes in Computer Science, 6928:585–592, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-27579-1_75.

Sieniutycz:2012:MSP

[884] Stanislaw Sieniutycz. Modeling and simulation of power yield in thermal, chemical and electrochemical systems: Fuel cell case. Lecture Notes in Computer Science, 6928:593–600, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_76.

Herrmann:2012:PCC

[885] Stefan Herrmann, Andreas Attenberger, and Klaus Buchenrieder. Prostheses control with combined near-infrared and myoelectric signals. Lecture Notes in Computer Science, 6928:601–608, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27579-1_77.

Quiroga:2012:HFG

[886] Lisandro Quiroga and Eckehard Schnieder. Heuristic forecasting of geometry deterioration of high speed railway tracks. Lecture Notes in Computer Science, 6928:609–616, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_78.

Halas:2012:ANC

[887] Miroslav Halás and Vladimír Žilka. Application of noninteracting control problem to coupled tanks. Lecture Notes in Computer Science, 6928:617–624, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27579-1_79.

Cabrera:2012:DLC

[888] Francisco Cabrera, Víctor Araña, and Carlos Barrera. Description of a low-cost radio-frequency system to detect hydrocarbons. Lecture Notes in Computer Science, 6928:625–632, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27579-1_80.

Anonymous:2012:BMv

[889] Anonymous. Back matter. Lecture Notes in Computer Science, 6928: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27579-1/1.

Anonymous:2012:FMy

[890] Anonymous. Front matter. Lecture Notes in Computer Science, 6928: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27579-1/1.

Klint:2012:TOS

[891] Paul Klint, Bert Lisser, and Atze van der Ploeg. Towards a one-stop-shop for analysis, transformation and visualization of software. Lecture Notes in Computer Science, 6940:1–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28830-2_1/.

Seibel:2012:DLC

[892] Andreas Seibel, Regina Hebig, Stefan Neumann, and Holger Giese. A dedicated language for context composition and execution of true blackbox model transformations. Lecture Notes in Computer Science, 6940: 19–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28830-2_2/.

deJonge:2012:ALP

[893] Maartje de Jonge and Eelco Visser. An algorithm for layout preservation in refactoring transformations. Lecture Notes in Computer Science, 6940:40–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28830-2_3/.

Tairas:2012:CDE

[894] Robert Tairas and Jordi Cabot. Cloning in DSLs: Experiments with OCL. Lecture Notes in Computer Science, 6940:60–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28830-2_4/.

Joncheere:2012:UMW

[895] Niels Joncheere and Ragnhild Van Der Straeten. Uniform modularization of workflow concerns using unify. Lecture Notes in Computer Science, 6940:77–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28830-2_5/.

Sutton:2012:DCL

[896] Andrew Sutton and Bjarne Stroustrup. Design of concept libraries for C++. Lecture Notes in Computer Science, 6940:97–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28830-2_6/.

Nishimori:2012:JTB

[897] Taketoshi Nishimori and Yasushi Kuno. Join token-based event handling: a comprehensive framework for game programming. Lecture Notes in Computer Science, 6940:119–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28830-2_7/.

Tombelle:2012:RPS

[898] Christophe Tombelle, Gilles Vanwormhoudt, and Emmanuel Renaux. Reusing pattern solutions in modeling: a generic approach based on a role language. Lecture Notes in Computer Science, 6940:139–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28830-2_8/.

Paulheim:2012:AIE

[899] Heiko Paulheim, Daniel Oberle, and Roland Plendl. An architecture for information exchange based on reference models. Lecture Notes in Computer Science, 6940:160–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28830-2_9/.

Kling:2012:MDQ

[900] Wolfgang Kling, Frédéric Jouault, and Dennis Wagelaar. MoScript: a DSL for querying and manipulating model repositories. Lecture Notes in Computer Science, 6940:180–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28830-2_10/.

Vermolen:2012:RCM

[901] Sander D. Vermolen, Guido Wachsmuth, and Eelco Visser. Reconstructing complex metamodel evolution. Lecture Notes in Computer Science, 6940:201–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28830-2_11/.

Czarnecki:2012:DVM

[902] Krzysztof Czarnecki. Designing variability modeling languages. Lecture Notes in Computer Science, 6940:222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-28830-2_12.

Stappers:2012:FDS

[903] Frank P. M. Stappers, Sven Weber, and Michel A. Reniers. Formalizing a domain specific language using SOS: An industrial case study. Lecture Notes in Computer Science, 6940:223–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28830-2_13/.

Erwig:2012:SF

[904] Martin Erwig and Eric Walkingshaw. Semantics first! Lecture Notes in Computer Science, 6940:243–262, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28830-2_14/.

Kaminski:2012:IAG

[905] Ted Kaminski and Eric Van Wyk. Integrating attribute grammar and functional programming language features. Lecture Notes in Computer Science, 6940:263–282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28830-2_15/.

Basten:2012:PFD

[906] Hendrikus J. S. Basten and Jurgen J. Vinju. Parse forest diagnostics with Dr. Ambiguity. Lecture Notes in Computer Science, 6940:283– 302, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28830-2_16/.

Basten:2012:ADS

[907] Hendrikus J. S. Basten, Paul Klint, and Jurgen J. Vinju. Ambiguity detection: Scaling to scannerless. Lecture Notes in Computer Science, 6940:303–323, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28830-2_17/.

Fischer:2012:CCF

[908] Bernd Fischer, Ralf Lämmel, and Vadim Zaytsev. Comparison of contextfree grammars based on parsing generated test data. Lecture Notes in Computer Science, 6940:324–343, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28830-2_18/.

Hills:2012:RLR

[909] Mark Hills, Paul Klint, and Jurgen J. Vinju. RLSRunner: Linking Rascal with K for program analysis. Lecture Notes in Computer Science, 6940:344–353, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28830-2_19/.

Nilsson:2012:MOO

[910] Anders Nilsson and Görel Hedin. Metacompiling OWL ontologies. Lecture Notes in Computer Science, 6940:354–366, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28830-2_20/.

Wider:2012:TCB

[911] Arif Wider. Towards combinators for bidirectional model transformations in Scala. Lecture Notes in Computer Science, 6940:367–377, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
28830-2_21.

Wojciechowski:2012:TFC

[912] Paweł T. Wojciechowski. Typed first-class communication channels and mobility for concurrent scripting languages. Lecture Notes in Computer Science, 6940:378–387, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-28830-2_22.

Anonymous:2012:BMw

[913] Anonymous. Back matter. Lecture Notes in Computer Science, 6940: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28830-2/1.

Anonymous:2012:FMz

[914] Anonymous. Front matter. Lecture Notes in Computer Science, 6940: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28830-2/1.

vonOheimb:2012:AFS

[915] David von Oheimb and Sebastian Mödersheim. ASLan++ — a formal security specification language for distributed systems. Lecture Notes in Computer Science, 6957:1–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25271-6_1/.

Chevalier:2012:OUS

[916] Yannick Chevalier, Mohamed Anis Mekki, and Michaël Rusinowitch. Orchestration under security constraints. Lecture Notes in Computer Science, 6957:23–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25271-6_2/.

Groza:2012:CPS

[917] Bogdan Groza and Marius Minea. Customizing protocol specifications for detecting resource exhaustion and guessing attacks. Lecture Notes in Computer Science, 6957:45–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25271-6_3/.

Beckert:2012:IUS

[918] Bernhard Beckert, Thorsten Bormer, and Vladimir Klebanov. Improving the usability of specification languages and methods for annotationbased verification. Lecture Notes in Computer Science, 6957:61–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25271-6_4/.

Bubel:2012:PSS

[919] Richard Bubel, Reiner Hähnle, and Ran Ji. Program specialization via a software verification tool. Lecture Notes in Computer Science, 6957:80– 101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25271-6_5/.

Petre:2012:MBA

[920] Luigia Petre, Kaisa Sere, and Leonidas Tsiopoulos. Model–based analysis tools for component synthesis. Lecture Notes in Computer Science, 6957:102–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25271-6_6/.

Silva:2012:SEC

[921] Renato Silva and Michael Butler. Shared event composition/ decomposition in Event-B. Lecture Notes in Computer Science, 6957:

122–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-25271-6_7/.

Johnsen:2012:ACL

[922] Einar Broch Johnsen, Reiner Hähnle, Jan Schäfer, Rudolf Schlatte, and Martin Steffen. ABS: a core language for abstract behavioral specification. Lecture Notes in Computer Science, 6957:142–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25271-6_8/.

Lienhardt:2012:CMA

[923] Michaël Lienhardt, Ivan Lanese, Mario Bravetti, Davide Sangiorgi, and Gianluigi Zavattaro. A component model for the ABS language. Lecture Notes in Computer Science, 6957:165–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-25271-6_9/.

Schaefer:2012:CAV

[924] Ina Schaefer, Dilian Gurov, and Siavash Soleimanifard. Compositional algorithmic verification of software product lines. Lecture Notes in Computer Science, 6957:184–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25271-6_10/.

Clarke:2012:VMA

[925] Dave Clarke, Radu Muschevici, José Proença, Ina Schaefer, and Rudolf Schlatte. Variability modelling in the ABS language. Lecture Notes in Computer Science, 6957:204–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25271-6_11/.

Hansen:2012:AVE

[926] Helle Hvid Hansen, Jeroen Ketema, Bas Luttik, MohammadReza Mousavi, and Jaco van de Pol. Automated verification of executable UML models. Lecture Notes in Computer Science, 6957:225–250, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25271-6_12/.

Snook:2012:VUM

[927] Colin Snook, Vitaly Savicks, and Michael Butler. Verification of UML models by translation to UML–B. Lecture Notes in Computer Science, 6957:251–266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25271-6_13/.

Baresi:2012:TUB

[928] Luciano Baresi, Angelo Morzenti, Alfredo Motta, and Matteo Rossi. Towards the UML-based formal verification of timed systems. Lecture Notes in Computer Science, 6957:267–286, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25271-6_14/.

Svenningsson:2012:GFM

[929] Rickard Svenningsson, Henrik Eriksson, Jonny Vinter, and Martin Törngren. Generic fault modelling for fault injection. Lecture Notes in Computer Science, 6957:287–296, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25271-6_15/.

Donaldson:2012:TTC

[930] Alastair F. Donaldson, Nannan He, Daniel Kroening, and Philipp Rümmer. Tightening test coverage metrics: a case study in equivalence checking using k-induction. Lecture Notes in Computer Science, 6957:297–315, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25271-6_16/.

Agut:2012:HCI

[931] Damian Nadales Agut, Bert van Beek, Harsh Beohar, Pieter Cuijpers, and Jasper Fonteijn. The hierarchical compositional interchange format. Lecture Notes in Computer Science, 6957:316–335, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25271-6_17/.

David:2012:AMC

[932] Alexandre David, Jacob Deleuran Grunnet, Jan Jakob Jessen, and Kim Guldstrand Larsen. Application of model-checking technology to controller synthesis. Lecture Notes in Computer Science, 6957:336–351, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25271-6_18/.

David:2012:TRT

[933] Alexandre David, Kim Guldstrand Larsen, Shuhao Li, Marius Mikucionis, and Brian Nielsen. Testing real-time systems under uncertainty. Lecture Notes in Computer Science, 6957:352–371, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-25271-6_19/.

Hartmanns:2012:MCS

[934] Arnd Hartmanns. Model-checking and simulation for stochastic timed systems. Lecture Notes in Computer Science, 6957:372–391, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25271-6_20/.

Anonymous:2012:BMx

[935] Anonymous. Back matter. Lecture Notes in Computer Science, 6957: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-25271-6/1.

Anonymous:2012:FMba

[936] Anonymous. Front matter. Lecture Notes in Computer Science, 6957: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-25271-6/1.

Backes:2012:UIT

[937] Michael Backes, Cătălin Hriţcu, and Matteo Maffei. Union and intersection types for secure protocol implementations. Lecture Notes in Computer Science, 6993:1–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27375-9_1/.

Cortier:2012:SCP

[938] Véronique Cortier. Secure composition of protocols. Lecture Notes in Computer Science, 6993:29–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27375-9_2/.

Maurer:2012:CCN

[939] Ueli Maurer. Constructive cryptography — a new paradigm for security definitions and proofs. Lecture Notes in Computer Science, 6993:33– 56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27375-9_3/.

Backes:2012:GCP

[940] Michael Backes, Matteo Maffei, and Kim Pecina. G2C: Cryptographic protocols from goal-driven specifications. Lecture Notes in Computer Science, 6993:57–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27375-9_4/.

MBarka:2012:MLT

[941] Moez Ben MBarka, Francine Krief, and Olivier Ly. Modeling long-term signature validation for resolution of dispute. Lecture Notes in Computer Science, 6993:78–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27375-9_5/.

Dahl:2012:FAP

[942] Morten Dahl, Stéphanie Delaune, and Graham Steel. Formal analysis of privacy for anonymous location based services. Lecture Notes in Computer Science, 6993:98–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27375-9_6/.

deRuiter:2012:FAE

[943] Joeri de Ruiter and Erik Poll. Formal analysis of the EMV protocol suite. Lecture Notes in Computer Science, 6993:113–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27375-9_7/.

Guttman:2012:SGP

[944] Joshua D. Guttman. Security goals and protocol transformations. Lecture Notes in Computer Science, 6993:130–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27375-9_8/.

Huisman:2012:MCS

[945] Marieke Huisman and Henri-Charles Blondeel. Model-checking secure information flow for multi-threaded programs. Lecture Notes in Computer Science, 6993:148–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27375-9_9/.

Jacquemard:2012:MCR

[946] Florent Jacquemard, Étienne Lozes, and Ralf Treinen. Multiple congruence relations, first-order theories on terms, and the frames of the applied pi-calculus. Lecture Notes in Computer Science, 6993:166–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27375-9_10/.

Luo:2012:ACI

[947] Zhengqin Luo, Tamara Rezk, and Manuel Serrano. Automated code injection prevention for Web applications. Lecture Notes in Computer Science, 6993:186–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27375-9_11/.

Malladi:2012:SRC

[948] Sreekanth Malladi. Soundness of removing cancellation identities in protocol analysis under exclusive-OR. Lecture Notes in Computer Science, 6993:205–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27375-9_12/.

Anonymous:2012:BMy

[949] Anonymous. Back matter. Lecture Notes in Computer Science, 6993: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27375-9/1.

Anonymous:2012:FMbb

[950] Anonymous. Front matter. Lecture Notes in Computer Science, 6993: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27375-9/1.

Harman:2012:SBS

[951] Mark Harman and Phil McMinn. Search based software engineering: Techniques, taxonomy, tutorial. Lecture Notes in Computer Science, 7007:1–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25231-0_1/.

Juristo:2012:RSE

[952] Natalia Juristo and Omar S. Gómez. Replication of software engineering experiments. Lecture Notes in Computer Science, 7007:60–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25231-0_2/.

Morandi:2012:FRS

[953] Benjamin Morandi and Sebastian Nanz. A formal reference for SCOOP. Lecture Notes in Computer Science, 7007:89–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25231-0_3/.

Braione:2012:IST

[954] Pietro Braione, Giovanni Denaro, and Mauro Pezzè. On the integration of software testing and formal analysis. Lecture Notes in Computer Science, 7007:158–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25231-0_4/.

Wei:2012:BCG

[955] Yi Wei, Bertrand Meyer, and Manuel Oriol. Is branch coverage a good measure of testing effectiveness? Lecture Notes in Computer Science, 7007:194–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25231-0_5/.

Anonymous:2012:BMz

[956] Anonymous. Back matter. Lecture Notes in Computer Science, 7007: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-25231-0/1.

Anonymous:2012:FMbc

[957] Anonymous. Front matter. Lecture Notes in Computer Science, 7007: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-25231-0/1.

Anonymous:2012:FMbd

[958] Anonymous. Front matter. Lecture Notes in Computer Science, 7010: 1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29326-9/1/1.

Parikh:2012:WSS

[959] Rohit Parikh and Marc Pauly. What is social software? Lecture Notes in Computer Science, 7010:3–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29326-9_1/.

Dowding:2012:FRS

[960] Keith Dowding and Martin van Hees. Freedom, rights and social software. Lecture Notes in Computer Science, 7010:14–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29326-9_2/.

Dimitri:2012:TDT

[961] Nicola Dimitri and Jan van Eijck. Time discounting and time consistency. Lecture Notes in Computer Science, 7010:29–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29326-9_3/.

Anonymous:2012:FMbe

[962] Anonymous. Front matter. Lecture Notes in Computer Science, 7010: 39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29326-9/2/1.

vanDitmarsch:2012:LL

[963] Hans van Ditmarsch, Jan van Eijck, Floor Sietsma, and Yanjing Wang. On the logic of lying. Lecture Notes in Computer Science, 7010:41– 72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29326-9_4/.

Segerberg:2012:SBR

[964] Krister Segerberg. Strategies for belief revision. Lecture Notes in Computer Science, 7010:73–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29326-9_5/.

vanBenthem:2012:PS

[965] Johan van Benthem. In praise of strategies. Lecture Notes in Computer Science, 7010:96–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29326-9_6/.

Anonymous:2012:FMbf

[966] Anonymous. Front matter. Lecture Notes in Computer Science, 7010: 117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29326-9/3/1.

vanEijck:2012:PCU

[967] Jan van Eijck. Perception and change in update logic. Lecture Notes in Computer Science, 7010:119–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29326-9_7/.

Dunin-Keplicz:2012:AAE

[968] Barbara Dunin-Keplicz and Andrzej Szałas. Agents in approximate environments. Lecture Notes in Computer Science, 7010:141–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29326-9_8/.

Gardenfors:2012:CCD

[969] Peter Gärdenfors. The cognitive and communicative demands of cooperation. Lecture Notes in Computer Science, 7010:164–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29326-9_9/.

Dunin-Keplicz:2012:LVT

[970] Barbara Dunin-Keplicz and Rineke Verbrugge. A logical view on teamwork. Lecture Notes in Computer Science, 7010:184–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29326-9_10/.

Anonymous:2012:BMba

[971] Anonymous. Back matter. Lecture Notes in Computer Science, 7010: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29326-9/1.

Anonymous:2012:FMbg

[972] Anonymous. Front matter. Lecture Notes in Computer Science, 7010: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29326-9/1.

Roth:2012:ICE

[973] Holger R. Roth, Thomas E. Hampshire, Jamie R. McClelland, Mingxing Hu, and Darren J. Boone. Inverse consistency error in the registration of prone and supine images in CT colonography. Lecture Notes in Computer Science, 7029:1–7, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28557-8_1/.

Cai:2012:DEE

[974] Wenli Cai, June-Goo Lee, Se Hyung Kim, and Hiroyuki Yoshida. Dualenergy electronic cleansing for artifact-free visualization of the colon in fecal-tagging CT colonography. Lecture Notes in Computer Science, 7029:8–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28557-8_2/.

Panjwani:2012:CAP

[975] Neil Panjwani, Marius George Linguraru, Joel G. Fletcher, and Ronald M. Summers. Computer-aided polyp detection for laxativefree CT colonography. Lecture Notes in Computer Science, 7029:18– 26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28557-8_3/.

Nappi:2012:CPR

[976] Janne J. Näppi, Daniele Regge, and Hiroyuki Yoshida. Comparative performance of random forest and support vector machine classifiers for detection of colorectal lesions in CT colonography. Lecture Notes in Computer Science, 7029:27–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28557-8_4/.

Vu:2012:IEC

[977] Hai Vu, Tomio Echigo, Keiko Yagi, Hirotoshi Okazaki, Yasuhiro Fujiwara, and Yasushi Yagi. Image-enhanced capsule endoscopy preserving the original color tones. Lecture Notes in Computer Science, 7029:35– 43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28557-8_5/.

Li:2012:NRM

[978] Zhang Li, Matthan W. A. Caan, Manon L. Ziech, Japp Stoker, and Lucas J. van Vliet. 3D non-rigid motion correction of free-breathing abdominal DCE–MRI data. Lecture Notes in Computer Science, 7029: 44–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28557-8_6/.

Fang:2012:SLT

[979] Ruogu Fang, Ramin Zabih, Ashish Raj, and Tsuhan Chen. Segmentation of liver tumor using efficient global optimal tree metrics graph cuts. Lecture Notes in Computer Science, 7029:51–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28557-8_7/.

Nappi:2012:EDC

[980] Janne J. Näppi, Daniele Regge, and Hiroyuki Yoshida. Ensemble detection of colorectal lesions for CT colonography. Lecture Notes in Computer Science, 7029:60–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28557-8_8/.

Nappi:2012:ADC

[981] Janne J. Näppi, Stefan Gryspeerdt, Philippe Lefere, Michael Zalis, and Hiroyuki Yoshida. Automated detection of colorectal lesions in noncathartic CT colonography. Lecture Notes in Computer Science, 7029: 68–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28557-8_9/.

Bernal:2012:IVO

[982] Jorge Bernal, Javier Sánchez, and Fernando Vilariño. Integration of valley orientation distribution for polyp region identification in colonoscopy. Lecture Notes in Computer Science, 7029:76–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28557-8_10/.

Nappi:2012:ADD

[983] Janne J. Näppi, Dushyant V. Sahani, Joel G. Fletcher, and Hiroyuki Yoshida. Automated detection and diagnosis of Crohn’s Disease in CT enterography. Lecture Notes in Computer Science, 7029:84–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28557-8_11/.

Jiang:2012:QEL

[984] Hongtao Jiang, Wei Qi, Qimei Liao, Haitao Zhao, Wei Lei, Li Guo, and Hongbing Lu. Quantitative evaluation of lumbar disc herniation based on MRI image. Lecture Notes in Computer Science, 7029:91– 98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28557-8_12/.

Mandava:2012:LTS

[985] Rajeswari Mandava, Lee Song Yeow, Bhavik Anil Chandra, and Ong Kok Haur. Liver tumor segmentation using kernel-based FGCM and PGCM. Lecture Notes in Computer Science, 7029:99–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28557-8_13/.

Erdt:2012:IDI

[986] Marius Erdt, Cristina Oyarzun Laura, Klaus Drechsler, Stefano De Beni, and Luigi Solbiati. Improving diagnosis and intervention: a complete approach for registration of liver CT data. Lecture Notes in Computer Science, 7029:108–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28557-8_14/.

Freiman:2012:AIN

[987] Moti Freiman, Stephan D. Voss, and Simon Keith Warfield. Abdominal images non-rigid registration using local-affine diffeomorphic demons. Lecture Notes in Computer Science, 7029:116–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28557-8_15/.

Steger:2012:FFI

[988] Sebastian Steger and Georgios Sakas. FIST: Fast interactive segmentation of tumors. Lecture Notes in Computer Science, 7029:125–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28557-8_16/.

Laura:2012:IRL

[989] Cristina Oyarzun Laura, Klaus Drechsler, Marius Erdt, Matthias Keil, and Matthias Noll. Intraoperative registration for liver tumor ablation. Lecture Notes in Computer Science, 7029:133–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28557-8_17/.

Pace:2012:SGD

[990] Danielle F. Pace, Marc Niethammer, and Stephen R. Aylward. Sliding geometries in deformable image registration. Lecture Notes in Computer Science, 7029:141–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28557-8_18/.

Drechsler:2012:SPV

[991] Klaus Drechsler and Cristina Oyarzun Laura. Simulation of portal vein clamping and the impact of safety margins for liver resection planning. Lecture Notes in Computer Science, 7029:149–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28557-8_19/.

Chen:2012:LSC

[992] Yufei Chen, Weidong Zhao, Qidi Wu, Zhicheng Wang, and Jinyong Hu. Liver segmentation in CT images for intervention using a graph-cut based model. Lecture Notes in Computer Science, 7029:157–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28557-8_20/.

Zhao:2012:AFG

[993] Xin Zhao and Arie Kaufman. Anatomical feature-guided mutual information registration of multimodal prostate MRI. Lecture Notes in Computer Science, 7029:165–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-28557-8_21.

Okada:2012:AMO

[994] Toshiyuki Okada, Marius George Linguraru, Yasuhide Yoshida, and Masatoshi Hori. Abdominal multi-organ segmentation of CT images based on hierarchical spatial modeling of organ interrelations. Lecture Notes in Computer Science, 7029:173–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-28557-8_22.

Oda:2012:OSA

[995] Masahiro Oda, Teruhisa Nakaoka, Takayuki Kitasaka, Kazuhiro Furukawa, and Kazunari Misawa. Organ segmentation from 3D abdominal CT images based on Atlas selection and graph cut. Lecture Notes in Computer Science, 7029:181–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-28557-8_23.

Pamulapati:2012:LSA

[996] Vivek Pamulapati, Aradhana Venkatesan, Bradford J. Wood, and Marius George Linguraru. Liver segmental anatomy and analysis from vessel and tumor segmentation via optimized graph cuts. Lecture Notes in Computer Science, 7029:189–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-28557-8_24.

Linguraru:2012:LTS

[997] Marius George Linguraru, William J. Richbourg, Jeremy M. Watt, and Vivek Pamulapati. Liver and tumor segmentation and analysis from CT of diseased patients via a generic affine invariant shape parameterization and graph cuts. Lecture Notes in Computer Science, 7029:198–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
28557-8_25.

Preiswerk:2012:BFE

[998] Frank Preiswerk, Patrik Arnold, Beat Fasel, and Philippe C. Cattin. A Bayesian framework for estimating respiratory liver motion from sparse measurements. Lecture Notes in Computer Science, 7029:207–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
28557-8_26.

Koek:2012:SAS

[999] Marcel Koek, Frederico Bastos Goncalves, Don Poldermans, and Wiro Niessen. Semi–automated subcutaneous and visceral adipose tissue quantification in computed tomography. Lecture Notes in Computer Science, 7029:215–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-28557-8_27.

Vera:2012:CEM

[1000] Sergio Vera, Debora Gil, Agnès Borràs, Xavi Sánchez, and Frederic Pérez. Computation and evaluation of medial surfaces for shape representation of abdominal organs. Lecture Notes in Computer Science, 7029:223–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-28557-8_28.

Kiriyanthan:2012:DPR

[1001] Silja Kiriyanthan, Ketut Fundana, and Philippe C. Cattin. Discontinuity preserving registration of abdominal MR images with apparent sliding organ motion. Lecture Notes in Computer Science, 7029:231–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
28557-8_29.

Berendsen:2012:SCI

[1002] Floris F. Berendsen, Uulke A. van der Heide, Thomas R. Langerak, and Alexis N. T. J. Kotte. Segmentation of cervical images by intersubject registration with a statistical organ model. Lecture Notes in Computer Science, 7029:240–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-28557-8_30.

Sahba:2012:NMC

[1003] Farhang Sahba. A new method for contour determination of the prostate in ultrasound images. Lecture Notes in Computer Science, 7029:248–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
28557-8_31.

Suzuki:2012:AMO

[1004] Miyuki Suzuki, Marius George Linguraru, Ronald M. Summers, and Kazunori Okada. Analyses of missing organs in abdominal multi-organ segmentation. Lecture Notes in Computer Science, 7029:256–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
28557-8_32.

Ho:2012:GMP

[1005] Harvey Ho, Adam Bartlett, and Peter Hunter. Geometric modelling of patient-specific hepatic structures using cubic Hermite elements. Lecture Notes in Computer Science, 7029:264–271, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-28557-8_33.

Anonymous:2012:BMbb

[1006] Anonymous. Back matter. Lecture Notes in Computer Science, 7029: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28557-8/1.

Anonymous:2012:FMbh

[1007] Anonymous. Front matter. Lecture Notes in Computer Science, 7029: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28557-8/1.

Eades:2012:KS

[1008] Peter Eades, Seok-Hee Hong, and Kazuo Misue. Kozo Sugiyama 1945– 2011. Lecture Notes in Computer Science, 7034:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/accesspage/chapter/10.1007/978-3-642-25878-
7_1.

Eppstein:2012:CHD

[1009] David Eppstein and Joseph A. Simons. Confluent Hasse diagrams. Lecture Notes in Computer Science, 7034:2–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-25878-7_2/.

Alamdari:2012:POR

[1010] Soroush Alamdari and Therese Biedl. Planar open rectangle-of-influence drawings with non-aligned frames. Lecture Notes in Computer Science, 7034:14–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25878-7_3/.

Alam:2012:PCR

[1011] Muhammad Jawaherul Alam, Therese Biedl, Stefan Felsner, and Michael Kaufmann. Proportional contact representations of planar graphs. Lecture Notes in Computer Science, 7034:26–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-25878-7_4/.

Durocher:2012:EPT

[1012] Stephane Durocher, Debajyoti Mondal, Rahnuma Islam Nishat, and Md. Saidur Rahman. Embedding plane 3-trees in R^{2} and R^{3}. Lecture Notes in Computer Science, 7034:39–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-25878-7_5/.

DiGiacomo:2012:OPS

[1013] Emilio Di Giacomo, Fabrizio Frati, Radoslav Fulek, Luca Grilli, and Marcus Krug. Orthogeodesic point-set embedding of trees. Lecture Notes in Computer Science, 7034:52–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25878-7_6/.

Dujmovic:2012:PSS

[1014] Vida Dujmovic, William Evans, Sylvain Lazard, William Lenhart, and Giuseppe Liotta. On point-sets that support planar graphs. Lecture Notes in Computer Science, 7034:64–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-25878-7_7/.

Angelini:2012:SPS

[1015] Patrizio Angelini, Giuseppe Di Battista, Michael Kaufmann, and Tamara Mchedlidze. Small point sets for simply-nested planar graphs. Lecture Notes in Computer Science, 7034:75–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-25878-7_8/.

vanWijk:2012:GV

[1016] Jarke J. van Wijk. Graph visualization. Lecture Notes in Computer Science, 7034:86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-25878-7_9.

Chimani:2012:APM

[1017] Markus Chimani and Carsten Gutwenger. Advances in the planarization method: Effective multiple edge insertions. Lecture Notes in Computer Science, 7034:87–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25878-7_10/.

Brandes:2012:QCS

[1018] Ulrik Brandes and Martin Mader. A quantitative comparison of stressminimization approaches for offline dynamic graph drawing. Lecture Notes in Computer Science, 7034:99–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-25878-7_11/.

Cornelsen:2012:ABM

[1019] Sabine Cornelsen and Andreas Karrenbauer. Accelerated bend minimization. Lecture Notes in Computer Science, 7034:111–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25878-7_12/.

Nguyen:2012:TEN

[1020] Quan Nguyen, Seok-Hee Hong, and Peter Eades. TGI–EB: a new framework for edge bundling integrating topology, geometry and importance. Lecture Notes in Computer Science, 7034:123–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25878-7_13/.

Pupyrev:2012:ERO

[1021] Sergey Pupyrev, Lev Nachmanson, Sergey Bereg, and Alexander E. Holroyd. Edge routing with ordered bundles. Lecture Notes in Computer Science, 7034:136–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25878-7_14/.

Eades:2012:RAC

[1022] Peter Eades and Giuseppe Liotta. Right angle crossing graphs and 1-planarity. Lecture Notes in Computer Science, 7034:148–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25878-7_15/.

Halupczok:2012:PBP

[1023] Immanuel Halupczok and André Schulz. Pinning balloons with perfect angles and optimal area. Lecture Notes in Computer Science, 7034:154– 165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25878-7_16/.

Evans:2012:APD

[1024] William Evans, Emden R. Gansner, Michael Kaufmann, Giuseppe Liotta, and Henk Meijer. Approximate proximity drawings. Lecture Notes in Computer Science, 7034:166–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25878-7_17/.

Brunel:2012:GGG

[1025] Edith Brunel, Andreas Gemsa, Marcus Krug, Ignaz Rutter, and Dorothea Wagner. Generalizing geometric graphs. Lecture Notes in Computer Science, 7034:179–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25878-7_18/.

DiBattista:2012:HVR

[1026] Giuseppe Di Battista, Claudio Squarcella, and Wolfgang Nagele. How to visualize the K-root name server (demo). Lecture Notes in Computer Science, 7034:191–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25878-7_19/.

Dumas:2012:ORL

[1027] Maxime Dumas, Michael J. McGuffin, Jean-Marc Robert, and Marie-Claire Willig. Optimizing a radial layout of bipartite graphs for a tool visualizing security alerts. Lecture Notes in Computer Science, 7034:203– 214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25878-7_20/.

Greffard:2012:VCD

[1028] Nicolas Greffard, Fabien Picarougne, and Pascale Kuntz. Visual community detection: an evaluation of 2D, 3D perspective and 3D stereoscopic displays. Lecture Notes in Computer Science, 7034:215–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25878-7_21.

Burch:2012:EPD

[1029] Michael Burch, Corinna Vehlow, Natalia Konevtsova, and Daniel Weiskopf. Evaluating partially drawn links for directed graph edges. Lecture Notes in Computer Science, 7034:226–237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25878-7_22.

Rote:2012:RPG

[1030] Günter Rote. Realizing planar graphs as convex polytopes. Lecture Notes in Computer Science, 7034:238–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25878-7_23.

Kornaropoulos:2012:OOD

[1031] Evgenios M. Kornaropoulos and Ioannis G. Tollis. Overloaded orthogonal drawings. Lecture Notes in Computer Science, 7034:242–253, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25878-7_24.

Mukkamala:2012:DCG

[1032] Padmini Mukkamala and Dömötör Pálvölgyi. Drawing cubic graphs with the four basic slopes. Lecture Notes in Computer Science, 7034:254–265, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25878-7_25.

Suk:2012:QPG

[1033] Andrew Suk. k-quasi-planar graphs. Lecture Notes in Computer Science, 7034:266–277, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25878-7_26.

Pach:2012:MCN

[1034] János Pach and Géza Tóth. Monotone crossing number. Lecture Notes in Computer Science, 7034:278–289, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25878-7_27.

Cano:2012:UBC

[1035] Javier Cano, Csaba D. Tóth, and Jorge Urrutia. Upper bound constructions for untangling planar geometric graphs. Lecture Notes in Computer Science, 7034:290–295, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25878-7_28.

Aichholzer:2012:TCA

[1036] Oswin Aichholzer, Wolfgang Aigner, Franz Aurenhammer, and Kateřina Čech Dobiášová. Triangulations with circular arcs. Lecture Notes in Computer Science, 7034:296–307, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-25878-7_29.

Duncan:2012:PPA

[1037] Christian A. Duncan, David Eppstein, Michael T. Goodrich, and Stephen G. Kobourov. Planar and poly-arc Lombardi drawings. Lecture Notes in Computer Science, 7034:308–319, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25878-7_30.

Chernobelskiy:2012:FDL

[1038] Roman Chernobelskiy, Kathryn I. Cunningham, Michael T. Goodrich, and Stephen G. Kobourov. Force-directed Lombardi-style graph drawing. Lecture Notes in Computer Science, 7034:320–331, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25878-7_31.

Pach:2012:EGA

[1039] János Pach. Every graph admits an unambiguous bold drawing. Lecture Notes in Computer Science, 7034:332–342, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25878-7_32.

Fulek:2012:ACD

[1040] Radoslav Fulek, Michael J. Pelsmajer, Marcus Schaefer, and Daniel Štefankovič. Adjacent crossings do matter. Lecture Notes in Computer Science, 7034:343–354, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25878-7_33.

Sarkar:2012:LDD

[1041] Rik Sarkar. Low distortion Delaunay embedding of trees in hyperbolic plane. Lecture Notes in Computer Science, 7034:355–366, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25878-7_34.

Bannister:2012:HAC

[1042] Michael J. Bannister and David Eppstein. Hardness of approximate compaction for nonplanar orthogonal graph drawings. Lecture Notes in Computer Science, 7034:367–378, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25878-7_35.

Angelini:2012:MDG

[1043] Patrizio Angelini, Walter Didimo, Stephen Kobourov, Tamara Mchedlidze, and Vincenzo Roselli. Monotone drawings of graphs with fixed embedding. Lecture Notes in Computer Science, 7034:379–390, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25878-7_36.

Frati:2012:PNU

[1044] Fabrizio Frati, Radoslav Fulek, and Andres J. Ruiz-Vargas. On the page number of upward planar directed acyclic graphs. Lecture Notes in Computer Science, 7034:391–402, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25878-7_37.

Kaufmann:2012:UPS

[1045] Michael Kaufmann, Tamara Mchedlidze, and Antonios Symvonis. Upward point set embeddability for convex point sets is in P. Lecture Notes in Computer Science, 7034:403–414, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25878-7_38.

Auer:2012:CPU

[1046] Christopher Auer, Christian Bachmaier, Franz Josef Brandenburg, and Andreas Gleißner. Classification of planar upward embedding. Lecture Notes in Computer Science, 7034:415–426, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25878-7_39.

Binucci:2012:UPT

[1047] Carla Binucci and Walter Didimo. Upward planarity testing of embedded mixed graphs. Lecture Notes in Computer Science, 7034:427–432, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25878-7_40.

Argyriou:2012:CPR

[1048] Evmorfia N. Argyriou, Michael A. Bekos, Michael Kaufmann, and Antonios Symvonis. Combining problems on RAC drawings and simultaneous graph drawings. Lecture Notes in Computer Science, 7034:433–434, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25878-7_41.

Bachmaier:2012:OGA

[1049] Christian Bachmaier, Franz Josef Brandenburg, Philip Effinger, and Carsten Gutwenger. The open graph archive: a community-driven effort. Lecture Notes in Computer Science, 7034:435–440, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25878-7_42.

Fink:2012:DGV

[1050] Martin Fink, Jan-Henrik Haunert, Tamara Mchedlidze, Joachim Spoerhase, and Alexander Wolff. Drawing graphs with vertices at specified positions and crossings at large angles. Lecture Notes in Computer Science, 7034:441–442, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25878-7_43.

Holy:2012:VCD

[1051] Lukas Holy and Premek Brada. Viewport for component diagrams. Lecture Notes in Computer Science, 7034:443–444, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25878-7_44.

Milea:2012:SPP

[1052] Tal Milea, Okke Schrijvers, Kevin Buchin, and Herman Haverkort. Shortest-paths preserving metro maps. Lecture Notes in Computer Science, 7034:445–446, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-25878-7_45.

Zelina:2012:CNW

[1053] Remus Zelina, Sebastian Bota, Siebren Houtman, Jaap Jan van Assen, and Bas Hattink. Challenger, a new way to visualize data. Lecture Notes in Computer Science, 7034:447–448, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25878-7_46.

Duncan:2012:GDC

[1054] Christian A. Duncan, Carsten Gutwenger, Lev Nachmanson, and Georg Sander. Graph drawing contest report. Lecture Notes in Computer Science, 7034:449–455, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-25878-7_47.

Anonymous:2012:BMbc

[1055] Anonymous. Back matter. Lecture Notes in Computer Science, 7034: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-25878-7/1.

Anonymous:2012:FMbi

[1056] Anonymous. Front matter. Lecture Notes in Computer Science, 7034: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-25878-7/1.

Wastlund:2012:ECM

[1057] Erik Wästlund, Julio Angulo, and Simone Fischer-Hübner. Evoking comprehensive mental models of anonymous credentials. Lecture Notes in Computer Science, 7039:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27585-2_1/.

Heupel:2012:TUI

[1058] Marcel Heupel and Dogan Kesdogan. Towards usable interfaces for proof based access rights on mobile devices. Lecture Notes in Computer Science, 7039:15–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27585-2_2/.

Milutinovic:2012:CHA

[1059] Milica Milutinovic, Koen Decroix, Vincent Naessens, and Bart De Decker. Commercial home assistance (eHealth) services. Lecture Notes in Computer Science, 7039:28–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27585-2_3/.

Biedermann:2012:DCW

[1060] Sebastian Biedermann and Stefan Katzenbeisser. Detecting computer worms in the cloud. Lecture Notes in Computer Science, 7039:43–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27585-2_4/.

Bangerter:2012:ESI

[1061] Endre Bangerter, Stefan Bühlmann, and Engin Kirda. Efficient and stealthy instruction tracing and its applications in automated malware analysis: Open problems and challenges. Lecture Notes in Computer Science, 7039:55–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27585-2_5/.

Szydlowski:2012:CDA

[1062] Martin Szydlowski, Manuel Egele, Christopher Kruegel, and Giovanni Vigna. Challenges for dynamic analysis of iOS applications. Lecture Notes in Computer Science, 7039:65–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27585-2_6/.

Minier:2012:EEC

[1063] Marine Minier and Raphael C.-W. Phan. Energy-efficient cryptographic engineering paradigm. Lecture Notes in Computer Science, 7039:78– 88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27585-2_7/.

Tondel:2012:TSM

[1064] Inger Anne Tøndel and Åsmund Ahlmann Nyre. Towards a similarity metric for comparing machine-readable privacy policies. Lecture Notes in Computer Science, 7039:89–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27585-2_8/.

Bussard:2012:APP

[1065] Laurent Bussard and Ulrich Pinsdorf. Abstract privacy policy framework: Addressing privacy problems in SOA. Lecture Notes in Computer Science, 7039:104–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27585-2_9/.

Asghar:2012:FDC

[1066] Muhammad Rizwan Asghar and Giovanni Russello. Flexible and dynamic consent-capturing. Lecture Notes in Computer Science, 7039:119– 131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27585-2_10/.

Gross:2012:TUC

[1067] Stephan Groß and Alexander Schill. Towards user centric data governance and control in the cloud. Lecture Notes in Computer Science, 7039:132–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27585-2_11/.

Asghar:2012:SDP

[1068] Muhammad Rizwan Asghar, Mihaela Ion, Giovanni Russello, and Bruno Crispo. Securing data provenance in the cloud. Lecture Notes in Computer Science, 7039:145–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27585-2_12/.

Anonymous:2012:BMbd

[1069] Anonymous. Back matter. Lecture Notes in Computer Science, 7039: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27585-2/1.

Anonymous:2012:FMbj

[1070] Anonymous. Front matter. Lecture Notes in Computer Science, 7039: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27585-2/1.

Prakash:2012:FMM

[1071] Arun Prakash and Zoltán Theisz. Formal methods for modeling, refining and verifying autonomic components of computer networks. Lecture Notes in Computer Science, 7050:1–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28525-7_1/.

Sahin:2012:SOA

[1072] Cem Şafak Şahin, M. Ümit Uyar, and Stephen Gundry. Self organization for area coverage maximization and energy conservation in mobile ad hoc networks. Lecture Notes in Computer Science, 7050:49–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28525-7_2/.

Phan:2012:DID

[1073] Cong-Vinh Phan. Data intensive distributed computing in data aware self-organizing networks. Lecture Notes in Computer Science, 7050:74– 107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28525-7_3/.

Boonma:2012:AEB

[1074] Pruet Boonma and Junichi Suzuki. Accelerated evolution: a biologicallyinspired approach for augmenting self-star properties in wireless sensor networks. Lecture Notes in Computer Science, 7050:108–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28525-7_4/.

Vassev:2012:DAP

[1075] Emil Vassev and Serguei A. Mokhov. Developing autonomic properties for distributed pattern-recognition systems with ASSL. Lecture Notes in Computer Science, 7050:130–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28525-7_5/.

Manzalini:2012:ANI

[1076] Antonio Manzalini and Nermin Brgulja. Autonomic nature-inspired eco-systems. Lecture Notes in Computer Science, 7050:158–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28525-7_6/.

Halle:2012:LAD

[1077] Sylvain Hallé, Roger Villemaire, and Omar Cherkaoui. A logical approach to data-aware automated sequence generation. Lecture Notes in Computer Science, 7050:192–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28525-7_7/.

Anonymous:2012:BMbe

[1078] Anonymous. Back matter. Lecture Notes in Computer Science, 7050: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28525-7/1.

Anonymous:2012:FMbk

[1079] Anonymous. Front matter. Lecture Notes in Computer Science, 7050: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28525-7/1.

Frey:2012:CHC

[1080] Gerhard Frey and Ernst Kani. Correspondences on hyperelliptic curves and applications to the discrete logarithm. Lecture Notes in Computer Science, 7053:1–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25261-7_1/.

Nivre:2012:BBD

[1081] Joakim Nivre. Bare-bones dependency parsing. Lecture Notes in Computer Science, 7053:20–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25261-7_2/.

Seredynski:2012:SSS

[1082] Marcin Seredynski and Pascal Bouvry. Solving soft security problem in MANETs using an evolutionary approach. Lecture Notes in Computer Science, 7053:33–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25261-7_3/.

Ellwart:2012:CSD

[1083] Damian Ellwart, Piotr Szczuko, and Andrzej Czyżewski. Camera sabotage detection for surveillance systems. Lecture Notes in Computer Science, 7053:45–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25261-7_4/.

Poniszewska-Maranda:2012:IAC

[1084] Aneta Poniszewska-Maranda. Implementation of access control model for distributed information systems using usage control. Lecture Notes in Computer Science, 7053:54–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25261-7_5/.

Bernard:2012:BTT

[1085] Nicolas Bernard and Franck Leprévost. Beyond TOR: The TrueNyms Protocol. Lecture Notes in Computer Science, 7053:68–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25261-7_6/.

Varrette:2012:SSD

[1086] Sébastien Varrette, Benoît Bertholon, and Pascal Bouvry. A signature scheme for distributed executions based on control flow analysis. Lecture Notes in Computer Science, 7053:85–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-25261-7_7/.

Kordy:2012:CAA

[1087] Barbara Kordy, Marc Pouly, and Patrick Schweitzer. Computational aspects of attack–defense trees. Lecture Notes in Computer Science, 7053:103–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25261-7_8/.

Priemuth-Schmid:2012:ASV

[1088] Deike Priemuth-Schmid. Attacks on simplified versions of K2. Lecture Notes in Computer Science, 7053:117–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-25261-7_9/.

Mielniczuk:2012:MSL

[1089] Jan Mielniczuk and Paweł Teisseyre. Model selection in logistic regression using p-values and greedy search. Lecture Notes in Computer Science, 7053:128–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25261-7_10/.

Plewczynski:2012:LTM

[1090] Dariusz Plewczynski. Landau theory of meta-learning. Lecture Notes in Computer Science, 7053:142–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25261-7_11/.

Czajkowski:2012:MTD

[1091] Marcin Czajkowski, Marek Grześ, and Marek Kretowski. Multi-test decision trees for gene expression data analysis. Lecture Notes in Computer Science, 7053:154–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25261-7_12/.

Stanczyk:2012:RBA

[1092] Urszula Stańczyk. Rule-based approach to computational stylistics. Lecture Notes in Computer Science, 7053:168–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-25261-7_13/.

Raciborski:2012:DEH

[1093] Mikołaj Raciborski, Krzysztof Trojanowski, and Piotr Kaczyński. Differential evolution for high scale dynamic optimization. Lecture Notes in Computer Science, 7053:180–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25261-7_14/.

Puzniakowski:2012:TOI

[1094] Tadeusz Puźniakowski and Marek A. Bednarczyk. Towards an OpenCL implementation of ‘genetic algorithms’ on GPUs. Lecture Notes in Computer Science, 7053:190–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25261-7_15/.

Pinel:2012:EAP

[1095] Frédéric Pinel, Grégoire Danoy, and Pascal Bouvry. Evolutionary algorithm parameter tuning with sensitivity analysis. Lecture Notes in Computer Science, 7053:204–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25261-7_16/.

Cudek:2012:IRS

[1096] Paweł Cudek, Wiesław Paja, and Mariusz Wrzesień. Image recognition system for diagnosis support of melanoma skin lesion. Lecture Notes in Computer Science, 7053:217–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25261-7_17/.

Wieczorkowska:2012:PUR

[1097] Alicja A. Wieczorkowska, Miron B. Kursa, Elżbieta Kubera, and Radosław Rudnicki. Playing in unison in the random forest. Lecture Notes in Computer Science, 7053:226–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25261-7_18/.

Chojnacki:2012:SIB

[1098] Szymon Chojnacki and Mieczysław A. Kłopotek. Scale invariant bipartite graph generative model. Lecture Notes in Computer Science, 7053:240– 250, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25261-7_19/.

Sydow:2012:IDL

[1099] Marcin Sydow, Krzysztof Ciesielski, and Jakub Wajda. Introducing diversity to log-based query suggestions to deal with underspecified user queries. Lecture Notes in Computer Science, 7053:251–264, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25261-7_20/.

Ciesielski:2012:WBD

[1100] Krzysztof Ciesielski, Piotr Borkowski, Mieczysław A. Kłopotek, and Krzysztof Trojanowski. Wikipedia-based document categorization. Lecture Notes in Computer Science, 7053:265–278, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25261-7_21.

Wroblewska:2012:PEP

[1101] Alina Wróblewska and Marcin Woliński. Preliminary experiments in Polish dependency parsing. Lecture Notes in Computer Science, 7053:

279–292, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25261-7_22.

Broda:2012:EMA

[1102] Bartosz Broda, Roman Kurc, Maciej Piasecki, and Radosław Ramocki. Evaluation method for automated Wordnet expansion. Lecture Notes in Computer Science, 7053:293–306, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25261-7_23.

Kobylinski:2012:MCA

[1103] Łukasz Kobyliński. Mining class association rules for word sense disambiguation. Lecture Notes in Computer Science, 7053:307–317, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25261-7_24.

Pohl:2012:OBM

[1104] Aleksander Pohl. An ontology-based method for an efficient acquisition of relation extraction training and testing examples. Lecture Notes in Computer Science, 7053:318–331, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25261-7_25.

Marcinczuk:2012:RSF

[1105] Michał Marcińczuk, Michał Stanek, Maciej Piasecki, and Adam Musiał. Rich set of features for proper name recognition in Polish texts. Lecture Notes in Computer Science, 7053:332–344, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25261-7_26.

Hajnicz:2012:SBM

[1106] Elżbieta Hajnicz. Similarity-based method of detecting diathesis alternations in semantic valence dictionary of Polish verbs. Lecture Notes in Computer Science, 7053:345–358, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25261-7_27.

Sniatowski:2012:CPM

[1107] Tomasz Śniatowski and Maciej Piasecki. Combining Polish morphosyntactic taggers. Lecture Notes in Computer Science, 7053:359–369, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25261-7_28.

Degorski:2012:TLP

[1108] Łukasz Degórski. Towards the lemmatisation of Polish nominal syntactic groups using a shallow grammar. Lecture Notes in Computer Science, 7053:370–378, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25261-7_29.

Junczys-Dowmunt:2012:SSW

[1109] Marcin Junczys-Dowmunt and Arkadiusz Szał. SyMGiza++: Symmetrized word alignment models for statistical machine translation. Lecture Notes in Computer Science, 7053:379–390, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25261-7_30.

Wawer:2012:HOA

[1110] Aleksander Wawer and Krzysztof Sakwerda. How opinion annotations and ontologies become objective? Lecture Notes in Computer Science, 7053:391–400, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25261-7_31.

Anonymous:2012:BMbf

[1111] Anonymous. Back matter. Lecture Notes in Computer Science, 7053: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-25261-7/1.

Anonymous:2012:FMbl

[1112] Anonymous. Front matter. Lecture Notes in Computer Science, 7053: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-25261-7/1.

Anonymous:2012:FMbm

[1113] Anonymous. Front matter. Lecture Notes in Computer Science, 7054: 1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28038-2/1/1.

Michalik:2012:TDC

[1114] Bartosz Michalik and Jerzy Nawrocki. Towards decision centric repository of architectural knowledge. Lecture Notes in Computer Science, 7054:3–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28038-2_1/.

Mederly:2012:CMB

[1115] Pavol Mederly, Marián Lekavý, Marek Závodský, and Pavol Návrat. Construction of messaging-based enterprise integration solutions using AI planning. Lecture Notes in Computer Science, 7054:16–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28038-2_2/.

Hnatkowska:2012:IAB

[1116] Bogumiła Hnatkowska and Krzysztof Kasprzyk. Integration of application business logic and business rules with DSL and AOP. Lecture Notes in Computer Science, 7054:30–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28038-2_3/.

Menkyna:2012:AOC

[1117] Radoslav Menkyna and Valentino Vranić. Aspect-oriented change realization based on multi-paradigm design with feature modeling. Lecture Notes in Computer Science, 7054:40–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28038-2_4/.

Cote:2012:ECB

[1118] Isabelle Côté, Maritta Heisel, and Jeanine Souquières. On the evolution of component-based software. Lecture Notes in Computer Science, 7054: 54–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28038-2_5/.

Gorawski:2012:QPU

[1119] Marcin Gorawski and Aleksander Chrószcz. Query processing using negative and temporal tuples in stream query engines. Lecture Notes in Computer Science, 7054:70–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28038-2_6/.

Bedla:2012:SSJ

[1120] Mariusz Bedla and Krzysztof Sapiecha. Scalable store of Java objects using range partitioning. Lecture Notes in Computer Science, 7054:84– 93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28038-2_7/.

Anonymous:2012:FMbn

[1121] Anonymous. Front matter. Lecture Notes in Computer Science, 7054: 95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28038-2/2/1.

Samolej:2012:HBM

[1122] Sławomir Samolej and Tomasz Szmuc. HTCPNs-based modelling and evaluation of dynamic computer cluster reconfiguration. Lecture Notes in Computer Science, 7054:97–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28038-2_8/.

Rak:2012:PAI

[1123] Tomasz Rak and Jan Werewka. Performance analysis of interactive Internet systems for a class of systems with dynamically changing offers. Lecture Notes in Computer Science, 7054:109–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28038-2_9/.

Cerina-Berzina:2012:ISD

[1124] Jana Ceriņa-Bērziņa, Jānis Bičevskis, and Ģirts Karnītis. Information systems development based on visual domain specific language BiLingva. Lecture Notes in Computer Science, 7054:124–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28038-2_10/.

Poernomo:2012:MNR

[1125] Iman Poernomo and Timur Umarov. A mapping from normative requirements to event-B to facilitate verified data-centric business process management. Lecture Notes in Computer Science, 7054:136–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28038-2_11/.

Rebout:2012:CGR

[1126] Maxime Rebout, Louis Féraud, Lionel Marie-Magdeleine, and Sergei Soloviev. Computations in graph rewriting: Inductive types and pullbacks in DPO approach. Lecture Notes in Computer Science, 7054:150– 163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28038-2_12/.

Wozna-Szczesniak:2012:BMC

[1127] Bożena Woźna-Szcześniak. Bounded model checking for the existential part of real-time CTL and knowledge. Lecture Notes in Computer Science, 7054:164–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28038-2_13/.

Felkner:2012:SRB

[1128] Anna Felkner and Krzysztof Sacha. The semantics of role-based trust management languages. Lecture Notes in Computer Science, 7054:179– 189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28038-2_14/.

Anonymous:2012:FMbo

[1129] Anonymous. Front matter. Lecture Notes in Computer Science, 7054: 191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28038-2/3/1.

Bicevska:2012:AST

[1130] Zane Bičevska. Applying of smart technologies: Evaluation of effectiveness. Lecture Notes in Computer Science, 7054:193–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28038-2_15/.

Ferber:2012:RCC

[1131] Marvin Ferber, Sascha Hunold, Björn Krellner, Thomas Rauber, and Thomas Reichel. Reducing the class coupling of legacy code by a metricsbased relocation of class members. Lecture Notes in Computer Science, 7054:202–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28038-2_16/.

Bluemke:2012:DTJ

[1132] Ilona Bluemke and Artur Rembiszewski. Dataflow testing of Java programs with DFC. Lecture Notes in Computer Science, 7054:215–228, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28038-2_17/.

Derezinska:2012:OOT

[1133] Anna Derezińska and Anna Szustek. Object-oriented testing capabilities and performance evaluation of the C# mutation system. Lecture Notes in Computer Science, 7054:229–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28038-2_18/.

Khalilian:2012:BCT

[1134] Alireza Khalilian and Saeed Parsa. Bi-criteria test suite reduction by cluster analysis of execution profiles. Lecture Notes in Computer Science, 7054:243–256, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28038-2_19/.

Anonymous:2012:BMbg

[1135] Anonymous. Back matter. Lecture Notes in Computer Science, 7054: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28038-2/1.

Anonymous:2012:FMbp

[1136] Anonymous. Front matter. Lecture Notes in Computer Science, 7054: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28038-2/1.

Gong:2012:KNF

[1137] Zheng Gong, Svetla Nikova, and Yee Wei Law. KLEIN: a new family of lightweight block ciphers. Lecture Notes in Computer Science, 7055: 1–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25286-0_1/.

Engels:2012:HLA

[1138] Daniel Engels, Markku-Juhani O. Saarinen, and Peter Schweitzer. The Hummingbird-2 lightweight authenticated encryption algorithm. Lecture Notes in Computer Science, 7055:19–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-25286-0_2/.

Pendl:2012:ECC

[1139] Christian Pendl, Markus Pelnar, and Michael Hutter. Elliptic curve cryptography on the WISP UHF RFID tag. Lecture Notes in Computer Science, 7055:32–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25286-0_3/.

Barenghi:2012:EFL

[1140] Alessandro Barenghi, Cédric Hocquet, and David Bol. Exploring the feasibility of low cost fault injection attacks on sub-threshold devices through an example of a 65nm AES implementation. Lecture Notes in Computer Science, 7055:48–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25286-0_4/.

Kasper:2012:SCA

[1141] Timo Kasper, David Oswald, and Christof Paar. Side-channel analysis of cryptographic RFIDs with analog demodulation. Lecture Notes in Computer Science, 7055:61–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25286-0_5/.

Kardas:2012:NRD

[1142] Süleyman Kardaş, Mehmet Sabir Kiraz, and Muhammed Ali Bingöl. A novel RFID distance bounding protocol based on physically unclonable functions. Lecture Notes in Computer Science, 7055:78–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25286-0_6/.

Abyaneh:2012:SAT

[1143] Mohammad Reza Sohizadeh Abyaneh. Security analysis of two distancebounding protocols. Lecture Notes in Computer Science, 7055:94–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25286-0_7/.

Amariucai:2012:ATB

[1144] George T. Amariucai, Clifford Bergman, and Yong Guan. An automatic, time-based, secure pairing protocol for passive RFID. Lecture Notes in Computer Science, 7055:108–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25286-0_8/.

Chai:2012:BSP

[1145] Qi Chai and Guang Gong. BUPLE: Securing passive RFID communication through physical layer enhancements. Lecture Notes in Computer Science, 7055:127–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25286-0_9/.

Fernandez-Mir:2012:SRA

[1146] Albert Fernàndez-Mir and Rolando Trujillo-Rasua. A scalable RFID authentication protocol supporting ownership transfer and controlled delegation. Lecture Notes in Computer Science, 7055:147–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25286-0_10/.

Elkhiyaoui:2012:RRO

[1147] Kaoutar Elkhiyaoui, Erik-Oliver Blass, and Refik Molva. ROTIV: RFID ownership transfer with issuer verification. Lecture Notes in Computer Science, 7055:163–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25286-0_11/.

Batina:2012:HEB

[1148] Lejla Batina, Stefaan Seys, Dave Singelée, and Ingrid Verbauwhede. Hierarchical ECC-based RFID authentication protocol. Lecture Notes in Computer Science, 7055:183–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25286-0_12/.

Anonymous:2012:BMbh

[1149] Anonymous. Back matter. Lecture Notes in Computer Science, 7055: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-25286-0/1.

Anonymous:2012:FMbq

[1150] Anonymous. Front matter. Lecture Notes in Computer Science, 7055: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-25286-0/1.

Wobcke:2012:WCA

[1151] Wayne Wobcke, Nirmit Desai, Frank Dignum, Aditya Ghose, and Srinivas Padmanabhuni. What can agent-based computing offer serviceoriented architectures, and vice versa? Lecture Notes in Computer Science, 7057:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25920-3_1/.

Billiau:2012:SNR

[1152] Graham Billiau, Chee Fon Chang, and Aditya Ghose. SBDO: a new robust approach to dynamic distributed constraint optimisation. Lecture Notes in Computer Science, 7057:11–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-25920-3_2/.

Mailler:2012:SDC

[1153] Roger Mailler and Jacob Graves. Solving distributed CSPs using dynamic, partial centralization without explicit constraint passing. Lecture Notes in Computer Science, 7057:27–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-25920-3_3/.

Doherty:2012:DTS

[1154] Patrick Doherty, Fredrik Heintz, and David Landén. A distributed task specification language for mixed-initiative delegation. Lecture Notes in Computer Science, 7057:42–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25920-3_4/.

Smith:2012:ANA

[1155] Melanie Smith, Sandip Sen, and Roger Mailler. Adaptive and nonadaptive distribution functions for DSA. Lecture Notes in Computer Science, 7057:58–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25920-3_5/.

Khanna:2012:MBS

[1156] Sankalp Khanna, Timothy Cleaver, Abdul Sattar, David Hansen, and Bela Stantic. Multiagent based scheduling of elective surgery. Lecture Notes in Computer Science, 7057:74–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-25920-3_6/.

Sugawara:2012:EAD

[1157] Toshiharu Sugawara, Kensuke Fukuda, Toshio Hirotsu, and Satoshi Kurihara. Effect of alternative distributed task allocation strategy based on local observations in contract net protocol. Lecture Notes in Computer Science, 7057:90–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25920-3_7/.

Savarimuthu:2012:GBS

[1158] Sharmila Savarimuthu, Martin Purvis, Bastin Tony Roy Savarimuthu, and Maryam Purvis. Gossip-based self-organising open agent societies. Lecture Notes in Computer Science, 7057:105–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25920-3_8/.

Le:2012:ANM

[1159] Thao P. Le, Timothy J. Norman, and Wamberto Vasconcelos. Adaptive negotiation in managing wireless sensor networks. Lecture Notes in Computer Science, 7057:121–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25920-3_9/.

Al-Jaljouli:2012:NSM

[1160] Raja Al-Jaljouli and Jemal Abawajy. Negotiation strategy for mobile agent-based e-negotiation. Lecture Notes in Computer Science, 7057:137– 151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25920-3_10/.

Grove:2012:ACB

[1161] Frank Grove, Sandip Sen, and Oly Mistry. Adaptive choice of behavior and protocol parameters. Lecture Notes in Computer Science, 7057:152– 165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25920-3_11/.

Okimoto:2012:EDV

[1162] Tenda Okimoto, Atsushi Iwasaki, and Makoto Yokoo. Effect of DisCSP variable-ordering heuristics in scale-free networks. Lecture Notes in Computer Science, 7057:166–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25920-3_12/.

Hindriks:2012:MAP

[1163] Koen V. Hindriks, Wietske Visser, and Catholijn M. Jonker. Multiattribute preference logic. Lecture Notes in Computer Science, 7057:181– 195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25920-3_13/.

Hindriks:2012:ESP

[1164] Koen V. Hindriks, M. Birna van Riemsdijk, and Catholijn M. Jonker. An empirical study of patterns in agent programs. Lecture Notes in Computer Science, 7057:196–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25920-3_14/.

Vo:2012:MLR

[1165] Duc-An Vo, Alexis Drogoul, Jean-Daniel Zucker, and Tuong-Vinh Ho. A modelling language to represent and specify emerging structures in agent-based model. Lecture Notes in Computer Science, 7057:212–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25920-3_15/.

Nakajima:2012:MMB

[1166] Yuu Nakajima, Shohei Yamane, and Hiromitsu Hattori. Multi-model based simulation platform for urban traffic simulation. Lecture Notes in Computer Science, 7057:228–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25920-3_16/.

Taillandier:2012:GSP

[1167] Patrick Taillandier, Duc-An Vo, Edouard Amouroux, and Alexis Drogoul. GAMA: a simulation platform that integrates geographical information data, agent-based modeling and multi-scale control. Lecture Notes in Computer Science, 7057:242–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25920-3_17/.

Seghrouchni:2012:ADA

[1168] Amal El Fallah Seghrouchni, Andrei Olaru, Nga Thi Thuy Nguyen, and Diego Salomone. Ao dai: Agent oriented design for ambient intelligence. Lecture Notes in Computer Science, 7057:259–269, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25920-3_18/.

Mistry:2012:PAT

[1169] Oly Mistry and Sandip Sen. Probabilistic approaches to tag recommendation in a social bookmarking network. Lecture Notes in Computer Science, 7057:270–287, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25920-3_19/.

Landen:2012:CTA

[1170] David Landén, Fredrik Heintz, and Patrick Doherty. Complex task allocation in mixed-initiative delegation: a UAV case study. Lecture Notes in Computer Science, 7057:288–303, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25920-3_20/.

Sindlar:2012:ABI

[1171] Michal Sindlar and John-Jules Meyer. Affordance-based intention recognition in virtual spatial environments. Lecture Notes in Computer Science, 7057:304–319, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-25920-3_21.

Furuhata:2012:RMU

[1172] Masabumi Furuhata. A robust multi-unit ascending-price auction with complementarities against strategic manipulation. Lecture Notes in Computer Science, 7057:320–335, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25920-3_22.

Godfrey:2012:MAC

[1173] W. Wilfred Godfrey and Shivashankar B. Nair. Mobile agent cloning for servicing networked robots. Lecture Notes in Computer Science, 7057:336–339, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25920-3_23.

Balsdon:2012:TDA

[1174] Quintin J. Balsdon and Elize M. Ehlers. Towards distributing agent intelligence: Using decentralized software services for the creation of complex problem modelling. Lecture Notes in Computer Science, 7057:340–354, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25920-3_24.

Sen:2012:ATC

[1175] Onkur Sen and Sandip Sen. Averting the tragedy of the commons by adapting aspiration levels. Lecture Notes in Computer Science, 7057: 355–370, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25920-3_25.

Koch:2012:RAA

[1176] Fernando Koch, Frank Dignum, Marcel Hiel, and Huib Aldewereld. The role of agents in adaptive service oriented architectures. Lecture Notes in Computer Science, 7057:371–386, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25920-3_26.

Dam:2012:ABD

[1177] Hoa Khanh Dam and Aditya Ghose. Agent-based development for business processes. Lecture Notes in Computer Science, 7057:387–393, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25920-3_27.

Grove:2012:TMA

[1178] Frank Grove and Sandip Sen. TwitAg: a multi-agent feature selection and recommendation framework for Twitter. Lecture Notes in Computer Science, 7057:394–397, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25920-3_28.

Caillou:2012:AMA

[1179] Philippe Caillou. Automated multi-agent simulation generation and validation. Lecture Notes in Computer Science, 7057:398–412, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25920-3_29.

Nguyen:2012:IEB

[1180] Ngoc Doanh Nguyen, Patrick Taillandier, Alexis Drogoul, and Pierre Auger. Inferring equation-based models from agent-based models: a case study in competition dynamics. Lecture Notes in Computer Science, 7057:413–427, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25920-3_30.

Chu:2012:TMP

[1181] Thanh-Quang Chu, Alexis Drogoul, Alain Boucher, and Jean-Daniel Jucker. Towards a methodology for the participatory design of agentbased models. Lecture Notes in Computer Science, 7057:428–442, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25920-3_31.

Nakagawa:2012:FVT

[1182] Hiroyuki Nakagawa, Nobukazu Yoshioka, Akihiko Ohsuga, and Shinichi Honiden. A framework for validating task assignment in multiagent systems using requirements importance. Lecture Notes in Computer Science, 7057:443–458, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25920-3_32.

Cheah:2012:TKP

[1183] WaiShiang Cheah, Leon Sterling, and Kuldar Taveter. Task knowledge patterns reuse in multi-agent systems development. Lecture Notes in Computer Science, 7057:459–474, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25920-3_33.

Lee:2012:EAA

[1184] Shih chiang Lee, Gu yuan Lin, Wan rong Jih, Chi-Chia Huang, and Jane Yung jen Hsu. Energy-aware agents for detecting nonessential appliances. Lecture Notes in Computer Science, 7057:475–486, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25920-3_34.

Lopez:2012:MEM

[1185] Beatriz Lopez, Albert Pla, David Daroca, Luis Collantes, Sara Lozano, and Joaquim Meléndez. Medical equipment maintenance support with service-oriented multi-agent services. Lecture Notes in Computer Science, 7057:487–498, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25920-3_35.

Dam:2012:AOA

[1186] Hoa Khanh Dam and Aditya Ghose. An agent-oriented approach to service analysis and design. Lecture Notes in Computer Science, 7057:

499–510, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25920-3_36.

Wobcke:2012:ABM

[1187] Wayne Wobcke and Adam Dunn. Agent-based modelling for risk assessment of routine clinical processes. Lecture Notes in Computer Science, 7057:511–522, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25920-3_37.

Solomonides:2012:HSP

[1188] Tony Solomonides. Healthgrids, the SHARE project, medical data and agents: Retrospect and prospect. Lecture Notes in Computer Science, 7057:523–534, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25920-3_38.

Khanna:2012:IAS

[1189] Sankalp Khanna, Abdul Sattar, Justin Boyle, David Hansen, and Bela Stantic. An intelligent approach to surgery scheduling. Lecture Notes in Computer Science, 7057:535–550, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25920-3_39.

Billiau:2012:UDA

[1190] Graham Billiau, Chee Fon Chang, Aditya Ghose, and Alexis Andrew Miller. Using distributed agents for patient scheduling. Lecture Notes in Computer Science, 7057:551–560, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25920-3_40.

Miller:2012:SAC

[1191] Alexis Andrew Miller and Fiona Hegi-Johnson. Software agents in clinical workflow, clinical guidelines and clinical trial medicine. Lecture Notes in Computer Science, 7057:561–574, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25920-3_41.

Taillandier:2012:UBT

[1192] Patrick Taillandier, Edouard Amouroux, Duc An Vo, and Ana-Maria Olteanu-Raimond. Using belief theory to formalize the agent behavior: Application to the simulation of avian flu propagation. Lecture Notes in Computer Science, 7057:575–587, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25920-3_42.

Vu:2012:CBA

[1193] Quang-Anh Nguyen Vu, Benoit Gaudou, Richard Canal, Salima Hassas, and Frédéric Armetta. A cluster-based approach for disturbed, spatialized, distributed information gathering systems. Lecture Notes in Computer Science, 7057:588–603, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25920-3_43.

Minh:2012:SED

[1194] Le Van Minh, Carole Adam, Richard Canal, Benoit Gaudou, and Ho Tuong Vinh. Simulation of the emotion dynamics in a group of agents in an evacuation situation. Lecture Notes in Computer Science, 7057:604–619, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25920-3_44.

Gil-Quijano:2012:BUC

[1195] Javier Gil-Quijano, Thomas Louail, and Guillaume Hutzler. From biological to urban cells: Lessons from three multilevel agent-based models. Lecture Notes in Computer Science, 7057:620–635, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25920-3_45.

Bui:2012:MAB

[1196] The Duy Bui, Duc Hai Ngo, and Cong Tran. Multi-agent based simulation of traffic in Vietnam. Lecture Notes in Computer Science, 7057: 636–648, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25920-3_46.

Anonymous:2012:BMbi

[1197] Anonymous. Back matter. Lecture Notes in Computer Science, 7057: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-25920-3/1.

Anonymous:2012:FMbr

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-25920-3/1.

Aghaee:2012:EMT

[1199] Saeed Aghaee and Cesare Pautasso. An evaluation of mashup tools based on support for heterogeneous mashup components. Lecture Notes in Computer Science, 7059:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27997-3_1/.

Liu:2012:ACD

[1200] Dong Liu, Ning Li, Carlos Pedrinaci, Jacek Kopecký, Maria Maleshkova, and John Domingue. An approach to construct dynamic service mashups using lightweight semantics. Lecture Notes in Computer Science, 7059: 13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27997-3_2/.

Tietz:2012:TBR

[1201] Vincent Tietz, Gregor Blichmann, Stefan Pietschmann, and Klaus Meißner. Task-based recommendation of mashup components. Lecture Notes in Computer Science, 7059:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27997-3_3/.

Chudnovskyy:2012:ITS

[1202] Olexiy Chudnovskyy, Frank Weinhold, Hendrik Gebhardt, and Martin Gaedke. Integration of telco services into enterprise mashup applications. Lecture Notes in Computer Science, 7059:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27997-3_4/.

Wilson:2012:OUI

[1203] Scott Wilson, Florian Daniel, Uwe Jugel, and Stefano Soi. Orchestrated user interface mashups using W3C widgets. Lecture Notes in Computer Science, 7059:49–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27997-3_5/.

Lautamaki:2012:CDE

[1204] Janne Lautamäki and Tommi Mikkonen. Cross-domain embedding for vaadin applications. Lecture Notes in Computer Science, 7059:62–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27997-3_6/.

Bellido:2012:WLB

[1205] Jesus Bellido, Rosa Alarcon, and Cristian Sepulveda. Web linkingbased protocols for guiding RESTful M2M interaction. Lecture Notes in Computer Science, 7059:74–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27997-3_7/.

Kochman:2012:BTR

[1206] Sebastian Kochman, Paweł T. Wojciechowski, and Miłosz Kmieciak. Batched transactions for RESTful Web services. Lecture Notes in Computer Science, 7059:86–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27997-3_8/.

Herbert:2012:SMP

[1207] Matthias Herbert, Tobias Thieme, Jan Zibuschka, and Heiko Roßnagel. Secure mashup-providing platforms — implementing encrypted wiring. Lecture Notes in Computer Science, 7059:99–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27997-3_9/.

Bozzon:2012:CFL

[1208] Alessandro Bozzon, Marco Brambilla, Emanuele Della Valle, and Piero Fraternali. A conceptual framework for linked data exploration. Lecture Notes in Computer Science, 7059:109–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27997-3_10/.

Cohen:2012:SRE

[1209] Marcelo Cohen and Daniel Schwabe. Support for reusable explorations of linked data in the semantic Web. Lecture Notes in Computer Science, 7059:119–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27997-3_11/.

Ko:2012:GSC

[1210] Han-Gyu Ko and In-Young Ko. Generation of semantic clouds based on linked data for efficient multimedia semantic annotation. Lecture Notes in Computer Science, 7059:127–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27997-3_12/.

Helou:2012:OBS

[1211] Mamoun Abu Helou. Ontology based segmentation of geo-referenced queries. Lecture Notes in Computer Science, 7059:135–144, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/10.1007/978-3-642-27997-3_
13/.

Durao:2012:SSB

[1212] Frederico Durao, Peter Dolog, Martin Leginus, and Ricardo Lage. Sim-Spectrum: a similarity based spectral clustering approach to generate a tag cloud. Lecture Notes in Computer Science, 7059:145–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27997-3_14/.

Keller:2012:GAP

[1213] Matthias Keller and Martin Nussbaumer. Graph access pattern diagrams (GAP–D): Towards a unified approach for modeling navigation over hierarchical, linear and networked structures. Lecture Notes in Computer Science, 7059:155–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27997-3_15/.

Morales-Chaparro:2012:DDU

[1214] Rober Morales-Chaparro, Juan C. Preciado, and Fernando Sánchez-Figueroa. Data-driven and user-driven multidimensional data visualization. Lecture Notes in Computer Science, 7059:159–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27997-3_16/.

Nebeling:2012:CAA

[1215] Michael Nebeling and Moira C. Norrie. Context-aware and adaptive Web interfaces: a crowdsourcing approach. Lecture Notes in Computer Science, 7059:167–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27997-3_17/.

Biewald:2012:MMH

[1216] Lukas Biewald. Massive multiplayer human computation for fun, money, and survival. Lecture Notes in Computer Science, 7059:171– 176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27997-3_18/.

Jayakanthan:2012:ECS

[1217] Ranganathan Jayakanthan and Deepak Sundararajan. Enterprise crowdsourcing solution for software development in an outsourcing organization. Lecture Notes in Computer Science, 7059:177–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27997-3_19/.

Achilleos:2012:MDF

[1218] Achilleas Achilleos, Georgia M. Kapitsaki, and George A. Papadopoulos. A model-driven framework for developing Web service oriented applications. Lecture Notes in Computer Science, 7059:181–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27997-3_20/.

Eickhoff:2012:DEW

[1219] Christoph Eickhoff, Nina Geiger, Marcel Hahn, and Albert Zündorf. Developing enterprise Web applications using the story driven modeling approach. Lecture Notes in Computer Science, 7059:196–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27997-3_21.

Zhang:2012:AOM

[1220] Gefei Zhang and Matthias Hölzl. Aspect-oriented modeling of Web applications with hiLA. Lecture Notes in Computer Science, 7059:211–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27997-3_22.

Escott:2012:MDW

[1221] Eban Escott, Paul Strooper, Paul King, and Ian J. Hayes. Modeldriven Web form validation with UML and OCL. Lecture Notes in Computer Science, 7059:223–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27997-3_23.

Rodriguez-Echeverria:2012:MLW

[1222] Roberto Rodríguez-Echeverría, José María Conejero, Pedro J. Clemente, and Juan C. Preciado. Modernization of legacy Web applications into rich Internet applications. Lecture Notes in Computer Science, 7059:236–250, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27997-3_24.

Polillo:2012:QMW

[1223] Roberto Polillo. Quality models for Web [2.0] sites: a methodological approach and a proposal. Lecture Notes in Computer Science, 7059:251–265, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27997-3_25.

Orehovacki:2012:EQU

[1224] Tihomir Orehovački, Andrina Granić, and Dragutin Kermek. Exploring the quality in use of Web 2.0 applications: The case of mind mapping services. Lecture Notes in Computer Science, 7059:266–277, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27997-3_26.

Urbieta:2012:DCI

[1225] Matias Urbieta, Maria Jose Escalona, Esteban Robles Luna, and Gustavo Rossi. Detecting conflicts and inconsistencies in Web application requirements. Lecture Notes in Computer Science, 7059:278–288, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27997-3_27.

Pandurino:2012:SCC

[1226] Andrea Pandurino, Davide Bolchini, Luca Mainetti, and Roberto Paiano. Streamlining complexity: Conceptual page re-modeling for rich Internet applications. Lecture Notes in Computer Science, 7059:289–301, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27997-3_28.

Braunschweig:2012:FGB

[1227] Katrin Braunschweig, Maik Thiele, and Wolfgang Lehner. A flexible graph-based data model supporting incremental schema design and evolution. Lecture Notes in Computer Science, 7059:302–306, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27997-3_29.

Kalchgruber:2012:PPL

[1228] Peter Kalchgruber. ProLD: Propagate linked data. Lecture Notes in Computer Science, 7059:307–311, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27997-3_30.

Katz:2012:CRD

[1229] Philipp Katz and Alexander Schill. Causal relation detection for activities from heterogeneous sources. Lecture Notes in Computer Science, 7059:312–316, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27997-3_31.

Maly:2012:XDV

[1230] Jakub Malý and Martin Nečaský. XML document versioning, revalidation and constraints. Lecture Notes in Computer Science, 7059:317–321, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27997-3_32.

Mather:2012:ROP

[1231] Neil Mather and Samia Oussena. A reuse-oriented product-line method for enterprise Web applications. Lecture Notes in Computer Science, 7059:322–326, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27997-3_33.

Firmenich:2012:FAC

[1232] Sergio Firmenich, Gustavo Rossi, Silvia Gordillo, and Marco Winckler. A flexible architecture for client-side adaptation. Lecture Notes in Computer Science, 7059:327–331, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27997-3_34.

Chaisatien:2012:AMA

[1233] Prach Chaisatien, Korawit Prutsachainimmit, and Takehiro Tokuda. Applications of mobile application interface description language MAIDL. Lecture Notes in Computer Science, 7059:332–336, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27997-3_35.

Eberius:2012:DSL

[1234] Julian Eberius, Maik Thiele, and Wolfgang Lehner. A domain-specific language for do-it-yourself analytical mashups. Lecture Notes in Computer Science, 7059:337–341, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27997-3_36.

Fayzrakhmanov:2012:IEW

[1235] Ruslan R. Fayzrakhmanov. Information extraction from Web pages based on their visual representation. Lecture Notes in Computer Science, 7059:342–346, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27997-3_37.

Aghaee:2012:EUP

[1236] Saeed Aghaee and Cesare Pautasso. End-user programming for Web mashups. Lecture Notes in Computer Science, 7059:347–351, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27997-3_38.

Motti:2012:MDC

[1237] Vivian Genaro Motti and Jean Vanderdonckt. Multi-dimensional context-aware adaptation for Web applications. Lecture Notes in Computer Science, 7059:352–354, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27997-3_39.

Abel:2012:EPS

[1238] Fabian Abel and Geert-Jan Houben. Engineering the personal social semantic Web. Lecture Notes in Computer Science, 7059:355–356, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27997-3_40.

Maleshkova:2012:AUW

[1239] Maria Maleshkova, Carlos Pedrinaci, Dong Liu, and Guillermo Alvaro. Automating the use of Web APIs through lightweight semantics. Lecture Notes in Computer Science, 7059:357–358, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27997-3_41.

Lew:2012:IQU

[1240] Philip Lew and Luis Olsina. Improving quality in use of Web applications in a systematic way. Lecture Notes in Computer Science, 7059:359–360, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27997-3_42.

Anonymous:2012:BMbj

[1241] Anonymous. Back matter. Lecture Notes in Computer Science, 7059: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27997-3/1.

Anonymous:2012:FMbs

[1242] Anonymous. Front matter. Lecture Notes in Computer Science, 7059: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27997-3/1.

Wicker:2012:LMM

[1243] Andrew W. Wicker and Jon Doyle. Leveraging multiple mechanisms for information propagation. Lecture Notes in Computer Science, 7068: 1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-27216-5_1.

Harb:2012:CSM

[1244] Hassan Harb, F. Jordan Srour, and Neil Yorke-Smith. A case study in model selection for policy engineering: Simulating maritime customs. Lecture Notes in Computer Science, 7068:3–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27216-5_2/.

Fridman:2012:TQR

[1245] Natalie Fridman, Gal A. Kaminka, and Avishay Zilka. Towards qualitative reasoning for policy decision support in demonstrations. Lecture Notes in Computer Science, 7068:19–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27216-5_3/.

Botti:2012:RMD

[1246] Vicente Botti, Antonio Garrido, Adriana Giret, and Pablo Noriega. The role of MAS as a decision support tool in a water-rights market. Lecture Notes in Computer Science, 7068:35–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27216-5_4/.

Dechesne:2012:UCD

[1247] Francien Dechesne, Virginia Dignum, and Yao-Hua Tan. Understanding compliance differences between legal and social norms: The case of smoking ban. Lecture Notes in Computer Science, 7068:50–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27216-5_5/.

Morris:2012:MCM

[1248] Alexis Morris, William Ross, and Mihaela Ulieru. Modelling culture in multi-agent organizations. Lecture Notes in Computer Science, 7068:65– 79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27216-5_6/.

Pedell:2012:SAB

[1249] Sonja Pedell, Tim Miller, Leon Sterling, Frank Vetere, and Steve Howard. Substantiating agent-based quality goals for understanding socio-technical systems. Lecture Notes in Computer Science, 7068:80– 95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27216-5_7/.

Ghorbani:2012:ADF

[1250] Amineh Ghorbani, Virginia Dignum, and Gerard Dijkema. An analysis and design framework for agent-based social simulation. Lecture Notes in Computer Science, 7068:96–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27216-5_8/.

Pedell:2012:BAB

[1251] Sonja Pedell and Leon Sterling. The benefits of agent-based motivation models in policy formulation and implementation. Lecture Notes in Computer Science, 7068:113–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27216-5_9/.

Hessler:2012:APA

[1252] Axel Hessler, Benjamin Hirsch, Tobias Küster, and Sahin Albayrak. AgentStore — a pragmatic approach to agent reuse. Lecture Notes in Computer Science, 7068:128–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27216-5_10/.

Nunes:2012:DAB

[1253] Ingrid Nunes, Michael Luck, Simone Diniz Junqueira Barbosa, and Simon Miles. Dynamically adapting BDI agents based on high-level user specifications. Lecture Notes in Computer Science, 7068:139–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27216-5_11/.

Schumann:2012:ECS

[1254] René Schumann. Engineering coordination: Selection of coordination mechanisms. Lecture Notes in Computer Science, 7068:164–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27216-5_12/.

Frantz:2012:AAA

[1255] Christopher Frantz, Mariusz Nowostawski, and Martin K. Purvis. Augmenting Android with AOSE principles for enhanced functionality reuse in mobile applications. Lecture Notes in Computer Science, 7068:187– 211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27216-5_13/.

Hindriks:2012:INI

[1256] Koen Hindriks, Mark A. Neerincx, and Mirek Vink. The iCat as a natural interaction partner. Lecture Notes in Computer Science, 7068:212– 231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27216-5_14/.

Sklar:2012:DHF

[1257] Elizabeth Sklar, A. Tuna Ozgelen, J. Pablo Munoz, Joel Gonzalez, and Mark Manashirov. Designing the HRTeam Framework: Lessons learned from a rough-and-ready human/multi-robot team. Lecture Notes in Computer Science, 7068:232–251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27216-5_15/.

Scerri:2012:FDM

[1258] Paul Scerri, Balajee Kannan, Pras Velagapudi, Kate Macarthur, Peter Stone, and Matt Taylor. Flood disaster mitigation: a real-world challenge problem for multi-agent unmanned surface vehicles. Lecture Notes in Computer Science, 7068:252–269, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27216-5_16/.

Sadeh-Or:2012:AFL

[1259] Eran Sadeh-Or and Gal A. Kaminka. AnySURF: Flexible local features computation. Lecture Notes in Computer Science, 7068:270–271, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-27216-5_17.

Agmon:2012:RNW

[1260] Noa Agmon, Yehuda Elmaliah, Yaron Mor, and Oren Slor. Robot navigation with weak sensors. Lecture Notes in Computer Science, 7068:272– 276, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27216-5_18/.

Mesbah:2012:ILT

[1261] Anousha Mesbah and Prashant Doshi. Individual localization and tracking in multi-robot settings with dynamic landmarks. Lecture Notes in Computer Science, 7068:277–280, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27216-5_19/.

Keidar:2012:FFD

[1262] Matan Keidar, Eran Sadeh-Or, and Gal A. Kaminka. Fast frontier detection for robot exploration. Lecture Notes in Computer Science, 7068:281– 294, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27216-5_20/.

Calliess:2012:LAM

[1263] Jan-P. Calliess, Daniel Lyons, and Uwe D. Hanebeck. Lazy auctions for multi-robot collision avoidance and motion control under uncertainty. Lecture Notes in Computer Science, 7068:295–312, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27216-5_21.

Wang:2012:MRP

[1264] Wenjie Wang and Wooi-Boon Goh. Multi-robot path planning with the spatio-temporal A∗ algorithm and its variants. Lecture Notes in Computer Science, 7068:313–329, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27216-5_22.

Dasgupta:2012:AMR

[1265] Prithviraj Dasgupta, Ke Cheng, and Bikramjit Banerjee. Adaptive multirobot team reconfiguration using a policy-reuse reinforcement learning approach. Lecture Notes in Computer Science, 7068:330–345, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27216-5_23.

Korsah:2012:BOC

[1266] G. Ayorkor Korsah, Anthony Stentz, and M. Bernardine Dias. Bounded optimal constrained coordination with delay penalties and location choice. Lecture Notes in Computer Science, 7068:346–349, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27216-5_24.

Imoto:2012:MHB

[1267] Tomoko Imoto, Shin’ya Nakano, and Tomoyuki Higuchi. Modeling human behavior selection under environmental subsidy policy by multiagent simulation. Lecture Notes in Computer Science, 7068:350–358, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27216-5_25.

Cheng:2012:TMS

[1268] Shih-Fen Cheng and Thi Duong Nguyen. TaxiSim: a multiagent simulation platform for evaluating taxi fleet operations. Lecture Notes in Computer Science, 7068:359–360, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27216-5_26.

Saito:2012:PAB

[1269] Masaya M. Saito, Seiya Imoto, Rui Yamaguchi, Satoru Miyano, and Tomoyuki Higuchi. Parallel agent-based simulator for influenza pandemic. Lecture Notes in Computer Science, 7068:361–370, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27216-5_27.

Anh:2012:HMM

[1270] Nguyen Thi Ngoc Anh, Zucker Jean Daniel, Nguyen Huu Du, Alexis Drogoul, and Vo Duc An. A hybrid macro-micro pedestrians evacuation model to speed up simulation in road networks. Lecture Notes in Computer Science, 7068:371–383, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27216-5_28.

Kobayashi:2012:UAB

[1271] Tomomi Kobayashi, Satoshi Takahashi, Masaaki Kunigami, Atsushi Yoshikawa, and Takao Terano. A unified agent-based model to analyze organizational deviation and Kaizen activities. Lecture Notes in Computer Science, 7068:384–395, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27216-5_29.

Laclavik:2012:ABS

[1272] Michal Laclavík, Štefan Dlugolinský, Martin Šeleng, Marcel Kvassay, and Bernhard Schneider. Agent-based simulation platform evaluation in the context of human behavior modeling. Lecture Notes in Computer Science, 7068:396–410, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27216-5_30.

Such:2012:AIP

[1273] Jose M. Such, Agustin Espinosa, and Ana Garcia-Fornes. An agent infrastructure for privacy-enhancing agent-based E-commerce applications. Lecture Notes in Computer Science, 7068:411–425, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27216-5_31.

Centeno:2012:AAO

[1274] Roberto Centeno and Holger Billhardt. Auto-adaptation of open MAS through on-line modifications of the environment. Lecture Notes in Computer Science, 7068:426–427, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27216-5_32.

Sensoy:2012:CSW

[1275] Murat Şensoy, Wamberto W. Vasconcelos, and Timothy J. Norman. Combining semantic Web and logic programming for agent reasoning. Lecture Notes in Computer Science, 7068:428–441, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27216-5_33.

Alberola:2012:CAR

[1276] Juan M. Alberola, Vicente Julian, and Ana Garcia-Fornes. Cost-aware reorganization service for multiagent systems. Lecture Notes in Computer Science, 7068:442–456, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27216-5_34.

Criado:2012:DAE

[1277] Natalia Criado, Estefania Argente, Pablo Noriega, and Vicent Botti. A distributed architecture for enforcing norms in open MAS. Lecture Notes in Computer Science, 7068:457–471, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27216-5_35.

Sensoy:2012:ESA

[1278] Murat Şensoy. Evolving semantics for agent-based collaborative search. Lecture Notes in Computer Science, 7068:472–487, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-27216-5_36.

Frantz:2012:MAA

[1279] Christopher Frantz, Mariusz Nowostawski, and Martin K. Purvis. Microagents on Android: Interfacing agents with mobile applications. Lecture Notes in Computer Science, 7068:488–502, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-27216-5_37.

Oh:2012:IPN

[1280] Jean Oh, Felipe Meneguzzi, Katia Sycara, and Timothy J. Norman. Introduction to prognostic normative reasoning. Lecture Notes in Computer Science, 7068:503–504, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27216-5_38.

Anonymous:2012:BMbk

[1281] Anonymous. Back matter. Lecture Notes in Computer Science, 7068: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27216-5/1.

Anonymous:2012:FMbt

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27216-5/1.

Depeursinge:2012:OSW

[1283] Adrien Depeursinge, Hayit Greenspan, Tanveer Syeda, and Henning Müller. Overview of the second workshop on medical content–based retrieval for clinical decision support. Lecture Notes in Computer Science, 7075:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28460-1_1/.

Andre:2012:CBR

[1284] Barbara André, Tom Vercauteren, and Nicholas Ayache. Contentbased retrieval in endomicroscopy: Toward an efficient smart Atlas for clinical diagnosis. Lecture Notes in Computer Science, 7075:12–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28460-1_2/.

Rahman:2012:BIR

[1285] Md. Mahmudur Rahman, Sameer K. Antani, Dina Demner Fushman, and George R. Thoma. Biomedical image retrieval using multimodal context and concept feature spaces. Lecture Notes in Computer Science, 7075:24–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28460-1_3/.

Mata:2012:UME

[1286] Jacinto Mata, Mariano Crespo, and Manuel J. Maña. Using MeSH to expand queries in medical image retrieval. Lecture Notes in Computer Science, 7075:36–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28460-1_4/.

Pauly:2012:BID

[1287] Olivier Pauly, Diana Mateus, and Nassir Navab. Building implicit dictionaries based on extreme random clustering for modality recognition. Lecture Notes in Computer Science, 7075:47–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28460-1_5/.

Haas:2012:SBI

[1288] Sebastian Haas, René Donner, Andreas Burner, Markus Holzer, and Georg Langs. Superpixel-based interest points for effective bags of visual words medical image retrieval. Lecture Notes in Computer Science, 7075: 58–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28460-1_6/.

Foncubierta-Rodriguez:2012:UMV

[1289] Antonio Foncubierta-Rodríguez, Adrien Depeursinge, and Henning Müller. Using multiscale visual words for lung texture classification and retrieval. Lecture Notes in Computer Science, 7075:69–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28460-1_7/.

Vanegas:2012:HII

[1290] Jorge A. Vanegas, Juan C. Caicedo, Fabio A. González, and Eduardo Romero. Histology image indexing using a non-negative semantic embedding. Lecture Notes in Computer Science, 7075:80–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28460-1_8/.

Costa:2012:DDL

[1291] María Jimena Costa, Alexey Tsymbal, Matthias Hammon, and Alexander Cavallaro. A discriminative distance learning–based CBIR framework for characterization of indeterminate liver lesions. Lecture Notes in Computer Science, 7075:92–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28460-1_9/.

Safi:2012:CAD

[1292] Asad Safi, Maximilian Baust, Olivier Pauly, Victor Castaneda, and Tobias Lasser. Computer-aided diagnosis of pigmented skin dermoscopic images. Lecture Notes in Computer Science, 7075:105–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28460-1_10/.

Burner:2012:TBA

[1293] Andreas Burner, René Donner, Marius Mayerhoefer, Markus Holzer, and Franz Kainberger. Texture bags: Anomaly retrieval in medical images based on local 3D-texture similarity. Lecture Notes in Computer Science, 7075:116–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28460-1_11/.

Donner:2012:EFM

[1294] René Donner, Sebastian Haas, Andreas Burner, Markus Holzer, Horst Bischof, and Georg Langs. Evaluation of fast 2D and 3D medical image retrieval approaches based on image miniatures. Lecture Notes in Computer Science, 7075:128–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28460-1_12/.

Venkatraghavan:2012:SAA

[1295] Vikram Venkatraghavan and Sohan Ranjan. Semantic analysis of 3D anatomical medical images for sub-image retrieval. Lecture Notes in Computer Science, 7075:139–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28460-1_13/.

Anonymous:2012:BMbl

[1296] Anonymous. Back matter. Lecture Notes in Computer Science, 7075: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28460-1/1.

Anonymous:2012:FMbu

[1297] Anonymous. Front matter. Lecture Notes in Computer Science, 7075: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28460-1/1.

Zhou:2012:UDM

[1298] Zhi-Hua Zhou. Unlabeled data and multiple views. Lecture Notes in Computer Science, 7081:1–7, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28258-4_1/.

Plumpton:2012:OSS

[1299] Catrin O. Plumpton. Online semi-supervised ensemble updates for fMRI data. Lecture Notes in Computer Science, 7081:8–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28258-4_2/.

Esparza:2012:SSA

[1300] José Esparza, Stefan Scherer, and Friedhelm Schwenker. Studying self- and active-training methods for multi-feature set emotion recognition. Lecture Notes in Computer Science, 7081:19–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28258-4_3/.

Loog:2012:SSL

[1301] Marco Loog. Semi-supervised linear discriminant analysis using moment constraints. Lecture Notes in Computer Science, 7081:32–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28258-4_4/.

Yoshiyama:2012:MRM

[1302] Kazuki Yoshiyama and Akito Sakurai. Manifold-regularized minimax probability machine. Lecture Notes in Computer Science, 7081:42–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28258-4_5/.

Castelli:2012:SUC

[1303] Ilaria Castelli and Edmondo Trentin. Supervised and unsupervised cotraining of adaptive activation functions in neural nets. Lecture Notes in Computer Science, 7081:52–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28258-4_6/.

Castelli:2012:SUW

[1304] Ilaria Castelli and Edmondo Trentin. Semi-unsupervised weighted maximum-likelihood estimation of joint densities for the co-training of adaptive activation functions. Lecture Notes in Computer Science, 7081: 62–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28258-4_7/.

Fausser:2012:SSK

[1305] Stefan Faußer and Friedhelm Schwenker. Semi-supervised kernel clustering with sample-to-cluster weights. Lecture Notes in Computer Science, 7081:72–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28258-4_8/.

Smith:2012:HRL

[1306] Simón C. Smith and J. Michael Herrmann. Homeokinetic reinforcement learning. Lecture Notes in Computer Science, 7081:82–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28258-4_9/.

Soullard:2012:IRH

[1307] Yann Soullard and Thierry Artieres. Iterative refinement of HMM and HCRF for sequence classification. Lecture Notes in Computer Science, 7081:92–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28258-4_10/.

Lausser:2012:UPL

[1308] Ludwig Lausser, Florian Schmid, and Hans A. Kestler. On the utility of partially labeled data for classification of microarray data. Lecture Notes in Computer Science, 7081:96–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28258-4_11/.

Muller:2012:MIM

[1309] Andreas Müller and Sven Behnke. Multi-instance methods for partially supervised image segmentation. Lecture Notes in Computer Science, 7081:110–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28258-4_12/.

Hillebrand:2012:SST

[1310] Matthias Hillebrand, Christian Wöhler, Ulrich Kreßel, and Franz Kummert. Semi-supervised training set adaption to unknown countries for traffic sign classifiers. Lecture Notes in Computer Science, 7081:120– 127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28258-4_13/.

Trentin:2012:CCP

[1311] Edmondo Trentin, Luca Lusnig, and Fabio Cavalli. Comparison of combined probabilistic connectionist models in a forensic application. Lecture Notes in Computer Science, 7081:128–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28258-4_14/.

Schels:2012:CES

[1312] Martin Schels, Markus Kächele, David Hrabal, Steffen Walter, and Harald C. Traue. Classification of emotional states in a Woz scenario exploiting labeled and unlabeled bio-physiological data. Lecture Notes in Computer Science, 7081:138–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28258-4_15/.

Cooper:2012:USO

[1313] Cameron Cooper and Robert Kilmer. Using self organizing maps to find good comparison universities. Lecture Notes in Computer Science, 7081:148–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28258-4_16/.

Popescu:2012:SWP

[1314] Doru Anastasiu Popescu. Sink Web pages in Web application. Lecture Notes in Computer Science, 7081:154–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28258-4_17/.

Anonymous:2012:BMbm

[1315] Anonymous. Back matter. Lecture Notes in Computer Science, 7081: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28258-4/1.

Anonymous:2012:FMbv

[1316] Anonymous. Front matter. Lecture Notes in Computer Science, 7081: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28258-4/1.

Petriu:2012:MBP

[1317] Dorina C. Petriu. Model-based performance analysis of serviceoriented systems. Lecture Notes in Computer Science, 7083:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-25264-8_1.

LeSergent:2012:SCF

[1318] Thierry Le Sergent. SCADE: a comprehensive framework for critical system and software engineering. Lecture Notes in Computer Science, 7083:2–3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-25264-8_2.

Reed:2012:SBR

[1319] Rick Reed. SDL-2010: Background, rationale, and survey. Lecture Notes in Computer Science, 7083:4–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25264-8_3/.

Perrotin:2012:TRT

[1320] Maxime Perrotin, Eric Conquet, Julien Delange, André Schiele, and Thanassis Tsiodras. TASTE: a real-time software engineering tool-chain overview, status, and future. Lecture Notes in Computer Science, 7083: 26–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25264-8_4/.

Svendsen:2012:SSM

[1321] Andreas Svendsen, Øystein Haugen, and Birger Møller-Pedersen. Synthesizing software models: Generating train station models automatically. Lecture Notes in Computer Science, 7083:38–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25264-8_5/.

Hassine:2012:EEA

[1322] Jameleddine Hassine and Abdelouahed Gherbi. Exploring early availability requirements using use case maps. Lecture Notes in Computer Science, 7083:54–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25264-8_6/.

Nassiet:2012:PRU

[1323] Didier Nassiet, Yohan Livet, Marc Palyart, and David Lugato. Paprika: Rapid UI development of scientific dataset Editors for high performance computing. Lecture Notes in Computer Science, 7083:69–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25264-8_7/.

Baranov:2012:VCC

[1324] Sergey Baranov, Vsevolod Kotlyarov, and Thomas Weigert. Verifiable coverage criteria for automated testing. Lecture Notes in Computer Science, 7083:79–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25264-8_8/.

Wu-Hen-Chang:2012:NAM

[1325] Antal Wu-Hen-Chang, Gusztáv Adamis, Levente Erős, Gábor Kovács, and Tibor Csöndes. A new approach in model-based testing: Designing test models in TTCN-3. Lecture Notes in Computer Science, 7083:90– 105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25264-8_9/.

Mussa:2012:TMB

[1326] Mohamed Mussa and Ferhat Khendek. Towards a model based approach for integration testing. Lecture Notes in Computer Science, 7083:106– 121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25264-8_10/.

Fatima:2012:SIS

[1327] Urooj Fatima, Rolv Bræk, and Humberto Nicolás Castejón. Session initiation as a service. Lecture Notes in Computer Science, 7083:122– 137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25264-8_11/.

Sulistyo:2012:PPM

[1328] Selo Sulistyo and Andreas Prinz. PMG-pro: a model-driven development method of service-based applications. Lecture Notes in Computer Science, 7083:138–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25264-8_12/.

Kathayat:2012:MDF

[1329] Surya Bahadur Kathayat, Hien Nam Le, and Rolv Bræk. A modeldriven framework for component-based development. Lecture Notes in Computer Science, 7083:154–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25264-8_13/.

Cottenier:2012:SCT

[1330] Thomas Cottenier, Aswin van den Berg, and Thomas Weigert. Separation of concerns with transactional regions. Lecture Notes in Computer Science, 7083:168–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25264-8_14/.

Kramer:2012:RTS

[1331] Marc Krämer, Tobias Braun, Dennis Christmann, and Reinhard Gotzhein. Real-time signaling in SDL. Lecture Notes in Computer Science, 7083:186–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25264-8_15/.

Christmann:2012:PSS

[1332] Dennis Christmann, Philipp Becker, and Reinhard Gotzhein. Priority scheduling in SDL. Lecture Notes in Computer Science, 7083:202–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25264-8_16/.

Kraas:2012:MBF

[1333] Alexander Kraas. A model-based formalization of the textual notation for SDL–UML. Lecture Notes in Computer Science, 7083:218–232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25264-8_17/.

Fleurey:2012:SVC

[1334] Franck Fleurey, Øystein Haugen, Birger Møller-Pedersen, Andreas Svendsen, and Xiaorui Zhang. Standardizing variability — challenges and solutions. Lecture Notes in Computer Science, 7083:233–246, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25264-8_18/.

Palyart:2012:MAU

[1335] Marc Palyart, David Lugato, Ileana Ober, and Jean-Michel Bruel. MDE4HPC: An approach for using model-driven engineering in highperformance computing. Lecture Notes in Computer Science, 7083:247– 261, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25264-8_19/.

Anssi:2012:AVM

[1336] Saoussen Anssi, Sébastien Gérard, Stefan Kuntz, and François Terrier. AUTOSAR vs. MARTE for enabling timing analysis of automotive applications. Lecture Notes in Computer Science, 7083:262–275, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25264-8_20/.

Anonymous:2012:BMbn

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-25264-8/1.

Anonymous:2012:FMbw

[1338] Anonymous. Front matter. Lecture Notes in Computer Science, 7083: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-25264-8/1.

Pop:2012:ECS

[1339] Mihaela Pop, Maxime Sermesant, Tommaso Mansi, Eugene Crystal, Sudip Ghate, and Jatin Relan. EP challenge — STACOM’11: Forward approaches to computational electrophysiology using MRI-based models and in-vivo CARTO mapping in swine hearts. Lecture Notes in Computer Science, 7085:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28326-0_1/.

Relan:2012:PVE

[1340] Jatin Relan, Maxime Sermesant, Hervé Delingette, and Nicholas Ayache. Personalisation of a 3D ventricular electrophysiological model, using endocardial and epicardial contact mapping and MRI. Lecture Notes in Computer Science, 7085:14–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28326-0_2/.

Wang:2012:TES

[1341] Linwei Wang, Fady Dawoud, Ken C. L. Wong, Heye Zhang, Huafeng Liu, and Albert C. Lardo. Transmural electrophysiologic and scar imaging on porcine heart with chronic infarction. Lecture Notes in Computer Science, 7085:23–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28326-0_3/.

Tobon-Gomez:2012:MDC

[1342] Catalina Tobon-Gomez, Mathieu De Craene, Annette Dahl, and Stam Kapetanakis. A multimodal database for the 1^{st} cardiac motion analysis challenge. Lecture Notes in Computer Science, 7085:33–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28326-0_4/.

Wang:2012:ACM

[1343] Haiyan Wang, Wenzhe Shi, Xiahai Zhuang, Simon Duckett, KaiPin Tung, and Philip Edwards. Automatic cardiac motion tracking using both untagged and 3D tagged MR images. Lecture Notes in Computer Science, 7085:45–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28326-0_5/.

McLeod:2012:ILD

[1344] Kristin McLeod, Adityo Prakosa, Tommaso Mansi, Maxime Sermesant, and Xavier Pennec. An incompressible log-domain demons algorithm for tracking heart tissue. Lecture Notes in Computer Science, 7085:55– 67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28326-0_6/.

DeCraene:2012:TDF

[1345] Mathieu De Craene, Catalina Tobon-Gomez, Constantine Butakoff, and Nicolas Duchateau. Temporal diffeomorphic free form deformation (TDFFD) applied to motion and deformation quantification of tagged MRI sequences. Lecture Notes in Computer Science, 7085:68– 77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28326-0_7/.

Tautz:2012:MAQ

[1346] Lennart Tautz, Anja Hennemuth, and Heinz-Otto Peitgen. Motion analysis with quadrature filter based registration of tagged MRI sequences. Lecture Notes in Computer Science, 7085:78–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28326-0_8/.

Suinesiaputra:2012:LVS

[1347] Avan Suinesiaputra, Brett R. Cowan, J. Paul Finn, Carissa G. Fonseca, and Alan H. Kadish. Left ventricular segmentation challenge from cardiac MRI: a collation study. Lecture Notes in Computer Science, 7085: 88–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28326-0_9/.

Jolly:2012:ASM

[1348] Marie-Pierre Jolly, Christoph Guetter, Xiaoguang Lu, Hui Xue, and Jens Guehring. Automatic segmentation of the myocardium in cine MR images using deformable registration. Lecture Notes in Computer Science, 7085:98–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28326-0_10/.

Margeta:2012:LST

[1349] Ján Margeta, Ezequiel Geremia, Antonio Criminisi, and Nicholas Ayache. Layered spatio-temporal forests for left ventricle segmentation from 4D cardiac MRI data. Lecture Notes in Computer Science, 7085:109– 119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28326-0_11/.

Fahmy:2012:MSU

[1350] Ahmed S. Fahmy, Ahmed O. Al-Agamy, and Ayman Khalifa. Myocardial segmentation using contour-constrained optical flow tracking. Lecture Notes in Computer Science, 7085:120–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28326-0_12/.

Ulen:2012:OMR

[1351] Johannes Ulén, Petter Strandmark, and Fredrik Kahl. Optimization for multi-region segmentation of cardiac MRI. Lecture Notes in Computer Science, 7085:129–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28326-0_13/.

Truong:2012:ACB

[1352] Michael Truong, Thomas Gordon, Reza Razavi, Graeme Penney, and Kawal S. Rhode. Analysis of catheter-based registration with vesselradius weighting of 3D CT data to 2D X-ray for cardiac catheterisation procedures in a phantom study. Lecture Notes in Computer Science, 7085:139–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28326-0_14/.

Wang:2012:MCR

[1353] Vicky Y. Wang, Daniel B. Ennis, Brett R. Cowan, Alistair A. Young, and Martyn P. Nash. Myocardial contractility and regional work throughout the cardiac cycle using FEM and MRI. Lecture Notes in Computer Science, 7085:149–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28326-0_15/.

Lombaert:2012:VHC

[1354] Hervé Lombaert, Jean-Marc Peyrat, Laurent Fanton, Farida Cheriet, and Hervé Delingette. Variability of the Human Cardiac Laminar Structure. Lecture Notes in Computer Science, 7085:160–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28326-0_16/.

Dikici:2012:PRB

[1355] Engin Dikici and Fredrik Orderud. Polynomial regression based edge filtering for left ventricle tracking in 3D echocardiography. Lecture Notes in Computer Science, 7085:168–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28326-0_17/.

Shi:2012:MIG

[1356] Wenzhe Shi, Xiahai Zhuang, Robin Wolz, Duckett Simon, KaiPin Tung, and Haiyan Wang. A multi-image graph cut approach for cardiac image segmentation and uncertainty estimation. Lecture Notes in Computer Science, 7085:178–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28326-0_18/.

Rahimi:2012:TCF

[1357] Azar Rahimi, Hongda Mao, Pengcheng Shi, and Linwei Wang. Toward clinically-feasible noninvasive electrophysiological imaging: Investigating the impact of local anatomical details. Lecture Notes in Computer Science, 7085:188–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28326-0_19/.

Elen:2012:TST

[1358] An Elen, Jeroen Hermans, Hadewich Hermans, Frederik Maes, and Paul Suetens. A 3D+time spatio-temporal model for joint segmentation and registration of sparse cardiac cine MR image stacks. Lecture Notes in Computer Science, 7085:198–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28326-0_20/.

Lombaert:2012:SAH

[1359] Hervé Lombaert, Jean-Marc Peyrat, Laurent Fanton, Farida Cheriet, and Hervé Delingette. Statistical Atlas of human cardiac fibers: Comparison with abnormal hearts. Lecture Notes in Computer Science, 7085:207–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
28326-0_21.

Medrano-Gracia:2012:MLC

[1360] Pau Medrano-Gracia, David A. Bluemke, Brett R. Cowan, J. Paul Finn, and Carissa G. Fonseca. Maximum likelihood correction of shape bias arising from imaging protocol: Application to cardiac MRI. Lecture Notes in Computer Science, 7085:214–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-28326-0_22.

Mao:2012:VME

[1361] Hongda Mao, Linwei Wang, Ken C. L. Wong, Huafeng Liu, and Pengcheng Shi. Volumetric modeling electromechanics of the heart. Lecture Notes in Computer Science, 7085:224–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-28326-0_23.

Ardekani:2012:MSS

[1362] Siamak Ardekani, Aastha Jain, Saurabh Jain, Theodore P. Abraham, and Maria R. Abraham. Matching sparse sets of cardiac image crosssections using large deformation diffeomorphic metric mapping algorithm. Lecture Notes in Computer Science, 7085:234–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-28326-0_24.

Radau:2012:VVP

[1363] Perry E. Radau, Stefan Pintilie, Roey Flor, Labonny Biswas, and Samuel O. Oduneye. VURTIGO: Visualization platform for real-time, MRI-guided cardiac electroanatomic mapping. Lecture Notes in Computer Science, 7085:244–253, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-28326-0_25.

Karim:2012:VNM

[1364] Rashed Karim, Aruna Arujuna, Alex Brazier, Jaswinder Gill, and C. Aldo Rinaldi. Validation of a novel method for the automatic segmentation of left atrial scar from delayed-enhancement magnetic resonance. Lecture Notes in Computer Science, 7085:254–262, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-28326-0_26.

Becciu:2012:CME

[1365] Alessandro Becciu, Remco Duits, Bart J. Janssen, Luc M. J. Florack, and Hans C. van Assen. Cardiac motion estimation using covariant derivatives and Helmholtz decomposition. Lecture Notes in Computer Science, 7085:263–273, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-28326-0_27.

Zhang:2012:TDM

[1366] Zhijun Zhang, David J. Sahn, and Xubo Song. Temporal diffeomorphic motion analysis from echocardiographic sequences by using intensity transitivity consistency. Lecture Notes in Computer Science, 7085:274–284, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-28326-0_28.

Anonymous:2012:BMbo

[1367] Anonymous. Back matter. Lecture Notes in Computer Science, 7085: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28326-0/1.

Anonymous:2012:FMbx

[1368] Anonymous. Front matter. Lecture Notes in Computer Science, 7085: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28326-0/1.

Ghimire:2012:NTF

[1369] Deepak Ghimire and Joonwhoan Lee. Nonlinear transfer functionbased image detail preserving dynamic range compression for color image enhancement. Lecture Notes in Computer Science, 7087:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25367-6_1/.

Gil:2012:PAS

[1370] Jong In Gil, Seung Eun Jang, and Manbae Kim. 3D perception adjustment of stereoscopic images based upon depth map. Lecture Notes in Computer Science, 7087:13–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25367-6_2/.

Takahashi:2012:SRF

[1371] Keita Takahashi and Takeshi Naemura. Super-resolved free-viewpoint image synthesis using semi-global depth estimation and depth-reliability-based regularization. Lecture Notes in Computer Science, 7087:22–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25367-6_3/.

Kim:2012:HKS

[1372] Seung-Goo Kim, Moo K. Chung, and Seongho Seo. Heat kernel smoothing via Laplace–Beltrami eigenfunctions and its application to subcortical structure modeling. Lecture Notes in Computer Science, 7087:36– 47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25367-6_4/.

Lin:2012:SNI

[1373] Shang-Yen Lin and Yung-Chang Chen. SLAM and navigation in indoor environments. Lecture Notes in Computer Science, 7087:48–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25367-6_5/.

Muthukudage:2012:CBS

[1374] Jayantha Muthukudage and JungHwan Oh. Color based stool region detection in colonoscopy videos for quality measurements. Lecture Notes in Computer Science, 7087:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25367-6_6/.

Nguyen:2012:IME

[1375] Duc Dung Nguyen and Jae Wook Jeon. Improving motion estimation using image-driven functions and hybrid scheme. Lecture Notes in Computer Science, 7087:73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25367-6_7/.

Nguyen:2012:RTB

[1376] Thuy Tuong Nguyen and Jae Wook Jeon. Real-time background compensation for PTZ cameras using GPU accelerated and range-limited genetic algorithm search. Lecture Notes in Computer Science, 7087:85– 96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25367-6_8/.

Komai:2012:AVS

[1377] Yuto Komai, Yasuo Ariki, and Tetsuya Takiguchi. Audio-visual speech recognition based on AAM parameter and phoneme analysis of visual feature. Lecture Notes in Computer Science, 7087:97–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25367-6_9/.

Saracchini:2012:MSI

[1378] Rafael F. V. Saracchini and Jorge Stolfi. Multi-scale integration of slope data on an irregular mesh. Lecture Notes in Computer Science, 7087:109– 120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25367-6_10/.

Shin:2012:VVD

[1379] In-Yong Shin and Yo-Sung Ho. Virtual viewpoint disparity estimation and convergence check for real-time view synthesis. Lecture Notes in Computer Science, 7087:121–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25367-6_11/.

Yao:2012:SFI

[1380] Anbang Yao and Shan Yu. Spatial feature interdependence matrix (SFIM): a robust descriptor for face recognition. Lecture Notes in Computer Science, 7087:132–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25367-6_12/.

Chiang:2012:CDM

[1381] Jui-Chiu Chiang, Chun-Hung Chen, and Wen-Nung Lie. Coding of dynamic 3D mesh model for 3D video transmission. Lecture Notes in Computer Science, 7087:144–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25367-6_13/.

Hess-Flores:2012:RDB

[1382] Mauricio Hess-Flores and Daniel Knoblauch. Ray divergence-based bundle adjustment conditioning for multi-view stereo. Lecture Notes in Computer Science, 7087:153–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25367-6_14/.

Hosni:2012:TCD

[1383] Asmaa Hosni, Christoph Rhemann, and Michael Bleyer. Temporally consistent disparity and optical flow via efficient spatio-temporal filtering. Lecture Notes in Computer Science, 7087:165–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25367-6_15/.

Migita:2012:SFR

[1384] Tsuyoshi Migita and Kazuhiro Sogawa. Specular-free residual minimization for photometric stereo with unknown light sources. Lecture Notes in Computer Science, 7087:178–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25367-6_16/.

May:2012:AFP

[1385] Michael May, Martin Turner, and Tim Morris. Analysing false positives and 3D structure to create intelligent thresholding and weighting functions for SIFT features. Lecture Notes in Computer Science, 7087:190– 201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25367-6_17/.

Jawed:2012:VAS

[1386] Khurram Jawed and John Morris. Verging axis stereophotogrammetry. Lecture Notes in Computer Science, 7087:202–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25367-6_18/.

Wang:2012:MWF

[1387] Sheng Wang, Qiang Wu, Xiangjian He, and Wenjing Jia. More on weak feature: Self-correlate histogram distances. Lecture Notes in Computer Science, 7087:214–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25367-6_19/.

Hermann:2012:MLS

[1388] Simon Hermann, Anko Börner, and Reinhard Klette. Mid-level segmentation and segment tracking for long-range stereo analysis. Lecture Notes in Computer Science, 7087:224–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25367-6_20/.

Adluru:2012:AER

[1389] Nagesh Adluru, Moo K. Chung, and Nicholas T. Lange. Applications of epsilon radial networks in neuroimage analyses. Lecture Notes in Computer Science, 7087:236–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25367-6_21.

Kang:2012:RIS

[1390] Yousun Kang, Koichiro Yamaguchi, and Takashi Naito. Road image segmentation and recognition using hierarchical bag-of-textons method. Lecture Notes in Computer Science, 7087:248–256, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25367-6_22.

Ling:2012:SHS

[1391] Huo-Chong Ling and Raphael C.-W. Phan. On the security of a hybrid SVD–DCT watermarking method based on LPSNR. Lecture Notes in Computer Science, 7087:257–266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25367-6_23.

Heo:2012:IEC

[1392] Jin Heo and Yo-Sung Ho. Improved entropy coder in H.264/AVC for lossless residual coding in the spatial domain. Lecture Notes in Computer Science, 7087:267–276, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25367-6_24.

Yamada:2012:APE

[1393] Kentaro Yamada, Yusuke Sugano, and Takahiro Okabe. Attention prediction in egocentric video using motion and visual saliency. Lecture Notes in Computer Science, 7087:277–288, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25367-6_25.

May:2012:FME

[1394] Michael May, Martin Turner, and Tim Morris. FAW for multi-exposure fusion features. Lecture Notes in Computer Science, 7087:289–300, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25367-6_26.

Kang:2012:ESI

[1395] Yun-Suk Kang and Yo-Sung Ho. Efficient stereo image rectification method using horizontal baseline. Lecture Notes in Computer Science, 7087:301–310, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25367-6_27.

deSouza:2012:RTI

[1396] Rafael Henrique Castanheira de Souza. Real-time image mosaicing using non-rigid registration. Lecture Notes in Computer Science, 7087:311–322, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25367-6_28.

Pham:2012:AGI

[1397] Cuong Cao Pham, Synh Viet Uyen Ha, and Jae Wook Jeon. Adaptive guided image filtering for sharpness enhancement and noise reduction. Lecture Notes in Computer Science, 7087:323–334, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25367-6_29.

Matsui:2012:HSI

[1398] Shuhei Matsui and Hajime Nagahara. Half-sweep imaging for depth from defocus. Lecture Notes in Computer Science, 7087:335–347, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25367-6_30.

Yeh:2012:HAP

[1399] Mei-Chen Yeh and Jason Tai. A hierarchical approach to practical beverage package recognition. Lecture Notes in Computer Science, 7087: 348–357, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25367-6_31.

Sthitpattanapongsa:2012:EOT

[1400] Puthipong Sthitpattanapongsa and Thitiwan Srinark. An equivalent 3D Otsu’s thresholding method. Lecture Notes in Computer Science, 7087:358–369, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25367-6_32.

Xu:2012:HMT

[1401] Jianfeng Xu, Koichi Takagi, and Shigeyuki Sakazawa. Human motion tracking with monocular video by introducing a graph structure into Gaussian process dynamical models. Lecture Notes in Computer Science, 7087:370–383, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25367-6_33.

Lee:2012:DMS

[1402] Gyo-Yoon Lee and Yo-Sung Ho. Depth map up-sampling using random walk. Lecture Notes in Computer Science, 7087:384–394, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25367-6_34.

Hermann:2012:ENC

[1403] Simon Hermann and Reinhard Klette. Evaluation of a new coarseto-fine strategy for fast semi-global stereo matching. Lecture Notes in Computer Science, 7087:395–406, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25367-6_35.

Nakashima:2012:TAM

[1404] Ryo Nakashima, Keita Takahashi, and Takeshi Naemura. Theoretical analysis of multi-view camera arrangement and light-field superresolution. Lecture Notes in Computer Science, 7087:407–420, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25367-6_36.

Anonymous:2012:BMbp

[1405] Anonymous. Back matter. Lecture Notes in Computer Science, 7087: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-25367-6/1.

Anonymous:2012:FMby

[1406] Anonymous. Front matter. Lecture Notes in Computer Science, 7087: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-25367-6/1.

Park:2012:LIC

[1407] Jiho Park, Je Woo Kim, and Jechang Jeong. Lossless image coding based on inter-color prediction for ultra high definition image. Lecture Notes in Computer Science, 7088:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25346-1_1/.

Kim:2012:MAR

[1408] Yong-Hwan Kim, Jiho Park, and Je-Woo Kim. Multithreading architecture for real-time MPEG-4 AVC/H.264 SVC decoder. Lecture Notes in Computer Science, 7088:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25346-1_2/.

Yoon:2012:FMD

[1409] Da-Hyun Yoon and Yo-Sung Ho. Fast mode decision algorithm for depth coding in 3D video systems using H.264/AVC. Lecture Notes in Computer Science, 7088:25–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25346-1_3/.

Aranda:2012:IDB

[1410] Ramón Aranda and Mariano Rivera. Improved diffusion basis functions fitting and metric distance for brain axon fiber estimation. Lecture Notes in Computer Science, 7088:36–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25346-1_4/.

Zou:2012:AMD

[1411] Ruobing Zou, Oscar C. Au, Lin Sun, and Sijin Li. An adaptive motion data storage reduction method for temporal predictor. Lecture Notes in Computer Science, 7088:48–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25346-1_5/.

Pham:2012:LVB

[1412] Cuong Cao Pham, Synh Viet Uyen Ha, and Jae Wook Jeon. A local variance-based bilateral filtering for artifact-free detail- and edgepreserving smoothing. Lecture Notes in Computer Science, 7088:60– 70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25346-1_6/.

Tae-o-sot:2012:IGD

[1413] Sarawut Tae o sot and Akinori Nishihara. Iterative gradient-driven patch-based inpainting. Lecture Notes in Computer Science, 7088:71– 81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25346-1_7/.

Nosaka:2012:FEB

[1414] Ryusuke Nosaka, Yasuhiro Ohkawa, and Kazuhiro Fukui. Feature extraction based on co-occurrence of adjacent local binary patterns. Lecture Notes in Computer Science, 7088:82–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-25346-1_8/.

Wang:2012:NIC

[1415] Dong Wang, Weijia Jia, Guiqing Li, and Yunhui Xiong. Natural image composition with inhomogeneous boundaries. Lecture Notes in Computer Science, 7088:92–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25346-1_9/.

Seto:2012:DEL

[1416] Hiroyuki Seto and Tomoyuki Taguchi. Directional eigentemplate learning for sparse template tracker. Lecture Notes in Computer Science, 7088:104–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25346-1_10/.

Lin:2012:GIU

[1417] Shen-Ju Lin, Chung-Lin Huang, and Shih-Chung Hsu. Gender identification using feature patch-based Bayesian classifier. Lecture Notes in Computer Science, 7088:116–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25346-1_11/.

Chen:2012:MOT

[1418] Ke-Yin Chen, Chung-Lin Huang, and Shih-Chung Hsu. Multiple objects tracking across multiple non-overlapped views. Lecture Notes in Computer Science, 7088:128–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25346-1_12/.

Yang:2012:FHP

[1419] Zhuo Yang and Sei ichiro Kamata. Fast hypercomplex polar Fourier analysis for image processing. Lecture Notes in Computer Science, 7088:141– 148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25346-1_13/.

Du:2012:CLP

[1420] Weiwei Du, Shiya Mori, and Nobuyuki Nakamori. Colorization by landmark pixels extraction. Lecture Notes in Computer Science, 7088:149– 156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25346-1_14/.

Nguyen:2012:FBN

[1421] Tuan-Anh Nguyen and Min-Cheol Hong. Filtering-based noise estimation for denoising the image degraded by Gaussian noise. Lecture Notes in Computer Science, 7088:157–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25346-1_15/.

Habed:2012:CMC

[1422] Adlane Habed, Tarik Elamsy, and Boubakeur Boufama. Combining Mendonça–Cipolla self-calibration and scene constraints. Lecture Notes in Computer Science, 7088:168–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25346-1_16/.

Imaizumi:2012:KDS

[1423] Shoko Imaizumi, Masaaki Fujiyoshi, and Hitoshi Kiya. A key derivation scheme for hierarchical access control to JPEG 2000 coded images. Lecture Notes in Computer Science, 7088:180–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25346-1_17/.

Carrasco:2012:BMU

[1424] Miguel Carrasco and Domingo Mery. Bifocal matching using multiple geometrical solutions. Lecture Notes in Computer Science, 7088:192– 203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25346-1_18/.

Sim:2012:DHC

[1425] Jae-Young Sim. Digital hologram compression using correlation of reconstructed object images. Lecture Notes in Computer Science, 7088:204– 214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25346-1_19/.

Li:2012:PIS

[1426] Ke-Chun Li, Hong-Ren Su, and Shang-Hong Lai. Pedestrian image segmentation via shape-prior constrained random walks. Lecture Notes in Computer Science, 7088:215–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25346-1_20/.

Zhu:2012:NRC

[1427] Jiangying Zhu, Mei Yu, Qiaoyan Zheng, and Zongju Peng. A novel rate control algorithm for H.264/AVC based on human visual system. Lecture Notes in Computer Science, 7088:227–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25346-1_21.

Yang:2012:BID

[1428] Hao-Liang Yang, Yen-Hao Chiao, and Po-Hao Huang. Blind image deblurring with modified Richardson–Lucy deconvolution for ringing artifact suppression. Lecture Notes in Computer Science, 7088:240–251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25346-1_22.

Wang:2012:QEH

[1429] Ren-Jie Wang, Yan-Ting Jiang, and Jiunn-Tsair Fang. Quality estimation for H.264/SVC inter-layer residual prediction in spatial scalability. Lecture Notes in Computer Science, 7088:252–261, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-25346-1_23.

Kimura:2012:EID

[1430] Ryohei Kimura, Noriko Takemura, and Yoshio Iwai. Extracting interval distribution of human interactions. Lecture Notes in Computer Science, 7088:262–273, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25346-1_24.

Geng:2012:FML

[1431] Haokun Geng, James Russell, and Bok-Suk Shin. A flexible method for localisation and classification of footprints of small species. Lecture Notes in Computer Science, 7088:274–286, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25346-1_25.

Madrigal:2012:LRM

[1432] Francisco Madrigal and Mariano Rivera. Learning and regularizing motion models for enhancing particle filter-based target tracking. Lecture Notes in Computer Science, 7088:287–298, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25346-1_26.

Su:2012:CMI

[1433] Hong-Ren Su and Shang-Hong Lai. CT–MR image registration in 3D K-space based on Fourier moment matching. Lecture Notes in Computer Science, 7088:299–310, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25346-1_27.

Chew:2012:STR

[1434] S. W. Chew, R. Rana, P. Lucey, and S. Lucey. Sparse temporal representations for facial expression recognition. Lecture Notes in Computer Science, 7088:311–322, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25346-1_28.

Daribo:2012:DCC

[1435] Ismael Daribo, Ryo Furukawa, and Ryusuke Sagawa. Dynamic compression of curve-based point cloud. Lecture Notes in Computer Science, 7088:323–334, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25346-1_29.

Chen:2012:RDM

[1436] Hsiao-Wei Chen and Shang-Hong Lai. Recovering depth map from video with moving objects. Lecture Notes in Computer Science, 7088:335–346, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25346-1_30.

DinhQuoc:2012:IAE

[1437] Khanh DinhQuoc, Xiem HoangVan, and Byeungwoo Jeon. An iterative algorithm for efficient adaptive GOP size in transform domain Wyner– Ziv video coding. Lecture Notes in Computer Science, 7088:347–358, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25346-1_31.

Shakeri:2012:RZW

[1438] Mahsa Shakeri and Mansour Jamzad. A robust zero-watermark copyright protection scheme based on DWT and image normalization. Lecture Notes in Computer Science, 7088:359–370, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25346-1_32.

Oh:2012:MVV

[1439] Kwan-Jung Oh, Jaejoon Lee, and Du-Sik Park. Multi-view video coding based on high efficiency video coding. Lecture Notes in Computer Science, 7088:371–380, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-25346-1_33.

Lin:2012:ICB

[1440] Guo-Shiang Lin, Han-Wen Liu, and Wei-Chih Chen. 2D to 3D image conversion based on classification of background depth profiles. Lecture Notes in Computer Science, 7088:381–392, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-25346-1_34.

Wang:2012:SMR

[1441] Junwei Wang, Yu Zhou, Xiang Bai, and Wenyu Liu. Shape matching and recognition using group-wised points. Lecture Notes in Computer Science, 7088:393–404, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-25346-1_35.

Anonymous:2012:BMbq

[1442] Anonymous. Back matter. Lecture Notes in Computer Science, 7088: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-25346-1/1.

Anonymous:2012:FMbz

[1443] Anonymous. Front matter. Lecture Notes in Computer Science, 7088: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-25346-1/1.

Terziyan:2012:GUE

[1444] Vagan Terziyan. Global understanding environment: Towards selfmanaged Web of everything. Lecture Notes in Computer Science, 7096: 1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-27916-4_1.

Fugini:2012:SSR

[1445] MariaGrazia Fugini, Stefano Pinardi, and Claudia Raibulet. Smart solutions for risk prevention through analysis of people movements. Lecture Notes in Computer Science, 7096:3–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27916-4_2/.

Pantsar-Syvaniemi:2012:SSA

[1446] Susanna Pantsar-Syväniemi, Jarkko Kuusijärvi, and Eila Ovaska. Supporting situation-awareness in smart spaces. Lecture Notes in Computer Science, 7096:14–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27916-4_3/.

Stenudd:2012:MUM

[1447] Sakari Stenudd. A model for using machine learning in smart environments. Lecture Notes in Computer Science, 7096:24–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27916-4_4/.

Suomalainen:2012:FSD

[1448] Jani Suomalainen. Flexible security deployment in smart spaces. Lecture Notes in Computer Science, 7096:34–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27916-4_5/.

Niezen:2012:UST

[1449] Gerrit Niezen, Bram van der Vlist, Jun Hu, and Loe Feijs. Using semantic transformers to enable interoperability between media devices in a ubiquitous computing environment. Lecture Notes in Computer Science, 7096:44–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27916-4_6/.

Steenhuyse:2012:TKO

[1450] Maarten Steenhuyse, Jeroen Hoebeke, Ann Ackaert, Ingrid Moerman, and Piet Demeester. TV-kiosk: An open and extensible platform for the wellbeing of an ageing population. Lecture Notes in Computer Science, 7096:54–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27916-4_7/.

Li:2012:DEA

[1451] Geng Li, Yuping Zhao, Bingli Jiao, and Timo Korhonen. Design of easy access Internet browsing system for elderly people based on Android. Lecture Notes in Computer Science, 7096:64–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27916-4_8/.

Zhou:2012:EPS

[1452] Jiehan Zhou, Xiang Su, Mika Ylianttila, and Jukka Riekki. Exploring pervasive service computing opportunities for pursuing successful ageing. Lecture Notes in Computer Science, 7096:73–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27916-4_9/.

Korhonen:2012:ACR

[1453] Timo Korhonen, Xirui Wang, Shuo Liu, Christos Karaiskos, and Yuping Zhao. Alert calls in remote health: Cultural adaptation and usability inspection for China. Lecture Notes in Computer Science, 7096: 83–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27916-4_10/.

Merilahti:2012:EOP

[1454] Juho Merilahti, Juha Pärkkä, and Ilkka Korhonen. Estimating older people’s physical functioning with automated health monitoring technologies at home: Feature correlations and multivariate analysis. Lecture Notes in Computer Science, 7096:94–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27916-4_11/.

Mutafungwa:2012:AMF

[1455] Edward Mutafungwa. Applying MTC and femtocell technologies to the Continua Health Reference Architecture. Lecture Notes in Computer Science, 7096:105–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27916-4_12/.

Zhang:2012:DRT

[1456] Ye Zhang, Olli Martikainen, Petri Pulli, and Valeriy Naumov. Developing a real-time process data acquisition system for automatic process measurement. Lecture Notes in Computer Science, 7096:115–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27916-4_13/.

Pantsar-Syvaniemi:2012:ACA

[1457] Susanna Pantsar-Syväniemi. Adaptable context-aware micro-architecture. Lecture Notes in Computer Science, 7096:125–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27916-4_14/.

Oja:2012:UFC

[1458] Mika Oja and Jukka Riekki. Ubiquitous framework for creating and evaluating persuasive applications and games. Lecture Notes in Computer Science, 7096:133–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27916-4_15/.

Leppanen:2012:DDP

[1459] Teemu Leppänen and Jukka Riekki. Dynamic data processing middleware for sensor networks. Lecture Notes in Computer Science, 7096:141– 147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27916-4_16/.

Gilman:2012:SSM

[1460] Ekaterina Gilman and Jukka Riekki. Smart spaces: a metacognitive approach. Lecture Notes in Computer Science, 7096:148–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27916-4_17/.

Linden:2012:CFE

[1461] Tomas Lindén. A conceptual framework for enabling community-driven extensible, open and privacy-preserving ubiquitous computing networks. Lecture Notes in Computer Science, 7096:156–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27916-4_18/.

Salminen:2012:DCI

[1462] Arto Salminen. Doctoral colloquium: Integrating Web content into mashups on desktop and mobile devices. Lecture Notes in Computer Science, 7096:164–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27916-4_19/.

Polojarvi:2012:LSB

[1463] Mikko Polojärvi and Jukka Riekki. Lightweight service-based software architecture. Lecture Notes in Computer Science, 7096:172–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27916-4_20/.

Anonymous:2012:BMbr

[1464] Anonymous. Back matter. Lecture Notes in Computer Science, 7096: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27916-4/1.

Anonymous:2012:FMca

[1465] Anonymous. Front matter. Lecture Notes in Computer Science, 7096: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27916-4/1.

Sassi:2012:AQP

[1466] Minyar Sassi, Oussama Tlili, and Habib Ounelli. Approximate query processing for database flexible querying with aggregates. Lecture Notes in Computer Science, 7100:1–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28148-8_1/.

Vlachou:2012:MBS

[1467] Akrivi Vlachou and Christos Doulkeridis. Metric-based similarity search in unstructured peer-to-peer systems. Lecture Notes in Computer Science, 7100:28–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28148-8_2/.

Sabesan:2012:APQ

[1468] Manivasakan Sabesan and Tore Risch. Adaptive parallelization of queries to data providing Web service operations. Lecture Notes in Computer Science, 7100:49–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28148-8_3/.

Tian:2012:PBA

[1469] Yuan Tian, Haofen Wang, Wei Jin, and Yuan Ni. A pattern-based approach for efficient query processing over RDF data. Lecture Notes in Computer Science, 7100:70–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28148-8_4/.

Manna:2012:HSS

[1470] Marco Manna, Ermelinda Oro, and Massimo Ruffolo. The HiLϵX system for semantic information extraction. Lecture Notes in Computer Science, 7100:91–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28148-8_5/.

Hedeler:2012:DAF

[1471] Cornelia Hedeler, Khalid Belhajjame, and Lu Mao. DSToolkit: An architecture for flexible dataspace management. Lecture Notes in Computer Science, 7100:126–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28148-8_6/.

Atzeni:2012:TCM

[1472] Paolo Atzeni and Pierluigi Del Nostro. Temporal content management and Website modeling: Putting them together. Lecture Notes in Computer Science, 7100:158–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28148-8_7/.

Stahl:2012:HHD

[1473] Frederic Stahl and Mohamed Medhat Gaber. Homogeneous and heterogeneous distributed classification for pocket data mining. Lecture Notes in Computer Science, 7100:183–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28148-8_8/.

Frank:2012:IDM

[1474] Lars Frank and Rasmus Ulslev Pedersen. Integrated distributed/mobile logistics management. Lecture Notes in Computer Science, 7100:206– 221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28148-8_9/.

Anonymous:2012:BMbs

[1475] Anonymous. Back matter. Lecture Notes in Computer Science, 7100: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28148-8/1.

Anonymous:2012:FMcb

[1476] Anonymous. Front matter. Lecture Notes in Computer Science, 7100: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28148-8/1.

Anonymous:2012:FMcc

[1477] Anonymous. Front matter. Lecture Notes in Computer Science, 7103: 1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27609-5/1/1.

Stone:2012:IFU

[1478] Peter Stone. Intersections of the future: Using fully autonomous vehicles. Lecture Notes in Computer Science, 7103:3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-27609-5_1.

Sklar:2012:TMA

[1479] Elizabeth Sklar, Chipp Jansen, Jonathan Chan, and Michael Byrd. Toward a methodology for agent-based data mining and visualization. Lecture Notes in Computer Science, 7103:4–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27609-5_2/.

Chaimontree:2012:MAB

[1480] Santhana Chaimontree, Katie Atkinson, and Frans Coenen. A multiagent based approach to clustering: Harnessing the power of agents. Lecture Notes in Computer Science, 7103:16–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27609-5_3/.

Bhamra:2012:AED

[1481] G. S. Bhamra, A. K. Verma, and R. B. Patel. Agent enriched distributed association rules mining: a review. Lecture Notes in Computer Science, 7103:30–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27609-5_4/.

Moemeng:2012:OOM

[1482] Chayapol Moemeng, Can Wang, and Longbing Cao. Obtaining an optimal MAS configuration for agent-enhanced mining using constraint optimization. Lecture Notes in Computer Science, 7103:46–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27609-5_5/.

Corniglion:2012:TNA

[1483] Sébastien Corniglion and Nadine Tournois. Towards a numerical, agentbased, behaviour analysis: The case of tourism. Lecture Notes in Computer Science, 7103:58–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27609-5_6/.

Neri:2012:CSF

[1484] Filippo Neri. A comparative study of a financial agent based simulator across learning scenarios. Lecture Notes in Computer Science, 7103:86– 97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27609-5_7/.

Zhang:2012:ABC

[1485] Wei Zhang and Yuanfei Wang. Agent-based cluster analysis of tropical cyclone tracks in the Western North Pacific. Lecture Notes in Computer Science, 7103:98–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27609-5_8/.

Anonymous:2012:FMcd

[1486] Anonymous. Front matter. Lecture Notes in Computer Science, 7103: 115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27609-5/2/1.

Emele:2012:EDK

[1487] Chukwuemeka David Emele, Timothy J. Norman, Murat Şensoy, and Simon Parsons. Exploiting domain knowledge in making delegation decisions. Lecture Notes in Computer Science, 7103:117–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27609-5_9/.

Epstein:2012:DMS

[1488] Susan L. Epstein, Rebecca Passonneau, Tiziana Ligorio, and Joshua Gordon. Data mining to support human-machine dialogue for autonomous agents. Lecture Notes in Computer Science, 7103:132–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27609-5_10/.

Attar:2012:IWB

[1489] Vahida Attar, Prashant Chaudhary, Sonali Rahagude, Gaurish Chaudhari, and Pradeep Sinha. An instance-window based classification algorithm for handling gradual concept drifts. Lecture Notes in Computer Science, 7103:156–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27609-5_11/.

Brahmi:2012:TMB

[1490] Imen Brahmi, Sadok Ben Yahia, Hamed Aouadi, and Pascal Poncelet. Towards a multiagent-based distributed intrusion detection system using data mining approaches. Lecture Notes in Computer Science, 7103:173– 194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27609-5_12/.

Fiosins:2012:CPA

[1491] Maksims Fiosins, Jelena Fiosina, and Jörg P. Müller. Change point analysis for intelligent agents in city traffic. Lecture Notes in Computer Science, 7103:195–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27609-5_13/.

Wang:2012:MFA

[1492] Xiaofeng Wang, Wenjia Niu, Gang Li, Xinghua Yang, and Zhongzhi Shi. Mining frequent agent action patterns for effective multi-agent-based Web service composition. Lecture Notes in Computer Science, 7103:211– 227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27609-5_14/.

Chatzidimitriou:2012:EAI

[1493] Kyriakos C. Chatzidimitriou, Antonios C. Chrysopoulos, and Andreas L. Symeonidis. Enhancing agent intelligence through evolving reservoir networks for predictions in power stock markets. Lecture Notes in Computer Science, 7103:228–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27609-5_15/.

Kaur:2012:PAO

[1494] Preetinder Kaur, Madhu Goyal, and Jie Lu. Pricing analysis in online auctions using clustering and regression tree approach. Lecture Notes in Computer Science, 7103:248–257, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27609-5_16/.

Anonymous:2012:FMce

[1495] Anonymous. Front matter. Lecture Notes in Computer Science, 7103: 259, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27609-5/3/1.

Anonymous:2012:FMcf

[1496] Anonymous. Front matter. Lecture Notes in Computer Science, 7103: 259, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27609-5/3/1.

Kowalczyk:2012:IRS

[1497] W. Kowalczyk, Z. Szlávik, and M. C. Schut. The impact of recommender systems on item-, user-, and rating-diversity. Lecture Notes in Computer Science, 7103:261–287, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27609-5_17/.

Sobkowicz:2012:OFS

[1498] Pawel Sobkowicz, Michael Kaschesky, and Guillaume Bouchard. Opinion formation in the social web: Agent-based simulations of opinion convergence and divergence. Lecture Notes in Computer Science, 7103:288– 303, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27609-5_18/.

Christensen:2012:DDA

[1499] Dion Christensen, Henrik Ossipoff Hansen, and Jorge Pablo Cordero Hernandez. A data-driven approach for resource gathering in realtime strategy games. Lecture Notes in Computer Science, 7103:304– 315, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27609-5_19/.

Wu:2012:DCD

[1500] Zhiang Wu, Jie Cao, and Changjian Fang. Data cloud for distributed data mining via pipelined MapReduce. Lecture Notes in Computer Science, 7103:316–330, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27609-5_20/.

Arora:2012:SEI

[1501] Monika Arora, Uma Kanjilal, and Dinesh Varshney. Successful efficient and intelligent retrieval using analytic hierarchy process. Lecture Notes in Computer Science, 7103:331–343, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-27609-5_21.

Kularbphettong:2012:HSB

[1502] Kunyanuth Kularbphettong, Phayung Meesad, and Gareth Clayton. A hybrid system based on multi-agent systems in case of e-WeddingThailand. Lecture Notes in Computer Science, 7103:344–359, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
27609-5_22.

Anonymous:2012:BMbt

[1503] Anonymous. Back matter. Lecture Notes in Computer Science, 7103: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27609-5/1.

Anonymous:2012:FMcg

[1504] Anonymous. Front matter. Lecture Notes in Computer Science, 7103: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27609-5/1.

Anonymous:2012:FMch

[1505] Anonymous. Front matter. Lecture Notes in Computer Science, 7104: 1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28320-8/1/1.

Kim:2012:ERH

[1506] Hyoungnyoun Kim and Ji-Hyung Park. Evaluating the regularity of human behavior from mobile phone usage logs. Lecture Notes in Computer Science, 7104:3–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28320-8_1/.

Akehurst:2012:EIU

[1507] Joshua Akehurst, Irena Koprinska, Kalina Yacef, Luiz Pizzato, Judy Kay, and Tomasz Rej. Explicit and implicit user preferences in online dating. Lecture Notes in Computer Science, 7104:15–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28320-8_2/.

Tsai:2012:BLT

[1508] Flora S. Tsai. Blogger-link-topic model for blog mining. Lecture Notes in Computer Science, 7104:28–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28320-8_3/.

Wan:2012:RIA

[1509] Miao Wan, Arne Jönsson, Cong Wang, Lixiang Li, and Yixian Yang. A random indexing approach for Web user clustering and Web prefetching. Lecture Notes in Computer Science, 7104:40–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28320-8_4/.

Nguyen:2012:ERR

[1510] Thin Nguyen, Dinh Phung, Brett Adams, and Svetha Venkatesh. Emotional reactions to real-world events in social networks. Lecture Notes in Computer Science, 7104:53–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28320-8_5/.

Huang:2012:CPK

[1511] Yin-Fu Huang and Cin-Siang Ciou. Constructing personal knowledge base: Automatic key-phrase extraction from multiple-domain Web pages. Lecture Notes in Computer Science, 7104:65–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28320-8_6/.

Shie:2012:DVU

[1512] Bai-En Shie, Hui-Fang Hsiao, Philip S. Yu, and Vincent S. Tseng. Discovering valuable user behavior patterns in mobile commerce environments. Lecture Notes in Computer Science, 7104:77–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28320-8_7/.

Yiwen:2012:NMC

[1513] Wang Yiwen and Yao Min. A novel method for community detection in complex network using new representation for communities. Lecture Notes in Computer Science, 7104:89–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28320-8_8/.

Spiegel:2012:LPE

[1514] Stephan Spiegel, Jan Clausen, Sahin Albayrak, and Jérôme Kunegis. Link prediction on evolving data using tensor factorization. Lecture Notes in Computer Science, 7104:100–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28320-8_9/.

He:2012:PAI

[1515] Xianmang He, Yanghua Xiao, Yujia Li, Qing Wang, Wei Wang, and Baile Shi. Permutation anonymization: Improving anatomy for privacy preservation in data publication. Lecture Notes in Computer Science, 7104:111–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28320-8_10/.

Amphawan:2012:EMT

[1516] Komate Amphawan, Philippe Lenca, and Athasit Surarerks. Efficient mining top-k regular-frequent itemset using compressed tidsets. Lecture Notes in Computer Science, 7104:124–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28320-8_11/.

Li:2012:MSM

[1517] Hailin Li, Chonghui Guo, and Libin Yang. A method of similarity measure and visualization for long time series using binary patterns. Lecture Notes in Computer Science, 7104:136–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28320-8_12/.

Nhon:2012:BBC

[1518] Vo Le Quy Nhon and Duong Tuan Anh. A BIRCH-based clustering method for large time series databases. Lecture Notes in Computer Science, 7104:148–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28320-8_13/.

Ishikawa:2012:VCS

[1519] Masahiro Ishikawa. Visualizing cluster structures and their changes over time by two-step application of self-organizing maps. Lecture Notes in Computer Science, 7104:160–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28320-8_14/.

Denny:2012:ACM

[1520] Denny, Peter Christen, and Graham J. Williams. Analysis of cluster migrations using self-organizing maps. Lecture Notes in Computer Science, 7104:171–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28320-8_15/.

Anonymous:2012:FMci

[1521] Anonymous. Front matter. Lecture Notes in Computer Science, 7104: 183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28320-8/1.

Anonymous:2012:FMcj

[1522] Anonymous. Front matter. Lecture Notes in Computer Science, 7104: 183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28320-8/2/1.

Ivanescu:2012:CMR

[1523] Anca Maria Ivanescu, Marc Wichterich, and Thomas Seidl. ClasSi: Measuring ranking quality in the presence of object classes with similarity information. Lecture Notes in Computer Science, 7104:185–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28320-8_16/.

Estivill-Castro:2012:IES

[1524] Vladimir Estivill-Castro. The instance easiness of supervised learning for cluster validity. Lecture Notes in Computer Science, 7104:197–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28320-8_17/.

Lamirel:2012:NEU

[1525] Jean-Charles Lamirel, Pascal Cuxac, Raghvendra Mall, and Ghada Safi. A new efficient and unbiased approach for clustering quality evaluation. Lecture Notes in Computer Science, 7104:209–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28320-8_18/.

Hadzic:2012:SPF

[1526] Fedja Hadzic. A structure preserving flat data format representation for tree-structured data. Lecture Notes in Computer Science, 7104:221– 233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28320-8_19/.

Fan:2012:FAN

[1527] Jun Fan and Tiejun Huang. A fusion of algorithms in near duplicate document detection. Lecture Notes in Computer Science, 7104:234–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28320-8_20/.

Yang:2012:SIA

[1528] Guangfei Yang, Yanzhong Dang, Shingo Mabu, Kaoru Shimada, and Kotaro Hirasawa. Searching interesting association rules based on evolutionary computation. Lecture Notes in Computer Science, 7104:243–253, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
28320-8_21.

Surana:2012:EAM

[1529] Akshat Surana, R. Uday Kiran, and P. Krishna Reddy. An efficient approach to mine periodic-frequent patterns in transactional databases. Lecture Notes in Computer Science, 7104:254–266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-28320-8_22.

Wu:2012:ADC

[1530] Jianjun Wu, Li Wan, and Zeren Xu. Algorithms to discover complete frequent episodes in sequences. Lecture Notes in Computer Science, 7104:267–278, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-28320-8_23.

Garriga:2012:CUE

[1531] Joan Garriga. Certainty upon empirical distributions. Lecture Notes in Computer Science, 7104:279–290, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-28320-8_24.

Anonymous:2012:FMck

[1532] Anonymous. Front matter. Lecture Notes in Computer Science, 7104: 291, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28320-8/3/1.

Yuan:2012:MOT

[1533] Bo Yuan and Wenhuang Liu. A measure oriented training scheme for imbalanced classification problems. Lecture Notes in Computer Science, 7104:293–303, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-28320-8_25.

Wang:2012:SBA

[1534] Yi Wang, Cheqing Jin, Minqi Zhou, and Aoying Zhou. An SVM-based approach to discover MicroRNA precursors in plant genomes. Lecture Notes in Computer Science, 7104:304–315, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-28320-8_26.

Alam:2012:TRS

[1535] Shafiq Alam, Gillian Dobbie, and Patricia Riddle. Towards recommender system using particle swarm optimization based Web usage clustering. Lecture Notes in Computer Science, 7104:316–326, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-28320-8_27.

Pears:2012:WAR

[1536] Russel Pears and Yun Sing Koh. Weighted association rule mining using particle swarm optimization. Lecture Notes in Computer Science, 7104:327–338, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-28320-8_28.

Jiang:2012:UFS

[1537] Sheng yi Jiang and Lian xi Wang. An unsupervised feature selection framework based on clustering. Lecture Notes in Computer Science, 7104:339–350, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-28320-8_29.

Anonymous:2012:FMcl

[1538] Anonymous. Front matter. Lecture Notes in Computer Science, 7104: 351, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28320-8/4/1.

Zhang:2012:DRU

[1539] Nevin L. Zhang, Runsun Zhang, and Tao Chen. Discovery of regularities in the use of herbs in traditional Chinese medicine prescriptions. Lecture Notes in Computer Science, 7104:353–360, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-28320-8_30.

Detterer:2012:CCE

[1540] Dion Detterer and Paul Kwan. COW: a co-evolving memetic wrapper for herb-herb interaction analysis in TCM informatics. Lecture Notes in Computer Science, 7104:361–371, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-28320-8_31.

Zhang:2012:SAI

[1541] Lei Zhang, Qi ming Zhang, Yi guo Wang, and Dong lin Yu. Selecting an appropriate interestingness measure to evaluate the correlation between syndrome elements and symptoms. Lecture Notes in Computer Science, 7104:372–383, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-28320-8_32.

Poon:2012:IFR

[1542] Simon Poon, Zhe Luo, and Runshun Zhang. The impact of feature representation to the biclustering of symptoms-herbs in TCM. Lecture Notes in Computer Science, 7104:384–394, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-28320-8_33.

Anonymous:2012:FMcm

[1543] Anonymous. Front matter. Lecture Notes in Computer Science, 7104: 395, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28320-8/5/1.

Saravanan:2012:UMP

[1544] M. Saravanan, S. Shanthi, and S. Shalini. Usage of mobile phones for personalized healthcare solutions. Lecture Notes in Computer Science, 7104:397–407, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-28320-8_34.

Wang:2012:RLM

[1545] Boyu Wang, Feng Wan, Peng Un Mak, Pui In Mak, and Mang I. Vai. Robust learning of mixture models and its application on trial pruning for EEG signal analysis. Lecture Notes in Computer Science, 7104:408–419, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
28320-8_35.

Gu:2012:IAM

[1546] Wei Gu, Baijie Wang, and Xin Wang. An integrated approach to multicriteria-based health care facility location planning. Lecture Notes in Computer Science, 7104:420–430, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-28320-8_36.

Pechsiri:2012:MPK

[1547] Chaveevan Pechsiri, Sumran Painuall, and Uraiwan Janviriyasopak. Medicinal property knowledge extraction from herbal documents for supporting question answering system. Lecture Notes in Computer Science, 7104:431–443, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-28320-8_37.

Anonymous:2012:FMcn

[1548] Anonymous. Front matter. Lecture Notes in Computer Science, 7104: 445, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28320-8/6/1.

Anonymous:2012:FMco

[1549] Anonymous. Front matter. Lecture Notes in Computer Science, 7104: 445, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28320-8/6/1.

Zhang:2012:AEU

[1550] Yu Zhang. Age estimation using Bayesian process. Lecture Notes in Computer Science, 7104:447–458, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-28320-8_38.

Tseng:2012:SNI

[1551] Chi-Yao Tseng and Ming-Syan Chen. Significant node identification in social networks. Lecture Notes in Computer Science, 7104:459–470, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
28320-8_39.

Hsu:2012:IBP

[1552] Kuo-Wei Hsu and Jaideep Srivastava. Improving bagging performance through multi-algorithm ensembles. Lecture Notes in Computer Science, 7104:471–482, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-28320-8_40.

Zhang:2012:MTP

[1553] Chen Zhang and Jie Zhang. Mining tourist preferences with twicelearning. Lecture Notes in Computer Science, 7104:483–493, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
28320-8_41.

Liu:2012:TCS

[1554] Xu-Ying Liu and Zhi-Hua Zhou. Towards cost-sensitive learning for realworld applications. Lecture Notes in Computer Science, 7104:494–505, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
28320-8_42.

Anonymous:2012:BMbu

[1555] Anonymous. Back matter. Lecture Notes in Computer Science, 7104: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28320-8/1.

Anonymous:2012:FMcp

[1556] Anonymous. Front matter. Lecture Notes in Computer Science, 7104: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28320-8/1.

Meerwald:2012:ERW

[1557] Peter Meerwald and Andreas Uhl. An efficient robust watermarking method integrated in H.264/SVC. Lecture Notes in Computer Science, 7110:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28693-3_1/.

Yamada:2012:PBR

[1558] Takaaki Yamada and Isao Echizen. PC-based real-time video watermark embedding system independent of platform for parallel computing. Lecture Notes in Computer Science, 7110:15–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28693-3_2/.

Echizen:2012:IHM

[1559] Isao Echizen and Takayuki Yamada. IR hiding: Method for preventing illegal recording of videos based on differences in sensory perception between humans and devices. Lecture Notes in Computer Science, 7110: 34–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28693-3_3/.

Wu:2012:SWG

[1560] Hao tian Wu and Yiu ming Cheung. Secure watermarking on 3D geometry via ICA and orthogonal transformation. Lecture Notes in Computer Science, 7110:52–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28693-3_4/.

Cao:2012:MSC

[1561] Hong Cao and Alex C. Kot. Measuring the statistical correlation inconsistencies in mobile images for tamper detection. Lecture Notes in Computer Science, 7110:63–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28693-3_5/.

Sur:2012:SSU

[1562] Arijit Sur and Vignesh Ramanathan. Secure steganography using randomized cropping. Lecture Notes in Computer Science, 7110:82–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28693-3_6/.

Zhao:2012:SSM

[1563] Hong Zhao, Yun Q. Shi, and Nirwan Ansari. Steganography in streaming multimedia over networks. Lecture Notes in Computer Science, 7110:96– 114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28693-3_7/.

Anonymous:2012:BMbv

[1564] Anonymous. Back matter. Lecture Notes in Computer Science, 7110: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28693-3/1.

Anonymous:2012:FMcq

[1565] Anonymous. Front matter. Lecture Notes in Computer Science, 7110: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28693-3/1.

Dolev:2012:DMP

[1566] Shlomi Dolev. Dynamic multi-party computation forever for swarm and cloud computing and code obfuscation. Lecture Notes in Computer Science, 7111:1–3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28209-6_1/.

Kempkes:2012:LSO

[1567] Barbara Kempkes and Friedhelm Meyer auf der Heide. Local, selforganizing strategies for robotic formation problems. Lecture Notes in Computer Science, 7111:4–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28209-6_2/.

Abshoff:2012:LAA

[1568] Sebastian Abshoff, Andreas Cord-Landwehr, Bastian Degener, and Barbara Kempkes. Local approximation algorithms for the uncapacitated metric facility location problem in power-aware sensor networks. Lecture Notes in Computer Science, 7111:13–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28209-6_3/.

Bar-Noy:2012:MNL

[1569] Amotz Bar-Noy and Ben Baumer. Maximizing network lifetime on the line with adjustable sensing ranges. Lecture Notes in Computer Science, 7111:28–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28209-6_4/.

Cohen:2012:SFD

[1570] Asaf Cohen, Shlomi Dolev, and Guy Leshem. Sensor fusion: From dependence analysis via matroid bases to online synthesis. Lecture Notes in Computer Science, 7111:42–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28209-6_5/.

Du:2012:NDS

[1571] Jingzhe Du, Evangelos Kranakis, Oscar Morales Ponce, and Sergio Rajsbaum. Neighbor discovery in a sensor network with directional antennae. Lecture Notes in Computer Science, 7111:57–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28209-6_6/.

Eyal:2012:LLM

[1572] Ittay Eyal, Idit Keidar, and Raphael Rom. LiMoSense — live monitoring in dynamic sensor networks. Lecture Notes in Computer Science, 7111:

72–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28209-6_7/.

Johnson:2012:EIC

[1573] Matthew P. Johnson and Alexander Gutfraind. Evader interdiction and collateral damage. Lecture Notes in Computer Science, 7111:86–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28209-6_8/.

Li:2012:EAN

[1574] Meng Li, Yota Otachi, and Takeshi Tokuyama. Efficient algorithms for network localization using cores of underlying graphs. Lecture Notes in Computer Science, 7111:101–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28209-6_9/.

Lou:2012:MAI

[1575] Tiancheng Lou, Haisheng Tan, Yuexuan Wang, and Francis C. M. Lau. Minimizing average interference through topology control. Lecture Notes in Computer Science, 7111:115–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28209-6_10/.

Tseng:2012:BRS

[1576] Kuan-Chieh Robert Tseng and David Kirkpatrick. On barrier resilience of sensor networks. Lecture Notes in Computer Science, 7111:130–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28209-6_11/.

Yu:2012:DCP

[1577] Dongxiao Yu, Yuexuan Wang, Qiang-Sheng Hua, and Francis C. M. Lau. Distributed (Δ + 1)-coloring in the physical model. Lecture Notes in Computer Science, 7111:145–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28209-6_12/;   http:
//link.springer.com/content/pdf/10.1007/978-3-642-28209-6_12.

Alvarez:2012:CMD

[1578] Carme Àlvarez, Josep Díaz, Dieter Mitsche, and Maria Serna. Continuous monitoring in the dynamic sensor field model. Lecture Notes in Computer Science, 7111:161–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28209-6_13/.

Bar-Noy:2012:MCB

[1579] Amotz Bar-Noy, Prithwish Basu, Matthew P. Johnson, and Ram Ramanathan. Minimum-cost broadcast through varying-size neighborcast. Lecture Notes in Computer Science, 7111:173–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28209-6_14/.

Even:2012:RTV

[1580] Guy Even, Yaniv Fais, Moti Medina, Shimon (Moni) Shahar, and Alexander Zadorojniy. Real-time video streaming in multi-hop wireless static ad hoc networks. Lecture Notes in Computer Science, 7111:188–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28209-6_15/.

Even:2012:MHR

[1581] Guy Even, Yakov Matsri, and Moti Medina. Multi-hop routing and scheduling in wireless networks in the SINR model. Lecture Notes in Computer Science, 7111:202–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28209-6_16/.

Halldorsson:2012:WCA

[1582] Magnús M. Halldórsson and Pradipta Mitra. Wireless capacity with arbitrary gain matrix. Lecture Notes in Computer Science, 7111:215– 224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28209-6_17/.

Tonoyan:2012:COP

[1583] Tigran Tonoyan. On the capacity of oblivious powers. Lecture Notes in Computer Science, 7111:225–237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28209-6_18/.

Anonymous:2012:BMbw

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28209-6/1.

Anonymous:2012:FMcr

[1585] Anonymous. Front matter. Lecture Notes in Computer Science, 7111: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28209-6/1.

Cygan:2012:MCP

[1586] Marek Cygan, Marcin Pilipczuk, and Michał Pilipczuk. On multiway cut parameterized above lower bounds. Lecture Notes in Computer Science, 7112:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28050-4_1/.

Cygan:2012:PCF

[1587] Marek Cygan, Fedor V. Fomin, and Erik Jan van Leeuwen. Parameterized complexity of firefighting revisited. Lecture Notes in Computer Science, 7112:13–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28050-4_2/.

Jiang:2012:PCM

[1588] Minghui Jiang and Yong Zhang. Parameterized complexity in multipleinterval graphs: Domination. Lecture Notes in Computer Science, 7112: 27–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28050-4_3/.

Iwata:2012:FAD

[1589] Yoichi Iwata. A faster algorithm for dominating set analyzed by the potential method. Lecture Notes in Computer Science, 7112:41–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28050-4_4/.

Heggernes:2012:CGP

[1590] Pinar Heggernes, Pim van ’t Hof, and Benjamin Lévêque. Contracting graphs to paths and trees. Lecture Notes in Computer Science, 7112:55– 66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28050-4_5/.

Golovach:2012:IMD

[1591] Petr A. Golovach, Marcin Kamiński, and Daniël Paulusma. Increasing the minimum degree of a graph by contractions. Lecture Notes in Computer Science, 7112:67–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28050-4_6/.

Adler:2012:PDP

[1592] Isolde Adler and Stavros G. Kolliopoulos. Planar disjoint-paths completion. Lecture Notes in Computer Science, 7112:80–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28050-4_7/.

Damaschke:2012:SSS

[1593] Peter Damaschke. Sparse solutions of sparse linear systems: Fixedparameter tractability and an application of complex group testing. Lecture Notes in Computer Science, 7112:94–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28050-4_8/.

Golovnev:2012:NUB

[1594] Alexander Golovnev. New upper bounds for MAX-2-SAT and MAX-2-CSP w.r.t. the average variable degree. Lecture Notes in Computer Science, 7112:106–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28050-4_9/.

Kim:2012:IPA

[1595] Eun Jung Kim and Ryan Williams. Improved parameterized algorithms for above average constraint satisfaction. Lecture Notes in Computer Science, 7112:118–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28050-4_10/.

Jansen:2012:PKS

[1596] Bart M. P. Jansen and Stefan Kratsch. On polynomial kernels for structural parameterizations of odd cycle transversal. Lecture Notes in Computer Science, 7112:132–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28050-4_11/.

Bodlaender:2012:KBP

[1597] Hans L. Bodlaender and Bart M. P. Jansen. Kernel bounds for path and cycle problems. Lecture Notes in Computer Science, 7112:145–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28050-4_12/.

Cygan:2012:HLW

[1598] Marek Cygan, Daniel Lokshtanov, and Marcin Pilipczuk. On the hardness of losing width. Lecture Notes in Computer Science, 7112:159– 168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28050-4_13/.

Guo:2012:SAR

[1599] Jiong Guo, Iyad Kanj, and Stefan Kratsch. Safe approximation and its relation to kernelization. Lecture Notes in Computer Science, 7112:169– 180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28050-4_14/.

Hagerup:2012:SLT

[1600] Torben Hagerup. Simpler linear-time kernelization for planar dominating set. Lecture Notes in Computer Science, 7112:181–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28050-4_15/.

vanBevern:2012:LTC

[1601] René van Bevern, Sepp Hartung, and Frank Kammer. Linear-time computation of a linear problem kernel for dominating set on planar graphs. Lecture Notes in Computer Science, 7112:194–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28050-4_16/.

Broersma:2012:TCB

[1602] Hajo Broersma, Petr A. Golovach, and Viresh Patel. Tight complexity bounds for FPT subgraph problems parameterized by clique-width. Lecture Notes in Computer Science, 7112:207–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28050-4_17/.

Hvidevold:2012:FGD

[1603] Eivind Magnus Hvidevold and Sadia Sharmin. Finding good decompositions for dynamic programming on dense graphs. Lecture Notes in Computer Science, 7112:219–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28050-4_18/.

Lampis:2012:PMP

[1604] Michael Lampis. Parameterized maximum path coloring. Lecture Notes in Computer Science, 7112:232–245, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28050-4_19/.

Cygan:2012:CPV

[1605] Marek Cygan, Daniel Lokshtanov, and Marcin Pilipczuk. On cutwidth parameterized by vertex cover. Lecture Notes in Computer Science, 7112:246–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28050-4_20/.

Ganian:2012:TCB

[1606] Robert Ganian. Twin-Cover: Beyond vertex cover in parameterized algorithmics. Lecture Notes in Computer Science, 7112:259–271, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
28050-4_21.

Anonymous:2012:BMbx

[1607] Anonymous. Back matter. Lecture Notes in Computer Science, 7112: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28050-4/1.

Anonymous:2012:FMcs

[1608] Anonymous. Front matter. Lecture Notes in Computer Science, 7112: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28050-4/1.

Robinson:2012:CLS

[1609] Edward Robinson, Peter McBurney, and Xin Yao. Co-learning segmentation in marketplaces. Lecture Notes in Computer Science, 7113:1– 20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28499-1_1/.

Ammar:2012:RLT

[1610] Haitham Bou Ammar and Matthew E. Taylor. Reinforcement learning transfer via common subspaces. Lecture Notes in Computer Science, 7113:21–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28499-1_2/.

Kemmerich:2012:CMR

[1611] Thomas Kemmerich and Hans Kleine Büning. A convergent multiagent reinforcement learning approach for a subclass of cooperative stochastic games. Lecture Notes in Computer Science, 7113:37–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28499-1_3/.

Martinez-Gil:2012:MAR

[1612] Francisco Martinez-Gil, Miguel Lozano, and Fernando Fernández. Multiagent reinforcement learning for simulating pedestrian navigation. Lecture Notes in Computer Science, 7113:54–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28499-1_4/.

Hosseini:2012:LDK

[1613] Hadi Hosseini and Mihaela Ulieru. Leveraging domain knowledge to learn normative behavior: a Bayesian approach. Lecture Notes in Computer Science, 7113:70–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28499-1_5/.

Comanici:2012:BFD

[1614] Gheorghe Comanici and Doina Precup. Basis function discovery using spectral clustering and bisimulation metrics. Lecture Notes in Computer Science, 7113:85–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28499-1_6/.

Catteeuw:2012:HPL

[1615] David Catteeuw and Bernard Manderick. Heterogeneous populations of learning agents in the minority game. Lecture Notes in Computer Science, 7113:100–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28499-1_7/.

DeHauwere:2012:SSD

[1616] Yann-Michaël De Hauwere, Peter Vrancx, and Ann Nowé. Solving sparse delayed coordination problems in multi-agent reinforcement learning. Lecture Notes in Computer Science, 7113:114–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28499-1_8/.

Anonymous:2012:BMby

[1617] Anonymous. Back matter. Lecture Notes in Computer Science, 7113: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28499-1/1.

Anonymous:2012:FMct

[1618] Anonymous. Front matter. Lecture Notes in Computer Science, 7113: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28499-1/1.

Sekar:2012:PAC

[1619] Gautham Sekar and Bart Preneel. Practical attacks on a cryptosystem proposed in Patent WO/2009/066313. Lecture Notes in Computer Science, 7115:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27890-7_1/.

Kogure:2012:GSA

[1620] Jun Kogure, Noboru Kunihiro, and Hirosuke Yamamoto. Generalized security analysis of the random key bits leakage attack. Lecture Notes in Computer Science, 7115:13–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27890-7_2/.

Li:2012:IIA

[1621] Yanjun Li, Wenling Wu, and Lei Zhang. Improved integral attacks on reduced-round CLEFIA block cipher. Lecture Notes in Computer Science, 7115:28–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27890-7_3/.

Ohtahara:2012:PAF

[1622] Chiaki Ohtahara, Keita Okada, and Yu Sasaki. Preimage attacks on full-ARIRANG: Analysis of DM-mode with middle feed-forward. Lecture Notes in Computer Science, 7115:40–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27890-7_4/.

Dong:2012:KKD

[1623] Le Dong, Wenling Wu, Shuang Wu, and Jian Zou. Known-key distinguisher on round-reduced 3D block cipher. Lecture Notes in Computer Science, 7115:55–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27890-7_5/.

Lee:2012:IBS

[1624] Woomyo Lee, Jae Woo Seo, and Pil Joong Lee. Identity-based signcryption from identity-based cryptography. Lecture Notes in Computer Science, 7115:70–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27890-7_6/.

Yum:2012:OPE

[1625] Dae Hyun Yum, Duk Soo Kim, Jin Seok Kim, and Pil Joong Lee. Orderpreserving encryption for non-uniformly distributed plaintexts. Lecture Notes in Computer Science, 7115:84–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27890-7_7/.

Sakemi:2012:SDA

[1626] Yumi Sakemi, Tetsuya Izu, and Masahiko Takenaka. Solving a DLP with auxiliary input with the ρ-algorithm. Lecture Notes in Computer Science, 7115:98–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27890-7_8/.

Pan:2012:GNL

[1627] Yanbin Pan and Yingpu Deng. A general NTRU-like framework for constructing lattice-based public-key cryptosystems. Lecture Notes in Computer Science, 7115:109–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27890-7_9/.

Matsushita:2012:PPC

[1628] Tatsuyuki Matsushita, Shinji Yamanaka, and Fangming Zhao. A peer-to-peer content-distribution scheme resilient to key leakage. Lecture Notes in Computer Science, 7115:121–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27890-7_10/.

Kang:2012:RIE

[1629] Boojoong Kang, Hye Seon Kim, Ji Su Yang, and Eul Gyu Im. Rule indexing for efficient intrusion detection systems. Lecture Notes in Computer Science, 7115:136–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27890-7_11/.

Kim:2012:SDE

[1630] Chung-Hyo Kim, Moon-Seok Choi, and Seong-Ho Ju. Security data extraction from IEC 61850 ACSI models for network and system management. Lecture Notes in Computer Science, 7115:142–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27890-7_12/.

Hong:2012:LMB

[1631] Insung Hong, Jisung Byun, and Sehyun Park. Lightweight middlewarebased ZigBee security in building energy management system. Lecture Notes in Computer Science, 7115:151–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27890-7_13/.

Zhang:2012:MLB

[1632] Yingjun Zhang, Yang Zhang, and Kai Chen. A map-layer-based access control model. Lecture Notes in Computer Science, 7115:157–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27890-7_14/.

Ohtake:2012:AAH

[1633] Go Ohtake and Kazuto Ogawa. Application authentication for hybrid services of broadcasting and communications networks. Lecture Notes in Computer Science, 7115:171–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27890-7_15/.

Kiribuchi:2012:AMC

[1634] Naoto Kiribuchi, Ryo Kato, and Takashi Nishide. Accelerating multiparty computation by efficient random number bitwise-sharing protocols. Lecture Notes in Computer Science, 7115:187–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27890-7_16/.

Cho:2012:BBS

[1635] Kwantae Cho and Dong Hoon Lee. Biometric based secure communications without pre-deployed key for biosensor implanted in body sensor networks. Lecture Notes in Computer Science, 7115:203–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27890-7_17/.

Kim:2012:MPS

[1636] Myungsun Kim, Hyung Tae Lee, and Jung Hee Cheon. Mutual private set intersection with linear complexity. Lecture Notes in Computer Science, 7115:219–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27890-7_18/.

Sun:2012:APS

[1637] Chung Il Sun and Tae Ho Cho. Advanced path selection method for detection of false reports in statistical filtering based WSNs. Lecture Notes in Computer Science, 7115:232–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27890-7_19/.

Salonen:2012:ESP

[1638] Jarno Salonen. Evaluating the security and privacy of near field communication — case: Public transportation. Lecture Notes in Computer Science, 7115:242–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27890-7_20/.

Hegazy:2012:ERT

[1639] Islam Hegazy, Reihaneh Safavi-Naini, and Carey Williamson. Exploiting routing tree construction in CTP. Lecture Notes in Computer Science, 7115:256–270, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-27890-7_21.

Anonymous:2012:BMbz

[1640] Anonymous. Back matter. Lecture Notes in Computer Science, 7115: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27890-7/1.

Anonymous:2012:FMcu

[1641] Anonymous. Front matter. Lecture Notes in Computer Science, 7115: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27890-7/1.

Anonymous:2012:FMcv

[1642] Anonymous. Front matter. Lecture Notes in Computer Science, 7116: 1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29843-1/1/1.

Brezina:2012:SAS

[1643] Marian Brezina and Panayot S. Vassilevski. Smoothed aggregation spectral element agglomeration AMG: SA-ρ AMGe. Lecture Notes in Computer Science, 7116:3–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29843-1_1/.

Casas:2012:ASC

[1644] Eduardo Casas, Roland Herzog, and Gerd Wachsmuth. Approximation of sparse controls in semilinear elliptic equations. Lecture Notes in Computer Science, 7116:16–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29843-1_2/.

Hofreither:2012:NSF

[1645] Clemens Hofreither, Ulrich Langer, and Clemens Pechstein. A nonstandard finite element method based on boundary integral operators. Lecture Notes in Computer Science, 7116:28–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29843-1_3/.

Anonymous:2012:FMcw

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29843-1/2/1.

Efendiev:2012:RSS

[1647] Yalchin Efendiev, Juan Galvis, Raytcho Lazarov, and Joerg Willems. Robust solvers for symmetric positive definite operators and weighted Poincaré inequalities. Lecture Notes in Computer Science, 7116:43–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29843-1_4/.

Kraus:2012:ASC

[1648] J. Kraus. Additive Schur complement approximation for elliptic problems with oscillatory coefficients. Lecture Notes in Computer Science, 7116: 52–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29843-1_5/.

Anonymous:2012:FMcx

[1649] Anonymous. Front matter. Lecture Notes in Computer Science, 7116: 61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29843-1/3/1.

Bochev:2012:OBM

[1650] Pavel Bochev, Denis Ridzal, and Joseph Young. Optimization–based modeling with applications to transport: Part 1. abstract formulation. Lecture Notes in Computer Science, 7116:63–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29843-1_6/.

Young:2012:OBM

[1651] Joseph Young, Denis Ridzal, and Pavel Bochev. Optimization-based modeling with applications to transport: Part 2. the optimization algorithm. Lecture Notes in Computer Science, 7116:72–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29843-1_7/.

Ridzal:2012:OBM

[1652] Denis Ridzal, Joseph Young, Pavel Bochev, and Kara Peterson. Optimization-based modeling with applications to transport: Part 3. computational studies. Lecture Notes in Computer Science, 7116:81– 88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29843-1_8/.

Anonymous:2012:FMcy

[1653] Anonymous. Front matter. Lecture Notes in Computer Science, 7116: 89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29843-1/4/1.

Baier:2012:NMS

[1654] Robert Baier and Mirko Hessel von Molo. Newton’s method and secant method for set-valued mappings. Lecture Notes in Computer Science, 7116:91–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29843-1_9/.

Chernousko:2012:OCM

[1655] F. L. Chernousko. Optimal control of multibody systems in resistive media. Lecture Notes in Computer Science, 7116:99–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29843-1_10/.

Chryssoverghi:2012:CRP

[1656] I. Chryssoverghi, J. Coletsos, and B. Kokkinis. Classical and relaxed progressively refining discretization-optimization methods for optimal control problems defined by ordinary differential equations. Lecture Notes in Computer Science, 7116:106–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29843-1_11/.

Dimitrova:2012:ASU

[1657] Neli S. Dimitrova and Mikhail I. Krastanov. On the asymptotic stabilization of an uncertain bioprocess model. Lecture Notes in Computer Science, 7116:115–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29843-1_12/.

Filippova:2012:RSI

[1658] Tatiana F. Filippova and Oksana G. Matviychuk. Reachable sets of impulsive control system with cone constraint on the control and their estimates. Lecture Notes in Computer Science, 7116:123–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29843-1_13/.

Marigonda:2012:OMT

[1659] Antonio Marigonda and Giandomenico Orlandi. Optimal mass transportation-based models for neuronal fibers. Lecture Notes in Computer Science, 7116:131–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29843-1_14/.

Moser:2012:OCM

[1660] Elke Moser, Alexia Prskawetz, and Gernot Tragler. Optimal controls in models of economic growth and the environment. Lecture Notes in Computer Science, 7116:139–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29843-1_15/.

Poggiolini:2012:MTP

[1661] L. Poggiolini and G. Stefani. On the minimum time problem for Dodgem Car-like bang–singular extremals. Lecture Notes in Computer Science, 7116:147–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29843-1_16/.

Popchev:2012:PBN

[1662] Ivan Popchev, Petko Petkov, Mihail Konstantinov, and Vera Angelova. Perturbation bounds for the nonlinear matrix equation. Lecture Notes in Computer Science, 7116:155–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29843-1_17/.

Anonymous:2012:FMcz

[1663] Anonymous. Front matter. Lecture Notes in Computer Science, 7116: 163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29843-1/5/1.

Anonymous:2012:FMda

[1664] Anonymous. Front matter. Lecture Notes in Computer Science, 7116: 163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29843-1/5/1.

Angelova:2012:SAP

[1665] Maria Angelova and Tania Pencheva. Sensitivity analysis for the purposes of parameter identification of a S. cerevisiae fed-batch cultivation. Lecture Notes in Computer Science, 7116:165–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29843-1_18/.

Anghinolfi:2012:MAL

[1666] D. Anghinolfi, L. M. Gambardella, R. Montemanni, C. Nattero, M. Paolucci, and N. E. Toklu. A matheuristic algorithm for a largescale energy management problem. Lecture Notes in Computer Science, 7116:173–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29843-1_19/.

Atanassova:2012:GMM

[1667] Lilija Atanassova and Krassimir Atanassov. On a game-method for modelling with intuitionistic fuzzy estimations: Part 1. Lecture Notes in Computer Science, 7116:182–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29843-1_20/.

Atanassova:2012:GNA

[1668] Vassia Atanassova, Stefka Fidanova, Panagiotis Chountas, and Krassimir Atanassov. A generalized net with an ACO-algorithm optimization component. Lecture Notes in Computer Science, 7116:190–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29843-1_21.

Balabanov:2012:TSP

[1669] Todor Balabanov, Iliyan Zankinski, and Nina Dobrinkova. Time series prediction by artificial neural networks and differential evolution in distributed environment. Lecture Notes in Computer Science, 7116:198–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29843-1_22.

Duda:2012:DEA

[1670] Jerzy Duda and Iwona Skalna. Differential evolution applied to large scale parametric interval linear systems. Lecture Notes in Computer Science, 7116:206–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29843-1_23.

Espinar:2012:UCO

[1671] Javier Espinar, Carlos Cotta, and Antonio J. Fernández-Leiva. Usercentric optimization with evolutionary and memetic systems. Lecture Notes in Computer Science, 7116:214–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29843-1_24.

Fidanova:2012:IFE

[1672] Stefka Fidanova, Krassimir Atanassov, and Pencho Marinov. Intuitionistic fuzzy estimation of the ant colony optimization starting points. Lecture Notes in Computer Science, 7116:222–229, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29843-1_25.

Mucherino:2012:VNS

[1673] A. Mucherino, M. Fuchs, X. Vasseur, and S. Gratton. Variable neighborhood search for robust optimization and applications to aerodynamics. Lecture Notes in Computer Science, 7116:230–237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29843-1_26.

Ratuszniak:2012:PAD

[1674] Piotr Ratuszniak. Processor array design with the use of genetic algorithm. Lecture Notes in Computer Science, 7116:238–246, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29843-1_27.

Roeva:2012:HGA

[1675] Olympia Roeva. A hybrid genetic algorithm for parameter identification of bioprocess models. Lecture Notes in Computer Science, 7116:247–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29843-1_28.

Simian:2012:GFB

[1676] Dana Simian and Florin Stoica. A general frame for building optimal multiple SVM kernels. Lecture Notes in Computer Science, 7116:256–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29843-1_29.

Anonymous:2012:FMdb

[1677] Anonymous. Front matter. Lecture Notes in Computer Science, 7116: 265, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29843-1/6/1.

Brandiyska:2012:MTS

[1678] A. Brandiyska, K. Ganev, D. Syrakov, M. Prodanova, and N. Miloshev. Modeling of toxic substances in the atmosphere — risk analysis and emergency forecast. Lecture Notes in Computer Science, 7116:267–274, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29843-1_30.

Chervenkov:2012:SAI

[1679] Hristo Chervenkov. Some aspects of impact in the potential climate change on ozone pollution levels over Bulgaria from high resolution simulations. Lecture Notes in Computer Science, 7116:275–282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29843-1_31.

Georgiev:2012:NPI

[1680] Krassimir Georgiev, Tzvetan Ostromsky, and Zahari Zlatev. New parallel implementation of an air pollution computer model — performance study on an IBM Blue Gene/P computer. Lecture Notes in Computer Science, 7116:283–290, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29843-1_32.

Jordanov:2012:SHF

[1681] Georgi Jordanov, Jonathan D. Beezley, Nina Dobrinkova, Adam K. Kochanski, and Jan Mandel. Simulation of the 2009 Harmanli Fire (Bulgaria). Lecture Notes in Computer Science, 7116:291–298, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29843-1_33.

Liolios:2012:CAR

[1682] Konstantinos Liolios, Vassilios Tsihrintzis, Konstantinos Moutsopoulos, and Ivan Georgiev. A computational approach for remediation procedures in horizontal subsurface flow constructed wetlands. Lecture Notes in Computer Science, 7116:299–306, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29843-1_34.

Ostromsky:2012:PCS

[1683] Tzvetan Ostromsky, Ivan Dimov, Rayna Georgieva, and Zahari Zlatev. Parallel computation of sensitivity analysis data for the Danish Eulerian model. Lecture Notes in Computer Science, 7116:307–315, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29843-1_35.

SanJose:2012:ITD

[1684] R. San Jose, J. L. Perez, and R. M. Gonzalez. Implementation of two different shadow models into EULAG model: Madrid case study. Lecture Notes in Computer Science, 7116:316–323, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29843-1_36.

Syrakov:2012:MSA

[1685] D. Syrakov and M. Prodanova. Model simulation of air pollution due to April 2010 Iceland volcano eruption. Lecture Notes in Computer Science, 7116:324–332, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29843-1_37.

Tchorbadjieff:2012:ADQ

[1686] A. Tchorbadjieff. Automatic data quality control of environmental data. Lecture Notes in Computer Science, 7116:333–340, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29843-1_38.

Anonymous:2012:FMdc

[1687] Anonymous. Front matter. Lecture Notes in Computer Science, 7116: 341, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29843-1/1.

Haase:2012:CBE

[1688] Gundolf Haase, Martin Schanz, and Samar Vafai. Computing boundary element method’s matrices on GPU. Lecture Notes in Computer Science, 7116:343–350, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29843-1_39.

Shterev:2012:PAI

[1689] Kiril S. Shterev, Stefan K. Stefanov, and Emanouil I. Atanassov. A parallel algorithm with improved performance of finite volume method (SIMPLE–TS). Lecture Notes in Computer Science, 7116:351–358, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29843-1_40.

Weinbub:2012:TDH

[1690] Josef Weinbub, Karl Rupp, and Siegfried Selberherr. Towards distributed heterogeneous high-performance computing with ViennaCL. Lecture Notes in Computer Science, 7116:359–367, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29843-1_41.

Zinterhof:2012:HTS

[1691] Peter Zinterhof. High-throughput-screening of medical image data on heterogeneous clusters. Lecture Notes in Computer Science, 7116: 368–377, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29843-1_42.

Anonymous:2012:FMdd

[1692] Anonymous. Front matter. Lecture Notes in Computer Science, 7116: 379, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29843-1/8/1.

Popov:2012:PLS

[1693] P. Popov. Preconditioning of linear systems arising in finite element discretizations of the Brinkman equation. Lecture Notes in Computer Science, 7116:381–389, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-29843-1_43.

Anonymous:2012:FMde

[1694] Anonymous. Front matter. Lecture Notes in Computer Science, 7116: 391, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29843-1/9/1.

Bang:2012:BFH

[1695] Børre Bang, Lubomir T. Dechevsky, Arne Lakså, and Peter Zanaty. Blending functions for Hermite interpolation by beta-function B-splines on triangulations. Lecture Notes in Computer Science, 7116:393–401, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29843-1_44.

Dechevsky:2012:IMB

[1696] Lubomir T. Dechevsky, Jostein Bratlie, and Joakim Gundersen. Index mapping between tensor-product wavelet bases of different number of variables, and computing multivariate orthogonal discrete wavelet transforms on graphics processing units. Lecture Notes in Computer Science, 7116:402–410, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29843-1_45.

Dechevsky:2012:ICT

[1697] Lubomir T. Dechevsky and Georgi H. Georgiev. Interpolation of curvature and torsion using expo-rational B-splines. Lecture Notes in Computer Science, 7116:411–419, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29843-1_46.

Dechevsky:2012:HIU

[1698] Lubomir T. Dechevsky and Rumen Uluchev. Hermite interpolation using ERBS with trigonometric polynomial local functions. Lecture Notes in Computer Science, 7116:420–428, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29843-1_47.

Dechevsky:2012:TBF

[1699] Lubomir T. Dechevsky and Peter Zanaty. Triangular beta-function Bspline finite elements: Evaluation and graphical comparisons. Lecture Notes in Computer Science, 7116:429–436, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29843-1_48.

Anonymous:2012:FMdf

[1700] Anonymous. Front matter. Lecture Notes in Computer Science, 7116: 437, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29843-1/10/1.

Atanassov:2012:SSH

[1701] Emanouil I. Atanassov and Sofiya Ivanovska. Sensitivity study of Heston stochastic volatility model using GPGPU. Lecture Notes in Computer Science, 7116:439–446, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29843-1_49.

Filipovic:2012:MCS

[1702] Lado Filipovic and Siegfried Selberherr. A Monte Carlo simulator for non-contact mode atomic force microscopy. Lecture Notes in Computer Science, 7116:447–454, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29843-1_50.

Ivanovska:2012:NIU

[1703] Sofiya Ivanovska, A. Karaivanova, and N. Manev. Numerical integration using sequences generating permutations. Lecture Notes in Computer Science, 7116:455–463, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29843-1_51.

Sahpaski:2012:OII

[1704] Dragan Sahpaski, Ljupčo Pejov, and Anastas Misev. Optimization of intermolecular interaction potential energy parameters for Monte–Carlo and molecular dynamics simulations. Lecture Notes in Computer Science, 7116:464–471, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29843-1_52.

Schwaha:2012:PID

[1705] Philipp Schwaha, Mihail Nedjalkov, Siegfried Selberherr, and Ivan Dimov. Phonon-induced decoherence in electron evolution. Lecture Notes in Computer Science, 7116:472–479, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29843-1_53.

Trandafir:2012:SHI

[1706] Romică Trandafir and Cornel Resteanu. Study of human influenza’s spreading phenomenon. Lecture Notes in Computer Science, 7116: 480–486, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29843-1_54.

Anonymous:2012:FMdg

[1707] Anonymous. Front matter. Lecture Notes in Computer Science, 7116: 487, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29843-1/11/1.

Blaheta:2012:MSA

[1708] Radim Blaheta and Vojtěch Sokol. Multilevel solvers with aggregations for voxel based analysis of geomaterials. Lecture Notes in Computer Science, 7116:489–497, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29843-1_55.

Flaig:2012:HSM

[1709] Cyril Flaig and Peter Arbenz. A highly scalable matrix-free multigrid solver for μ FE analysis based on a pointer-less octree. Lecture Notes in Computer Science, 7116:498–506, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29843-1_56.

Linner:2012:APV

[1710] E. Linnér and R. Strand. Aliasing properties of voxels in threedimensional sampling lattices. Lecture Notes in Computer Science, 7116:507–514, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29843-1_57.

Vondrejc:2012:AFF

[1711] Jaroslav Vondřejc, Jan Zeman, and Ivo Marek. Analysis of a Fast Fourier Transform based method for modeling of heterogeneous materials. Lecture Notes in Computer Science, 7116:515–522, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29843-1_58.

Anonymous:2012:FMdh

[1712] Anonymous. Front matter. Lecture Notes in Computer Science, 7116: 523, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29843-1/12/1.

Anonymous:2012:FMdi

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29843-1/12/1.

Andreev:2012:PEI

[1714] A. B. Andreev and M. R. Racheva. Properties and estimates of an integral type nonconforming finite element. Lecture Notes in Computer Science, 7116:525–532, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29843-1_59.

Andreev:2012:QFE

[1715] A. B. Andreev and M. R. Racheva. Quadratic finite element approximation of a contact eigenvalue problem. Lecture Notes in Computer Science, 7116:533–540, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29843-1_60.

Blaheta:2012:OMC

[1716] Radim Blaheta, Rostislav Hrtus, Roman Kohut, and Ondřej Jakl. Optimization methods for calibration of heat conduction models. Lecture Notes in Computer Science, 7116:541–548, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29843-1_61.

Boyanova:2012:BPC

[1717] P. Boyanova, M. Do-Quang, and M. Neytcheva. Block-preconditioners for conforming and non-conforming FEM discretizations of the Cahn– Hilliard equation. Lecture Notes in Computer Science, 7116:549–557, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29843-1_62.

Kandilarov:2012:CTN

[1718] J. D. Kandilarov and D. Ševčovič. Comparison of two numerical methods for computation of American type of the floating strike Asian option. Lecture Notes in Computer Science, 7116:558–565, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29843-1_63.

Koleva:2012:KBA

[1719] Miglena N. Koleva and Lubin G. Vulkov. A kernel-based algorithm for numerical solution of nonlinear PDEs in finance. Lecture Notes in Computer Science, 7116:566–573, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29843-1_64.

Kosturski:2012:IEP

[1720] N. Kosturski, S. Margenov, and Y. Vutov. Improving the efficiency of parallel FEM simulations on voxel domains. Lecture Notes in Computer Science, 7116:574–581, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29843-1_65.

Kraus:2012:RTL

[1721] J. Kraus, M. Lymbery, and S. Margenov. On the robustness of two-level preconditioners for quadratic FE orthotropic elliptic problems. Lecture Notes in Computer Science, 7116:582–589, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29843-1_66.

Liolios:2012:CAE

[1722] Angelos Liolios, Konstantinos Chalioris, Asterios Liolios, and Stefan Radev. A computational approach for the earthquake response of cablebraced reinforced concrete structures under environmental actions. Lecture Notes in Computer Science, 7116:590–597, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29843-1_67.

Marinov:2012:IPS

[1723] Tchavdar T. Marinov and Rossitza Marinova. An inverse problem for the stationary Kirchhoff equation. Lecture Notes in Computer Science, 7116:598–605, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29843-1_68.

Martone:2012:ISM

[1724] Michele Martone, Marcin Paprzycki, and Salvatore Filippone. An improved sparse matrix-vector multiply based on recursive sparse blocks layout. Lecture Notes in Computer Science, 7116:606–613, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29843-1_69.

Mincsovics:2012:DDW

[1725] Miklós E. Mincsovics and Tamás L. Horváth. On the differences of the discrete weak and strong maximum principles for elliptic operators. Lecture Notes in Computer Science, 7116:614–621, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29843-1_70.

Olas:2012:AFP

[1726] Tomasz Olas and Roman Wyrzykowski. Adaptive FEM package with decentralized parallel adaptation of tetrahedral meshes. Lecture Notes in Computer Science, 7116:622–629, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29843-1_71.

Osintsev:2012:EST

[1727] D. Osintsev, A. Makarov, V. Sverdlov, and S. Selberherr. Efficient simulations of the transport properties of spin field-effect transistors built on silicon fins. Lecture Notes in Computer Science, 7116:630–637, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29843-1_72.

Tashev:2012:LSS

[1728] Tasho Tashev and Vladimir Monov. Large-scale simulation of uniform load traffic for modeling of throughput on a crossbar switch node. Lecture Notes in Computer Science, 7116:638–645, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29843-1_73.

Trapeznikova:2012:TPP

[1729] Marina Trapeznikova, Boris Chetverushkin, Natalia Churbanova, and Dmitrii Morozov. Two-phase porous media flow simulation on a hybrid cluster. Lecture Notes in Computer Science, 7116:646–653, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29843-1_74.

Valkov:2012:PGA

[1730] R. L. Valkov. Petrov–Galerkin analysis for a degenerate parabolic equation in zero-coupon bond pricing. Lecture Notes in Computer Science, 7116:654–661, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29843-1_75.

Wasielewska:2012:AGS

[1731] Katarzyna Wasielewska, Michał Drozdowicz, Paweł Szmeja, Maria Ganzha, and Marcin Paprzycki. Agents in Grid system — design and implementation. Lecture Notes in Computer Science, 7116:662–669, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29843-1_76.

Wyrzykowski:2012:UBG

[1732] Roman Wyrzykowski, Krzysztof Rojek, and Łukasz Szustak. Using Blue Gene/P and GPUs to accelerate computations in the EULAG model. Lecture Notes in Computer Science, 7116:670–677, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29843-1_77.

Anonymous:2012:BMca

[1733] Anonymous. Back matter. Lecture Notes in Computer Science, 7116: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29843-1/1.

Anonymous:2012:FMdj

[1734] Anonymous. Front matter. Lecture Notes in Computer Science, 7116: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29843-1/1.

Celino:2012:LKS

[1735] Irene Celino and Daniele Dell’Aglio. Linking knowledge for simulation learning. Lecture Notes in Computer Science, 7117:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25953-1_1/.

Foulonneau:2012:GEA

[1736] Muriel Foulonneau. Generating educational assessment items from linked open data: The case of DBpedia. Lecture Notes in Computer Science, 7117:16–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25953-1_2/.

Robinson:2012:ULD

[1737] Julien Robinson, Johann Stan, and Myriam Ribière. Using linked data to reduce learning latency for e-book readers. Lecture Notes in Computer Science, 7117:28–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25953-1_3/.

Zablith:2012:OLO

[1738] Fouad Zablith, Miriam Fernandez, and Matthew Rowe. The OU linked open data: Production and consumption. Lecture Notes in Computer Science, 7117:35–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25953-1_4/.

Przyjaciel-Zablocki:2012:RPQ

[1739] Martin Przyjaciel-Zablocki, Alexander Schätzle, Thomas Hornung, and Georg Lausen. RDFPath: Path query processing on large RDF graphs with MapReduce. Lecture Notes in Computer Science, 7117:50–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25953-1_5/.

Moore:2012:NMI

[1740] Joshua L. Moore, Florian Steinke, and Volker Tresp. A novel metric for information retrieval in semantic networks. Lecture Notes in Computer Science, 7117:65–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25953-1_6/.

Celino:2012:TBU

[1741] Irene Celino, Daniele Dell’Aglio, Emanuele Della Valle, Yi Huang, Tony Lee, and Seon-Ho Kim. Towards BOTTARI: Using stream reasoning to make sense of location-based micro-posts. Lecture Notes in Computer Science, 7117:80–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25953-1_7/.

Maynard:2012:ADP

[1742] Diana Maynard and Adam Funk. Automatic detection of political opinions in tweets. Lecture Notes in Computer Science, 7117:88–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25953-1_8/.

Skilters:2012:PPM

[1743] Jurǵis Šķilters, Monika Kreile, Uldis Bojārs, Inta Brikše, Jānis Pencis, and Laura Uzule. The pragmatics of political messages in Twitter communication. Lecture Notes in Computer Science, 7117:100–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25953-1_9/.

Mazumdar:2012:KDM

[1744] Suvodeep Mazumdar, Andrea Varga, Vita Lanfranchi, Daniela Petrelli, and Fabio Ciravegna. A knowledge dashboard for manufacturing industries. Lecture Notes in Computer Science, 7117:112–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25953-1_10/.

Damljanovic:2012:FIW

[1745] Danica Damljanovic, Milan Agatonovic, and Hamish Cunningham. FREyA: An interactive way of querying linked data using natural language. Lecture Notes in Computer Science, 7117:125–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25953-1_11/.

Beltran:2012:OBU

[1746] Victoria Beltran, Knarig Arabshian, and Henning Schulzrinne. Ontologybased user-defined rules and context-aware service composition system. Lecture Notes in Computer Science, 7117:139–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25953-1_12/.

Damljanovic:2012:RIF

[1747] Danica Damljanovic, Johann Petrak, Mihai Lupu, Hamish Cunningham, and Mats Carlsson. Random indexing for finding similar nodes within large RDF graphs. Lecture Notes in Computer Science, 7117:156–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25953-1_13/.

Dey:2012:RPP

[1748] Saumen Dey, Daniel Zinn, and Bertram Ludäscher. Reconciling provenance policy conflicts by inventing anonymous nodes. Lecture Notes in Computer Science, 7117:172–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25953-1_14/.

Stajner:2012:IRR

[1749] Tadej Štajner, Dunja Mladenić, and Marko Grobelnik. Information resource recommendation in knowledge processes. Lecture Notes in Computer Science, 7117:186–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25953-1_15/.

Verginadis:2012:SAR

[1750] Yiannis Verginadis, Ioannis Patiniotakis, Nikos Papageorgiou, and Roland Stuehmer. Service adaptation recommender in the event marketplace: Conceptual view. Lecture Notes in Computer Science, 7117:194– 201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25953-1_16/.

Cirlanaru:2012:APU

[1751] Mihai Cîrlănaru, Deyan Ginev, and Christoph Lange. Authoring and publishing units and quantities in semantic documents. Lecture Notes in Computer Science, 7117:202–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25953-1_17/.

Boero:2012:PIE

[1752] Riccardo Boero, Enrico Ferro, Michele Osella, Yannis Charalabidis, and Euripidis Loukis. Policy intelligence in the era of social computing: Towards a cross-policy decision support system. Lecture Notes in Computer Science, 7117:217–228, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25953-1_18/.

Hypponen:2012:XRS

[1753] Konstantin Hyppönen, Miika Alonen, Sami Korhonen, and Virpi Hotti. XHTML with RDFa as a semantic document format for CCTS modelled documents and its application for social services. Lecture Notes in Computer Science, 7117:229–240, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25953-1_19/.

Tsarouchis:2012:BSE

[1754] Christos Tsarouchis, Declan O’Sullivan, and David Lewis. Balancing system expressivity and user cognitive load in semantically enhanced policy modelling. Lecture Notes in Computer Science, 7117:241–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25953-1_20/.

Cano:2012:SPP

[1755] Amparo-Elizabeth Cano, Aba-Sah Dadzie, Victoria Uren, and Fabio Ciravegna. Sensing presence (PreSense) ontology: User modelling in the Semantic Sensor Web. Lecture Notes in Computer Science, 7117:253–268, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
25953-1_21.

Tao:2012:TTB

[1756] Ke Tao, Fabian Abel, Qi Gao, and Geert-Jan Houben. TUMS: Twitterbased user modeling service. Lecture Notes in Computer Science, 7117: 269–283, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-25953-1_22.

Anonymous:2012:BMcb

[1757] Anonymous. Back matter. Lecture Notes in Computer Science, 7117: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-25953-1/1.

Anonymous:2012:FMdk

[1758] Anonymous. Front matter. Lecture Notes in Computer Science, 7117: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-25953-1/1.

Basu:2012:RSD

[1759] Ananda Basu, Saddek Bensalem, and Marius Bozga. Rigorous system design: The BIP approach. Lecture Notes in Computer Science, 7119: 1–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25929-6_1/.

Bentley:2012:NBC

[1760] Peter J. Bentley. Natural Born computing. Lecture Notes in Computer Science, 7119:20–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25929-6_2/.

Chatterjee:2012:GMD

[1761] Krishnendu Chatterjee and Laurent Doyen. Games and Markov decision processes with mean-payoff parity and energy parity objectives. Lecture Notes in Computer Science, 7119:37–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-25929-6_3/.

Fey:2012:ASV

[1762] Görschwin Fey. Assessing system vulnerability using formal verification techniques. Lecture Notes in Computer Science, 7119:47–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25929-6_4/.

Renner:2012:ISQ

[1763] Renato Renner. Information security in a quantum world. Lecture Notes in Computer Science, 7119:57–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25929-6_5/.

Babka:2012:CMW

[1764] Vlastimil Babka and Petr Tůma. Computer memory: Why we should care what is under the hood. Lecture Notes in Computer Science, 7119: 63–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25929-6_6/.

Balodis:2012:FPF

[1765] Kaspars Balodis and Ilja Kucevalovs. Frequency prediction of functions. Lecture Notes in Computer Science, 7119:76–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25929-6_7/.

Barnat:2012:TAA

[1766] Jiří Barnat, Ivana Černá, and Jana Tůmová. Timed automata approach to verification of systems with degradation. Lecture Notes in Computer Science, 7119:84–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25929-6_8/.

Cmorik:2012:BOB

[1767] Roland Cmorik and Galina Jirásková. Basic operations on binary suffix-free languages. Lecture Notes in Computer Science, 7119:94–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25929-6_9/.

Klusacek:2012:EDR

[1768] Dalibor Klusáček and Hana Rudová. Efficient data representation of large job schedules. Lecture Notes in Computer Science, 7119:103–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25929-6_10/.

Krausova:2012:PFR

[1769] Monika Krausová. Prefix-free regular languages: Closure properties, difference, and left quotient. Lecture Notes in Computer Science, 7119:114– 122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25929-6_11/.

Krena:2012:NIH

[1770] Bohuslav Křena, Zdeněk Letko, and Tomáš Vojnar. Noise injection heuristics for concurrency testing. Lecture Notes in Computer Science, 7119:123–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25929-6_12/.

Matela:2012:LGO

[1771] Jiří Matela, Martin Šrom, and Petr Holub. Low GPU occupancy approach to fast arithmetic coding in JPEG2000. Lecture Notes in Computer Science, 7119:136–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25929-6_13/.

Mazgut:2012:UDR

[1772] Jakub Mažgut, Martina Paulinyová, and Peter Tiňo. Using dimensionality reduction method for binary data to questionnaire analysis. Lecture Notes in Computer Science, 7119:146–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-25929-6_14/.

Hlineny:2012:GMR

[1773] Petr Hliněný and Ondrej Moriš. Generalized maneuvers in route planning. Lecture Notes in Computer Science, 7119:155–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25929-6_15/.

Obdrzalek:2012:SBF

[1774] Jan Obdržálek, Jiří Slabý, and Marek Trtík. STANSE: Bug-finding framework for C programs. Lecture Notes in Computer Science, 7119:167–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25929-6_16/.

Sakellariou:2012:IFB

[1775] Christos Sakellariou and Peter J. Bentley. Introducing the FPGA-based hardware architecture of systemic computation (HAoS). Lecture Notes in Computer Science, 7119:179–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25929-6_17/.

vanderVegt:2012:PCH

[1776] Steven van der Vegt and Alfons Laarman. A parallel compact hash table. Lecture Notes in Computer Science, 7119:191–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-25929-6_18/.

Zilka:2012:FAP

[1777] Roman Žilka, Vashek Matyáš, and Libor Kyncl. Four authorization protocols for an electronic payment system. Lecture Notes in Computer Science, 7119:205–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25929-6_19/.

Anonymous:2012:BMcc

[1778] Anonymous. Back matter. Lecture Notes in Computer Science, 7119: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-25929-6/1.

Anonymous:2012:FMdl

[1779] Anonymous. Front matter. Lecture Notes in Computer Science, 7119: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-25929-6/1.

Avoine:2012:PCR

[1780] Gildas Avoine. Privacy challenges in RFID. Lecture Notes in Computer Science, 7122:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28879-1_1/.

Karakasidis:2012:FIS

[1781] Alexandros Karakasidis, Vassilios S. Verykios, and Peter Christen. Fake injection strategies for private phonetic matching. Lecture Notes in Computer Science, 7122:9–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28879-1_2/.

Matteucci:2012:DPD

[1782] Ilaria Matteucci, Marinella Petrocchi, Marco Luca Sbodio, and Luca Wiegand. A design phase for data sharing agreements. Lecture Notes in Computer Science, 7122:25–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28879-1_3/.

Parra-Arnau:2012:PPA

[1783] Javier Parra-Arnau, David Rebollo-Monedero, and Jordi Forné. A privacy-protecting architecture for collaborative filtering via forgery and suppression of ratings. Lecture Notes in Computer Science, 7122:42– 57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28879-1_4/.

Duncan:2012:CAI

[1784] Christian A. Duncan and Vir V. Phoha. On the complexity of aggregating information for authentication and profiling. Lecture Notes in Computer Science, 7122:58–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28879-1_5/.

Sen:2012:SPA

[1785] Jaydip Sen. Secure and privacy-aware searching in peer-to-peer networks. Lecture Notes in Computer Science, 7122:72–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28879-1_6/.

Safkhani:2012:SMA

[1786] Masoumeh Safkhani, Nasour Bagheri, Somitra Kumar Sanadhya, Majid Naderi, and Hamid Behnam. On the security of mutual authentication protocols for RFID systems: The case of Wei et al. ’s protocol. Lecture Notes in Computer Science, 7122:90–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28879-1_7/.

Biskup:2012:IPV

[1787] Joachim Biskup and Cornelia Tadros. Inference-proof view update transactions with minimal refusals. Lecture Notes in Computer Science, 7122:104–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28879-1_8/.

Pretschner:2012:RID

[1788] Alexander Pretschner, Enrico Lovat, and Matthias Büchler. Representationindependent data usage control. Lecture Notes in Computer Science, 7122:122–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28879-1_9/.

Tasidou:2012:UPP

[1789] Aimilia Tasidou and Pavlos S. Efraimidis. Using personal portfolios to manage customer data. Lecture Notes in Computer Science, 7122:141– 154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28879-1_10/.

Graa:2012:URE

[1790] Mariem Graa, Nora Cuppens-Boulahia, Fabien Autrel, Hanieh Azkia, and Frédéric Cuppens. Using requirements engineering in an automatic security policy derivation process. Lecture Notes in Computer Science, 7122:155–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28879-1_11/.

Avanesov:2012:WSV

[1791] Tigran Avanesov, Yannick Chevalier, Mohammed Anis Mekki, and Michaël Rusinowitch. Web services verification and prudent implementation. Lecture Notes in Computer Science, 7122:173–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28879-1_12/.

Idrees:2012:ESR

[1792] Muhammad Sabir Idrees, Gabriel Serme, Yves Roudier, and Anderson Santana De Oliveira. Evolving security requirements in multilayered service-oriented-architectures. Lecture Notes in Computer Science, 7122:190–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28879-1_13/.

Krautsevich:2012:RBA

[1793] Leanid Krautsevich, Fabio Martinelli, Charles Morisset, and Artsiom Yautsiukhin. Risk-based auto-delegation for probabilistic availability. Lecture Notes in Computer Science, 7122:206–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28879-1_14/.

Martinez-Garcia:2012:IRP

[1794] Carles Martínez-García, Guillermo Navarro-Arribas, and Joan Borrell. Intra-role progression in RBAC: An RPG-like access control scheme. Lecture Notes in Computer Science, 7122:221–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28879-1_15/.

Avanesov:2012:DOW

[1795] Tigran Avanesov, Yannick Chevalier, Mohammed Anis Mekki, and Michaël Rusinowitch. Distributed orchestration of Web services under security constraints. Lecture Notes in Computer Science, 7122:235–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28879-1_16/.

Hernandez-Castro:2012:KSS

[1796] Julio Cesar Hernandez-Castro, Pedro Peris-Lopez, and Jean-Philippe Aumasson. On the key schedule strength of PRESENT. Lecture Notes in Computer Science, 7122:253–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28879-1_17/.

Pujol-Gil:2012:TRM

[1797] Enric Pujol-Gil and Nikolaos Chatzis. A traffic regulation method based on MRA signatures to reduce unwanted traffic from compromised end-user machines. Lecture Notes in Computer Science, 7122:264–279, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28879-1_18/.

Fu:2012:NSA

[1798] Yulong Fu and Ousmane Kone. Network securing against threatening requests. Lecture Notes in Computer Science, 7122:280–294, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28879-1_19/.

Koukovini:2012:WCA

[1799] Maria N. Koukovini, Eugenia I. Papagiannakopoulou, and Georgios V. Lioudakis. A workflow checking approach for inherent privacy awareness in network monitoring. Lecture Notes in Computer Science, 7122:295– 302, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28879-1_20/.

Janicke:2012:CDD

[1800] Helge Janicke, Mohamed Sarrab, and Hamza Aldabbas. Controlling data dissemination. Lecture Notes in Computer Science, 7122:303–309, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
28879-1_21.

Hennebert:2012:FDS

[1801] Christine Hennebert and Vincent Berg. A framework of deployment strategy for hierarchical WSN security management. Lecture Notes in Computer Science, 7122:310–318, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-28879-1_22.

Anonymous:2012:BMcd

[1802] Anonymous. Back matter. Lecture Notes in Computer Science, 7122: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28879-1/1.

Anonymous:2012:FMdm

[1803] Anonymous. Front matter. Lecture Notes in Computer Science, 7122: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28879-1/1.

Grimaldo:2012:ASP

[1804] Francisco Grimaldo, Mario Paolucci, and Rosaria Conte. Agent simulation of peer review: The PR-1 model. Lecture Notes in Computer Science, 7124:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28400-7_1/.

Osman:2012:SRB

[1805] Nardine Osman, Jordi Sabater-Mir, Carles Sierra, and Jordi Madrenas-Ciurana. Simulating research behaviour. Lecture Notes in Computer Science, 7124:15–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28400-7_2/.

Yu:2012:ABM

[1806] Tongkui Yu and Shu-Heng Chen. Agent-based modeling of the prediction markets for political elections. Lecture Notes in Computer Science, 7124: 31–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28400-7_3/.

Vanek:2012:UMA

[1807] Ondřej Vaněk, Michal Jakob, Ondřej Hrstka, and Michal Pěchouček. Using multi-agent simulation to improve the security of maritime transit. Lecture Notes in Computer Science, 7124:44–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28400-7_4/.

Gaudou:2012:HDS

[1808] Benoit Gaudou, Andreas Herzig, Emiliano Lorini, and Christophe Sibertin-Blanc. How to do social simulation in logic: Modelling the segregation game in a dynamic logic of assignments. Lecture Notes in Computer Science, 7124:59–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28400-7_5/.

Manenti:2012:ABP

[1809] Lorenza Manenti, Sara Manzoni, Giuseppe Vizzari, Kazumichi Ohtsuka, and Kenichiro Shimura. An agent-based proxemic model for pedestrian and group dynamics: Motivations and first experiments. Lecture Notes in Computer Science, 7124:74–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28400-7_6/.

Xing:2012:VAB

[1810] Pengfei Xing, Michael Lees, Hu Nan, and T. Vaisagh Viswanthatn. Validation of agent-based simulation through human computation: An example of crowd simulation. Lecture Notes in Computer Science, 7124:90– 102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28400-7_7/.

Morvan:2012:OLS

[1811] Gildas Morvan, Alexandre Veremme, and Daniel Dupont. Observation of large-scale multi-agent based simulations. Lecture Notes in Computer Science, 7124:103–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28400-7_8/.

Parunak:2012:BAM

[1812] H. Van Dyke Parunak. Between agents and mean fields. Lecture Notes in Computer Science, 7124:113–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28400-7_9/.

Anonymous:2012:BMce

[1813] Anonymous. Back matter. Lecture Notes in Computer Science, 7124: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28400-7/1.

Anonymous:2012:FMdn

[1814] Anonymous. Front matter. Lecture Notes in Computer Science, 7124: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28400-7/1.

MacKinnon:2012:DSS

[1815] Lachlan M. MacKinnon, editor. Data Security and Security Data: 27th British National Conference on Databases, BNCOD 27, Dundee, UK, June 29–July 1, 2010. Revised Selected Papers, volume 6121 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25703-8 (print), 3-642-25704-6 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-25704-9.

Varbanescu:2012:CAI

[1816] Ana Lucia Varbanescu, Anca Molnos, and Rob van Nieuwpoort, editors. Computer Architecture: ISCA 2010 International Workshops A4MMC, AMAS-BT, EAMA, WEED, WIOSCA, Saint-Malo, France, June 19– 23, 2010, Revised Selected Papers, volume 6161 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-24321-5 (print), 3-642-24322-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-24322-6.

Kannan:2012:DEM

[1817] Rajkumar Kannan and Frederic Andres, editors. Data Engineering and Management: Second International Conference, ICDEM 2010, Tiruchirappalli, India, July 29–31, 2010. Revised Selected Papers, volume 6411 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27871-X (print), 3-642-27872-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-
642-27872-3.

Horimoto:2012:ANB

[1818] Katsuhisa Horimoto, Masahiko Nakatsui, and Nikolaj Popov, editors. Algebraic and Numeric Biology: 4th International Conference, ANB 2010, Hagenberg, Austria, July 31- August 2, 2010, Revised Selected Papers, volume 6479 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28066-8 (print), 3-642-28067-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-28067-2.

Kutulakos:2012:TTCa

[1819] Kiriakos N. Kutulakos, editor. Trends and Topics in Computer Vision: ECCV 2010 Workshops, Heraklion, Crete, Greece, September 10–11, 2010, Revised Selected Papers, Part I, volume 6553 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35748-2 (print), 3-642-35749-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-35749-7.

Kutulakos:2012:TTCb

[1820] Kiriakos N. Kutulakos, editor. Trends and Topics in Computer Vision: ECCV 2010 Workshops, Heraklion, Crete, Greece, September 10–11, 2010, Revised Selected Papers, Part II, volume 6554 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35739-3 (print), 3-642-35740-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-35740-4.

Beneventano:2012:APP

[1821] Domenico Beneventano, Zoran Despotovic, Francesco Guerra, Sam Joseph, Gianluca Moro, and Adrián Perreau de Pinninck, editors. Agents and Peer-to-Peer Computing: 7th International Workshop, AP2PC 2008, Estoril, Portugal, May 13, 2008 and 8th International Workshop, AP2PC 2009, Budapest, Hungary, May 11, 2009. Revised Selected Papers, volume 6573 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31808-8 (print), 3-642-31809-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-31809-2.

Collier:2012:PMA

[1822] Rem Collier, Jürgen Dix, and Peter Novák, editors. Programming Multi-Agent Systems: 8th International Workshop, ProMAS 2010, Toronto, ON, Canada, May 11, 2010. Revised Selected Papers, volume 6599 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28938-X (print), 3-642-28939-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-
642-28939-2.

Freitas:2012:BPM

[1823] Ana T. Freitas and Arcadi Navarro, editors. Bioinformatics for Personalized Medicine: 10th Spanish Symposium, JBI 2010, Torremolinos, Spain, October 27–29, 2010. Revised Selected Papers, volume 6620 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28061-7 (print), 3-642-28062-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-28062-7.

Bruckstein:2012:SSV

[1824] Alfred M. Bruckstein, Bart M. ter Haar Romeny, Alexander M. Bronstein, and Michael M. Bronstein, editors. Scale Space and Variational Methods in Computer Vision: Third International Conference, SSVM 2011, Ein-Gedi, Israel, May 29–June 2, 2011, Revised Selected Papers, volume 6667 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-24784-9 (print), 3-642-24785-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-24785-9.

Lacroix:2012:RDT

[1825] Zoé Lacroix and María Esther Vidal, editors. Resource Discovery: Third International Workshop, RED 2010, Paris, France, November 5, 2010, Revised Selected Papers, volume 6799 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27391-2 (print), 3-642-27392-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-27392-6.

Huang:2012:AICa

[1826] De-Shuang Huang, Yong Gan, Vitoantonio Bevilacqua, and Juan Carlos Figueroa, editors. Advanced Intelligent Computing: 7th International Conference, ICIC 2011, Zhengzhou, China, August 11–14, 2011. Revised Selected Papers, volume 6838 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-24727-X (print), 3-642-24728-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-24728-6.

Huang:2012:AICb

[1827] De-Shuang Huang, Yong Gan, Phalguni Gupta, and M. Michael Gromiha, editors. Advanced Intelligent Computing Theories and Applications. With Aspects of Artificial Intelligence: 7th International Conference, ICIC 2011, Zhengzhou, China, August 11–14, 2011, Revised Selected Papers, volume 6839 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25943-X (print), 3-642-25944-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-25944-9.

Huang:2012:BIC

[1828] De-Shuang Huang, Yong Gan, Prashan Premaratne, and Kyungsook Han, editors. Bio-Inspired Computing and Applications: 7th International Conference on Intelligent Computing, ICIC 2011, Zhengzhou,China, August 11–14. 2011, Revised Selected Papers, volume 6840 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-24552-8 (print), 3-642-24553-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-24553-4.

Moot:2012:LCGb

[1829] Richard Moot and Christian Retoré, editors. The Logic of Categorial Grammars: A Deductive Account of Natural Language Syntax and Semantics, volume 6850 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31554-2 (print), 3-642-31555-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). 297 (est.) pp. LCCN ???? URL http://www.

springerlink.com/content/978-3-642-31555-8.

Garcia-Alfaro:2012:FPS

[1830] Joaquin Garcia-Alfaro and Pascal Lafourcade, editors. Foundations and Practice of Security: 4th Canada-France MITACS Workshop, FPS 2011, Paris, France, May 12–13, 2011, Revised Selected Papers, volume 6888 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27900-7 (print), 3-642-27901-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-27901-0.

Jensen:2012:TPN

[1831] Kurt Jensen, Susanna Donatelli, and Jetty Kleijn, editors. Transactions on Petri Nets and Other Models of Concurrency V, volume 6900 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29071-X (print), 3-642-29072-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-
642-29072-5.

Boissonnat:2012:CSI

[1832] Jean-Daniel Boissonnat, Patrick Chenin, Albert Cohen, Christian Gout, Tom Lyche, Marie-Laurence Mazure, and Larry Schumaker, editors. Curves and Surfaces: 7th International Conference, Avignon, France, June 24–30, 2010, Revised Selected Papers, volume 6920 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27412-9 (print), 3-642-27413-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-27413-8.

Barbosa:2012:FAC

[1833] Luís Soares Barbosa and Markus Lumpe, editors. Formal Aspects of Component Software: 7th International Workshop, FACS 2010, Guimarães, Portugal, October 14–16, 2010, Revised Selected Papers, volume 6921 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27268-1 (print), 3-642-27269-X (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-27269-1.

Riano:2012:KRH

[1834] David Riaño, Annette ten Teije, and Silvia Miksch, editors. Knowledge Representation for Health-Care: AIME 2011 Workshop KR4HC 2011, Bled, Slovenia, July 2–6, 2011. Revised Selected Papers, volume 6924 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27696-2 (print), 3-642-27697-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-27697-2.

Moreno-Diaz:2012:CASa

[1835] Roberto Moreno-Díaz, Franz Pichler, and Alexis Quesada-Arencibia, editors. Computer Aided Systems Theory — EUROCAST 2011: 13th International Conference, Las Palmas de Gran Canaria, Spain, February 6–11, 2011, Revised Selected Papers, Part I, volume 6927 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27548-6 (print), 3-642-27549-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-27549-4.

Moreno-Diaz:2012:CASb

[1836] Roberto Moreno-Díaz, Franz Pichler, and Alexis Quesada-Arencibia, editors. Computer Aided Systems Theory — EUROCAST 2011: 13th International Conference, Las Palmas de Gran Canaria, Spain, February 6–11, 2011, Revised Selected Papers, Part II, volume 6928 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27578-8 (print), 3-642-27579-6 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-27579-1.

Sloane:2012:SLE

[1837] Anthony Sloane and Uwe Aßmann, editors. Software Language Engineering: 4th International Conference, SLE 2011, Braga, Portugal, July 3–4, 2011, Revised Selected Papers, volume 6940 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28829-4 (print), 3-642-28830-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-28830-2.

Aichernig:2012:FMC

[1838] Bernhard K. Aichernig, Frank S. de Boer, and Marcello M. Bonsangue, editors. Formal Methods for Components and Objects: 9th International Symposium, FMCO 2010, Graz, Austria, November 29 — December 1, 2010. Revised Papers, volume 6957 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25270-2 (print), 3-642-25271-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-25271-6.

Modersheim:2012:TSA

[1839] Sebastian Mödersheim and Catuscia Palamidessi, editors. Theory of Security and Applications: Joint Workshop, TOSCA 2011, Saarbrücken, Germany, March 31–April 1, 2011, Revised Selected Papers, volume 6993 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27374-2 (print), 3-642-27375-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-
642-27375-9.

Meyer:2012:ESE

[1840] Bertrand Meyer and Martin Nordio, editors. Empirical Software Engineering and Verification: International Summer Schools, LASER 2008-2010, Elba Island, Italy, Revised Tutorial Lectures, volume 7007 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25230-3 (print), 3-642-25231-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-25231-0.

vanEijck:2012:GAS

[1841] Jan van Eijck and Rineke Verbrugge, editors. Games, Actions and Social Software: Multidisciplinary Aspects, volume 7010 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29325-5 (print), 3-642-29326-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-29326-9.

Yoshida:2012:AIC

[1842] Hiroyuki Yoshida, Georgios Sakas, and Marius George Linguraru, editors. Abdominal Imaging. Computational and Clinical Applications: Third International Workshop, Held in Conjunction with MICCAI 2011, Toronto, ON, Canada, September 18, 2011, Revised Selected Papers, volume 7029 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28556-2 (print), 3-642-28557-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-28557-8.

vanKreveld:2012:GDI

[1843] Marc van Kreveld and Bettina Speckmann, editors. Graph Drawing: 19th International Symposium, GD 2011, Eindhoven, The Netherlands, September 21–23, 2011, Revised Selected Papers, volume 7034 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25877-8 (print), 3-642-25878-6 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-25878-7.

Camenisch:2012:OPN

[1844] Jan Camenisch and Dogan Kesdogan, editors. Open Problems in Network Security: IFIP WG 11.4 International Workshop, iNetSec 2011, Lucerne, Switzerland, June 9, 2011, Revised Selected Papers, volume 7039 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27584-2 (print), 3-642-27585-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-
642-27585-2.

Gavrilova:2012:TCS

[1845] Marina L. Gavrilova, C. J. Kenneth Tan, and Cong-Vinh Phan, editors. Transactions on Computational Science XV: Special Issue on Advances in Autonomic Computing: Formal Engineering Methods for Nature-Inspired Computing Systems, volume 7050 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28524-4 (print), 3-642-28525-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-28525-7.

Bouvry:2012:SII

[1846] Pascal Bouvry, Mieczysław A. Kłopotek, Franck Leprévost, Małgorzata Marciniak, Agnieszka Mykowiecka, and Henryk Rybiński, editors. Security and Intelligent Information Systems: International Joint Conferences, SIIS 2011, Warsaw, Poland, June 13–14, 2011, Revised Selected Papers, volume 7053 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25260-5 (print), 3-642-25261-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-25261-7.

Szmuc:2012:ASE

[1847] Tomasz Szmuc, Marcin Szpyrka, and Jaroslav Zendulka, editors. Advances in Software Engineering Techniques: 4th IFIP TC 2 Central and East European Conference on Software Engineering Techniques, CEE-SET 2009, Krakow, Poland, October 12–14, 2009. Revised Selected Papers, volume 7054 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28037-4 (print), 3-642-28038-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-28038-2.

Juels:2012:RSP

[1848] Ari Juels and Christof Paar, editors. RFID. Security and Privacy: 7th International Workshop, RFIDSec 2011, Amherst, USA, June 26–28, 2011, Revised Selected Papers, volume 7055 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25285-0 (print), 3-642-25286-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-25286-0.

Desai:2012:PPM

[1849] Nirmit Desai, Alan Liu, and Michael Winikoff, editors. Principles and Practice of Multi-Agent Systems: 13th International Conference, PRIMA 2010, Kolkata, India, November 12–15, 2010, Revised Selected Papers, volume 7057 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25919-7 (print), 3-642-25920-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-25920-3.

Harth:2012:CTW

[1850] Andreas Harth and Nora Koch, editors. Current Trends in Web Engineering: Workshops, Doctoral Symposium, and Tutorials, Held at ICWE 2011, Paphos, Cyprus, June 20–21, 2011. Revised Selected Papers, volume 7059 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27996-1 (print), 3-642-27997-X (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-27997-3.

Dechesne:2012:AAT

[1851] Francien Dechesne, Hiromitsu Hattori, Adriaan ter Mors, Jose Miguel Such, Danny Weyns, and Frank Dignum, editors. Advanced Agent Technology: AAMAS 2011 Workshops, AMPLE, AOSE, ARMS, DOCM 3 AS, ITMAS, Taipei, Taiwan, May 2–6, 2011. Revised Selected Papers, volume 7068 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27215-0 (print), 3-642-27216-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-27216-5.

Muller:2012:MCB

[1852] Henning Müller, Hayit Greenspan, and Tanveer Syeda-Mahmood, editors. Medical Content-Based Retrieval for Clinical Decision Support: Second MICCAI International Workshop, MCBR-CDS 2011, Toronto, ON, Canada, September 22, 2011, Revised Selected Papers, volume 7075 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28459-0 (print), 3-642-28460-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-28460-1.

Schwenker:2012:PSL

[1853] Friedhelm Schwenker and Edmondo Trentin, editors. Partially Supervised Learning: First IAPR TC3 Workshop, PSL 2011, Ulm, Germany, September 15–16, 2011, Revised Selected Papers, volume 7081 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28257-1 (print), 3-642-28258-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-28258-4.

Ober:2012:SIS

[1854] Iulian Ober and Ileana Ober, editors. SDL 2011: Integrating System and Software Modeling: 15th International SDL Forum Toulouse, France, July 5–7, 2011. Revised Papers, volume 7083 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25263-X (print), 3-642-25264-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-25264-8.

Camara:2012:SAC

[1855] Oscar Camara, Ender Konukoglu, Mihaela Pop, Kawal Rhode, Maxime Sermesant, and Alistair Young, editors. Statistical Atlases and Computational Models of the Heart. Imaging and Modelling Challenges: Second International Workshop, STACOM 2011, Held in Conjunction with MICCAI 2011, Toronto, ON, Canada, September 22, 2011, Revised Selected Papers, volume 7085 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28325-X (print), 3-642-28326-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-28326-0.

Ho:2012:AIVa

[1856] Yo-Sung Ho, editor. Advances in Image and Video Technology: 5th Pacific Rim Symposium, PSIVT 2011, Gwangju, South Korea, November 20–23, 2011, Proceedings, Part I, volume 7087 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25366-0 (print), 3-642-25367-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-25367-6.

Ho:2012:AIVb

[1857] Yo-Sung Ho, editor. Advances in Image and Video Technology: 5th Pacific Rim Symposium, PSIVT 2011, Gwangju, South Korea, November 20–23, 2011, Proceedings, Part II, volume 7088 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25345-8 (print), 3-642-25346-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-25346-1.

Rautiainen:2012:GPC

[1858] Mika Rautiainen, Timo Korhonen, Edward Mutafungwa, Eila Ovaska, Artem Katasonov, Antti Evesti, Heikki Ailisto, Aaron Quigley, Jonna Häkkilä, Natasa Milic-Frayling, and Jukka Riekki, editors. Grid and Pervasive Computing Workshops: International Workshops, S3E, HWTS, Doctoral Colloquium, Held in Conjunction with GPC 2011, Oulu, Finland, May 11–13, 2011. Revised Selected Papers, volume 7096 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27915-5 (print), 3-642-27916-3 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-27916-4.

Hameurlain:2012:TLS

[1859] Abdelkader Hameurlain, Josef Küng, and Roland Wagner, editors. Transactions on Large-Scale Data- and Knowledge-Centered Systems V, volume 7100 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28147-8 (print), 3-642-28148-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-28148-8.

Cao:2012:ADM

[1860] Longbing Cao, Ana L. C. Bazzan, Andreas L. Symeonidis, Vladimir I. Gorodetsky, Gerhard Weiss, and Philip S. Yu, editors. Agents and Data Mining Interaction: 7th International Workshop on Agents and Data Mining Interation, ADMI 2011, Taipei, Taiwan, May 2–6, 2011, Revised Selected Papers, volume 7103 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27608-3 (print), 3-642-27609-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-27609-5.

Cao:2012:NFA

[1861] Longbing Cao, Joshua Zhexue Huang, James Bailey, Yun Sing Koh, and Jun Luo, editors. New Frontiers in Applied Data Mining: PAKDD 2011 International Workshops, Shenzhen, China, May 24–27, 2011, Revised Selected Papers, volume 7104 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28319-5 (print), 3-642-28320-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-28320-8.

Shi:2012:TDH

[1862] Yun Q. Shi, editor. Transactions on Data Hiding and Multimedia Security VII, volume 7110 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28692-5 (print), 3-642-28693-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-28693-3.

Erlebach:2012:ASS

[1863] Thomas Erlebach, Sotiris Nikoletseas, and Pekka Orponen, editors. Algorithms for Sensor Systems: 7th International Symposium on Algorithms for Sensor Systems, Wireless Ad Hoc Networks and Autonomous Mobile Entities, ALGOSENSORS 2011, Saarbrücken, Germany, September 8–9, 2011, Revised Selected Papers, volume 7111 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28208-3 (print), 3-642-28209-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-28209-6.

Marx:2012:PEC

[1864] Dániel Marx and Peter Rossmanith, editors. Parameterized and Exact Computation: 6th International Symposium, IPEC 2011, Saarbrücken, Germany, September 6–8, 2011. Revised Selected Papers, volume 7112 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28049-8 (print), 3-642-28050-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-28050-4.

Vrancx:2012:ALA

[1865] Peter Vrancx, Matthew Knudson, and Marek Grześ, editors. Adaptive and Learning Agents: International Workshop, ALA 2011, Held at AAMAS 2011, Taipei, Taiwan, May 2, 2011, Revised Selected Papers, volume 7113 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28498-1 (print), 3-642-28499-X (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-28499-1.

Jung:2012:ISA

[1866] Souhwan Jung and Moti Yung, editors. Information Security Applications: 12th International Workshop, WISA 2011, Jeju Island, Korea, August 22–24, 2011. Revised Selected Papers, volume 7115 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27889-2 (print), 3-642-27890-6 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-27890-7.

Lirkov:2012:LSS

[1867] Ivan Lirkov, Svetozar Margenov, and Jerzy Waśniewski, editors. Large-Scale Scientific Computing: 8th International Conference, LSSC 2011, Sozopol, Bulgaria, June 6–10, 2011, Revised Selected Papers, volume 7116 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29842-7 (print), 3-642-29843-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-
642-29843-1.

Garcia-Castro:2012:SWE

[1868] Raúl García-Castro, Dieter Fensel, and Grigoris Antoniou, editors. The Semantic Web: ESWC 2011 Workshops: ESWC 2011 Workshops, Heraklion, Greece, May 29–30, 2011, Revised Selected Papers, volume 7117 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25952-9 (print), 3-642-25953-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-
642-25953-1.

Kotasek:2012:MEM

[1869] Zdeněk Kotásek, Jan Bouda, Ivana Černá, Lukáš Sekanina, Tomáš Vojnar, and David Antoš, editors. Mathematical and Engineering Methods in Computer Science: 7th International Doctoral Workshop, MEMICS 2011, Lednice, Czech Republic, October 14–16, 2011, Revised Selected Papers, volume 7119 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25928-6 (print), 3-642-25929-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-25929-6.

Garcia-Alfaro:2012:DPM

[1870] Joaquin Garcia-Alfaro, Guillermo Navarro-Arribas, Nora Cuppens-Boulahia, and Sabrina de Capitani di Vimercati, editors. Data Privacy Management and Autonomous Spontaneus Security: 6th International Workshop, DPM 2011, and 4th International Workshop, SETOP 2011, Leuven, Belgium, September 15–16, 2011, Revised Selected Papers, volume 7122 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28878-2 (print), 3-642-28879-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-28879-1.

Villatoro:2012:MAB

[1871] Daniel Villatoro, Jordi Sabater-Mir, and Jaime Simão Sichman, editors. Multi-Agent-Based Simulation XII: International Workshop, MABS 2011, Taipei, Taiwan, May 2–6, 2011, Revised Selected Papers, volume 7124 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28399-3 (print), 3-642-28400-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-
642-28400-7.